A Bibliography of Publications in ACM SIGPLAN Notices, 2000–2009

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.84

Title word cross-reference

2 [1630]. 3 [1092, 1344, 1682, 1695, 248]. < [1230]. > [1230]. ^{F} [2067, 871]. ^{T M} [179, 175, 177]. Δ [1505]. Γ [1615]. k [2088]. λ [877, 960]. m [722]. N [1510, 2293, 522]. Π [1924, 2434]. Π0 [1630]. T [1364].

-calculus [722, 960]. -complete [1630]. -pin [1344]. -queen [522]. -STL [27]. -synchronous [1510]. -valued [248]. -way [2293].

.NET [1236, 2119, 1204, 865, 285, 973, 1348].

/garbage [1919].

’00 [2478, 33, 2480, 118, 119]. ’01 [211].

16-bit [855].

2 [128, 1280]. 2000 [36, 327]. 2001 [59, 58, 537, 274, 60]. 2003 [958, 1122]. 2008 [2060, 2005, 2164]. 21364 [674]. 21st [271]. 28th [2482, 54].

6 [1715]. 64 [233, 364, 204].

’99 [16, 17].

= [534].

abandoned [1032]. ABCD [90]. abduction [2193]. Abstract [1190, 18, 985, 370, 926, 2292, 2362, 2365, 2363, 2442, 499, 2366, 2367, 2295, 2358, 2075, 2443, 19, 1591, 500, 984, 2440, 2359, 408, 744, 1582, 1946, 2297, 2360, 1826, 2446, 2400, 1615, 253, 2441, 2199, 1192, 2364, 1211, 827, 1958, 1590, 266, 2368, 1196, 2444, 2296, 419, 894, 890, 2361, 2445, 755, 2294, 1957, 1200, 2293]. Abstraction [304, 1640, 2438, 854, 501, 489, 876, 1481, 1333, 1932, 1324, 844, 1183, 1262, 2318, 1318, 1626]. abstraction-safe [876]. Abstractions [303, 989, 1741, 1584, 1071, 1785, 1786, 2180, 1456, 1337, 1069]. abstracts [1534, 2060, 2005, 1608, 2164]. Accelerating [1970, 2416, 2241]. Accelerator [1701]. accelerators [2275, 2263, 2350]. Acceptability [968]. Acceptability-oriented [968]. Access [1638, 704, 1973, 499, 68, 2045, 743, 888, 1675, 635, 1036]. accessed [78]. accesses [76]. accessing [667]. accidents [148]. accommodate [1549]. accommodating [920]. account [1936]. accounting [1334, 2231]. accumulating [1635, 641]. Accuracy [222, 816, 1376]. Accurate [1979, 772, 1688, 336, 1572, 1687, 1793, 2050, 346, 358]. Accurately [1, 2, 1041, 1039, 1040]. achieving [1386]. Ackermann [931]. ACM [54, 539, 319, 30, 52, 1486, 1539, 2485, 824, 2480]. ACM-SIGACT [54]. ACM/SIGPLAN [2485]. acme [1155, 1209, 1402]. acoustic [2072]. acquisition [1314]. Across [105, 341, 1481, 1883, 1091]. action [1062]. actions [2052, 2169]. ActionScript [2359]. Activation [205]. active [2363, 677, 768]. Activities [317]. Acute [1427]. ad [1392, 1962, 2382, 1388, 1733]. ad-hoc [1388]. Ada [1003, 1001]. ADAPT [340]. adaptability [599]. Adaptable [1345, 107, 625, 703]. adaptation [563, 337, 1643]. adaptations [668]. Adapters [27, 1179, 1260]. Adapting [1647, 857, 896, 1991]. adaption [2449]. A}daptive [2480, 506, 377, 159, 236, 769, 371, 1980, 474, 1108, 1402, 637, 893, 1281, 2048, 673, 770, 839, 921, 2334, 2134, 340, 1077, 1572, 373]. adaptivity [1971]. Adding [1474, 1864, 1000, 2430, 1466]. additions [1548]. address [667, 2239, 1978, 1274]. Addressing [2149, 1837, 2355, 1240]. ADDS [1493]. adjusting [1558, 1953, 2301, 2092, 2185]. administration [1225]. ado [1928]. Adoption [554, 1336, 1124]. Advanced [1156, 140, 1109, 1742, 832]. Advancements [1224]. advances [1963]. advantage [757]. adventure [155]. Adventures [1509]. advice [1527]. advisor [1710]. affine [724, 341]. affinity [1088]. again [650]. age [2145, 2151, 453]. agent [1345, 172, 1238]. agent-oriented [172]. Agents [1345, 1234]. aggregates [1841]. aggregation [911, 2461]. aggressive [1528]. Agile [2436, 1218, 1176, 1485]. ahead [1818, 1011]. ahead-of-time [1818]. AI [1300]. AJA [1345]. Ajax [2406]. AJEER [1165, 1221]. Alexander [2436]. algebra [1003, 171, 997, 1211]. Algebraic [1635, 424, 2214, 716, 1455, 1430, 1514, 1264]. Algorithm [1573, 396, 1487, 1545, 767, 1683, 758, 1549, 557, 850, 1111, 797, 1565, 2251, 1046, 1048, 848, 1759, 1964, 141, 142, 1090, 691]. Algorithmic [420, 2302]. Algorithms [101, 88, 1940, 607, 2170, 1478, 669, 674, 174, 768, 1579, 313, 732, 729]. Alias [701, 928, 1562, 2029, 1078, 1943]. aliased [1897]. Aliasing [309, 804, 1450, 1048, 1904, 1749]. alignment [1074]. Allocation [226, 229, 193, 350, 101, 548, 2425, 783, 2055, 609, 681, 1034, 1037, 1017, 1511, 1019, 1113, 2056, 1745, 1762, 745, 1378, 1841, 1415, 1070, 1529, 2026, 733, 759, 1380, 608, 152, 682, 1090, 725, 1027, 1033, 856, 1091, 1112, 1381]. Allocator [194, 1567, 770]. Allocators [296, 230]. allow [2433]. almost [1634, 2209]. along [1183, 1262]. Alpha [48, 674]. AlphaServer [184]. ALT [1230]. Alternatives [190]. Ambient [246]. ambients [265, 490]. ambiguity [1549]. ambiguous [521]. amnesia [1142]. analogy [1919]. analyses [971, 1738, 2419, 2356, 559, 1008, 508, 1342, 2470, 1615, 2469, 560, 2030]. Analysis [305, 469, 1357, 1038, 312, 290, 287, 309, 128, 2122, 590, 406, 2163, 207, 517, 289, 478, 547, 2481, 2423, 1558, 842, 2240, 808, 1055, 2047, 1293, 709, 2365, 2370, 484, 2363, 1366, 1450, 802, 1651, 723, 1023, 1217, 2193, 728, 1538, 1947, 626, 903, 68, 1285, 1811, 930, 2035, 1833, 758, 1029, 2399, 1709, 2118, 148, 1937, 66, 1562, 2031, 813, 628, 637, 1902, 84, 1623, 254, 1798, 91, 741, 638, 530, 1340, 2114, 1791, 586, 1339, 2032, 2331, 1790, 1581, 2414, 1338, 1872, 1241, 1793, 2188, 865, 1131, 2467, 2338, 513, 344, 1750, 1761, 655, 2308, 1065]. analysis [1826, 1747, 2029, 1066, 2284, 330, 1664, 780, 704, 519, 1755, 486, 795, 1792, 76, 817, 529, 171, 1383, 1650, 2468, 879, 1938, 1513, 2400, 1505, 1736, 1592, 253, 997, 2120, 1769, 1536, 878, 1161, 2462, 452, 2167, 1577, 2329, 250, 2046, 67, 441, 497, 2322, 78, 2466, 1356, 331, 1597, 1854, 1473, 2107, 1032, 1457, 1583, 1052, 1411, 72, 2337, 2274, 1413, 1883, 136, 2306, 1981, 1855, 343, 288, 1077, 2011, 1770, 1078, 1737, 1838, 694, 1943, 1079, 1594, 803, 540, 358, 356, 357]. analyzable [696]. analyzer [809]. analyzers [1279]. Analyzing [487, 2123, 2317, 1832, 1591, 1542, 1590]. anatomy [1426]. Ancestor [1185]. anchored [1480]. andante [1234]. Android [2358]. ands [1916]. ANI [1250]. annealed [1485]. Annotated [100, 783, 1797, 441, 517]. Annotation [6, 2117, 696]. annotations [2423, 701, 2307, 300]. Annual [54]. anomalies [1557, 1036]. Anomaly [2225, 1691, 1421, 1762]. Anomaly-based [2225]. ANSI [2321]. ANSI-C [2321]. answer [1521]. ant [1793]. anthropology [2436]. antipatterns [1261]. any [1301]. anything [2242]. AOP [863, 1220, 1239, 1468, 966]. APE.FORTH [706]. aperiodic [1814]. API [2288, 1371]. apples [1903]. Application [1288, 179, 613, 2253, 1218, 908, 1228, 1834, 454, 1169, 1784, 1475, 452, 2389, 1764, 72, 1799, 1182, 1259, 2057, 1862, 1396, 691]. Application-aware [2253]. Application-level [1288, 908, 1784]. Applications [377, 37, 194, 195, 376, 302, 2477, 858, 38, 683, 2133, 463, 1217, 845, 2238, 1999, 697, 1386, 946, 1387, 1557, 1828, 2332, 1254, 1158, 1215, 868, 2250, 1896, 1488, 1858, 2281, 1093, 658, 2401, 2313, 2406, 2469, 1331, 593, 2462, 926, 682, 1574, 1526, 849, 2404, 2274, 2117, 2306, 1986, 343, 1843, 1770, 2276, 755, 1588, 2425, 333]. apply [1130]. Applying [935, 759, 757, 658]. Approach [24, 46, 185, 402, 1218, 782, 938, 2363, 723, 2202, 2133, 180, 845, 1228, 2330, 1484, 1612, 1968, 1658, 711, 1358, 1359, 2225, 1332, 794, 1752, 1319, 1435, 1959, 1424, 1542, 1761, 883, 686, 1214, 2424, 1753, 2205, 1179, 1260, 2290, 2160, 567, 2299, 2191, 2314, 166, 599, 2257, 1449, 1885, 366]. approaches [970]. approximate [1048]. approximating [2230]. approximation [1722]. arbitrarily [341]. arbitrary [915]. arbitration [674]. Archipelago [1978]. architect [1245]. Architectural [2476, 200, 201, 2248, 199, 2280, 2221, 101, 2423, 1977, 1460, 1689, 1461]. Architecture [1292, 656, 184, 363, 192, 202, 207, 474, 204, 1155, 1209, 903, 1687, 1601, 2050, 1684, 688, 1202, 2355, 1025, 379, 1693, 1097, 2436, 2247, 461, 1671, 1463, 1694, 756, 833, 899, 613, 1781, 552, 1607, 1253]. architectures [1821, 1227, 1294, 607, 1683, 1387, 1278, 1074, 679, 766, 604, 857, 848, 2280, 1462, 2342, 605, 608, 2241, 1094, 1605, 1835, 2294, 351, 362]. Archive [42, 1250]. archJava [1155, 1209]. ArchMatE [1460]. Area [1605, 333]. Aristotelian [178]. arithmetical [269]. arithmetics [1592]. Arity [2445]. Arity-generic [2445]. ARM [2292, 1114, 598, 2355]. ARMI [921]. Array [2464, 421, 1088, 1819, 90, 2045, 898, 915, 1340, 814, 341, 936, 78, 1086]. arrays [2056, 1273, 2037, 630, 1272, 832, 1185]. Arrows [423, 2452, 2375]. ART [965]. Artemis [1557]. artifact [1175]. artificial [1858]. arts [2142, 1203]. asparagus [1121]. Aspect [218, 479, 480, 1647, 1452, 1917, 1167, 596, 1468, 1909, 1214, 932, 1453, 2160, 1670, 2396, 1966, 1163]. aspect-oriented [480, 1452, 1917, 596, 1909, 1214, 932, 1453, 2160, 1670]. AspectC2C [1966]. AspectCOOL [480]. aspectJ [1221, 1474, 1377, 1717, 1653, 1232, 692, 1165, 693]. aspectJ-enabled [1221, 1165]. Aspects [33, 2131, 2295, 1917, 1243, 1757, 518, 1229, 693, 880, 894]. Aspectual [1453]. ASPLOS [2476]. ASPLOS-IX [2476]. assemblies [2409]. Assembly [417, 1772, 735, 1448, 1584, 597, 1522, 1771, 418, 1146]. Assertion [247, 306]. assertions [2319, 600]. Assessing [1739, 1261]. Assessment [791, 2408]. Assignment [298, 607, 1111, 852, 1099]. assignments [66]. assist [1706]. assistance [172]. assistant [1628, 1498, 2383]. Assisted [231, 2273, 716, 2348, 1416]. Associated [1313, 1447, 1454]. Associating [1523]. associations [2461]. assurance [1129, 1488]. ASSURE [2223]. ASTs [386, 387]. asymmetric [2268, 2241]. Asynchronous [980, 1096, 311, 2278, 2177, 1949, 1750, 505, 1416]. Asynchrony [375]. ATOM [1052]. AtomCaml [1433]. Atomic [2252, 2113, 2034, 2169, 690, 1524, 2006, 1657]. atomicity [2269, 776, 821, 991, 2033, 1675, 2222, 1433, 2110]. Atomizer [991]. ATOMOS [1550]. ATS [2440]. attacks [2363, 1526]. Attribute [2396]. attribution [2337]. audit [2064]. auditing [1570]. augment [1451]. augmented [528]. augmenting [855]. AURA [2064, 2360, 2369]. authentication [986]. Author [209, 316, 2475, 427, 2058]. authorization [2064]. Authors [30, 389, 390, 43, 52, 99, 124, 216, 125, 217, 279, 324, 476, 280, 325, 477, 544, 583, 621, 545, 584, 622, 962, 57]. Auto [1561, 2314]. auto-parallelization [2314]. Auto-vectorization [1561]. Autolocker [1524]. automata [263, 1554, 521]. Automate [8, 1228]. Automated [908, 761, 2206, 1813, 1342, 1820, 2225, 1385, 2424, 2001, 1809]. Automatic [1018, 1013, 2012, 304, 1707, 1569, 1020, 937, 601, 530, 745, 1378, 2180, 1780, 422, 2329, 1779, 1942, 358, 1830, 1265, 829, 1931, 532, 2016, 2051, 92, 1962, 1362, 1581, 1424, 1788, 900, 2261, 1852, 2245, 2211, 882, 494, 2210, 1795, 2223, 1799]. Automatically [1387, 811, 2401, 1769, 660, 104, 1767, 1279]. Automating [1470]. automotive [1222]. autonomic [910]. autonomous [164, 337]. available [337]. AVIO [1675]. avionics [842, 922]. avoidance [2190]. Avoiding [262]. Award [537]. Awards [120, 31, 537]. Aware [348, 193, 2288, 1602, 1208, 1789, 2338, 1102, 603, 2426, 2342, 1604, 1407, 725, 2109, 2253, 1814]. awareness [2422]. away [602]. Awesome [1909]. Axiomatic [1352]. axiomatization [884].

Back [1941, 1006, 810, 650, 495, 1498, 754]. back-end [810, 1498]. Back-propagation [1006]. Backtracking [2255, 1442, 146, 1134]. Backtracking-based [2255]. backward [516]. bad [1903, 955]. Balanced [1051, 856]. Balancing [1091, 2255, 945, 333, 373]. bananas [890]. bandwidth [1700, 1287]. bank [1111]. Banked [350]. Banks [351]. Barrier [547, 2187, 1901, 694]. barriers [775, 2293]. Barron [1851]. base [250]. Based [2484, 298, 108, 355, 37, 41, 225, 361, 473, 319, 357, 1089, 1819, 2258, 948, 2176, 38, 1247, 2133, 820, 2178, 342, 2358, 1834, 1968, 1691, 1658, 1019, 911, 818, 1703, 2052, 1076, 984, 1500, 66, 2056, 2225, 454, 408, 786, 715, 1398, 1584, 81, 979, 1154, 1870, 2339, 2025, 577, 2203, 2239, 1338, 2255, 954, 864, 2332, 1618, 364, 2325, 1375, 459, 1080, 1102, 1488, 563, 1465, 1271, 1852, 2353, 1513, 688, 1698, 182, 1568, 2424, 258, 2426, 1404, 2462, 1606, 1680, 1760, 1327, 1261, 1753, 770, 1097, 1390, 562, 657, 1679, 1437, 1238, 970, 1257, 592, 2130]. based [183, 334, 567, 1583, 819, 174, 2314, 166, 1280, 2011, 2257, 2361, 1449, 1078, 1094, 1412, 2287, 1360, 360, 2301]. bases [996]. Basic [2103, 1489]. basis [1754]. batteries [2219]. BC [2478]. BDD [1080]. BDD-based [1080]. BDDs [802]. be [1389, 515, 1355, 1893, 22, 23, 749]. beans [955]. Beautiful [2390]. Bebop [2474]. became [1248]. Bee [1819]. before [1184]. Behavior [312, 207, 1973, 1787, 1169, 2352, 660, 1045]. behavioral [934]. behaviour [751, 1161, 1832]. Bell [1677]. Beltway [566]. benchmark [1092]. benchmarking [1651, 1233]. benchmarks [1651, 2446, 1986]. benefit [2232]. benefits [1989, 362]. benign [1769]. Berlin [539]. best [749]. Better [1997, 1553, 1508]. between [1361, 2149, 524, 2323, 1238, 670]. Beyond [966, 2448, 2259, 1153]. bi [2193, 1332, 2197, 247]. bi-abduction [2193]. bi-directional [1332]. bias [2058]. biased [1657]. Bibliography [100, 517]. bidirectional [88, 1632]. Bidirectionalization [1852, 2183]. Bidwidth [72]. big [2240]. Bigraphs [721]. binaries [1114, 503, 359]. Binary [112, 2337, 107, 113, 1102, 563, 1775, 1765, 270, 1078]. Binders [402, 27, 891]. Binding [409, 655, 1616, 2384]. binding-time [655]. bindings [145, 2329]. biology [1193, 1925]. Biorthogonality [2381]. Bisimulation [490, 983, 1318]. bisimulations [1506]. Bit [522, 1677, 2248, 855, 1391, 1099]. bit-differential [1099]. bit-encoding [1677]. bit-streaming [1391]. Bit-vector [522]. Bitwidth [725]. biXid [1632]. black [1776]. black-box [1776]. blind [2058, 520]. bloat [1894, 2333, 2335]. block [1610, 2170, 2175]. Blocking [341, 2170, 1746]. blocks [929, 2113]. BLOOM [2459]. blueprint [459]. Bluespec [436]. BNF [1540]. boilerplate [778, 1152, 1443, 1731]. Book [1609, 2382]. boolean [1341]. Boomerang [1961]. boosted [2291]. Bootloader [123, 122]. Bootstrapping [630, 2029]. border [1863]. Boston [2480]. both [208]. BOTS [1606]. Bottleneck [1217]. bottlenecks [1837]. bound [1418, 530, 2331, 1086]. Boundaries [105]. bounded [2311, 872]. Bounding [492, 1059, 1808, 1287]. bounds [90, 78]. box [1776, 1543]. Boxes [890]. Boxy [1637]. brady [1248]. BRAIN.FORTH [1300]. branch [2246, 1979, 816, 1376]. Branches [26, 25]. Breadth [141]. Breadth-first [141]. breakpoints [1042]. breaks [294]. brick [1271]. brick-based [1271]. bridging [524]. broken [2129]. browser [1740]. browsers [131]. Bruijn [1629]. brute [1358]. BTB [2246]. buddy [1418]. BuddyCache [683]. buffer [2246, 806, 1673]. buffers [1101]. Bug [805, 1997, 2225, 1368, 1998]. bugs [1691, 1201, 1308, 2222]. build [2238, 2279]. Building [371, 929, 1208, 1870, 559, 1783, 1896, 1383, 1650, 1272, 1162, 1052, 2343, 790]. built [624]. built-in [624]. bulk [1657]. bunch [1248]. burned [1570]. bus [1274]. buts [1916]. Bypass [1604]. Bytecode [299, 661, 1315, 964, 2000, 2359, 2105, 653]. bytes [1727].

c [1849, 1166, 2020, 304, 2410, 823, 2013, 826, 114, 699, 1977, 806, 1251, 290, 1063, 1167, 1659, 963, 2301, 807, 309, 1573, 115, 139, 277, 278, 1592, 2320, 1556, 2111, 2321, 917, 1128, 22, 23, 1520, 102, 283, 27, 1764, 866, 828, 1883, 1086, 365, 284, 1662, 1006, 1966, 1943]. C# [1907, 1423]. C-based [2301]. c-rules [1849]. C-SAW [1167]. C/C [2410, 2111]. Cache [228, 312, 1571, 1407, 197, 1816, 572, 611, 1673, 673, 867, 753, 751, 1404, 1106, 493, 2046, 756, 1107, 1411, 1413, 1843, 1412, 362]. Cache-conscious [1571]. caches [673, 1844, 1980, 1838]. Caching [89, 2100, 773, 2239, 769, 675, 1698]. Calculating [750]. calculations [568]. Calculator [1351, 1190]. calculus [1638, 1507, 439, 975, 1766, 1235, 259, 2081, 731, 1772, 2169, 1518, 1846, 748, 1742, 635, 779, 702, 498, 722, 1000, 2317, 270, 960, 1617, 257]. Calendar [382, 318, 466, 32, 53, 94, 121, 210, 275, 538]. Call [211, 54, 36, 539, 540, 426, 1880, 319, 885, 960, 1875, 1845, 1015, 1959, 2192, 2415, 1854, 1137, 173, 174, 2346, 654]. Call-by-name [960, 1959, 885]. Call-by-Need [426]. Call-by-value [885, 960, 1875, 2192]. Call-pattern [1880]. call-stack [1845]. call/cc [654]. calling [1886, 1529, 1572]. calls [1372, 89, 867, 2400, 1900, 161]. Cambridge [2476]. caml [2392, 1453, 1950]. Can [37, 1893, 2362, 38, 2413, 690, 1187, 1969, 1508, 385, 2129, 1002, 981]. Canada [2478, 2479]. cannot [1389]. canonical [1143]. capabilities [2081]. capable [909]. capacity [2024, 2364]. Capo [2226]. card [2069, 1109]. Carlo [253]. Cartesian [1851]. Case [234, 7, 2288, 169, 1826, 2105, 1492, 1927, 613]. cases [1636]. cast [134]. casts [1152]. Catch [2362, 927]. categorization [459]. caught [1863]. Causal [2375]. cause [726]. causes [1894]. CBEA [1825]. CC [909, 2201, 654]. CCSL [2357]. CCured [812, 496]. CDuce [873]. CEAL [2301]. cell [2391]. cellML [1873]. center [1972]. centred [883]. Centric [197, 873, 1197, 1307, 2353, 754, 2299].

Century [271]. certification [2176, 1498, 2392, 710]. Certified [1773, 1522, 1628, 1772, 647, 970, 503]. Certifying [2022, 71, 1810, 1058, 560]. CFA [1615, 2088, 1505]. CFGs [1356]. CFL [1084, 250]. CFL-reachability [250]. CGCExplorer [1809]. chains [1408, 244]. Chair [380, 93]. Challenge [2446]. Challenges [656, 1205, 338, 329, 1573, 2355, 1825, 103, 367]. Chalmers [2150]. Chameleon [2334]. Change [479, 2466, 1464, 167, 1205, 252]. changeable [1890]. Channel [982, 1741, 1511, 1566, 2364]. channels [1866, 844]. chaos [2297]. characteristics [1998]. characterizing [660]. Charting [414]. charts [687]. chat [1207]. Cheap [143]. Check [189, 196, 2319, 1538, 574]. checked [1480]. checker [991, 2033, 1859, 1713]. CheckFence [1768]. Checking [804, 1372, 1302, 2007, 2078, 2020, 1897, 2178, 1768, 488, 955, 2195, 1663, 950, 573, 1322, 1515, 2040, 1067, 717, 1008, 1964, 1016, 2039, 258, 2128, 2066, 527, 1086, 69, 2172]. checkpointing [908, 1288, 1626]. checks [90, 835, 837, 1996, 1795]. Chez [1614]. chip [1973, 1511, 1834, 1800, 1682, 673, 591, 2272, 1980]. chips [2279, 1803, 1695]. choice [2302]. churn [2107]. CIL [2000]. Cilk [2280]. Circular [1354]. CISC [308]. citation [898, 895, 900, 901, 896, 899]. Cl [1819]. CLA [309]. CLASE [2218]. clashes [1890]. Class [328, 938, 2449, 1469, 1636, 1313, 568, 1518, 176, 1641, 2216, 1239, 1443, 459, 2089, 2174, 1439, 1958, 2209, 177, 1433, 2403, 1724, 79, 2450, 2396, 2090, 2472, 710, 666]. Class-is-type [328]. classages [1465]. Classbox [1464]. Classbox/J [1464]. Classes [44, 1321, 935, 165, 1723, 1888, 1641, 459, 646, 615, 2323, 1764, 1212, 2429, 2388, 45]. Classical [2197]. classification [572]. classifiers [720]. classifying [1769]. classroom [495]. client [1357, 1818, 560]. Clock [2053, 1289]. Clock-directed [2053]. Cloning [1078, 789, 825, 1611, 1792]. Cloning-based [1078]. Close [276]. Closing [188, 2382, 187]. Closure [404, 2076]. Closures [1125, 1724]. Clowns [1951]. cluster [342, 1484, 911, 841]. cluster-based [342, 911]. clustered [604, 605, 1835]. clusters [2238, 909, 1395]. CML [1748]. CMP [1697, 1290]. CMPs [1993]. Co [2480, 898, 2359, 1909, 1734]. co-array [898]. co-inductive [1734]. co-logic [2359]. co-weaving [1909]. coalescence [852]. coalescing [1821, 2347, 555, 2027]. coallocation [1571]. coarse [2056, 1685, 850, 848, 2342, 1605]. coarse-grained [2056, 1685, 848, 2342, 1605]. coated [1612]. Cobol [542]. Code [298, 301, 1010, 189, 1836, 349, 395, 309, 1376, 366, 1117, 1477, 1382, 2292, 2464, 929, 2176, 2053, 1406, 1236, 1773, 1999, 892, 712, 2195, 1585, 853, 647, 699, 854, 561, 1599, 601, 1448, 1584, 2119, 1459, 267, 1793, 2351, 580, 1397, 1762, 1302, 1826, 967, 597, 604, 1047, 855, 1053, 2417, 2175, 1157, 1969, 2432, 606, 496, 1522, 593, 1043, 2427, 2123, 1600, 170, 1666, 2201, 1159, 2371, 919, 1521, 1173, 2326, 1981, 1986, 69, 1006, 1146, 74, 1110, 1108, 610, 1587]. code-based [2176]. code-copying [2123]. code-size [606]. CodeBricks [929]. Codebugger [28]. codesign [613]. coercions [2404]. Coercive [731]. coffee [294]. cognitive [2432]. Coherence [1276, 912, 1268, 907]. coherent [2247]. cohesion [1612]. Cohort [370]. coinductive [535]. Collaboration [14]. collaborative [683, 1895, 856]. collapsing [1408]. collected [1719, 358]. Collection [231, 222, 233, 232, 230, 227, 1028, 1414, 947, 935, 566, 953, 2009, 1059, 1049, 463, 85, 1578, 1285, 930, 837, 786, 1919, 565, 652, 772, 1379, 1472, 954, 1759, 1415, 1615, 1824, 839, 682, 747, 785, 768, 1579, 2244, 1885]. collections [2295, 2334, 2253]. Collector [228, 297, 239, 238, 2319, 948, 294, 838, 742, 988, 2009, 86, 1419, 464, 763, 1976, 1353]. Collectors [234, 492, 760, 2206, 1810, 2010, 1809, 260]. Colored [249]. Coloring [1034, 1017, 607, 1019, 556, 1090, 2262]. COM [702]. combination [2189, 457]. Combinatorial [1708]. Combinators [1332, 1731]. Combining [1820, 1431, 349, 1582, 565, 1784, 151, 270, 1480, 667, 2114, 1186, 1815]. come [263]. comfort [1914]. command [1526]. commercial [2233, 2389, 1862]. Committee [958, 317, 120, 317, 31, 381]. Committing [2267]. commodity [1968, 2313, 1996, 1272, 2230]. Common [285, 746, 1757, 1265, 973, 2453]. Communicating [543, 1325]. Communication [190, 352, 191, 1988, 255, 679, 909, 2353, 698, 921, 256, 343, 1830, 2036, 670, 2387]. communication-centric [2353]. communication-exposed [679]. Commutative [2394, 2375]. Commutativity [2240]. Compact [227, 262, 1682]. Compaction [349, 767, 1405, 1580, 1976]. Compactly [918, 1751]. Comparability [2262]. Comparative [108, 939, 1063, 674]. Comparing [2217, 2296, 2298]. Comparison [1247, 1672, 586, 865, 1065, 1593, 2469, 2264, 799]. comparisons [2391]. compatibility [741]. compatible [2320]. compensations [1330]. competing [2345]. competitive [2046]. compilable [636]. Compilation [189, 1544, 117, 1116, 1315, 1911, 1654, 523, 2023, 711, 1402, 135, 116, 1073, 2121, 586, 744, 1022, 2105, 1491, 1102, 1802, 1492, 603, 615, 1655, 614, 970, 592, 72, 819, 784, 450]. Compile [801, 798, 1007, 1001]. Compile-time [801, 798, 1001]. compile/optimize [1007]. compiled [1386, 651, 909]. Compiler [551, 1552, 377, 2273, 1035, 903, 1511, 1834, 2045, 205, 743, 611, 76, 1603, 1284, 1844, 315, 365, 313, 204, 670, 391, 858, 2302, 1014, 965, 2381, 2000, 1010, 1566, 1831, 1772, 898, 71, 1990, 1881, 1601, 1021, 1060, 679, 587, 2068, 220, 1818, 797, 2282, 864, 175, 943, 261, 669, 1026, 242, 867, 509, 2345, 2344, 2261, 811, 1498, 817, 2281, 1712, 1528, 70, 1058, 2054, 1655, 2321, 379, 1213, 1390, 1128, 1148, 160, 567, 2272, 800, 451, 819, 1054, 613, 571, 1396, 1607, 1594, 790, 2141]. Compiler-assisted [2273]. Compiler-Directed [205, 1511, 1834, 743, 611, 1603, 1566, 2272]. Compiler-managed [1844]. compiler/architecture [613]. Compilers [387, 11, 42, 386, 2023, 1111, 2042, 1533, 494, 749]. Compiling [1417, 1865, 874, 2092, 606, 834, 653, 578, 1718, 1050, 1118, 1825]. complement [1852]. Complementing [1409]. Complete [1057, 2405, 2229, 1147, 2031, 2033, 884, 2320, 1630, 1734, 2088]. completeness [1729]. complex [1902, 449, 1463]. Complexity [48, 1822, 728, 1937, 2179, 1855]. Component [539, 378, 62, 1889, 369, 922, 1488, 1661, 1462, 1456, 1508, 2125, 1606, 1261, 560, 899, 1083, 368]. component-client [560]. components [1166, 1208, 1687, 1398, 1776, 453, 685, 1622, 1272, 1231, 1264]. Composable [636, 1397, 1864]. Composing [1393, 296, 155, 508, 642, 1909, 2213]. composite [157]. composition [1126, 2133, 1330, 1243, 1565, 1468, 1642, 458, 2316, 1575, 801]. Compositional [2193, 420, 1721, 1106, 1315, 1634, 1535, 1529]. Compositionality [531]. comprehensive [1971, 1998, 2265]. Comprehensively [1690]. Compression [299, 1406, 949, 1800, 561, 65, 757]. Compressor [1580]. compromise [447]. computability [728]. Computation [1697, 205, 1558, 1953, 1269, 150, 1320, 1478, 2301, 507, 2185, 2384, 1686, 449, 1011, 1390, 534, 1427, 256, 1928, 1077]. Computational [264, 147, 2179, 344, 1955, 1352, 1211, 625, 2194]. computational-environment [625]. Computations [409, 1788, 2280, 2294]. Computer [20, 716, 182, 2139, 2143, 21, 2220, 2061, 245, 2377, 1671, 183, 517]. computer-vision [2377]. Computing [34, 273, 376, 1531, 632, 35, 1484, 338, 1687, 329, 664, 2251, 2275, 868, 345, 1102, 1763, 2221, 2049, 337, 968, 2290, 274]. Cona [1229]. Concatenate [631]. concept [808]. concepts [2140, 1189, 2155, 1520, 2158, 2159, 2162, 1659]. Conceptual [48, 1216]. concern [1183, 1262, 1168]. concern-dimensions [1183, 1262]. concerns [1240, 169, 1664, 2130]. concise [521, 1945]. concrete [1192]. concurrency [2013, 158, 1925, 906, 1784, 1998, 2289, 1097, 2387]. Concurrent [282, 231, 378, 232, 238, 2307, 1028, 947, 2412, 723, 2295, 1768, 2398, 626, 2035, 1627, 1369, 1448, 2414, 766, 1580, 2154, 1929, 1181, 564, 2010, 990, 2235, 2397, 2008, 1624, 2019, 2296, 2298, 1579, 1809, 2018, 1976, 1353, 248, 1146, 1110, 1626, 1132]. Conditional [1749, 2011, 1593, 25]. conditionals [1916]. Conditions [26, 2386, 262, 2473]. Conference [2476, 2477, 211, 318, 466, 468, 467, 470, 469, 539, 32, 53, 94, 121, 210, 275, 34, 538, 2485, 33, 16, 2478, 2479, 2002]. Confessions [1918]. confidence [1564]. Configuration [371, 1240, 1171]. configurations [676]. confined [455]. confinement [499, 955, 940]. confirmed [727]. conflict [803]. conflicting [2267]. Conflicts [619]. conformance [2423, 560]. congruences [490]. connected [1766]. Connecting [1336]. connection [1084]. Connectivity [954, 762]. Connectivity-based [954]. conquer [2184, 1780, 333]. Conscientious [1667, 1917]. conscious [1571, 1802, 493, 592]. consequences [2152]. Conservation [868]. Conservative [230, 492, 1049, 760]. conserving [616]. Considerations [204]. Considered [546, 446, 1764, 2371]. consistency [683, 1768, 1268, 900]. consistent [742, 1329, 600, 1929]. constant [1024]. Constrained [2126, 2043, 949, 1573, 1388, 1117, 1108]. ConstrainedJava [1170]. constraint [2032, 1454, 1084, 688, 1606]. constraint-based [688, 1606]. constraint/CFL [1084]. constraints [616, 1074, 84, 2315, 1797, 2180, 606, 2067, 441, 502, 933, 1523]. construct [1574]. Constructing [24, 1257, 2213]. construction [965, 1406, 220, 1435, 494, 1854, 174]. constructions [731]. constructor [783, 2388]. constructors [2420, 737]. Constructs [473, 2168, 1542, 1739, 2111, 1867]. Consumption [346, 548, 838, 2043, 1566, 764]. containment [1437]. Contaminated [85]. Content [411, 1935, 678]. content-directed [678]. Contention [335, 2109, 2265]. Contention-aware [2109]. Contest [50, 2099]. Context [1335, 1729, 2453, 1565, 2460, 1886, 84, 807, 1067, 2029, 1792, 2470, 879, 1808, 1577, 1583, 1182, 1259, 1078, 1572]. context-insensitive [2470]. Context-oriented [2453]. Context-sensitive [1565, 2460, 84, 807, 2029, 1792, 1577, 1583, 1078]. Contextual [1929, 2461]. Contification [403]. Continual [1277]. continuation [684, 730]. Continuations [1444, 1575, 1133, 1050, 1959, 884, 1865, 2092, 2401, 131, 754, 2403]. continued [1865]. Contract [438, 2172]. contraction [1819, 341]. Contracts [634, 2450, 1147, 1948, 1467, 155, 1513, 1229]. Control [2331, 2400, 1638, 1384, 499, 2181, 448, 1648, 1584, 1864, 2182, 888, 597, 1593, 662, 1109, 1974, 354, 1178, 754, 2304, 1032, 1426, 1734, 730, 1855, 457, 2190, 1289, 1823, 353]. Control-flow [2331, 2400, 1032]. controllable [907]. Controlling [838, 463, 2297, 2307, 1464, 1378]. conventions [1529]. convergent [896]. conversion [2076]. Cool [667]. Cool-Mem [667]. cooperation [1891]. cooperative [2171, 28]. coordinated [1972]. coordinating [2294]. Coordination [218, 2036]. copies [2335]. coproducts [642]. Copy [2027, 199, 2186, 555, 832, 952]. Copy-on-write [2186]. copying [988, 2426, 2123, 952]. Coq [1848]. CORBA [163, 375, 372, 922, 686, 374]. core [1638, 439, 1235, 2143, 1829, 1104, 2279, 2259, 1994, 2280, 2260, 2348, 1908, 2241, 1781, 2425]. cores [1697, 2257]. Cork [1719]. Corner [211, 318, 466, 468, 467, 470, 469, 32, 53, 94, 121, 210, 275, 34, 538, 33, 16]. correct [2113, 2327, 170, 256, 1809, 1449]. correcting [1767]. Correctness [1579, 971, 2381, 259, 1141, 509, 811]. Correctness-preserving [1579]. correlation [1577, 2011]. Correspondence [23, 401, 483, 1350]. CoSMIC [1240]. Cost [301, 425, 528, 1075, 987, 1073, 1845, 91, 1593, 2185, 2313, 336, 1678, 1107]. Cost-augmented [528]. cost-driven [1073]. count [981]. counter [1687, 1985]. counterexample [726]. counters [2463]. Counting [240, 953, 987, 464, 1615]. coupling [2115]. Course [551, 320, 321, 2141, 2147, 2156]. coverage [1679]. covering [850]. CPI [1687]. CPS [535, 148, 2403]. CPS-transform [2403]. CPU [1816, 797]. CPUs [1984]. crafts [1203]. Creating [682, 162, 170, 691]. creation [2416, 1448]. CRF [156]. criteria [1714]. critical [809, 2392, 2241]. cross [1471, 1175, 586, 943, 2036]. cross-artifact [1175]. cross-platform [943]. cross-run [1471]. cross-runtime [2036]. Crosscutting [1964, 1240, 2130, 1756]. cryptographic [2176]. Cryptographically [1954, 1955]. Cryptography [200, 201]. CS [2147]. CS1 [823, 963, 1185]. CS2 [823, 963]. CSSV [806]. CTrigger [2222]. CTRL [1230]. Curing [549]. Curricula [273, 274, 35, 34]. Curricular [320]. curriculum [2140, 2136, 2137, 2138, 2152, 2160, 59, 58, 60]. curry [1130]. cursor [2218]. curve [1283]. curves [2230]. custom [681]. customer [1206]. Customization [1035]. Customized [404, 2416, 1383, 1650, 1052]. cut [1451, 1072]. cycle [2231, 2050, 1759, 2342]. cycle-accurate [2050]. Cycles [233]. Cyclic [1934, 1610]. cyclically [1766]. cyclone [577, 777]. CZ [2407].

D [1218, 1092, 1682, 1271, 1695]. D-SPTF [1271]. DaCapo [1651]. dag [527]. dance [1210]. DART [1385, 396]. Dart-Zobel [396]. Data [836, 15, 303, 349, 13, 2080, 247, 109, 197, 2214, 1055, 2020, 2307, 181, 2442, 2053, 1961, 695, 1361, 1768, 572, 2238, 137, 2165, 570, 1571, 1840, 2045, 911, 1836, 2438, 678, 1937, 1663, 915, 937, 743, 628, 2097, 1902, 1856, 1392, 1495, 1962, 714, 1685, 1118, 2070, 2297, 2325, 895, 1455, 675, 2087, 1989, 1046, 745, 1378, 574, 1841, 533, 2281, 1725, 2310, 1951, 336, 1987, 2242, 1769, 1181, 1430, 1561, 916, 595, 1812, 733, 493, 1514, 917, 1972, 1844, 1560, 497, 1127, 1795, 872, 2019, 2272, 801, 1922, 1701, 1523, 755, 2038, 1496]. data-cache [572]. Data-flow [2080, 1055, 2053, 2165]. Data-Intensive [13, 2238]. data-layout [497]. data-parallel [2297]. data-race [2310]. data-sets [895]. data-types [872]. Database [221, 2391]. datacenter [1742]. Dataflow [2035, 2474, 2051, 508, 1342, 1684, 593, 2001]. datapath [1100]. datapath-width [1100]. datarace [575, 2035]. datatype [2445, 737]. datatype-generic [2445]. Datatypes [424, 639, 1922, 784, 2395]. day [2145]. days [516, 542, 961]. DBT [2355]. Dead [323, 1487, 1545, 672, 322, 1347]. dead-instruction [672]. deadlock [1766, 1181, 2190]. deadlocks [695, 1132, 2308, 1812]. Dealing [360, 361]. Debug [2417]. debuggability [1655]. debuggers [392]. Debugging [808, 420, 2353, 5, 2464, 484, 88, 1132, 2225, 1135, 2417, 1720, 1910, 599, 28]. decentralized [1271, 2304]. Decidability [1500]. decidable [2405]. decide [1082]. Deciding [781, 2088]. decision [2315, 1078, 359]. declarations [1480]. Declarative [1653, 2295, 13, 1858, 33, 2419, 2080, 1243, 138]. Declaring [950]. Decoding [359]. decomposition [1072]. decompression [853]. Deconstructing [1273]. decoupling [1276]. deductive [254]. Deep [2102]. DeepJava [1893]. deeply [1832]. def [2461]. def-use [2461]. Defaults [424]. defect [1693, 1678]. defects [2108]. deferred [164]. Deficiencies [963, 823]. defined [2431]. Definition [243, 1653, 1365, 792, 861, 1493, 1908]. Definitional [960]. defragmenting [1418]. defunctionalization [977]. Defunctionalized [2074]. degree [2055]. delay [2048, 673, 1605]. delayed [1899]. Delegation [2130]. Delegation-based [2130]. Delimited [1616, 1864, 2182, 1959, 884, 2403]. Delphi [327]. delta [2236]. Demand [287, 1457, 1943, 90, 637, 1554, 2239, 1842, 769, 685]. demand-based [2239]. Demand-Driven [287, 1457, 1943, 637]. demands [2260]. demonstrated [1173]. Demystifying [1382, 2424]. Denali [579]. Denotational [326, 391]. dense [2263]. density [1290, 2431]. Dependable [1219]. Dependence [478, 2258, 1023, 626, 1030, 1488, 1408, 1110, 244]. dependence-based [2258]. dependences [1361]. dependencies [89, 1242, 2215, 1692, 2429]. Dependency [881, 1638, 1463, 1140]. Dependency-style [881]. Dependent [1888, 2198, 529, 1494, 2083, 1944, 2374, 982]. Dependently [417, 418]. deployed [2108]. Deployment [539, 955, 1240, 1661, 1586]. deployment-time [955]. derivation [1852, 2245, 1579]. Derive [104]. derived [1845]. Deriving [1232, 146, 1458, 560, 1590, 2018, 1279]. Descent [24]. describing [612]. description [436, 1601, 1495, 1725, 1097, 1005]. Design [551, 355, 190, 1645, 184, 692, 285, 669, 221, 2485, 186, 2054, 2111, 595, 1223, 2431, 16, 1159, 27, 365, 398, 1713, 474, 204, 197, 2478, 480, 1820, 2133, 2246, 68, 500, 961, 915, 1895, 1154, 1983, 1399, 1752, 797, 1689, 600, 2215, 664, 1042, 617, 778, 2418, 1992, 171, 505, 594, 2432, 1742, 1058, 1404, 141, 1351, 17, 824, 1427, 397, 567, 1908, 919, 1960, 1613, 2283, 1669, 1191, 1530, 348, 373]. design-space [1820]. Designed [2135]. designer [2135]. Designing [375, 226, 372, 182, 183, 1468, 475, 1609]. designs [168, 922, 1183, 1262]. Desktop [195]. Destruct [877]. Destruct-time [877]. Destructors [740]. Details [859]. detect [758]. Detecting [2473, 2268, 2108, 1674, 806, 2308, 1675, 1261]. Detection [628, 1164, 1677, 672, 1811, 1691, 575, 2035, 937, 1807, 81, 2309, 1281, 1719, 2042, 2310, 1576, 1749, 916, 917, 1577, 1753, 1036, 517, 1287, 2314, 1341, 443, 1369]. detector [807, 1367]. determination [1415]. deterministic [2411, 980, 2227, 1627, 2373, 2226, 521, 2228, 2316]. Deus [1004]. develop [625, 2392]. developer [2002, 1517]. developers [1882, 2159]. Developing [376, 413, 1530]. development [1218, 1651, 1176, 1628, 2440, 1614, 1205, 1206, 1203, 591, 1172, 1097, 1064, 1862, 353]. device [1095, 2351, 829]. Devices [314, 1596, 1419, 771, 1388, 1560, 1117, 833, 1984]. DEVIL [792, 360, 361]. Devirtualizable [1286]. devirtualization [175]. DFTL [2239]. d’Hanoï [2372]. diagnosis [1384]. diagrams [1610, 2175, 2394, 1078]. dialect [181]. diamonds [2407]. dictionary [1406, 65]. DieHard [1563]. different [1183, 1262]. Differential [1381, 1099]. differentiation [532, 2390, 882]. Diffusion [240]. digital [1250]. Dijkstra [1766, 707, 1000]. Dilbert [879]. dimensional [1970, 1380]. dimensions [2368, 1183, 1262]. Dirac [882]. direct [654]. Directed [205, 234, 689, 975, 2053, 2187, 1511, 1566, 1834, 678, 743, 1385, 1874, 611, 579, 556, 1857, 1291, 1603, 562, 2403, 2008, 103, 2272, 1054]. Directing [2452]. directional [66, 1332]. Directions [192]. directly [1612]. dirt [1962]. Disambiguation [290]. disciplining [2162]. disconnect [2149]. Discovering [724, 2037]. discovery [761, 571]. discrete [875, 2190]. discrimination [2070]. Discriminative [870]. Discussion [118, 119, 2137]. DISE [853]. disjoint [745]. disjointness [700]. Disjunctive [886]. disk [1272]. disks [1291]. Dispatch [47, 576, 165, 1363, 2132]. dispatching [1482, 691, 729]. Dispersing [1986]. dissecting [1951]. distance [1361, 813]. distances [750]. distinctions [1641]. distinguish [2364]. Distributed [377, 128, 376, 1384, 845, 912, 1831, 2286, 1640, 1329, 1859, 1240, 1478, 1397, 1535, 1254, 1783, 1263, 345, 944, 2256, 335, 336, 1867, 841, 2205, 1272, 722, 266, 1427, 1219, 919, 339, 343, 1530, 348]. distributed-memory [1478]. distribution [1868, 2051, 2096, 1398, 1271, 693]. distributions [498, 237]. DITTO [1795]. divide [2184, 1780, 333]. divide-and-conquer [2184, 1780, 333]. DML [425]. DMP [2227]. DNA [1967]. Do [320, 321, 1743, 1610, 690, 1187, 2396, 2213]. documentation [2382]. does [623, 1127]. doing [2242]. Domain [393, 11, 4, 100, 1095, 1392, 2147, 1241, 1565, 1896, 1172, 2049, 1238, 1493, 1173, 1289]. domain-driven [1172]. domain-independent [1565]. Domain-Specific [393, 11, 100, 1095, 1392, 2147, 1896, 1493, 1173]. domains [1591, 1946, 2224]. Dominance [1356, 82]. dominant [2426]. dominant-thread-based [2426]. dominated [673]. Dominators [403, 82]. Don’t [1135, 602, 2170]. DOOM [581]. double [2058]. double-blind [2058]. doubling [2248]. Down [406, 1420, 1549, 2213]. Downgrading [1326]. downtime [1588]. downward [2184]. Draft [58, 273, 274, 273]. drive [1911]. Driven [543, 287, 1126, 1189, 2045, 1073, 637, 1105, 1158, 1215, 1803, 1202, 1172, 841, 1445, 1801, 1985, 1457, 1993, 2314, 528, 1832, 1923, 799, 1258, 1943]. drivers [1095, 2245, 829]. driving [527]. DSL [2082, 2069, 2377, 9]. DSLs [1174]. DSP [1831, 363, 313]. DSPs [602, 603]. dual [885]. dualising [2197]. duality [150]. dumb [1021]. during [590]. Dutch [999]. DVSleak [1815]. DyC’ [1061, 110, 111]. Dyn}amic [2480, 112, 877, 672, 570, 1840, 1076, 205, 116, 941, 1322, 310, 236, 286, 1101, 771, 2469, 406, 1430, 452, 1720, 337, 117, 2110, 2299, 107, 106, 2244, 1283, 2258, 2365, 2370, 644, 63, 2273, 2367, 1911, 2000, 1654, 1602, 2115, 64, 853, 113, 1448, 991, 2033, 2309, 2421, 2339, 1721, 2414, 1376, 1975, 2308, 1719, 1616, 300, 1383, 1757, 1974, 1876, 1070, 1556, 1929, 1775, 916, 444, 1622, 1490, 2107, 1765, 2368, 79, 1328, 451, 1275, 983, 2448, 450, 798, 2455, 1075, 1564, 1416, 366]. Dynamically [1035, 863, 472, 2465, 815, 615, 2454, 2235, 1756, 920, 729]. dynamically-scoped [1756]. Dynamically-Typed [1035]. dynamics [1176]. Dynamo [113, 2480, 63, 112, 118, 119].

E-path PRE [627]. eagerness [143]. Early [2233, 664, 1901]. ease [1386]. easier [1645]. easy [627, 1201, 1308]. echo [1406]. eclipse [1164, 1870, 1156, 1220, 1165, 1221, 1162, 2416]. eclipse-based [1870]. Economics [2471]. ECOSystem [666]. EDF [1596, 616, 1416, 1815]. Edge [109, 1683, 1278]. edit [1194]. Editor [51, 2218, 2212, 963]. Editors [959]. Edsger [707]. education [2142, 1148]. effect [2411, 700, 1029, 821, 1525]. effectful [1856]. Effective [2379, 1601, 196, 569, 1788, 1576, 80, 101, 2274, 1594, 1116, 2273, 831, 1081, 1031, 887, 2310, 1847, 2306, 1075]. Effectiveness [943, 223]. effects [1336, 147, 854, 2297, 780, 574, 662, 1929, 2194, 1434, 730]. Efficiency [342, 1971, 667, 2009, 1802, 1408]. Efficient [2078, 235, 446, 2365, 2370, 2127, 2278, 1648, 88, 375, 198, 303, 607, 575, 835, 1663, 1599, 254, 1798, 748, 1022, 2256, 1025, 47, 521, 1181, 1011, 917, 2128, 295, 2236, 571, 2028, 548, 1396, 445, 732, 333, 1552, 2108, 1049, 1361, 2238, 1912, 902, 758, 1402, 1902, 2309, 2391, 176, 1419, 2179, 1111, 1346, 664, 2251, 1682, 1785, 1096, 2345, 1688, 1759, 1964, 2271, 593, 2228, 698, 1680, 2427, 2067, 497, 2110, 1993, 1086, 2213, 846, 1395, 1099, 1572, 691]. Efficiently [1689, 2333, 750, 1690, 1274]. Eighth [2483]. EJB [697]. Elected [380]. electronic [568, 1093]. elegance [1191]. eliminate [1901]. Eliminating [1641, 1657, 2346, 90, 2237]. Elimination [196, 1124, 672, 627, 569, 335, 1115]. Emacs [406]. EMBARC [1111]. Embed [1120]. Embedded [355, 352, 1609, 353, 12, 347, 378, 11, 61, 1586, 361, 858, 842, 847, 2047, 2082, 2043, 1820, 845, 1831, 711, 1591, 1113, 616, 1403, 794, 1240, 2354, 1419, 580, 1131, 2048, 1818, 1828, 600, 1103, 2042, 2050, 1699, 591, 2349, 1698, 1592, 594, 2054, 2377, 1522, 1839, 2392, 595, 354, 841, 844, 1279, 1590, 839, 1117, 1408, 834, 849, 1086, 1830, 1832, 1395, 856, 830, 852, 1112, 1099, 2343, 843, 346, 360].

Embedding [624, 2451]. emergency [1417]. Emerging [656, 898, 2425]. Empirical [2408, 1267, 799]. employing [1697, 2239]. Emulation [186, 1681, 2351]. enable [1682]. Enabled [376, 1165, 1221]. enablers [1827]. Enabling [1831, 2118, 1827, 665, 1386, 1286, 2281]. enactment [684]. Encapsulating [455, 1375]. encapsulation [736, 700, 1525, 1904]. Encoding [2360, 2369, 1956, 1677, 176, 890, 445, 522]. encodings [104]. encounters [276]. end [810, 1498, 834]. Energy [858, 616, 664, 346, 603, 592, 1395, 856, 548, 667, 2043, 1566, 1419, 2048, 797, 1828, 2463, 1682, 1291, 1102, 1803, 1802, 1100, 668, 666, 1814]. energy-aware [1102]. Energy-balanced [856]. Energy-conscious [592]. Energy-conserving [616]. energy-delay [2048]. Energy-efficient [664, 1395]. enforcement [132, 2166]. Enforcing [291, 2114, 1774, 1562]. Engine [2352, 306, 2474]. Engineering [468, 467, 469, 1926, 41, 1499, 4, 2481, 1166, 1247, 1569, 169, 155, 1266, 2153, 824]. engineers [104]. Enhance [370]. Enhanced [2042]. Enhancing [1933, 1263, 855, 671, 2422]. enough [2211]. ENSEMBLE [352]. ensuring [2235]. enter [1130]. enterprise [590, 1742, 1272, 1259, 862, 1182]. entropia [338]. Environment [1505, 720, 1357, 1864, 788, 1244, 772, 1673, 162, 2417, 1763, 625, 518, 919, 1168, 1781]. Environmental [1314]. environments [949, 1329, 169, 2430]. Epi [1917]. Epi-aspects [1917]. equalities [724]. Equality [1630, 2191, 2078, 1319, 951]. equivalence [2076, 2327]. Erlang [709, 2386, 2072, 710]. errata [1920]. Error [2322, 2362, 1807, 1776, 2430, 1806, 1287, 1341, 1360]. Errors [420, 726, 1903, 937, 2344, 1992, 1475, 870, 1767, 1805]. Escape [289, 1863, 331, 288]. ESP [558, 314]. essay [2436]. essence [863, 1050, 1245, 718, 1526]. Essential [1373]. Establishing [1899]. Esterel [2357, 1699, 1598]. estimated [1045]. Estimating [764]. estimation [2179, 336, 1605]. ESys.Net [1103]. eta [1319]. Ethernet [909]. eval [1130]. eval/apply [1130]. evaluated [867]. Evaluating [2460, 190, 788, 1867, 374, 1896, 2470, 1686]. Evaluation [2484, 319, 923, 525, 144, 975, 1499, 1875, 68, 92, 893, 143, 2182, 2220, 1884, 2121, 924, 1061, 922, 1131, 797, 1022, 669, 1784, 524, 2054, 1445, 747, 599]. evaluators [131]. Event [1189, 2052, 1158, 1215, 844, 562, 1832, 1923]. event-based [562]. Event-driven [1189, 1923]. event-processing [2052]. Eventrons [1574]. Events [226, 416, 1625, 2071, 456, 1784]. EventScript [2052]. eventually [1743]. ever [51]. Everything [787, 546]. everywhere [1822, 2441]. evolution [447, 2435, 1258]. evolutionary [1359]. evolvable [944]. evolve [2433]. Evolving [677, 2393]. Ex [1004]. Exact [312]. Example [1197, 1307]. Examples [1187]. Excel [883]. Exception [638, 1424, 1346, 1542, 1761, 444, 824, 1480]. Exceptions [311, 927, 439, 886, 87, 1480]. Exclusion [157, 1931]. executable [600, 1826, 687, 1861]. Executing [2427]. execution [2258, 463, 1278, 1398, 1557, 1015, 1828, 1826, 330, 669, 2017, 2061, 1969, 1686, 657, 972, 2371, 1275, 1993, 2241, 2236, 2350, 2109, 2028, 1835, 1805]. executions [918]. Executive [317, 958, 120, 31]. exercise [141]. exhaustive [2210]. existential [2199, 872]. EXOCHI [1781]. exotasks [1817]. ExoVM [1799]. expanders [1643]. expanding [2162]. Experience [2402, 1868, 2072, 1873, 2399, 2097, 1870, 2069, 2380, 2094, 2377, 2392, 2389, 1860, 1849, 1862, 163, 2233, 1252]. Experiences [2150, 790, 1827, 664, 1586]. experiment [480, 1544, 1508]. experimental [1558, 1422, 652, 1280]. Explaining [2030]. Explanation [420]. explicit [1477, 1472, 2305]. explicitly [2460, 658, 2264]. exploitation [1680]. Exploiting [1973, 912, 303, 1685, 2282, 906, 75, 510, 1706, 1416]. exploration [2437, 1820, 594, 567, 2457]. exploratory [1460]. explorer [1169]. exploring [1689]. explosion [262]. explosions [1119]. exponential [262]. exponentially [727]. exposed [679, 1981]. exposing [2222]. EXPRESS [826]. expression [1136, 979, 130, 251]. expressions [1057, 2052, 1613]. Expressive [2014, 647, 475, 1428, 1889]. EXPTIME [2088]. Extended [573, 370, 499, 408, 1196, 419, 894, 1200]. Extender [440]. Extending [930, 2470, 185, 1489, 1226]. extensibility [1553, 1887, 945]. Extensible [1956, 1636, 1881, 1850, 424, 2412, 1653, 1443, 639, 2116, 866]. Extension [8, 576, 1990, 1105, 569, 1080, 872, 1850, 1966]. extension-oriented [1990]. Extensional [975]. Extensionality [246]. extensions [1820, 1907, 1909]. Exterminator [1767]. External [1839, 635]. externalization [1228]. extracted [1159, 1231]. Extracting [1358, 1359, 1737]. extraction [2423, 2101]. Extreme [241, 961, 1267, 1485, 1249].

F [871, 1857, 1437]. F90 [1611]. FAB [1272]. Facile [315]. facilitates [1189]. Facilities [515]. FaCSim [2050]. factor [2425]. factories [1190]. Factorization [913, 914]. factory [1172]. fad [1206]. failures [2030]. Fair [2039, 2256, 1324]. families [2323]. family [1248]. farewell [51]. fashion [1206]. fashioned [453]. Fast [1548, 923, 200, 555, 1512, 1554, 1042, 1081, 2454, 1445, 691, 953, 2009, 201, 2238, 1060, 1793, 309, 2050, 1130, 2245, 612, 354, 952, 567, 1110]. Faster [1876, 1597, 1922, 2347, 22, 23, 2429].

Fastest [437]. FastTrack [2309]. Fault [1771, 208, 1384, 1813, 1679]. Fault-tolerant [1771]. faults [1991]. faulty [1617]. FC [277, 278]. feasible [1913]. featherweight [1056, 940]. feature [596, 1252]. features [166]. FEB [2285]. Feedback [2187, 1874, 1105, 1993, 1923, 1416, 1587, 689, 616, 2430, 103, 1815]. Feedback-directed [2187, 689, 103]. Feedback-driven [1993]. fetch [1837, 2352, 661]. Few [308]. Field [305, 91, 1592, 549]. Field-sensitive [1592]. fields [1942]. Fifteen [1633]. fifth [2479]. File [350, 1837, 1604, 2322]. files [1831, 2344, 2262]. filter [1361]. filtering [2247]. Final [654, 2458, 1942]. finalizers [740]. finance [1253]. financial [155]. find [2335]. Finding [1116, 2386, 167, 1201, 1308, 831, 1475, 1157]. Fine [1467, 1836, 2304, 2018, 339]. fine-grain [1836, 339]. Fine-grained [1467, 2304, 2018]. finger [1848]. Fingerprinting [1287]. finite [1708]. First [1724, 938, 1384, 1636, 1143, 652, 734, 2216, 1239, 2089, 2174, 1439, 141, 1958, 1433, 2403, 1184, 747, 2450, 502, 2396, 2090, 83, 666, 539]. First-class [1724, 938, 1636, 2216, 2089, 2174, 1439, 1958, 1433, 2403, 2450, 2396, 2090]. first-order [1143, 734, 502]. fisheye [1613]. Fits [229, 2094]. fix [2129]. fixed [2079, 1098, 1732, 2317, 2395]. fixed-point [2317]. fixpoint [1507]. Flapjax [2406]. flash [2238, 2239, 1842, 1589, 189, 187, 188]. Flask [2093]. Flattening [1119, 783]. flaws [1475]. flexibility [1480]. Flexible [2041, 198, 2174, 1482, 1554, 1656, 630, 2343]. FlexRAM [905]. flight [1817]. Floating [1, 2, 1041, 1039, 1040]. Floating-Point [2, 1041]. Flow [2367, 553, 109, 1055, 1502, 2365, 2370, 964, 1384, 2053, 1648, 2178, 1330, 727, 2165, 1831, 1811, 1658, 2303, 628, 84, 2356, 2080, 1954, 2331, 2188, 807, 491, 1501, 2360, 2369, 2029, 597, 662, 1109, 2470, 2305, 879, 2024, 2400, 1615, 1736, 1987, 1861, 133, 512, 250, 2304, 2208, 1597, 1779, 2368, 1032, 1277, 1275, 2229, 1855, 2335]. flow-based [2178, 1658]. flow-insensitive [2470]. Flow-sensitive [2367, 553, 2188, 807, 1501]. flows [1507, 1955]. flux [264, 2062]. Fly [239, 1382, 948, 1697, 86, 717, 464, 917, 1036, 1985, 2396]. focus [554]. Focusing [2200, 1957]. foldl [2371]. foldr [2371]. Folklore [727]. Footprint [596]. force [1358]. foreign [1372]. forest [2251]. Form [298, 1822]. Formal [2176, 163, 1394, 1498, 861, 1927, 2277, 1289, 1711, 392, 1926, 1653, 601, 1963, 590, 1661, 792, 505, 640, 177, 1097]. Formalising [1477]. formalism [748]. formalism-only [748]. Formalization [973, 702]. formalizing [1907]. formally [2380]. Format [226]. forms [886, 1143, 1489]. Forth [1486, 481, 1009, 1311, 1349, 122, 123, 97, 39, 20, 40, 276, 241, 322, 471, 323, 581, 706, 618, 541, 788, 860, 1120, 1300, 1486, 1420, 1355, 1710, 1539, 1758, 214, 215, 514]. Fortran [898, 787, 1122]. forward [1733]. forward-mode [1733]. Found [58]. foundation [1133, 2178, 2147, 979, 1958]. foundational [735]. Foundations [2013, 1952, 1330, 2148, 2383, 468]. founded [994, 1504]. fountain [1984]. Fourier [437, 1060]. FPGA [2353, 567, 1605]. FPGA-based [2353, 567]. FPH [2090]. Fractional [593]. fragmentation [838]. fragments [929, 1645]. Frame [1521, 841]. Framework [301, 355, 360, 361, 307, 106, 398, 1644, 2402, 1912, 984, 1482, 2399, 1113, 1073, 2084, 1340, 2189, 2414, 1656, 624, 1706, 330, 2261, 1102, 817, 1661, 1001, 1810, 2077, 1775, 172, 2392, 1213, 1261, 452, 703, 497, 1257, 1889, 397, 1765, 79, 451, 1555, 1182, 1259, 1348, 843]. frameworks [500, 454, 1645, 1005]. FranTk [138]. Free [2443, 978, 2388, 1474, 442, 774, 1070, 878, 133, 657, 1804, 631, 2183, 1581]. Free-Me [1581]. FreeBSD [122, 123]. freedom [1766, 2311, 2007]. Freon [1681]. Frequency [1814, 1974, 1574, 1289]. Frequency-aware [1814]. Frequent [197]. Fresh [402, 974]. freshman [2376, 1191]. FreshML [891]. friendly [1871]. frontier [82]. FRP [415]. fruit [1420]. FSE [468]. Full [2038, 1987, 2321, 1173]. Fully [136, 1962]. fun [2376]. Function [789, 825, 404, 1851, 1611, 1372, 89, 867, 931, 2400, 970, 1137]. function-based [970]. Functional [2479, 399, 421, 2081, 1149, 1863, 2079, 2372, 139, 277, 278, 640, 504, 882, 385, 2095, 402, 83, 506, 1868, 1633, 1441, 2062, 2443, 1937, 1452, 1537, 2096, 2097, 1853, 1871, 143, 2376, 1150, 2373, 633, 1859, 149, 2391, 643, 734, 2150, 155, 2086, 2215, 1442, 868, 2093, 1725, 1453, 1603, 629, 751, 2377, 1430, 1760, 2389, 2374, 151, 1854, 2085, 2371, 1138, 1613, 2388, 1849, 134, 1144, 129, 2038, 1626, 1608]. functionally [1449]. Functioning [404]. functions [1450, 1634, 863, 634, 883, 1635, 1443, 1852, 639, 1327, 1590, 1779, 2066, 641]. functors [1621]. fundamental [2162]. fundamentals [2432]. Fusion [421, 1879, 1451, 1635, 1732, 605, 641]. future [1941, 684, 1195, 1310, 2447]. futures [1479]. fuzzing [2025].

GADTs [1952, 1618, 2405]. gain

[1002, 2294]. gains [1491]. Galois [1129]. game [1517, 1191, 2252]. games [1320, 879]. gap [524]. Garbage [228, 297, 239, 786, 234, 231, 1379, 222, 233, 232, 238, 230, 227, 2319, 948, 294, 838, 742, 1414, 947, 988, 566, 953, 2009, 1059, 492, 1049, 463, 85, 1578, 1285, 837, 86, 1419, 1919, 565, 652, 2206, 772, 1472, 954, 1719, 464, 1415, 1810, 1615, 1824, 2010, 763, 839, 682, 747, 785, 768, 1579, 260, 1353, 1885, 358]. garbage-collected [1719, 358]. gates [2229]. GC [2319, 1536, 2426, 564, 839, 952, 223]. GCC [586, 1762]. GCspy [703]. Geiger [1673]. genAWeave [1167]. General [360, 361, 230, 1728, 873, 570, 2084, 1995, 1286, 1810, 1213, 2158, 1701]. general-purpose [873, 570, 2084, 1995, 1701]. generalised [1451, 1152]. generalization [1192, 933, 1942]. Generalized [2044, 1455, 1008, 2430, 1430, 1444, 1514, 1090]. generalized-LR [2430]. Generalizing [1833]. generate [919]. generated [2465, 2430, 867]. generates [1780]. Generating [847, 1600, 359, 1945, 1110, 262]. Generation [376, 1410, 422, 391, 1707, 2053, 1010, 1999, 1020, 1599, 1601, 1962, 1721, 2351, 1397, 655, 967, 604, 1053, 2175, 2329, 996, 1173, 1830, 74, 244, 829, 1253, 366]. Generational [238, 1414, 86, 1976]. Generative [1423, 409, 1726, 1214, 919]. generativity [1429]. generator [1532, 219, 1360]. Generic [411, 2070, 597, 1178, 1660, 410, 2395, 723, 1912, 976, 939, 1659, 1872, 257, 1573, 778, 1443, 881, 2077, 2125, 2217, 1257, 142, 1373, 2429, 2445]. genericity [938]. Generics [1063, 1151, 285, 2124, 1454, 1573, 1180, 973]. Geometric [422]. Geometry [1752]. Germany [539]. get [2413]. getter [623]. getting [566, 91, 1863, 1570]. giotto [1397, 353]. GIPSY [1213]. Global [349, 998, 202, 876, 1027, 548, 1507, 1205, 557, 2282, 1567, 2006, 1568, 1988, 668, 725, 339]. Global-Scale [202]. globally [1329, 600]. Go [2335, 2145, 602, 1969, 890]. goal [579]. goal-directed [579]. goes [51]. Going [706]. Goldilocks [1789]. gone [516, 542, 961]. Good [551, 1743, 2149, 2432]. Goodness [1714]. Gordon [2238]. GOTO [97, 98]. GPCE [1166, 1175, 1168, 1173]. GPGPU [2261]. gprof [1015]. GPUs [2281, 1701]. GraalBench [1092]. Grace [2410]. grain [1836, 850, 339]. grained [2056, 1685, 1467, 848, 2342, 2304, 2018, 1605]. Grammar [299, 2025, 1358, 1359, 1714, 1540, 1179, 1260].

Grammar-based [2025]. Grammars [3, 979, 2316, 527, 2396]. granting [1882]. granular [594]. Granule [1200, 1309]. Granule-oriented [1200, 1309]. graph [1017, 607, 1015, 1478, 850, 2027, 556, 2427, 1257, 1854, 1090, 1924, 174, 1264, 2262, 2343]. graph-based [1257]. graph-coloring [1090]. graphic [2067]. graphics [1092, 2287]. graphs [2078, 2041, 727, 2044, 1030, 2251, 1841, 1110]. grasshopper [1793]. great [1204, 1863]. greenfoot [1186]. Grid [793]. gridlock [566]. Gröbner [996]. Groovy [2449]. group [714]. groups [574]. Growing [526]. growth [463]. GS320 [184]. guarantee [2196, 2311]. Guaranteeing [2352]. guarantees [2468]. Guarded [737, 1811]. Guava [181]. guest [2245]. GUI [138, 1713]. guidance [1396]. guide [140, 1543]. Guided [105, 1578, 561, 2421, 1323, 598, 1043]. guiding [1245]. Guru [2444]. guts [129].

H [552]. Hacking [1920]. hammer [546]. Hancock [15]. Handling [1421, 2362, 1357, 1424, 1346, 1542, 1761, 444, 824]. Hard [1251, 2003, 1992, 379, 1233, 1574]. Hardbound [1977]. hardness [493]. Hardware [205, 231, 196, 361, 186, 1985, 1112, 2269, 1672, 1089, 436, 1217, 1697, 1285, 711, 1596, 2233, 2022, 2356, 1118, 2414, 2413, 1706, 1586, 1969, 2313, 2226, 751, 1996, 1404, 2263, 668, 1801, 756, 151, 567, 849, 1981, 2457, 1416, 1609, 360]. hardware-assist [1706]. Hardware-Assisted [231, 1416]. Hardware-managed [1112]. hardware-supported [1969]. hardware/software [849]. harm [1187]. Harmful [546, 1769, 1764, 2371]. Harmless [1527, 446]. hash [876]. hashing [1754]. haskell [2217, 2208, 2082, 2212, 2366, 2398, 153, 1132, 154, 1873, 2219, 1870, 2068, 1435, 1880, 2216, 881, 311, 2378, 2209, 416, 138, 2389, 713, 2296, 2298, 2396, 2213, 2090, 2172, 2294]. haven’t [2467]. head [1652]. healing [2223]. health [1894, 1238]. Heap [949, 226, 766, 295, 2319, 463, 1691, 762, 734, 1690, 1378, 1792, 1720, 703, 764, 1473, 1589, 2457]. heap-based [1691]. HeapMD [1691]. Heaps [224, 1418, 765, 1337]. Heat [1290]. Heat-and-run [1290]. heavyweight [1775]. help [520]. Helper [1280]. helping [1371]. here [2441]. heterogeneous [1821, 1829, 1994, 594, 1238, 2336, 1830, 1781, 1069]. Heuristics [108, 1034, 1082, 2042, 1938, 749, 800]. hidden [1674]. HIDE [1274]. hiding [992, 2222]. Hierarchical [1824, 2423, 847, 818, 2391, 242, 2455, 1496]. hierarchically [639]. hierarchies [1956, 2449, 2264, 77]. High [296, 190, 2144, 291, 1440, 895, 1930, 473, 315, 340, 1700, 683, 912, 1386, 2380, 791, 1129, 907, 612, 1767, 1867, 1679, 2290, 1427, 1574, 1892, 1993, 1083, 2294]. high- [1679]. high-bandwidth [1700]. high-frequency [1574]. high-gain [2294]. High-Level [291, 473, 2144, 1440, 1930, 340, 912, 612, 1867, 1427, 1083]. High-Performance [296, 315, 683, 2380, 2290, 1993]. High-Speed [190]. high-throughput [1892]. Higher [1032, 1450, 2075, 2379, 2051, 738, 634, 530, 2192, 2204, 1506, 1375, 1130, 1736, 2124, 646, 1866, 1733, 1958, 722, 1854, 1900, 1637, 960, 890, 982, 1957]. Higher-order [1032, 1450, 2075, 2379, 2051, 738, 634, 530, 2204, 1506, 1375, 1130, 1736, 646, 1866, 1733, 1958, 722, 1854, 1900, 960, 890, 982, 1957]. higher-rank [1637]. highly [2320]. Hilbert [879]. historical [394]. History [414, 41]. HM [872]. HMF [2089]. Hoard [194]. Hoare [1619]. HOAS [1629]. HOAS/de [1629]. hoc [1392, 1962, 2382, 1388]. HOIST [1279]. HOL [1499]. holistic [794, 2290, 2314]. homomorphism [2184]. honors [2147]. hope [1301]. Hosting [2457]. Hot [203, 570, 1571, 408, 624, 1570, 2042].

HPC [620, 2284]. HPM [1911].

HPM-sampling [1911]. HSSM [552]. HTM [2232]. HTML [2465, 748]. Hybrid [1702, 1515, 347, 1492, 61, 916, 62, 847, 1629, 1544]. HydroJ [944]. hypervisor [2245].

I/O [1627, 2376]. IA [233, 364, 204]. IA-64 [233, 364, 204]. ICFP [2099, 2002, 50]. ICSE [467]. I’d [1188]. IDE [1870, 1760]. ideal [974]. idempotency [330, 1142]. Idempotent [2254]. Identification [1264, 555, 1265]. Identifying [2393, 1691]. IDEs [2416]. idiom [1706]. idle [2237]. If [1188, 546, 2362]. IFIP [539]. IFIP/ACM [539]. ifs [1916]. ii [2015]. Ilea [1883]. ILP [364]. ILP-based [364]. imbalances [1901]. Immix [2009]. immutability [1466]. Impact [2043, 1671, 2115, 148, 1267, 2466, 824, 764, 830]. Impala [910]. Imperative [1953, 1450, 2379, 554, 1506, 2083, 1490, 1449, 2328]. Implement [416, 1757]. Implementation [532, 412, 285, 2321, 16, 9, 27, 2478, 1819, 446, 480, 1730, 2278, 2366, 68, 853, 915, 2233, 1983, 654, 651, 692, 797, 1042, 900, 1784, 1058, 2111, 17, 1960, 2283, 1713, 1605, 2485]. implementations [1499, 1741, 2408, 1954, 87, 2296, 2298]. implemented [1389]. Implementing [925, 239, 2147, 1174, 625, 2403, 693, 1138, 1155, 1209, 1644, 1656, 1529]. implications [1992, 2221]. Implicit [1878, 1874]. Implicitly [2073]. Implicitly-threaded [2073]. Importance [290]. Important [30, 52]. impredicativity [1637]. improbable [1344]. Improved [1776, 74, 1487, 1545, 1885]. Improvement [426, 1635, 610]. Improving [1055, 1471, 1037, 220, 1131, 2048, 1216, 1975, 156, 904, 2058, 1615, 208, 572, 1376, 261, 1378, 662, 1655, 800]. impulses [882]. in-memory [2427]. inadequate [328]. included [2219]. Inclusion [396]. inclusive [2140, 135]. incoherence [1276].

Incommunicado [698]. incompatibilities [2393]. incomplete [1235]. inconsistency [1807]. Increasing [675, 507, 1408]. Incremental [449, 1493, 972, 729, 767, 632, 1580, 564]. incrementalisation [2100]. Incrementalization [1481, 931, 1795]. Incrementalized [288, 289]. independence [2198, 499]. independent [1010, 1565, 932, 685, 1179, 1260, 921]. Index [209, 316, 2475, 427]. indexed [1432, 714, 533, 268]. indexing [2381, 2028]. indices [78]. Indirect [240, 816, 1975]. individual [1581]. Inducing [1082]. Induction [29]. Inductive [1856, 535, 2248, 1947, 1790, 1734]. inductively [1445]. Indus [1531]. Industrial [2270, 365, 2399, 591, 1144]. industrial-strength [2399]. industry [1253]. inection [1089]. inefficiencies [2122]. inexpressivity [1729]. inference [2127, 1087, 1807, 1898, 1744, 565, 869, 1067, 1618, 2089, 2174, 2305, 1439, 1524, 2424, 249, 1720, 133, 512, 1425, 1514, 1778, 1437, 2067, 1779, 2405, 527, 2129, 1942, 1637]. Inferred [2415]. Inferring [783, 2034, 1904, 804, 2117]. influence [2364, 131]. Influences [824]. Influential [960]. Information [30, 389, 390, 43, 52, 99, 124, 216, 125, 217, 279, 324, 476, 280, 325, 477, 544, 583, 621, 545, 584, 622, 962, 57, 133, 512, 1502, 2365, 2370, 964, 2133, 2367, 912, 2303, 628, 1175, 2356, 1954, 91, 491, 2360, 2369, 868, 2305, 2024, 992, 2304, 2208, 764, 2368, 1275, 2166, 2229, 450, 1274]. information-flow [1954, 2208]. infrastructure [1390, 1128, 1148, 256, 1222, 1274]. infrastructures [587]. inheritance [1718, 2408, 454, 2407, 1932, 1184, 1662]. initial [244]. initialization [2173]. initiative [534]. Injecting [2140]. injection [1526, 1770]. Inlining [108, 789, 825, 92]. Input [1118]. insensitive [2470]. inspecting [815]. inspection [1538, 511, 1444]. instances [687]. instantiation [84, 172]. Instruction [1612, 20, 364, 346, 1684, 1820, 1569, 672, 1845, 2044, 104, 1554, 21, 1837, 2352, 1105, 1051, 857, 75, 1698, 1025, 1988, 1604, 1106, 1927, 73, 1838]. Instruction-coated [1612]. Instruction-Level [346, 1845]. instruction-set [1105]. instructions [1406, 1706, 598, 855, 603]. instrumentation [1236, 2153, 1383, 1775, 1401, 1740]. Instrumented [301]. integer [73]. integers [1548]. Integrated [1816, 355, 1700, 366, 5, 604, 1223, 2221, 2328]. Integrating [2356, 685, 936, 2134, 2314]. Integration [376, 746, 1096, 1409]. integrity [2165, 665, 791]. Intel [1104]. intelligence [1858]. Intelligent [49, 905]. intended [2428]. intensional [1935, 1213, 136]. Intensionality [246]. Intensive [13, 2238]. Intentional [1668]. Inter [609, 783, 1458, 1883]. inter-language [1883]. inter-object [1458]. inter-procedural [783]. Inter-task [609]. interact [942]. interacting [1747, 2120]. Interaction [1465, 369, 1186, 590, 534, 368]. Interaction-based [1465]. Interactive [195, 1177, 643, 752, 1237, 2379, 1244, 1861, 610, 2252]. interchange [1018]. interconnect [2049]. interdependent [2368]. interesting [1751]. Interface [1829, 1321, 675, 906, 2226, 785, 599, 1094, 1882]. interfaces [446, 1700, 1008, 2260, 1184]. interference [985, 1107]. interleaved [1561]. interleaving [1442, 1675]. intermediate [1316, 1718, 2456, 267, 744, 87, 1765, 1064]. intermittent [1991]. internal [459]. International [2476, 36, 539, 540, 2479, 467, 470, 2059]. Internet [338, 911]. Internet-scale [338]. interoperability [1467, 1238]. interoperable [635]. interposed [2442].

Interposition [862]. interpretation [500, 984, 985, 724, 998, 1339, 1826, 2400]. interpretation-based [984]. interpretations [1724]. interpreter [1873, 1762, 661]. Interpreters [42, 1612, 2074, 816, 924, 1582, 1946, 649, 960]. Interprocedural [1023, 1024, 1029, 741, 1030, 1750, 2101, 2474, 68, 91, 1339, 1048, 997, 452, 79]. interrupt [1400]. Interruptible [1519]. interrupts [2022, 2354]. Intersection [147, 1642, 1869]. Interval [357, 717]. Interval-Based [357]. Intriguing [482, 465]. intro [2147]. introduce [1191]. introducing [2341, 2104]. Introduction [2004, 708, 465]. introspection [2366, 866]. Introspective [1695]. intuition [1245]. invariance [1738]. invariant [1964, 996, 1795, 2472]. invariants [2442, 1794, 1899, 2331, 1675]. inverse [924]. inversion [1780, 754]. inverter [1362]. Inverting [754]. investigation [1267]. investigations [1766]. invitation [1210]. Invited [428, 532, 523, 140, 151]. Invokeinterface [446]. invoking [1885]. involving [2388]. IP [1700]. Ironman [273]. irreconcilable [1715]. irreducible [1356]. Irregular [363, 2045, 2250, 76, 608, 571, 2294]. ISMM [36, 470, 2059, 2060]. ISO [1249]. isolated [2187]. isolates [698]. Isolating [1664]. isolation [179, 2225, 805, 1368, 2235, 1774, 1891]. ISOLATOR [2235]. Isomorphisms [976, 732]. ISSTA [540]. issue [708, 465, 1835]. Issues [1825, 373, 1766, 283, 284]. Itanium [1280]. Itanium-2 [1280]. iTasks [1861]. iteration [801]. Iterative [1808, 878, 2015]. Iterative-free [878]. iterators [2063, 1519]. ITRON [599]. IX [2476]. iXj [1237]. IXP [796].

J [1464]. J& [1642]. J2EE [1204, 1225, 1233]. Jalapeño [159, 116, 117]. January [2482, 54, 2480, 319]. Jargons [4]. JastAdd [1881]. Java [34, 1477, 927, 446, 1321, 836, 439, 995, 1315, 689, 1817, 2041, 964, 440, 181, 294, 838, 1056, 1345, 576, 1665, 1357, 212, 213, 37, 1464, 448, 462, 1651, 2411, 1256, 1237, 442, 820, 463, 1217, 1301, 35, 1538, 282, 281, 949, 1754, 64, 165, 1646, 71, 835, 946, 179, 2118, 1537, 761, 2112, 705, 86, 239, 765, 941, 942, 1881, 1789, 1649, 81, 573, 2003, 1846, 1884, 2121, 1244, 1063, 1459, 2104, 1898, 1419, 692, 168, 460, 1169, 865, 1882, 1818, 623, 2413, 175, 943, 128, 1065, 690, 2122, 1266, 704, 791, 2042, 464, 1080, 1109, 2470]. Java [1904, 1650, 156, 2468, 1343, 453, 1546, 1576, 615, 444, 661, 1404, 8, 1812, 1660, 177, 770, 22, 23, 441, 80, 1889, 2116, 749, 834, 160, 223, 295, 764, 1795, 397, 833, 682, 2129, 1892, 1457, 1583, 747, 1908, 451, 819, 173, 1883, 1466, 1942, 166, 398, 1231, 1479, 2100, 2109, 478, 248, 161, 940, 2425, 1360]. Java-based [1360]. Java-like [995, 1315]. JavaBeans [590, 1182, 1259, 1231, 862]. Javana [1650]. Javari [1466]. JavaScript [2303, 1436, 2452, 1740]. Jazzing [1127]. Jeannie [1882]. Jedd [1080]. Jiazzi [453]. JICC4 [34, 35]. Jini [550]. JIT [61, 749]. JIVE [1244]. jMock [1154]. job [1205]. Jobscheduling [206]. Join [2103]. joining [920]. joining/leaving [920]. Joint [1828, 668]. jokers [1951]. Jolt [2107]. JOY [860]. JQuery [1157]. JRA [1161]. JSE [440]. jStar [2112]. JTL [1646]. JUDO [64]. jump [1863, 514, 625]. jumps [1975]. June [2478, 2481, 2483, 539]. jungle [1371]. Jungloid [1371]. just [969, 602, 2339, 864, 943, 2042, 1655, 451, 819, 175]. just-in-time [2339, 864, 943, 2042, 1655, 451, 819, 175]. JVM [159, 2043, 2412, 888, 653]. JVMs [1127, 1108].

Kahn [1510]. keep [1068]. Kendo [2228]. kernel [1405]. Key [200, 201]. Keynote [103]. KGOL [793]. kids [2376]. Kill [1071]. Kill-safe [1071]. kind [276, 2124]. Kingdom [2482]. KISS [1068]. Kleisli [129]. knots [2394]. know [2139]. knowledge [59, 1712, 1006, 793].

l2 [1816, 2230]. Lacking [58]. LALR [1011]. LAM [899]. LAM/MPI [899]. lambda [975, 1772, 1742, 498, 1617, 2383]. lambda-calculus [1742]. Laminar [2304]. land [620]. Language [377, 320, 15, 1035, 2358, 218, 705, 292, 6, 746, 956, 2168, 417, 247, 2215, 285, 314, 932, 14, 2485, 361, 8, 2451, 16, 315, 332, 313, 2478, 2302, 436, 480, 1126, 2140, 2136, 2137, 2138, 2437, 873, 321, 1531, 1550, 2416, 585, 1316, 1718, 1087, 1772, 826, 711, 1646, 2052, 1095, 735, 1500, 1452, 1365, 1977, 1714, 715, 950, 143, 1392, 1859, 2456, 939, 794, 530, 267, 1422, 559, 2301, 1535, 2064, 2192, 2430, 1158, 1215, 1632, 696, 1783, 2434, 676, 791, 900, 1896, 879, 1725, 1475, 1453, 1716, 1915, 1918, 2406, 245, 1252, 1742, 1867, 1508, 2433]. language [1327, 1812, 1771, 2383, 861, 1097, 87, 1428, 2158, 1179, 1260, 17, 824, 138, 2389, 2159, 1427, 2160, 1765, 2368, 520, 1489, 1373, 2134, 2001, 243, 1364, 1922, 1517, 1850, 1883, 2186, 1083, 1523, 1449, 418, 1530, 973, 2328, 1966, 793, 790, 360, 2100]. Language-based [2358, 1500, 1449]. Language-independent [932, 1179, 1260]. Language-specific [705, 2416]. Languages [2476, 2482, 54, 41, 58, 656, 126, 60, 46, 393, 2153, 2163, 1195, 1310, 391, 100, 2477, 2439, 995, 1315, 59, 1532, 1563, 2053, 2141, 526, 2142, 2143, 1330, 1316, 2144, 2074, 1640, 1562, 2145, 127, 438, 1495, 2146, 2147, 2084, 2421, 2339, 638, 1446, 2148, 1131, 1226, 2315, 1975, 766, 1719, 1516, 2151, 1375, 1363, 2154, 1130, 524, 629, 2126, 2433, 779, 1866, 649, 2167, 2454, 1178, 2156, 2157, 1490, 2161, 1493, 1138, 960, 2162, 552, 2455, 1005, 729, 1125, 33, 358]. Large [15, 1559, 809, 1361, 2023, 2473, 895, 751, 660, 1196, 1306, 1086]. Large-Scale [15, 2023]. larger [727]. latency [1051, 1287]. latent [2131, 1664]. later [1202]. lattice [1819]. lattice-based [1819]. Layer [352, 2239, 844]. layered [845]. layering [2455]. Layout [395, 2045, 1378, 733, 497, 1587]. laziness [1876]. Lazy [957, 1875, 229, 1537, 2182, 489, 1047, 2061, 1733, 102, 144, 987, 143, 2373, 633, 1022, 751, 2389, 2326, 2210]. LCTES [211, 381]. LCTES’02 [384]. lead [2184]. leads [1000]. leads-to [1000]. Leak [2243, 1677, 1811, 628, 1281, 807, 1719]. leakage [1602, 857, 1815, 1274]. leakage-aware [1602]. leaks [2106, 2333]. learned [163, 1032]. Learning [39, 40, 1998, 1816, 1189, 1220, 1187, 896, 2134, 800, 2314, 2257]. learnt [2291]. leaving [920]. left [1549, 1951]. legacy [496, 2117]. legends [2447]. lenses [1961, 2098]. Less [203, 952]. lessons [163, 516, 542, 961, 141, 1032]. let [1439]. let-polymorphism [1439]. Letter [51, 380, 93, 963, 381]. Letters [959]. letting [2240]. Level [291, 195, 346, 7, 473, 410, 908, 1288, 912, 1829, 189, 2195, 2144, 699, 1440, 1845, 942, 2022, 1403, 1167, 2239, 1762, 1072, 2284, 857, 906, 1893, 75, 1784, 658, 2353, 612, 1930, 1944, 1867, 897, 1972, 1427, 1765, 2422, 1083, 339, 340, 77]. levels [1183, 1262]. leveraging [1290]. Lexer [56, 55]. Lexical [1066, 519]. liberal [2142]. liberated [1893]. liberty [1083, 1094]. liberty-based [1094]. Libraries [6, 1878, 2217]. Library [278, 2218, 1469, 1389, 2398, 515, 867, 2418, 277, 1211, 2329, 2208, 921]. LIFE [2352, 1191]. lifecycle [590]. lifetime [237]. lifetimes [2340]. Lifting [1478, 1946]. light [2208]. light-weight [2208]. Lighting [413]. Lightweight [2307, 1937, 2207, 1732, 940, 1256, 154, 956, 2310, 2107, 1212, 1850]. like [995, 1315, 1516, 1731, 2120, 1866, 1212, 641, 546]. Limitations [426, 1357]. limiting [2425]. limits [1971, 1894, 798]. line [67, 1127]. Linear [795, 2194, 1029, 554, 633, 2070, 171, 2180, 997, 2263, 996, 142, 2007, 419]. linear-time [142]. linearizable [2018]. Lines [309, 1793, 1162]. linguistic [1332, 1659]. Link [1114, 1622, 1027]. Link-time [1114]. linked [1504, 2296, 2298, 2038]. linked-list [2296, 2298]. linking [1587]. Linnaean [2151]. Linux [1868, 1405, 2096]. Liquid [2021]. Lisp [2453, 395, 1362, 1516, 406]. Lisp-like [1516]. list [2296, 2298]. Lists [102, 1879, 1504]. LiteRace [2310]. Live [37, 2347, 358, 38, 555, 1915, 1568, 1536]. Live-range [2347, 555, 1568]. liveness [2177, 764]. Living [1915, 1914]. LL [219]. Load [373, 572, 2255, 1842, 1901, 333]. loaded [615]. loading [177, 79]. Local [988, 2196, 804, 1507, 886, 2127, 765, 1342, 688, 249, 1404, 1337, 668]. Locality [303, 1282, 1722, 197, 2016, 1578, 813, 261, 1046, 1680, 682, 1496]. localization [1813]. localizing [726]. locate [870]. locating [2333, 1805]. locations [1338]. Lock [1745, 690, 774, 2203, 2006, 1070, 2311, 657]. lock-based [2203, 657]. lock-free [774, 1070, 657]. lock-freedom [2311]. locking [1657]. locks [1056, 2034, 690, 334]. LOCKSMITH [1577]. LogGPS [344]. Logic [1736, 306, 487, 1502, 1499, 1334, 928, 1934, 2197, 2178, 781, 1335, 1729, 1228, 1933, 2443, 2359, 1145, 717, 509, 529, 504, 1333, 1932, 1138, 1727, 248, 981]. Logic-flow [1736]. logical [782, 1450, 1398, 1936, 1946, 1613]. Logics [246]. logistic [1654]. logTM [1704]. London [2482, 54]. long [953]. longer [2291, 1676]. look [1011, 974]. look-ahead [1011]. Lookahead [2352]. Looking [1199, 1304]. looks [546]. Loop [188, 351, 605, 1347, 1018, 2055, 2273, 1487, 1545, 1394, 187, 996, 1426]. Loop-dead [1347]. looping [1739]. Loops [312, 363, 26, 848, 341, 25, 82, 1804, 1380, 1002, 2057, 77, 74]. loose [1512]. Lost [1907]. love [2291, 1199, 1304]. Low [291, 1281, 2294, 742, 1612, 2195, 699, 1440, 1270, 2022, 91, 1969, 2313, 1944, 1844, 1679, 834, 1678]. low-cost [91, 2313, 1678]. low-end [834]. Low-Level [291, 2195, 1440, 2022, 1944]. Low-overhead [1281, 1679]. Low-pain [2294]. low-power [1270]. LR [2430, 1541]. LTL [1256]. Lustre [2341, 845].

MA [2476]. Macaulay [1012]. Mace [1783]. Machine [97, 62, 107, 1816, 2292, 1471, 1647, 1010, 447, 816, 1021, 98, 580, 1673, 1007, 663, 2468, 827, 896, 461, 2201, 800, 2314, 2257, 69, 179, 1404, 1812, 833, 747, 552]. Machine-Adaptable [107]. machine-independent [1010]. machine-learning [2314]. Machines [308, 1559, 1940, 1674, 942, 1031, 1286, 514, 1598, 2123, 1924, 2455]. Macintosh [1004]. macro [552]. macro-architecture [552]. MacroML [409]. Macros [409, 526, 636]. made [1402, 627, 891]. main [1291]. mainstream [1454, 1517]. maintaining [1329]. maintenance [2464, 1286, 1179, 1260]. MaJIC [578]. major [1728]. majors [2139]. Make [142, 705, 1645, 1135, 1142]. Making [1913, 2458, 2216, 1792, 1130, 1276]. malloc [774]. malware [1674, 1753]. manage [1868, 1290, 1463]. managed [1844, 2264, 2244, 1112]. Management [36, 470, 2059, 347, 398, 858, 2012, 820, 2295, 1840, 786, 577, 1681, 1472, 596, 2332, 771, 1291, 1603, 1171, 1100, 1972, 397, 510, 2265, 1182, 1259, 2011, 2253, 1283, 1594]. Manager [225, 2426]. managing [910, 1980, 666]. mandatory [1900]. Manipulation [2484, 319, 716, 1168]. Manticore [2073]. manual [589, 897]. Many [1610, 623, 2280]. many-core [2280]. manycore [2457]. map [631]. Maple [1012]. Mapping [168, 1699, 2257, 2288, 1834, 848, 2314, 1976, 1412, 755]. mappings [2239].

Mar [1920]. Marine [1193]. mark [948, 2009, 1285, 952]. Mark-copy [952]. mark-region [2009]. mark-sweep [1285]. Marketing [2156]. marking [2401]. markup [1493]. Mars [706]. Marshalling [877, 876]. mashups [1871]. Masked [2173, 1548, 1955]. Massachusetts [2480]. masses [1151]. massively [1064]. Matching [405, 923, 1474, 2178, 714, 251, 2200, 1491, 944, 2211, 521, 1850]. Maté [663]. math [1203, 867]. mathematics [875, 752]. MATLAB [578, 814, 7]. matrices [913, 914, 332]. Matrix [1488]. Matrix-based [1488]. matter [1494]. matters [1250, 1143]. maximizing [1242]. Maximum [2232, 142]. maximum-weightsum [142]. may [263, 749]. Maya [576]. McErlang [1859]. MDA [1183, 1262]. MDA-levels [1183, 1262]. me [2362, 1951, 2213, 1581]. meanings [1779]. Means [473, 2193]. measurement [2337, 2274]. Measuring [2364, 1107, 1974]. mechanical [1530]. mechanism [678, 1346, 1239]. mechanisms [1468, 1756]. Mechanized [1629, 2075, 1735]. Mechanizing [1446]. MEDASL [1158, 1215]. Meeting [958]. Mem [667]. members [2442]. memoization [719, 2387]. memories [2040, 186]. Memory [36, 470, 1293, 194, 296, 229, 837, 290, 2059, 225, 230, 358, 1931, 2269, 1552, 667, 1559, 681, 1563, 2366, 2013, 1677, 2106, 2202, 820, 1288, 2295, 1361, 1768, 2341, 2238, 949, 1811, 607, 1840, 1703, 1705, 758, 835, 2453, 1702, 2286, 761, 2227, 2233, 2312, 786, 905, 913, 914, 1478, 577, 1919, 2203, 1551, 2168, 1281, 807, 1656, 1472, 1111, 1842, 2332, 1719, 1417, 2251, 2122, 1051, 771, 1291, 76, 335, 156, 757, 2349, 1343, 907, 2313, 1070, 1704, 336, 2320, 1692, 1839, 1767, 2333, 2426, 2321, 595, 733, 2427, 759, 78, 2264, 152, 510, 2265, 1589]. memory [919, 339, 2283, 2011, 1976, 2036, 680, 1353, 1367, 1676, 77, 1607, 1283, 2252]. memory-constrained [949]. memory-reference [835]. memory-safe [2321]. MEMS [182, 183]. MEMS-based [182, 183]. Mercury [1681]. Merge [1994]. merging [1394, 1541]. Merlin [2305]. Message [543, 47, 488, 766, 1748]. Message-Driven [543]. message-passing [488, 1748]. messages [1806]. Meta [892, 189, 1160, 648, 800, 585, 1446, 1629, 713, 1226]. meta-language [585]. Meta-level [189]. Meta-programming [892, 1160, 648, 713]. meta-reasoning [1629]. meta-theory [1446]. metadata [1089]. MetaEdit [1173]. metamorphic [526]. metaOCaml [1174]. metaphor [2435]. metatheory [1926, 1735]. Method [1126, 1654, 1365, 1241, 2153, 162, 253, 173, 450]. Method-specific [1654]. Methodology [1198, 1305, 5]. Methods [290, 626, 254, 623, 505, 245, 640, 1900, 270, 1289]. Metric [48]. metrics [167, 941]. metronome [838, 1414]. Micro [1459]. microarchitectural [942, 1688, 1971]. microcontroller [1344, 1401]. microdrivers [1983]. microfluidics [2012]. microkernel [2380]. MicroMED [1710]. Microphase [1885]. microprocessor [1678]. microprocessors [2259, 1289]. Microsoft [1204]. Middle [367]. Middleware [128, 164, 376, 371, 373, 1263, 910, 1183, 1262, 367]. middleware-specific [1183, 1262]. Migrating [1254]. Migration [862, 1469, 1697, 2348, 1396]. migrations [1588]. MILAN [355]. mile [2458]. Million [309]. millions [1793]. Min [1072]. Min-cut [1072]. mind [516, 2155, 2162]. mindstorms [471]. minimal [2232, 1969]. minimise [2055]. minimization [2087, 244]. Minimizing [1536, 2429, 1353, 548, 1588]. minimum [2251]. Mining [485, 895, 2281, 1371, 1778, 1666, 1163]. Minneapolis [2477, 36]. Minnesota [36]. mirrors [1467]. miss [2230, 1283]. Misses [228, 572]. missing [1622]. mistakes [1135, 1998]. Mitosis [1390]. Mixed [534, 2234, 1532, 2417]. mixed-environment [2417]. Mixed-initiative [534]. Mixed-mode [2234]. Mixin [2091, 454, 1439]. mixin-based [454]. Mixing [2105]. mixins [1235]. MJ [946]. ML [2398, 404, 2091, 2071, 871, 1053, 1735, 1725, 1866, 512, 1425, 2067, 2397, 1622, 407, 1054]. ML-like [1866]. MLF [1438, 1857]. MMU [1840]. MN [2477]. Mobile [255, 298, 1234, 1218, 1092, 1388, 256, 982, 1588, 1218]. Mobile-Code [298]. Mobile-D [1218]. mobility [1477, 1827]. Mock [1246, 1154]. mockObject [1212]. mockObject-like [1212]. modal [1728, 1729, 1516]. mode [1733, 2234]. Model [355, 375, 2040, 1242, 1202, 2171, 1432, 2258, 1728, 2041, 1126, 1147, 2013, 2178, 2295, 488, 826, 1510, 1663, 1188, 1329, 1322, 1859, 1320, 1063, 922, 717, 1641, 344, 1994, 1003, 156, 1343, 612, 1352, 2039, 2424, 593, 2015, 1097, 2128, 2336, 1900, 920, 1485, 1222, 974, 1713, 799, 1258, 1496, 346]. model-driven [799, 1258]. Modeling [1155, 1209, 1176, 1158, 1215, 1093, 2199, 1162, 1461, 1999, 1963, 1216, 1689, 1103, 162, 1688, 1173, 1083, 1530]. Modelling [1627]. Models [378, 474, 847, 1294, 2202, 1768, 488, 1829, 1460, 1599, 1925, 2154, 900, 1965, 1463, 1921, 2264, 237, 1134, 843]. modern [1728]. Modest [549, 271]. modifying [1773]. Modular [2442, 1897, 68, 1628, 1723, 454, 1448, 1584, 2175, 1513, 639, 1746, 1900, 1180, 2053, 957, 1933, 165, 1791, 1553, 2430, 676, 2401, 945, 2180, 2361, 1626, 39, 40]. modularity [2175]. modularizing [2130]. module [946, 2091, 140, 1908]. Modules [411, 266, 410, 738, 1877, 408, 1439, 2199, 1620, 1621, 2450]. Modulo [2340, 2342]. MoJo [615]. molecular [1967]. monad [146, 1442]. Monadic [1139, 1451, 10, 712, 145, 2207, 1784, 1547]. Monads [632, 9, 2216, 1639, 642, 2394, 498].

Mondrian [680]. monitor [2413].

monitoring [1089, 263, 743, 1557, 1673, 2353, 1546, 1801, 756]. monitors [1913, 1217]. monotonic [1602]. Monte [253]. Monte-Carlo [253]. Montréal [2479]. MOP [1912]. morally [1512]. mornin [1188]. Morphable [362]. MorphJ [2014]. morphology [1149]. Morton [913, 914, 332]. Morton-order [332]. Morton-ordered [913, 914]. Mostly [947, 774, 238, 690]. Mostly-Concurrent [238]. motion [1047, 2326]. motivate [2154]. mouse [1920]. moving [2435]. MPI [908, 909, 899, 2277, 2276]. m}pilation [2480]. MPIWiz [2276]. MPSoC [1830]. MR [1012]. MST [607]. Much [1928, 2250]. mud [699]. Multi [352, 409, 2260, 1137, 224, 755, 1345, 2278, 1829, 2279, 267, 1323, 1101, 1516, 1893, 1699, 1994, 1716, 594, 2300, 1331, 1988, 841, 1972, 1380, 2348, 1993, 2241, 1781, 2257, 77, 2425, 691, 803]. multi-agent [1345]. multi-cluster [841]. Multi-core [2260, 1829, 2279, 1994, 2348, 2241, 1781, 2425]. multi-cores [2257]. multi-dimensional [1380]. multi-dispatching [691]. multi-granular [594]. multi-interfaces [2260]. multi-language [267, 1716]. multi-level [1893, 1972, 77]. Multi-objective [755]. multi-place [2278]. multi-process [1323]. Multi-Processor [352]. Multi-return [1137]. Multi-Stage [409]. multi-staged [1516]. multi-task [1101]. Multi-Threaded [224, 1699, 2300, 1988, 1993, 803]. multi-tier [1331]. multicluster [818]. multicode [1224]. multicore [2273, 1995, 2414, 2332, 2017, 2349, 2378, 1991, 2234, 2294]. multicores [2350]. multidimensional [169, 2015]. MultiJava [165, 945]. multimedia [75, 1839, 593]. multimethods [1363]. Multiparadigm [221]. Multiparty [218, 1949]. multiplayer [2252]. Multiple [1905, 687, 2132, 576, 1357, 1718, 165, 2408, 587, 1909, 1491, 2407, 2120, 2263, 87, 2368, 1138, 1662, 1289, 1607]. multiple-dispatch [576]. multiprocessing [2227]. Multiprocessor [186, 2292, 294, 1059, 1682, 2256, 2226, 2201]. multiprocessors [1028, 1973, 1834, 902, 907, 2264, 2272, 1980]. Multitasking [447]. Multithreaded [194, 232, 207, 2020, 2410, 575, 991, 2033, 904, 1546, 1808, 917, 331, 2274, 1781, 2109, 1091]. Multithreading [206, 903, 777, 2228, 1680, 1280]. multivariate [1733]. musical [1234]. must [1749]. Mutable [247]. mutandis [1328]. mutation [1986]. Mutatis [1328]. mutator [2009]. mutators [1810]. mutual [1931]. mutually [2395]. MxYacc [1532]. Mython [2451].

nail [546]. Naked [475]. name [969, 1890, 1959, 885, 960]. nameplate [1441]. names [255, 648, 1726]. NAND [1842]. nanopass [1148]. nanoscale [1693, 2221]. narrowing [1445, 528]. narrowing-driven [1445, 528]. Native [112, 1882, 113]. natural [629]. naturalistic [966]. navigate [1371]. NB [2285]. NB-FEB [2285]. NDL [1095]. near [1595]. near-optimal [1595]. necessity [648]. Need [426, 1610, 2146]. needs [591]. NePalTM [2283]. nesC [794]. Nested [312, 44, 2266, 137, 341, 1704, 1642, 1255, 2283, 45]. nests [2273]. net [242]. netlists [2095]. Network [656, 49, 365, 263, 1700, 342, 1386, 796, 675, 906, 1784, 1803, 2024, 2245, 337, 1094, 755, 1091, 1594]. network-on-chips [1803]. Networked [192, 329, 794, 856]. Networks [190, 1511, 1510, 1270, 1785, 663, 1802, 2093, 1388]. neural [242, 49]. new-age [453]. Newlines [55, 56]. Newly [380]. News [120, 317, 31]. Next [376, 1495, 1184, 1517]. Next-Generation [376]. NEXTGEN [1889]. NGMF [1257]. NIAM [48]. Ninth [2476]. NIPL [326]. nitrO [827]. NixOS [2096]. NNPS [49]. No [969, 1916, 1972, 2209]. NoC [1566, 1834]. node [1286]. Nomadic [256]. Non [153, 2385, 996, 804, 1484, 607, 893, 950, 2373, 985, 638, 2170, 1939, 507, 673, 1746]. non-aliasing [804]. non-blocking [2170, 1746]. non-deterministic [2373]. non-interference [985]. Non-linear [996]. non-null [950]. non-orthogonal [607]. Non-parametric [2385]. non-size [507]. Non-stop [153]. non-strict [893, 638]. non-termination [1939]. non-uniform [1484, 673]. nondestructive [2078]. noninterference [1500, 1326, 2361]. Nonintrusive [1401, 294]. nonlinearity [1142]. nonstandard [1724]. normal [886]. normalisation [975]. normalization [2105]. Notation [1895, 423, 62, 1540]. Notes [1194, 1303, 969]. nothing [1879]. Notices [30, 52, 1486, 1539]. Notification [372]. November [2476]. Null [196, 1903, 950]. NUMA [2426]. NUMA-aware [2426]. Numbering [1143, 557, 998, 141, 1751]. Numbers [1, 2, 1041, 1039, 1040, 1121]. numeric [1340]. numerical [171].

O [1627, 2376]. Object

[2477, 235, 158, 1035, 281, 24, 865, 128, 376, 951, 47, 1171, 473, 458, 29, 546, 474, 443, 1089, 1502, 1126, 582, 683, 1531, 957, 736, 180, 1484, 1316, 2023, 1087, 575, 1895, 1460, 92, 2408, 950, 1899, 438, 1177, 741, 176, 2114, 220, 1581, 1186, 1454, 624, 1346, 1975, 1455, 164, 617, 1263, 219, 1363, 1755, 162, 590, 944, 1481, 2409, 171, 936, 449, 1458, 1211, 1181, 2126, 635, 2427, 2173, 178, 2466, 2107, 2461, 1589, 237, 2457, 768, 1523, 270, 1643, 1264, 1265, 328, 790, 1125, 1587, 803, 1421, 327]. Object-Models [474]. Object-Oriented [1035, 281, 47, 473, 29, 546, 2477, 158, 1171, 458, 1502, 582, 1484, 1316, 2023, 1087, 575, 1895, 1460, 2408, 950, 438, 1177, 220, 1454, 624, 1346, 1975, 1455, 617, 219, 944, 936, 1181, 2126, 178, 2466, 1523, 1264, 1265, 790, 1125, 803]. objective [2392, 755]. Objects [543, 519, 1184, 1235, 1897, 980, 2341, 456, 1623, 1154, 1246, 1846, 455, 762, 1243, 815, 686, 1003, 684, 2120, 1536, 157, 475, 1255, 1652, 768, 2018, 862]. oblivious [1754]. observational [2076, 1895]. Observe [388]. obviously [2242]. OCaml [2399, 1631]. OceanStore [202]. ocsigen [2402]. October [2477, 36, 958]. Off [95, 96, 67, 2269, 568, 1757, 1127]. Off-and-On [95, 96]. Off-line [67, 1127]. off-the-shelf [2269, 1757]. Offline [1800, 1161]. offloading [1077]. offset [1099]. old [453]. old-fashioned [453]. Older [747, 652]. Older-first [747]. omission [1805]. omniscient [1910]. on-chip [1511, 1800, 673]. on-demand [1554, 685]. On-the-Fly [239, 717, 1036, 1382, 948, 1697, 86, 464, 917, 1985]. one [2146, 630]. one-sided [630]. Online [689, 525, 1570, 1801, 1677, 1286, 2230, 2236, 1289]. only [2292, 2362, 2365, 2363, 2442, 2366, 2367, 2295, 2358, 2443, 898, 2440, 2359, 2297, 748, 2360, 2446, 895, 900, 2441, 2364, 901, 896, 2368, 899, 2444, 2296, 2361, 2445, 546, 2294, 2293]. onto [1118, 848, 1699]. ontology [1198, 1305]. onward [1312]. OO [1933]. OOLACA [1211]. OoLALA [171]. OOP [1220]. OOPAL [936]. OOPSLA [1166, 1168, 482, 1312, 465]. OOPSLA/GPCE [1166, 1168]. Open [472, 2266, 165, 2199, 2066, 1196, 1306, 243, 828, 1222]. open-nested [2266]. opening [1724]. OpenMP [2282, 2261]. OpenUH [2282]. Operating [2476, 2440, 207, 609, 1968, 1290, 1435, 2224, 662, 1671, 1984, 666, 793]. operation [818]. Operational [1716, 144, 2202, 259, 1365, 1930, 2205, 1662]. operationalization [1493]. operationally [1639]. operations [758, 1340, 1118, 2297, 690, 1657, 343]. operator [915, 1000]. operators [2181]. Opportunities [2259, 329, 798]. Optimal [308, 604, 73, 2347, 1595, 2077, 366]. optimisation [1403]. Optimising [1377]. Optimistic [1821, 893, 1786, 1989, 452, 1473]. O}ptimization [2480, 105, 128, 1016, 307, 106, 2455, 1013, 1293, 159, 689, 63, 1487, 1545, 2187, 810, 1578, 2045, 568, 1114, 92, 1828, 814, 300, 831, 1081, 1595, 867, 1491, 795, 2344, 2261, 533, 2375, 1001, 1528, 2469, 1430, 444, 661, 595, 1100, 2015, 562, 2451, 2110, 1407, 1347, 1765, 103, 79, 800, 1224, 451, 2191, 1555, 1981, 340, 799, 670, 1814]. optimizations [836, 64, 1709, 1020, 1061, 2282, 943, 676, 2327, 509, 1570, 2345, 811, 817, 1988, 497, 1801, 510, 2230, 1927, 339, 830, 852]. optimize [912, 1007]. optimized [1600]. Optimizer [112, 2016, 113, 1779]. Optimizing [1035, 911, 387, 1611, 816, 372, 6, 1551, 368, 405, 1053, 931, 374, 1560, 1115, 1843, 2057, 369, 1607, 2428, 386, 2023, 330, 857, 1026, 1046, 70, 1054, 2262, 1587, 367]. optional [1611]. Oracle [258]. Oracle-based [258]. Orbit [1026]. orchestrated [1284]. Orchestrating [2017]. order [1450, 2075, 2379, 2051, 738, 634, 1322, 1143, 530, 734, 2048, 2192, 2204, 1506, 1375, 1130, 1736, 646, 625, 1866, 1733, 1958, 722, 1854, 1900, 1032, 502, 960, 890, 332, 982, 1957]. ordered [913, 914]. ordering [1774]. orderings [1595]. Oregon [319, 2439]. organization [2136, 242]. Organizing [2371, 2224, 1693]. Oriented [1035, 281, 24, 46, 47, 473, 29, 546, 2477, 1227, 1502, 480, 582, 1531, 158, 1484, 1316, 2023, 1087, 575, 2453, 1990, 1452, 1895, 1460, 1787, 2408, 950, 438, 1917, 1177, 220, 1241, 596, 1454, 624, 1346, 1975, 1455, 617, 1909, 219, 1214, 1363, 1755, 932, 944, 171, 1453, 936, 1171, 1211, 1181, 2126, 172, 458, 178, 968, 2466, 2160, 1670, 1523, 2396, 1264, 1265, 1200, 1309, 1253, 790, 1125, 803]. origin [1903]. orthogonal [607]. Other [290, 2432].

Ott [1847]. out-of-order [2048]. outsourcing [1205]. Overcoming [27, 103]. overflows [806]. overhead [742, 1281, 1679, 952]. overlapped [2340]. overload [1400]. overloading [504, 645]. Overshadow [1968]. overview [2439, 1014, 626, 1122, 2459]. own [787]. Ownership [695, 736, 820, 700, 1906, 2266, 1905, 1375, 1178, 1660]. OXenstored [2391].

P [1924]. package [2449, 1012]. packages [161]. packet [1387, 2353, 1843]. packets [263]. packing [841]. Packrat [633]. PADL [33]. PADS [1392, 1725]. PADS/ML [1725]. Page [193, 1970, 1703, 2239, 2086, 1839, 754, 1283]. page-based [1703]. page-centric [754]. page-level [2239]. Pages [213]. paging [1379, 1842]. pain [2294]. Palm [618]. Panel [111, 118, 116, 113, 114, 119, 2270]. paper [1534, 2060, 2005, 1608, 2148, 2164]. Papers [211, 54, 36, 319, 960]. PARADIGM [1230, 828, 1923]. Paradise [2082]. paradoxical [1670]. Parallel [2483, 2249, 297, 338, 543, 1021, 329, 2409, 2397, 2258, 383, 144, 767, 2411, 2295, 2248, 2000, 1999, 137, 515, 620, 859, 822, 915, 1610, 2084, 905, 918, 1478, 2285, 2297, 344, 1580, 2418, 910, 658, 1780, 2377, 564, 2311, 337, 2085, 2371, 919, 920, 1928, 1976, 1064, 2287, 2253]. Parallelism [388, 543, 195, 515, 1836, 2073, 913, 914, 1685, 1874, 2284, 857, 1786, 1989, 2250, 75, 25, 1701, 2314, 2283, 2257, 2109]. Parallelization [2284, 2240, 2273, 1023, 902, 1787, 1073, 1788, 897, 2314]. parallelizer [2016]. Parallelizing [2313, 1996]. parameter [1635]. Parameterized [1804, 410, 442, 2327]. parameterizing [985]. parameters [1611, 641]. Parametric [1301, 2075, 411, 1317, 1085, 1462, 166, 356, 1077, 1729, 1658, 1833, 2385, 890]. parametricity [2385, 1140]. Parametrization [101]. parametrized [2214]. ParaModel [1999]. parser [1532, 219, 1541, 1360]. Parsers [24, 889, 2213]. Parsing [18, 56, 95, 1121, 979, 8, 400, 619, 1554, 633, 1549, 748, 2430]. part [2143, 1486, 1539, 2015, 702]. Partial [2484, 934, 319, 450, 2387, 923, 525, 975, 1124, 2118, 627, 1322, 924, 922, 1131, 524, 2211, 1445, 1679, 1437, 1159, 1115]. partial-order [1322]. Participation [539, 540]. partition [2189]. partitioned [1844, 351]. partitioning [818, 1836, 1387, 2070, 1989, 341, 849, 851]. party [1468]. Pascal [589, 327]. Pascal-2000 [327]. pass [2055, 563]. passing [488, 766, 1748, 730]. passive [768]. Past [390, 125, 217, 280, 325, 477, 545, 584, 622, 2427]. PASTE [469]. PASTE’01 [2481]. Path [1794, 203, 1370, 529, 302, 1055, 2474, 1754, 1141, 1076, 1683, 558, 2031, 1887, 1085, 2077, 2415, 1751]. path-based [1076]. path-sensitive [2474, 558, 2031]. path PRE [627]. paths [68, 1798, 1751]. Pattern [714, 2087, 405, 923, 1634, 692, 251, 1880, 2434, 2200, 1491, 778, 944, 2211, 521, 2125, 1223, 2431, 1850]. Pattern-matching [714, 521]. Patterns [414, 320, 18, 56, 95, 400, 1354, 1366, 739, 1136, 2045, 619, 1399, 1459, 874, 2211, 521, 2103, 919, 1503, 1191, 571, 1461]. pause [760]. pay [1127]. PDG [2361]. PDG-based [2361]. Pearl [2183, 1441, 633, 149, 1863, 2079, 2372, 155, 1442, 1951, 2388, 1928, 134]. pearls [1633]. pedagogy [2153, 2155]. peephole [1707, 1020]. Peer [345]. Pensieve [900]. people [2432]. PEPM [1534]. PEPM’02 [319, 2484]. Per-thread [2231]. perceived [1974]. Perfecting [617]. Performance [296, 697, 1845, 195, 1169, 128, 370, 1291, 315, 208, 1471, 683, 2009, 1217, 1386, 1687, 1884, 2121, 1472, 2048, 2413, 1975, 1065, 895, 2380, 855, 1378, 1570, 2153, 1688, 753, 757, 1974, 907, 1261, 2123, 1801, 2290, 653, 1993, 2296, 2298, 2337, 2274, 1843, 1885]. periodic [2055, 1596, 1814]. Perl [1715]. Permission [1334, 1375]. Permission-based [1375]. permissive [2362]. permutations [1410, 1560]. perplexed [140]. persistence [693, 1737, 2101]. Persistent [202, 957, 518, 1589]. perspective [1827, 2086, 617, 591, 1517]. perspectives [2270]. PetaBricks [2302]. Petascale [2275]. Phantom [2246]. Phantom-BTB [2246]. phase [831, 1081, 1595, 1901, 1282]. Phased [840]. phasing [1878]. Phobia [549]. Phoenix [920]. phones [1092]. PHP [2186]. Pi [257, 2065]. Pi-calculus [257]. PicoServer [1682]. PICSEL [1974]. pict [256]. picture [2240, 2435]. Pilot [618]. pin [1344, 1652, 1383]. pinning [1980]. PIP [1248]. pipeline [1685]. pipelined [1387, 1828, 1025, 674, 1380, 74]. pipelines [2097, 612, 1678, 1277]. Pipelining [363, 26, 1031, 25, 1409]. piping} [612]. PIROL [169]. PL [2141, 2152, 2155]. PL.8 [1014]. PL/Compiler [2141]. place [2278, 775]. Placement [2056, 1376, 2271, 493]. Placement-and-routing-based [2056]. places [2222, 1199, 1304]. planning [1248]. PLAS [2163, 2164]. platform [943, 1944, 625, 2336, 921, 1280]. platform-dependent [1944]. platforms [2017, 2263, 1830, 2425]. playing [2069]. PLDI [16, 17]. Pluggable [1591, 1468, 1644, 1008]. Point [1, 2, 1041, 1039, 1040, 1732, 2317]. pointcuts [1717]. Pointer [287, 2467, 196, 306, 331, 289, 68, 1709, 66, 1793, 2188, 2029, 1048, 1876, 1592, 288, 1078, 1079]. pointers [1038, 1522, 78, 1764, 1521]. Points [802, 290, 441, 2419, 728, 903, 2079, 1792, 2470, 2469, 1901, 67, 2223, 1457, 1583, 2395]. Points-to [802, 290, 441, 2419, 903, 1792, 2470, 2469, 67, 1457, 1583]. policies [1393, 2367, 2341, 1326, 2046, 2166, 1395]. Polish [889]. PolyAML [1452]. PolyD [1482]. polyhedral [2016, 2015]. polymer [1393]. Polymorphic [1315, 977, 1503, 1452, 1516, 250, 2403]. Polymorphism [2086, 1619, 1301, 1317, 611, 2089, 2174, 1439, 646, 1462, 166, 2090, 890]. polynomial [558, 976, 1549]. Pool [374, 745, 1378]. POPL [2482, 495]. POPL’01 [54]. POPL’02 [319]. popular [2149]. portability [342]. Portable [448, 686, 1817, 2278, 580, 864, 2417]. Portland [319]. Position [2148]. positioning [1043]. Positive [2192]. Possibilities [426]. Post [2055, 1491, 563, 2151]. Post-compilation [1491]. post-Linnaean [2151]. Post-pass [2055, 563]. postmodern [1194, 1303]. potency [1142]. potential [1710, 2469, 362]. Potpourri [1295]. ’power [1972, 1296, 348, 347, 193, 1099, 2292, 1973, 2238, 1511, 1104, 1270, 1290, 871, 1688, 2237, 1192, 2065, 1604, 1972, 1844, 1408, 1993, 1191]. Power-Aware [348]. Power-efficient [1099, 2238, 1993]. PowerEpsilon [391]. powerful [2330, 633]. PowerNap [2237]. PPoPP’01 [2483]. PQ [445]. PQ-encoding [445]. PQL [1475]. Practical [400, 1811, 18, 56, 95, 619, 1556, 33, 1624, 173, 2016, 1124, 2112, 554, 1557, 2458, 2148, 2206, 807, 778, 1792, 2226, 1178, 2304, 142, 1543, 2277]. practice [1084, 753, 2132, 747, 2343]. Practices [2483, 1267]. Practicing [64]. Pre [236, 669, 1390]. pre-computation [1390]. pre-execution [669]. Pre-Tenuring [236]. preallocation [741]. Preceding [319]. Precise [1339, 1755, 997, 1086, 575, 2309, 2179, 89, 1941, 1770, 1945]. precisely [2333]. precision [1411, 1401, 1855]. precomputation [563]. precondition [2363, 1766]. preconditions [2330]. Predicate [304, 501, 1324, 1778, 2318]. predicated [557]. Predicates [422]. predication [1975, 1284]. predictability [572]. predictable [580, 1328]. Predicting [813, 1045, 830]. Prediction [203, 1979, 2225, 816, 734, 1376, 1688, 662, 1679, 1282, 2244]. prediction-based [1679]. predictive [1689]. Predictor [1982]. preemption [617]. preemptive [2022, 1107, 1411]. Preference [556]. Preference-directed [556]. Preferential [1751]. Prefetch [1089]. prefetching [773, 1696, 1285, 570, 678, 815, 1284, 2272, 571, 1838, 1099]. prefix [1928]. Preliminary [2285]. prepaging [769]. presence [978, 1458, 79]. presentations [1150]. preservation [1640]. preserves [2076]. preserving [2023, 1772, 2068, 682, 1579, 260]. pressure [1823]. Pretenuring [462]. Preventing [1400, 695, 1812]. prevention [1181]. Price [414]. primacy [1185]. primitive [2285, 702]. primitives [1784, 2271, 1748, 1461]. Principal [995, 1439]. principle [2248, 252, 2224]. Principled [293, 1613, 1435]. Principles [2482, 2483, 551, 54, 1245, 2432, 83]. Print [2, 1041]. priorities [759]. prioritized [1413]. Priority [412, 1019, 1098]. priority-based [1019]. PRISM [1163]. privacy [1997]. proactive [1578, 862]. proactively [1885]. Probabilistic [1886, 1403, 109, 1563, 903, 1709, 1744, 253, 779, 1327]. probability [1767, 498]. Problem [44, 1332, 2467, 1242, 1351, 1630, 45, 522]. Problems [27, 2144, 2087, 2305, 142]. procedural [783]. procedure [2315, 1529, 1337, 1809, 1945]. procedures [723, 990]. Proceedings [2476, 2478, 2479, 2483, 2484, 536, 2480]. Process [105, 1003, 716, 1323, 869, 1145, 722]. processes [1325, 1196, 1306, 1485, 982, 357]. Processing [15, 2133, 2248, 585, 2052, 1387, 1021, 1392, 164, 2353, 1843, 856]. Processor [352, 315, 206, 365, 1820, 1602, 1270, 796, 1699, 612, 1280, 1091, 347]. processor-based [1280]. Processors [232, 208, 2273, 1831, 818, 1113, 2231, 1995, 2048, 2332, 1410, 904, 1698, 2348, 2057, 1823, 2262, 2287, 852, 1112, 1099, 1594, 346]. Procrastination [1602, 1098]. Producing [2242]. product [1730, 1248, 1851, 1162, 710]. Production [3, 1864, 1253]. Professional [317]. Profile [108, 1578, 561, 234, 2421, 105, 598, 1803, 1568, 1043, 562, 1227, 1471, 497, 1127, 2314, 450]. profile-analysis [497]. Profile-Based [108, 1568]. Profile-Directed [234, 562]. Profile-driven [1803, 2314]. Profile-Guided [105, 1578, 561, 2421]. Profiled [299]. profiler [1015]. profiles [1055, 2345, 1045]. Profiling [203, 231, 109, 295, 849, 1281, 1470, 951, 2415, 2085, 1751, 2335, 1572]. Program [2481, 2484, 1166, 1744, 2032, 655, 533, 2468, 232, 1965, 2318, 319, 106, 302, 2460, 701, 2240, 1507, 1973, 2437, 971, 1934, 2178, 925, 1628, 500, 114, 148, 558, 743, 813, 254, 1362, 918, 1476, 2179, 1145, 1813, 1072, 2215, 2308, 261, 2327, 1941, 252, 805, 1964, 65, 1383, 1650, 115, 1475, 1938, 1528, 1686, 2469, 1331, 878, 2167, 1720, 131, 560, 1597, 1694, 660, 520, 2422, 1052, 1275, 2337, 1701, 1555, 340, 1045, 846, 1711, 2028, 1110, 851, 1075, 1258, 1547, 1088, 2030, 469]. Programmable [314, 186, 2012, 2279, 906, 771, 1984, 1094]. programmer [2086]. Programming [2476, 2478, 2482, 2483, 54, 41, 58, 2142, 2143, 1035, 516, 542, 961, 2099, 126, 905, 60, 1268, 46, 366, 676, 49, 1639, 2485, 278, 245, 1388, 473, 2167, 2163, 402, 50, 16, 2336, 2159, 1391, 1848, 29, 472, 2162, 546, 391, 2477, 2479, 506, 2439, 59, 399, 1817, 2041, 1345, 2140, 2136, 2137, 2138, 2402, 1559, 2410, 582, 1868, 1636, 695, 1189, 2295, 1550, 1999, 137, 585, 892, 1431, 1386, 338, 2144, 1990, 1452, 2074, 2359, 1537, 1977, 705, 1423, 1714, 1197, 1307, 1483, 1853, 1871, 715, 554, 2145, 127, 1329, 2373, 2146, 1864, 1160, 1859, 939, 2391, 1135]. programming [1659, 1995, 2285, 643, 1446, 2148, 2150, 596, 1573, 2064, 1952, 2086, 1455, 1785, 2151, 1893, 791, 932, 778, 1267, 2154, 900, 1498, 1109, 1994, 1465, 1858, 966, 2409, 2093, 1187, 1915, 139, 277, 1918, 524, 2406, 936, 1742, 629, 1620, 648, 1929, 1522, 1430, 1194, 1303, 1763, 2433, 385, 1213, 1760, 1958, 2383, 2454, 1178, 861, 754, 178, 2156, 2157, 2158, 2217, 824, 2264, 2374, 151, 1427, 713, 2160, 891, 2161, 1373, 2134, 1196, 1306, 2422, 1574, 1892, 1670, 2001, 1922, 1517, 920, 2007, 1485, 2396, 960, 83, 1781, 1849, 2445, 1449, 1144, 73, 1967, 2395, 2294, 1005]. programming [1200, 1309, 1608, 17]. Programs [304, 282, 281, 388, 290, 425, 224, 478, 487, 927, 1502, 836, 709, 2278, 723, 442, 908, 1288, 488, 626, 570, 2379, 575, 1132, 2443, 2035, 154, 1743, 1358, 1359, 2118, 2051, 1073, 942, 1369, 893, 2022, 2080, 2473, 991, 2033, 1403, 2177, 1177, 1685, 2037, 734, 2282, 1750, 1542, 1761, 1880, 2446, 840, 2284, 1266, 2204, 1506, 2017, 1755, 795, 2092, 335, 529, 1716, 1736, 1592, 253, 1546, 1780, 1808, 2083, 2300, 751, 1331, 2054, 1181, 2311, 917, 990, 562, 2235, 657, 1445, 2466, 331, 2008, 1854, 1407, 152, 1900, 1624, 1391, 1708, 2085, 2317]. programs [2429, 2350, 1086, 2447, 2277, 1231, 571, 2109, 1367, 248, 2328, 1626, 2428, 803]. Progress [2311, 1210, 2440, 2203, 2331]. progressive [1567]. project [484, 824, 1234, 900]. projects [965, 1267]. prolific [510]. Prolog [753, 225]. promote [1238]. promotion [1732]. prompts [1133]. Proof [1323, 1628, 1500, 1498, 2383, 1662, 2361, 2328]. Proof-guided [1323]. proofing [1754]. proofs [2176, 971, 1934, 2379, 1585, 989, 1342, 1624]. propagation [1024, 1454, 1992, 2322, 174, 832, 1006]. propagation-based [174]. properties [2319, 1663, 2037, 1504, 1904, 1255, 248, 1069, 1146]. property [2195]. Proposal [549, 271, 756]. Proposed [59, 60, 1907]. proprietary [1986]. prospects [2447]. protected [1003]. Protecting [1525, 1690, 1274]. protection [2269, 1968, 1117, 1678, 680]. Protocol [189]. Protocols [291, 1196, 1306, 487, 2114]. prototype [909]. prototypes [1641]. Prototyping [393, 473, 354]. provably [810, 1809]. provide [1508]. Providing [2430]. Proving [1743, 1796, 2170, 1939, 2327, 509, 1431, 811]. Proxies [550, 1649]. Proxy [615]. Pruning [1564, 2243]. Pseudo [1212, 546]. Pseudo-classes [1212]. pspace [1936]. psychology [2432]. PTIDES [2343]. Public [549, 1225]. publication [2058]. PULSE [2386]. Pure [739, 2094]. Purely [2373, 2365, 2370, 1937, 2096]. purely-dynamic [2365, 2370]. purpose [873, 570, 2084, 1995, 1701]. Push [2348, 1130]. Push-assisted [2348]. push/enter [1130]. pushdown [1747]. put [385]. putting [775]. puzzle [531, 2026].

QoS [1208, 376]. QoS-aware [1208].

QoS-Enabled [376]. QR [914]. QuA [1208]. quadtree [913, 914]. quake [2252]. Qualified [1438]. qualifier [1898]. qualifiers [1374, 553, 1872]. qualities [1263]. quality [220, 1216, 1488, 2058]. quantification [1648]. quantified [1946]. Quantifying [303, 1472]. Quantitative [2024, 1320]. quantum [1325, 868, 1694]. Quartus [618]. quasi [160]. quasi-static [160]. queen [522]. queries [1141, 1476, 1085, 2077, 449, 1737]. Query [2100, 2393, 164, 1475, 2101, 129]. queue [334]. queue-based [334]. Queues [412]. Quick [612]. QuickCheck [154, 712, 2386]. Quicksilver [160]. Quotient [2098]. QVM [2108].

R6RS [1878]. Race [1067, 2008, 442, 2386, 1789, 81, 2473, 2309, 2310, 1576, 1749, 1181, 916, 917, 1577, 2007, 1676, 443]. race-free [442]. Raced [2345]. races [181, 695, 1769, 1812, 2268]. radio [1194]. RAIL [1236]. raising [871]. RAM [1800, 2346]. ran [516]. random [154, 1385, 724, 998, 1339, 2008]. randomized [2308]. range [2347, 555, 1016, 1568]. rank [1637]. Ranking [48]. Rapid [393, 594, 2430]. RapidMRC [2230]. rate [1602, 593, 2230]. rate-monotonic [1602]. rates [1415]. ratio [1283]. rational [946]. Raw [1764]. ray [2072, 2247]. re [1166, 1278, 913, 914]. re-engineering [1166]. re-execution [1278]. re-use [913, 914]. reachability [1084, 250]. reactis [1860]. Reactive [543, 1430, 2374, 83]. Read [1, 1039, 1040, 2213]. Real [1028, 297, 348, 186, 374, 415, 858, 1817, 838, 742, 1414, 820, 158, 1510, 812, 835, 761, 616, 1251, 1398, 1160, 1240, 580, 1101, 1098, 2308, 2446, 617, 1792, 162, 1998, 1415, 625, 1824, 2010, 839, 2389, 2348, 1574, 1107, 1045, 1832, 1353, 1348, 1814, 2343, 843, 347]. Real-Time [297, 348, 186, 374, 1028, 415, 858, 1817, 838, 742, 1414, 820, 158, 1510, 835, 761, 616, 1251, 1398, 1240, 580, 1101, 1098, 2446, 617, 162, 1415, 1824, 2010, 839, 2348, 1574, 1107, 1832, 1353, 1348, 1814, 2343, 843, 347]. realistic [806]. realities [1202]. Realization [1363]. realizing [1924]. reasoning [2442, 988, 2197, 1512, 2196, 1856, 1629, 1506]. rebiasing [1657]. Rebinding [877]. recipe [1351]. reclamation [1581]. recognition [979, 1706]. recognition-based [979]. recoloring [2027]. recommendations [2137, 1969]. reconciling [1715]. Reconfigurable [472, 1596, 2056, 850, 1118, 848, 2342, 2049, 1605]. Reconsidering [681]. reconstruction [1384, 597, 2454]. Recording [1692, 1676]. recoverable [2224]. Recovery [2224, 2430]. Recurrence [2342]. Recurrences [425]. Recursion [1366, 994, 1549, 1790, 2204, 1318, 2287, 77]. Recursive [24, 1429, 145, 149, 1620, 407, 478, 2214, 1940, 1877, 976, 1143, 1542, 2087, 533, 2317, 784, 737, 2395]. recursively [2077]. Recycle [233]. redesign [1285]. redistribution [336]. reduce [1973, 463, 300, 2344]. reducible [727]. Reducing [301, 228, 1566, 760, 1593, 2048, 2058]. reduction [1322, 651, 797, 2122, 1803, 1604, 1356, 1799, 1676, 1815]. reductions [1497]. redundancy [1124, 627, 1115]. redundant [1057, 758, 1680, 1679, 1182, 1259]. reeled [516]. Refactoring [1469, 1153, 933, 2289, 2433, 2001, 2117, 2102]. refactorings [1232, 167, 1156]. Reference [303, 240, 330, 953, 987, 589, 835, 1846, 2391, 464, 1466, 1088]. References [235, 1361, 1458]. Referentially [298]. Refinement [1583, 2331, 755, 1369]. Refinement-based [1583]. Refinement-violation [1369]. refinements [887]. Refining [1183, 1262]. Reflection [8, 1629, 2014, 1152, 934]. Reflections [213, 37, 212]. Reflective [1243, 686, 1239, 625, 166, 827]. reflexive [136]. Region [1087, 818, 577, 1338, 225, 2009, 820, 259, 786, 565, 770, 819, 2011]. Region-Based [225, 818, 577, 1338, 820, 786, 770, 819, 2011]. Regions [292, 761, 1139, 2207, 780, 78, 419]. Register [1017, 1019, 351, 350, 2026, 1380, 608, 101, 1033, 548, 1821, 783, 2055, 609, 1034, 1037, 1831, 607, 1113, 2056, 1837, 1762, 1567, 2344, 1604, 1090, 725, 1027, 1823, 2262, 244, 1091, 1112, 1381]. register-allocation [609]. register-reuse [244]. Registers [308, 101]. Regression [460, 1654, 1688]. regrouping [1088]. Regrowing [2433]. Regular [1136, 130, 251, 396, 2045, 2052, 2315, 874, 1085, 1503, 571]. regular/irregular [2045]. Regularly [1797]. regulated [1676]. reification [934]. reinforcement [2134]. Related [305, 1657]. Relating [1855, 1134, 1664]. Relational [1947, 1476, 221, 971, 1896, 1080]. relations [1500]. Relative [2046]. Relaxed [2202, 945, 1768, 1510, 1326]. release [2166].

Relevance [1938]. Reliability [1297, 1978, 671, 2234]. Reliable [190, 1785, 1346]. rely [2196]. rely-guarantee [2196]. remapping [915, 1839, 595]. remarks [394]. Reminiscences [960]. remodularization [2115, 685]. Remote [106, 805]. Removal [547, 1057, 80, 2107, 694]. removing [1182, 1259]. renaming [2116]. rendezvous [1001]. renewal [652]. renewal-older-first [652]. reorderings [801]. repair [937, 1902, 1360]. replacement [2046]. Replay [1655, 2121, 2226, 1769, 2276]. replication [335]. Report [958, 34, 2099, 123, 97, 39, 323, 541, 33, 16, 1009, 1311, 2138, 2402, 1868, 2072, 1873, 2399, 2097, 581, 706, 618, 1486, 1710, 1758, 1870, 214, 2069, 2380, 2094, 2377, 514, 2392, 2389, 1860, 1849, 1862]. reporting [1903, 1997, 1776]. repository [1471]. Representation [298, 499, 302, 2198, 892, 1895, 744, 1629, 1488, 1525, 1528, 2324, 1064]. Representations [303, 404, 1440, 1159]. Representing [479, 402, 918]. reproducible [2276]. request [1271, 1985]. requirement [1353]. Requirements [1348, 2357, 1247, 1664]. requires [1786]. rescue [2223]. Research [656, 2152, 824, 1163]. reservation [690]. reset [525, 654]. resiliency [2290]. resilient [1992]. Resistant [199, 851]. resizing [2464, 2048]. resolution [173]. resolving [1890]. Resource [513, 2444, 2043, 448, 2048, 780, 1388, 1117, 666, 1108]. resource-constrained [2043, 1388, 1117, 1108]. resourceful [1961]. resources [2197, 337, 920]. Response [273, 2058]. responsibility [1154]. responsibility-based [1154]. responsiveness [578]. RESTful [2401]. restricted [2041]. restructure [1612]. Restructuring [240]. Resultant [1012]. results [728, 2285]. Retargetable [586, 1113, 1111]. retargeting [1569]. Rethinking [2155]. retrieval [1175]. retrofitting [1968, 496]. retrospect [1032]. Retrospective [1040, 1041, 1035, 1061, 2167]. return [1137]. returns [2400]. reusable [1208, 2125, 1264]. Reuse [205, 2464, 1250, 2449, 969, 699, 813, 1118, 867, 1544, 1083, 328, 244]. reusing [1712]. revealed [149]. Reverse [1266, 1569, 104, 631]. reverse-engineering [1569]. reverse-engineers [104]. reversible [868]. review [1609]. reviewing [2058]. revisited [947, 1314, 2438, 1300, 1869, 1233, 1079]. Revisiting [1362, 1941]. Rewriting [299, 714, 1760]. Rice [1935]. right [1951]. rights [704]. rigorous [1884, 2121]. RISC [346, 20, 21]. RMI [1477, 1065, 397, 398]. robin [2256]. Robotics [10]. Robust [2047, 2412, 902, 2174, 839]. Robustness [842, 1480]. Role [486, 1206]. roles [1246, 1243]. rollback [1433]. rose [1819]. round [2256]. round-robin [2256]. router [676, 674]. Routine [967]. routing [2056]. rows [268]. Royal [2482]. RPC [677]. RTR [1676]. RTSJ [1251]. Rubicon [1153]. RUBiS [1233]. Ruby [2456]. rule [715, 2424]. rule-based [715]. rules [1714, 714, 1342, 1521, 1849]. run [1471, 743, 1290, 1061, 967, 1053, 801, 347]. run-time [743, 1061, 967, 1053, 801, 347]. running [408]. Runtime [285, 2378, 1473, 2423, 1552, 2108, 2449, 1256, 1912, 837, 1369, 1789, 1557, 2114, 746, 1165, 1221, 2281, 625, 1161, 2213, 2036, 2335, 973]. runtime-adaptable [625]. runtimes [2244].

§-calculus [748]. Safe

[2266, 298, 404, 409, 1826, 2300, 2374, 1479, 2410, 695, 820, 180, 265, 2341, 1071, 777, 2297, 2014, 1048, 876, 2245, 490, 1929, 496, 2321, 2125, 1866, 1598, 1574, 1328, 2011, 2036, 134]. SAFECode [1562]. Safely [170, 1408]. Safer [2119]. SafeTSA [298]. Safety [69, 1563, 809, 259, 1977, 837, 786, 1372, 2468, 2320, 2392, 1662, 248, 1069, 1146]. safety-critical [809, 2392]. salesman [1918]. SALSA [472]. sample [1666, 1412]. sample-based [1412]. Sampling [235, 198, 1911, 1845, 805, 2310, 1327]. satisfiability [1341]. saturation [2191]. save [2346]. Saving [955, 1511]. SAW [1167]. SCADE [845]. SCADE/lustre [845]. Scala [1508]. scalability [697, 343]. Scalable [194, 375, 1278, 84, 372, 347, 1368, 505, 1070, 1456, 1910, 334, 1411, 174, 1341, 647, 2031, 1887, 2029, 1784, 2256, 1642, 1606]. scalar [670]. Scale [15, 202, 1559, 2023, 338, 660]. scaling [1816, 1602, 1566, 1101, 1974, 67, 592, 798, 1416, 1815]. scannerless [2430]. Scavenging [293]. scenario [1965]. Schedulability [841, 2047, 1596]. Schedulability-driven [841]. Schedule [307, 1082]. scheduled [601]. scheduler [1569, 2109]. Scheduling [364, 370, 191, 2273, 2341, 1602, 1683, 616, 2084, 1101, 1098, 840, 1051, 617, 1031, 2256, 1684, 1025, 2342, 1988, 1604, 896, 839, 1411, 2340, 1927, 73, 1923, 1416, 1815]. Schemas [48, 2393]. Scheme [1608, 1842, 653, 152, 2272, 1960, 1862, 1412, 1614, 12, 1026]. schemes [2204]. school [2439]. science [2139, 2143, 1203]. scientific [2464, 2057]. SCOOP [901]. Scope [406, 1464, 1426, 2448]. scoped [863, 761, 1756, 1643]. SCOPES’02 [384]. SCR [1600]. Scrap [1441, 778, 1152, 1443, 1731]. Scratch [413]. Scratchpad [1841, 1840]. Scribble [2382]. Scripting [869, 2412, 2421, 520, 1489]. scripts [2447]. Sealed [161]. sealing [983]. seamless [1647, 1827, 1588]. Search [412, 1674, 1842, 1595, 688, 594, 1809, 1231]. search-while-load [1842]. searches [1081]. Searching [1132, 1806]. Second [309, 1253, 1143]. second-order [1143]. secrecy [487]. section [2241, 465]. sections [2034, 335, 1524]. Secure [298, 265, 1741, 1275, 255, 2362, 964, 491]. Security [1299, 44, 2163, 487, 1393, 2358, 1954, 1501, 1978, 1739, 1475, 245, 1996, 2167, 2208, 132, 45, 1740]. see [2240]. seeing [170]. seL4 [2380]. Selection [234, 2485, 2044, 1554, 460, 1565, 598, 2345, 2334, 934]. Selective [719, 667, 1278, 2239, 2403]. Self [1035, 2420, 1558, 1953, 1773, 2301, 2092, 2185, 1693, 2324, 2223]. self-adjusting [1558, 1953, 2301, 2092, 2185]. self-healing [2223]. self-modifying [1773]. self-organizing [1693]. self-representation [2324]. Semantic [1374, 974, 1807, 1664, 792, 1779, 1908, 1005]. semanticist [1847]. Semantics [2484, 1931, 1717, 46, 2338, 326, 319, 391, 2292, 144, 1973, 986, 2366, 928, 2367, 259, 2075, 1610, 1365, 531, 2203, 2185, 879, 1716, 2006, 1930, 1753, 1337, 2201, 2130, 1662, 2428]. Semantics-aware [2338]. Semantics-Based [2484, 319, 1753]. Semi [2188, 2245, 1809]. semi-automated [1809]. semi-automatic [2245]. Semi-sparse [2188]. semiformal [1937]. sensing [2221]. sensitive [2460, 2474, 2367, 558, 2031, 84, 553, 2188, 807, 1501, 1565, 2029, 1792, 1592, 1577, 1583, 1078]. Sensor [656, 1270, 1785, 663, 1802, 910, 2093]. Sensors [192]. sentences [516]. Separation [992, 1333, 1932, 1334, 1934, 1933, 169, 1619, 1727, 1069]. September [2479]. seq} [978]. sequence [687]. sequences [1116, 831, 1081]. sequential [1073, 1478, 335, 2313, 1331, 1068, 1445, 1734]. serializability [1367]. serializable [2401]. Serialization [2258, 865, 686]. Server [213, 375, 370, 186, 1357, 1681, 675, 904, 2237, 1367, 2252]. Server-side [375]. servers [342, 564, 131, 1182, 1259]. Service [372, 1227, 614]. Services [550, 163, 986, 1948, 911, 1065, 1784, 1238, 1182, 1259, 1253]. services-oriented [1253]. Servlets [213, 37]. Session [429, 430, 431, 432, 433, 434, 435, 111, 116, 113, 119, 114, 1949, 2209]. set [1820, 1105, 1084, 1797, 1008, 1513, 1980]. set-based [1513]. sets [2258, 895, 75, 2469, 1011]. settings [798]. Sexy [1062]. Shangri [1386]. Shangri-La [1386]. Shape [1790, 1665, 2193, 1947, 1190, 1791, 1338]. SharC [2020]. Shared [2214, 1559, 2442, 1288, 2341, 912, 2286, 2227, 335, 907, 336, 1692, 1980, 339, 2036, 1367, 1607]. shared-memory [1367, 1607]. Sharing [2323, 2020, 2307, 180]. shelf [2269, 1757]. SHIFT [1230, 525, 654]. SHIM [1599]. shirt [2094]. shootout [1204]. short [1979, 1451]. Shortcut [641]. should [2139, 515, 1355]. shovels [1962]. side [375, 1029, 854, 2297, 574, 1434]. side-effect [1029]. side-effects [854, 2297, 1434]. sided [630]. Sifting [699]. SIGACT [54, 2482]. sighing [51]. sight [1301]. sign [569]. signal [1394]. signatures [1428, 1985, 1981]. SIGPLAN [30, 52, 2480, 1486, 1539, 319, 2478, 2479, 2481, 2482, 2483, 2484, 2136, 2137, 2138, 271, 120, 273, 537, 317, 2485, 31]. SIGPLAN-SIGACT [2482]. SIGPLANet [272]. SIGSOFT [2481, 824]. silent [659]. silicon [72]. SIMD [1074, 1410, 603, 1561, 1693, 1560, 1607]. Simple [971, 412, 1618, 221, 1437, 1316, 859, 822, 633, 2037, 2089, 1759, 1068, 891, 1212, 1485, 732]. simpler [1365, 1254]. simplest [749]. simplicity [1483, 1916]. simplify [897]. Simplifying [1497]. Simulated [188, 187]. simulating [1093]. Simulation [355, 188, 1193, 187, 1103, 162, 753, 594, 354, 1106, 2253]. simulator [1104, 2050, 1094]. Simulators [315]. Simulink [845]. Simultaneous [207, 206, 904]. Single [298, 2006, 2259, 1363, 1286, 87]. single-core [2259]. single-node [1286]. Singleton [2441]. Sites [13]. size [836, 507, 252, 2175, 606]. size-change [252]. sized [930]. sizes [2189]. skeleton [1999]. Sketching [1782, 2019, 1391, 1708]. skills [2158]. Slam [484]. slice [1680]. slice-based [1680]. slicer [1547]. Slices [232, 1390, 1564]. Slicing [927, 282, 281, 922, 5, 1123, 2460, 2458, 1030, 1370, 1964, 1965, 1597, 1777, 1711, 1075]. SlicK [1680]. Slideshow [1150]. sliding [948]. slightly [2371]. Slipstream [208]. Slot [1320]. Small [2141, 393, 1506, 1710, 1758, 1930, 141, 2210, 833, 61]. small-step [1930]. SmallCheck [2210]. smaller [2429]. Smalltalk [935, 283, 284, 1530]. Smart [550, 1021]. smartTools [1172]. Smell [1164]. smgn [393, 394]. smiles [51]. SMPs [185]. SMT [2231, 1290, 1941]. snippets [1159]. snoop [1973]. Snooping [185]. Snowbird [2481, 2483]. Snugglebug [2330]. Society [2482]. SODA [1241]. Soft [1044, 2344, 1287]. soft-error [1287]. SoftBound [2320]. SoftSig [1981]. Software [773, 468, 467, 540, 41, 1285, 2286, 291, 1787, 203, 1399, 363, 6, 1203, 286, 1031, 378, 199, 2435, 2349, 361, 26, 1698, 875, 3, 2481, 1672, 1552, 1155, 1209, 2240, 847, 2047, 484, 1665, 969, 809, 1176, 902, 1076, 1591, 2453, 984, 1663, 2225, 715, 2356, 501, 1322, 1917, 1206, 1667, 2354, 1776, 460, 1656, 169, 2471, 600, 665, 1096, 1129, 1267, 1488, 1992, 563, 591, 1586, 1245, 904, 2313, 25, 2226, 336, 2432, 1742, 1556, 1929, 258, 1171, 1944, 1642, 1462, 2228, 671, 1606, 1172, 1100, 1097, 1590, 2128, 1380, 1179, 1260, 824, 1463, 2159, 2223]. software [1668, 1409, 2265, 1162, 1328, 2299, 849, 1401, 1981, 1485, 1222, 1843, 1064, 1264, 28, 74, 1258, 843, 360, 357]. Software-based [1698, 563]. software-exposed [1981]. software-hardware [2226]. software-through-models [843]. solution [1103]. solved [2467]. solvers [1941]. Solving [2263, 2032, 2026, 142]. Soma [1535]. Some [2139, 1766, 823, 963, 394, 1533, 2157, 283, 284]. something [1743]. sophisticated [2419, 2146]. sorting [2070]. sorts [1420]. Sound [2031, 2116, 1770, 1147, 810, 2033, 1954, 884, 2173, 1889]. soundness [438, 1955, 1342, 2128]. Source [7, 339, 1819, 1820, 1237, 1403, 1762, 1302, 2105, 870, 1159, 2422, 1006]. source-code-level [1762]. Source-Level [7, 339, 1403, 2422]. source-to-source [1819, 1820, 1237]. Sources [25, 26, 164]. Space [1049, 568, 426, 295, 2085, 838, 1820, 2009, 1059, 492, 987, 1509, 734, 1978, 879, 594, 595, 952, 1115, 764, 567, 691]. Space-Efficient [295]. Space-time [568]. spaces [1689]. spanning [2251]. sparse [557, 2188, 2251]. Spatial [1269, 781, 1683, 1977, 895, 1686, 2320, 934]. special [708, 465, 1265]. specialisation [1880]. Specialization [1748, 1405, 2339, 1573, 926, 1473, 528]. specialize [1697]. specialized [560]. specializer [114, 115]. Specific [2416, 393, 11, 224, 100, 1654, 1095, 705, 1392, 2147, 559, 1896, 2049, 1493, 1183, 1262, 1173, 613]. Specification [2357, 13, 14, 62, 163, 1499, 2419, 1158, 1215, 588, 590, 2305, 177, 1778, 1364, 1083, 599, 1711]. specifications [1321, 485, 808, 716, 717, 1266, 1593, 1861, 1600, 972, 1123, 392]. specified [1900]. specify [574]. Specifying [1520, 1529]. spectroscopy [2284]. speculation [1072, 658, 897, 1284]. Speculative [658, 1100, 2293, 667, 903, 902, 1709, 1073, 143, 330, 2061, 563, 817, 671, 1390, 1115, 670]. Speed [190, 578, 2179]. spelling [1302]. spill [1382, 1822, 74]. Spilling [308, 1017]. spin [334]. Spinach [1094]. SPL [313]. Split [65]. Split-stream [65]. splitting [1568, 1088]. spot [2042]. spots [624]. spreading [1697]. Spreadsheet [5]. Spreadsheets [5]. SPTF [1271]. SQL [1543]. SSA} [2044, 1822, 1528]. SSA}-graphs [2044]. stability [1241]. stability-oriented [1241]. Stabilizers [1626]. Stack [511, 2287, 782, 750, 1538, 1845, 1584, 1746, 1444, 2346]. Stack-based [2287, 1584]. stacking [1682]. Stage [409, 413, 2082]. Staged [523, 2303, 422, 1529, 1061, 1516, 2093, 649, 494]. Staging [9]. standard [1189, 1735, 404, 407]. Standardization [599]. standards [861]. STARC [1902]. State [2198, 62, 1432, 2442, 1940, 912, 1329, 1375, 514, 1598, 1921, 1734]. State-dependent [2198]. statecharts [846, 1119]. stateful [2166]. stateless [678, 2039]. States [56, 1541, 55, 1813]. Static [2423, 298, 108, 2465, 572, 1538, 1807, 1623, 734, 814, 1415, 402, 1778, 1408, 132, 1617, 2172, 547, 803, 1717, 484, 609, 971, 809, 723, 2115, 758, 2399, 2118, 2408, 1902, 573, 2421, 1798, 741, 2114, 2179, 1581, 559, 1131, 888, 2014, 1755, 2469, 1576, 1749, 1106, 2329, 1279, 160, 2448, 1086, 1737, 694]. static-dynamic [2115]. Statically [1643, 667, 806, 1125]. statically-typed [1125]. stationary [1942]. Statistical [343, 1281, 1368]. Statistically [1884, 1813]. stealing [2254]. Stealth [1696]. Steering [381]. stencil [1788]. stencils [1782]. step [1432, 2381, 889, 1930]. step-indexed [1432]. step-indexing [2381]. stewed [542]. STL [27]. STM [2187, 2006, 2267, 1774]. Stochastic [498]. stock [1028, 751]. stop [153, 2291]. Storage [1298, 202, 182, 307, 852, 1013, 683, 1273, 814, 1271, 183, 677]. Storeless [928]. stores [957, 659, 1138]. Storing [1589]. story [1225, 1426]. Strachey [1851]. strands [1408]. strange [620]. Strap [541]. Strap-On [541]. strata [1692]. Strategies [374, 2020]. strategy [2140, 1248, 893, 2265]. Stratified [259, 1514, 1425]. Stream [1879, 1835, 570, 679, 1685, 840, 2017, 795, 65, 2054, 2247, 1407, 1892, 2350, 2057, 2262]. StreamFlex [1892]. streaming [1828, 1391]. StreamRay [2247]. streams [2248, 1571, 1879, 1995, 2079, 1630]. Streamware [1995]. strength [2399, 507]. Stretching [2312]. strict [893, 638].

Strictly [2419]. Stride [815, 571]. string [1961]. strings [923, 2122]. strong [2269, 683, 651, 2110]. strong-consistency [683]. strongly [2187]. Structural [411, 2104, 1171, 1255, 1083]. structure [2423, 568, 1663, 491, 2325, 673, 459, 1378, 1795, 1088]. Structured [960, 2218, 1752, 1952, 1924]. Structures [247, 407, 2307, 1038, 1937, 937, 714, 745, 533, 1951, 1621, 2019, 2038]. struggles [1972]. Stuck [367]. students [542, 2146, 1191]. Study [108, 2288, 1895, 939, 587, 652, 169, 2332, 175, 1492, 753, 1998, 674, 444, 2010, 653, 1927, 613]. style [1893, 881]. styles [1460]. subclasses [170]. subclassing [1421]. Subcubic [1940]. subeffect [1872]. subgroup [2276]. subroutines [1611]. subscripted [1037]. subset [2315]. substitution [67]. substitutions [1958]. substructural [1432]. subsumption [270]. subsystem [1840]. Subtext [1483]. subtypes [2214]. Subtyping [269, 1869, 2214, 731, 135, 1640, 149, 250, 502, 445]. success [1225, 1670]. sufficient [2211]. sufficiently [1361]. suggestions [1533]. suite [1092]. sum [870]. Summaries [2137, 1945]. Summarizing [990]. Summary [118, 119, 2138]. summer [2439]. sums [975]. superclass [170]. supercompilation [2192]. superoperators [2043]. superoptimizer [579]. superoptimizers [1707]. superperfect [1841]. superset [1530]. superword [75]. Support [2476, 377, 200, 205, 292, 199, 536, 858, 1552, 1469, 1647, 201, 2248, 903, 1977, 1460, 1963, 939, 1659, 956, 1783, 2280, 1415, 2378, 904, 1508, 1694, 1847, 1373, 1976, 1984, 332]. supported [1969]. Supporting [1704, 1258, 2449, 1154]. supports [87]. survey [1005]. suspicious [1813]. sustainable [2047]. SW [1830]. swapping [408]. SWAR [2248]. sweep [948, 1285]. switched [909]. SWL [1842]. SXML [588].

Symbiotic [206]. Symbolic

[970, 78, 1079, 1499, 687, 594, 2294].

symbolic-evaluation [1499]. Symmetric [200, 201, 165, 1966]. Symmetric-Key [200, 201]. Symposium [2482, 2483, 54, 36, 470, 540, 2059]. symptom [726]. synchronization [1056, 740, 1071, 2285, 344, 1242, 617, 658, 1524, 2271, 1901, 80, 1657, 1523]. synchronization-related [1657]. Synchronous [2341, 2053, 1510, 2051, 2175]. synchrony [1510]. Syncopation [1414]. Synergistic [2350]. synonyms [1447]. Syntactic [440, 148, 979, 1319, 1734]. Syntax [18, 402, 1360, 576, 526, 1653, 2075, 19, 1623, 1553, 1958, 890, 1006, 1957]. Synthesis [1321, 348, 1921, 1730, 1829, 1752, 1790, 593, 1097, 846]. Synthesizing [1598, 1606]. System [469, 1405, 192, 49, 109, 207, 413, 398, 1293, 1250, 1728, 63, 484, 1766, 2411, 442, 1999, 2165, 859, 822, 946, 2440, 2051, 738, 994, 1877, 2091, 2376, 821, 905, 2220, 1290, 850, 559, 1435, 922, 888, 257, 617, 1516, 1909, 242, 871, 1857, 662, 1992, 910, 753, 2281, 1650, 757, 1546, 1142, 1606, 1812, 1279, 1921, 503, 397, 1052, 1908, 1799, 1669, 1781, 1231, 1984, 129, 1264, 666, 610, 2343, 347, 1437]. system-specific [559]. System-wide [1405]. Systematic [500, 600, 1808, 244]. Systems [2476, 355, 352, 656, 1925, 353, 591, 61, 182, 47, 371, 472, 2477, 842, 1644, 2108, 2043, 1345, 739, 609, 1820, 1970, 1406, 1968, 1829, 1510, 1585, 1990, 1500, 616, 1710, 1758, 794, 1240, 1323, 744, 140, 1681, 1818, 1535, 1098, 2086, 1747, 1783, 791, 2061, 1103, 162, 944, 2050, 2224, 1994, 910, 1271, 2349, 1586, 2353, 594, 336, 1987, 1839, 1508, 595, 354, 1606, 1261, 475, 1861, 841, 844, 2263, 1279, 1238, 2205, 2128, 839, 1671, 2322, 1257, 183, 834, 2290, 1196, 1306, 567, 1107, 1924, 1219, 2230, 339, 2283, 1832, 1991, 1353, 1530, 856, 830, 843, 348]. Systems-on-chip [591].

Tables [227, 2086, 1533, 514, 755, 691]. TACO [473]. tag [2133]. tag-based [2133]. tagging [1876]. Tagless [649]. taint [2306]. TAJ [2306]. takes [1817]. TAL [810]. tales [2447]. Talk [103, 428, 532, 523, 140, 151]. tall [2447]. Taming [796]. Tamper [1754, 199, 851]. Tamper-proofing [1754]. Tamper-resistant [851]. Tangible [1853]. tangled [1157]. Tapping [1984]. target [2246]. targets [587]. Tartan [1686]. task [2041, 609, 2288, 1685, 1101, 906, 2418, 856, 2343]. task-level [906]. tasks [1596, 2348, 1814]. TBNF [1540]. TCB [1969]. TCC [1268]. TCP [1700, 1499]. TCP/IP [1700]. teach [2145, 2383, 2158, 2161]. Teaching [963, 2151, 823, 711, 2144, 2150, 2154, 2155, 2157]. Teams [1421]. Technical [401, 483, 1350, 23]. Technique [412, 1406, 2308, 2029, 261, 1031, 590, 1529, 475, 819, 1823, 347]. Techniques [2271, 473, 1672, 2288, 1647, 586, 587, 220, 596, 175, 348]. Technology [1035, 465, 482, 705, 1682]. telecom [755]. temperature [1681]. Template [473, 713, 2216]. Templates [864, 2449, 2318]. Templates-based [864]. Tempo [115, 114]. Temporal [1674, 808, 2178, 509, 934]. Temporally [659]. Ten [1486, 1539, 1436]. Tenth [2099]. Tenuring [236]. terminating [1442]. Termination [1585, 1854, 1934, 1796, 1448, 1939, 655, 252, 1324, 1901]. Test [388, 1596, 1721, 460]. Testing [540, 712, 396, 5, 3, 1499, 154, 2080, 1385, 1424, 1761, 1808, 2210, 2008, 1212, 1543]. tests [624, 1265, 445]. Tetris [1823]. text [1247, 2248]. textbooks [752]. textile [1093]. their [775, 1882, 459, 2375, 1810, 2222]. Them [416]. theorem [1935, 1431, 2184]. theorems [2443, 978, 2388]. theoretic [1133, 1958]. Theoretical [1330]. theorist [1742]. theory [2131, 1730, 1948, 647, 511, 1319, 1446, 887, 1619, 1944, 733, 1734, 645, 502, 2404, 880, 2190, 1264, 2343]. There [650, 1301, 2250, 2441]. thermal [1603]. Thin [1056, 1777]. things [2139, 516, 542, 707]. thinking [2249, 2158]. third [1468, 2184]. third-party [1468]. Thorn [2412]. thoughts [2157]. Thread [765, 195, 1791, 374, 224, 2041, 1796, 2231, 1448, 1072, 2284, 1096, 658, 1546, 2426, 897, 1409, 1594]. thread-level [1072, 2284, 658, 897]. Thread-local [765]. Thread-modular [1791]. Thread-Specific [224]. Threaded [224, 1612, 2073, 1699, 2300, 1988, 1993, 803]. threading [1390, 1679, 1993]. Threads [1389, 2171, 1979, 2022, 1784, 671, 1138, 1280, 670, 1091]. threats [1739]. three [2167, 2436]. three-year [2167]. threshold [617]. throughout [2140]. throughput [1828, 675, 1892]. throw [927]. thumb [598]. tier [1331]. TIL [1054]. tiled [1684, 1804]. Time [297, 348, 780, 186, 374, 839, 858, 2357, 1028, 1817, 838, 742, 1414, 1366, 877, 1059, 820, 463, 158, 955, 568, 1510, 1029, 835, 1937, 1509, 558, 1114, 761, 743, 616, 1251, 760, 1398, 633, 1549, 2339, 1240, 530, 1061, 2070, 580, 1818, 864, 1101, 943, 1098, 655, 1826, 2446, 967, 617, 300, 162, 1053, 2042, 1001, 1415, 1655, 1824, 2010, 2015, 2348, 142, 1115, 1722, 1574, 1107, 801, 451, 819, 1027, 1832, 415, 1353, 798, 1348, 1814, 2343, 843, 175, 843, 347]. time-analysis [1366]. time-bound [530]. time-portable [1817]. Time-triggered [839]. timebombs [1674]. timely [2300]. timeout [334]. times [682]. Timestamp [185]. Timestamped [302]. timing [1833, 356]. Tina [1344]. tiny [1422, 663]. TinyOS [1399]. Tired [1355]. TLC [1713]. TMS320C6000 [363]. Tokens [95, 96]. Tolerance [208, 1679]. tolerant [1693, 1771]. Tolerating [2106, 2268]. tomography [1987]. TOOL [1230, 41, 46, 186, 1345, 1256, 154, 1460, 806, 104, 1962, 1066, 354, 1223, 1847, 1860, 2001, 546, 28]. Tool-Oriented [46]. Tools [469, 2481, 532, 1228, 1646, 1827, 2382, 1383, 1650, 1757, 2433, 2392, 1097, 2422, 1052, 849]. toolsuite [1167]. top [1549, 2213]. top-down [1549, 2213]. topics [2131]. Topology [2288]. tour [2372]. Trace [709, 226, 2339, 1474, 1913, 2338, 1102]. Trace-based [2339, 1102]. TraceBack [1384]. traces [726, 1476]. Tracing [2354, 2072, 2247]. tracked [1338]. Tracking [1903, 2189, 664, 1712, 2368, 1275, 2229, 1283]. trade [568]. trade-off [568]. Tradeoffs [1705, 664]. trading [1978, 1950]. train [763]. traits [935, 1890, 1252]. TRANSACT [2005]. transaction [1829, 1789, 2353]. transaction-aware [1789]. transaction-level [2353]. Transactional [2269, 1625, 2071, 657, 1931, 1552, 2366, 1550, 1703, 1705, 2453, 1702, 2286, 2233, 2312, 2391, 1919, 2040, 2203, 1268, 2168, 1656, 2251, 2349, 2313, 1704, 2111, 2265, 1589, 2283, 2252]. Transactions [1891, 2266, 956, 1551, 2291, 1930, 2267]. Transactors [1329]. transducer [1575]. transfer [1906, 1590, 970]. transfers [1593]. Transform [437, 1060, 2403]. transformation [500, 148, 1632, 1852, 1331, 1694, 972, 1258]. Transformations [7, 1166, 2240, 1820, 971, 1237, 925, 1332, 508, 1342, 2424, 2422, 2117, 351]. transformers [146, 1442]. Transforming [924, 77]. transforms [1394]. Transition [1324, 1924]. transitions [721]. transitive [1676]. Translating [826, 1140, 535]. Translation [70, 107, 667, 1907, 2081, 1612, 2239, 2261, 1857, 1404, 1765, 1927, 166]. translation-based [1404]. translational [1540]. translator [1819, 283, 284]. Transparent [112, 1398, 63, 113, 2422, 2036, 1737, 2101].

Transport [844]. Trap [196]. traversal [457]. treat [999]. Tree [240, 1335, 1554, 1332, 1503, 1006]. tree-parsing [1554]. Trees [18, 2078, 781, 19, 2184, 1848, 359]. trends [1971, 367]. TRICK [1712]. Tridirectional [993]. triggered [839]. TRIPS [2220]. trivial [623]. trusted [665]. TTA [845]. tunable [2290]. tuning [1845]. tunneling [1065]. Turbocharging [2291]. TwinDrivers [2245]. Two [109, 410, 2082, 1970, 1167, 1893, 1928]. two-dimensional [1970]. Two-Edge [109]. Two-Level [410, 1167, 1893]. two-stage [2082]. Type [298, 2023, 420, 404, 408, 81, 409, 1898, 777, 222, 2325, 396, 1439, 646, 2462, 1866, 250, 2066, 268, 260, 134, 1956, 1644, 1728, 1133, 975, 739, 2127, 2411, 442, 1447, 2165, 930, 1374, 1772, 700, 2195, 647, 2051, 738, 994, 1429, 1723, 1877, 637, 821, 1515, 553, 1372, 2339, 1319, 2068, 869, 888, 491, 257, 1618, 1516, 1755, 1857, 2089, 2174, 1806, 887, 1619, 496, 1142, 870, 249, 2125, 1812, 733, 1958, 2454, 1720, 1425, 1514, 1437, 2067, 2128, 2403, 2420, 2405, 503, 2129, 1318, 1054, 933, 136, 2388, 1662, 2036, 755, 328]. Type-base [250]. Type-based [408, 81, 2325, 2462]. type-containment [1437]. type-directed [975, 2403, 1054]. type-error [1806]. Type-indexed [268]. type-inference [1437]. Type-preserving [2023, 260, 1772, 2068]. Type-Safe [404, 409, 777, 1866, 134, 496, 2125, 2036]. type-theoretic [1133, 1958]. typechecking [993, 639, 945, 2102]. Typed [2076, 2063, 1035, 135, 417, 2324, 784, 975, 1741, 1316, 1718, 735, 1562, 1954, 1375, 649, 1771, 2454, 977, 2404, 730, 1960, 785, 418, 729, 1125]. typeful [925, 892]. Types [488, 420, 1141, 776, 2204, 1142, 410, 1727, 487, 2214, 783, 535, 695, 736, 820, 1768, 137, 1313, 1658, 147, 554, 950, 1899, 1856, 976, 1143, 1798, 269, 455, 1949, 130, 1501, 1887, 1454, 1455, 2087, 876, 1438, 2174, 1494, 1592, 2441, 2199, 1906, 2083, 870, 1430, 2126, 1869, 1514, 2209, 2173, 2021, 2374, 266, 1062, 510, 872, 2194, 132, 1521, 974, 1503, 1637, 419, 982, 732]. Typestate [2120, 2113, 1897, 1008]. Typestate-like [2120]. typestates [1458]. Typing [644, 267, 782, 1044, 2408, 2421, 2104, 2444, 1617]. typings [995, 1439].

ubiquitous [1763, 1531]. UCC [1802]. UK [215]. Ulterior [953]. Ultimate [20, 1207, 2383, 21]. Ultra [309, 1678, 1270]. Ultra-fast [309]. UML [1227, 1247, 168, 1266, 1965, 461]. unanticipated [1890]. unbinding [1726]. Unbounded [1703]. uncertain [1051]. unchecked [1480]. uncomputation [1694]. uncooperative [772]. uncovering [1483, 1916]. undefined [1903]. underapproximation [1323]. underapproximation-widening [1323]. Undergraduate [320, 321, 2141, 2144, 2160]. Understanding [1665, 582, 762, 662, 1992, 1987, 1679, 152, 894, 701, 2438, 2146, 1169, 2469, 1720, 178]. undue [2364]. unfolding [1002]. Unification [66, 410, 1618]. Unification-based [66, 1618]. Unified [307, 2041, 1126, 1413]. Uniform [18, 19, 1649, 1484, 491, 673]. unify [1243]. Unifying [2195, 1226]. Unimo [1639]. union [1592]. uniprocessor [1413]. unique [2079]. uniqueness [1336]. unit [1212, 1265]. unit-testing [1212]. United [2482]. Units [60, 59, 1603, 1621]. universal [872, 1064]. universe [2384, 1906]. University [551]. unloading [1108]. unlocking [2109]. unrealized [1710]. unreliable [1329]. unrestricted [1555]. unrolling [2055]. unsafe [1563, 2119]. unshared [2070]. unsplitting [2347]. untangled [458]. until [516]. untrusted [258]. Update [877, 1335, 1332, 1802]. update-conscious [1802]. updates [2062, 715, 164, 2300, 2299, 1587]. Updating [286, 1556, 1929, 1328]. upgrade [2117]. upgrades [957]. upon [1327]. Upper [1418]. upward [2184]. urban [2447]. USA [319, 2477, 2481, 2483]. usage [2113, 492, 1872, 734, 513, 780, 1115]. Use [2154, 913, 914, 1276, 791, 1533, 614, 2461, 571]. used [1918]. useful [91]. User [1890, 1871, 907, 883, 1974]. user-centred [883]. User-changeable [1890]. User-controllable [907]. User-friendly [1871]. user-perceived [1974]. uses [1776, 1701]. Using [1911, 189, 231, 393, 300, 370, 574, 1757, 8, 232, 897, 1901, 2158, 1463, 9, 3, 919, 768, 1191, 2269, 2423, 1816, 2319, 1471, 1766, 802, 1868, 2178, 1217, 2238, 1654, 1754, 68, 1811, 1691, 2035, 1807, 2044, 84, 1645, 403, 2391, 724, 998, 1339, 1995, 1281, 309, 89, 1629, 1837, 1030, 1762, 596, 1641, 1101, 748, 766, 1682, 1220, 1664, 1593, 855, 2327, 1941, 1008, 1570, 590, 2256, 341, 533, 1964, 642, 156, 1876, 1475, 277, 278, 2313, 594, 1692, 1769, 1430, 2392, 2383, 896, 1778, 2329, 441, 1179, 1260, 996, 1722, 2223, 2318, 1701, 933]. using [1138, 2317, 1045, 1231, 450, 1078, 710, 73, 1264, 1341, 248, 1069, 2455, 1461, 1088, 2252, 1480]. Utah [2481, 2483]. utilization [742, 2262]. Utilizing [196]. uu [127, 126].

Valgrind [1775]. Validating [1255].

Validation

[411, 2465, 2225, 70, 1860, 1922, 2326, 2236]. validators [1927]. validity [781]. Value [198, 867, 197, 1903, 1875, 572, 1811, 557, 998, 2192, 1592, 885, 960, 670]. Value-Centric [197]. value-flow [1811]. valued [248]. values [255, 2210]. Vancouver [2478]. vanish [631]. variable [1404, 67, 852]. Variables [310, 1474, 1037, 687]. Variance [1738]. Variant [1887, 1658]. variants [511]. VDM [1963]. VDMTools [1963]. vEC [2463]. vector [522, 589]. Vectorization [1074, 1561]. Velodrome [2033]. Verifiable [2316, 1528]. Verification [1146, 2357, 964, 1256, 1912, 1933, 558, 2359, 2112, 601, 1448, 1584, 262, 501, 1744, 2206, 2471, 2446, 2325, 2204, 1941, 1746, 151, 1900, 2318, 1927, 2277, 2038]. Verified [2166, 2326]. verifier [2211]. verifiers [1628]. Verifying [2113, 2177, 1504, 2205, 248, 1069, 2380, 1369]. Verilog [12]. Verischemelog [12]. version [1929]. version-consistent [1929]. versioning [2427, 266]. versus [789, 825, 1251, 754]. vertical [1470]. Very [15, 1728, 895, 833, 1212]. via [235, 484, 1017, 607, 1611, 167, 299, 1131, 1689, 1342, 805, 1675, 1505, 1615, 371, 595, 2311, 1284, 1433, 527, 410, 1275, 1850, 2007, 1280, 2190, 1099]. view [1332, 1207, 1852]. views [948, 1440, 1613, 270]. violation [1369, 2222, 1367]. violations [1675]. Virgil [1652]. Virtual [1404, 1812, 833, 747, 1471, 1647, 1674, 447, 942, 1518, 816, 1641, 1673, 2463, 1007, 663, 1286, 2468, 2123, 461, 173, 1924, 1280, 1976, 2455, 179, 552]. virtualization [1672, 1982, 1968, 1705, 1671]. virtualization-based [1968]. virtualized [1970, 2246]. viruses [517]. visibility [1890]. vision [2377]. visitation [866]. Visitor [457, 2125]. VISTA [610, 2165]. visual [1999, 643, 1896, 2001, 2103, 1489]. Visualisation [1220, 1186, 703]. Visualising [763]. visualization [1177, 1244, 459, 1760]. visualizing [2097, 1987, 1261]. visually [520]. VLIW [363, 1831, 601, 604, 857, 1031, 1603, 605, 1823, 1835]. VLSI [1752, 505]. VM [2413, 2299, 1799]. VM-centric [2299]. voices [2436]. voltage [1816, 1602, 1566, 1101, 592, 1289, 798, 1416, 1815, 347]. voltage/frequency [1289]. volume [2012]. vs [773, 342, 1204, 187, 188, 1472, 2175, 1130, 1033]. vulnerabilities [1770]. VYRD [1369].

W [707]. wait [953]. walks [1970]. wall [2425]. WAN [683]. wanders [1742]. want [636]. watermarking [1076, 984]. way [1361, 787, 1157, 2293]. WCET [1597, 1838]. WCRT [1413]. Weak [235]. weakest [2363, 1766, 2330]. weakly [1562, 2006]. weaver [843]. Weaving [518, 1167, 1909]. web [2402, 1896, 2401, 1861, 2306, 37, 38, 41, 986, 1948, 1871, 13, 2332, 1065, 675, 131, 1238, 1526, 1770, 1862, 1253]. web-applications [1896]. Web-Based [37, 41, 38, 2332]. WebRB [1896]. weight [2208]. weighted [2256]. weightsum [142]. Welcome [51]. well [994, 1504]. well-founded [994, 1504]. were [516, 542, 707]. West [437]. where [385]. whether [1082]. while [1386, 1132, 1842]. white [1543]. white-box [1543]. whitebox [2025]. Whiteoak [2104]. Whole [302, 1248, 813, 261, 1686, 1555, 851, 1088]. whole-program [813, 261, 1555, 1088]. wide [1405, 1835, 333]. wide-area [333]. wide-issue [1835]. widening [1323]. width [1100]. Wild [2181, 1586]. wildlife [664]. will [1205, 22, 23]. window [1118]. windows [1033, 2165, 788]. WinHIPE [1760]. Winners [537]. wire [673]. wire-delay [673]. wireless [1785, 1802]. wishes [1882]. within [2429, 1264]. without [181, 294, 953, 447, 837, 404, 1863, 1379, 690, 1570, 2407, 2242, 170, 2340, 1119]. Witnessing [1434]. work [2440, 1198, 1305, 2254, 1142, 1536, 1861]. Workflow [684, 398, 397]. Working [539, 1847]. workloads [1993]. Workshop [2484, 33, 2480, 319, 2481, 2136, 2137, 2138, 2136, 2138, 2163]. workstations [337]. World [710, 955, 812, 329, 1160, 1792, 162, 1998, 2389]. World-class [710]. worms [517]. worrying [2291]. worship [1185]. worst [1826]. worst-case [1826]. Write [2077, 547, 694, 775, 2186]. Writing [2432, 2389]. written [2468]. wrong [1135, 2242, 1199, 1304]. WSAgent [1238].

X [872]. X10 [1484, 2289]. x86 [1672, 2336, 2201]. x86-CC [2201]. XDuce [1631]. XIRC [1175]. XMem [2036]. XML [873, 2063, 2062, 1141, 2393, 1798, 130, 251, 1317, 748, 1632, 718, 1922]. XML-centric [873]. XML/HTML [748]. Xoc [1990]. XP [1206]. XPath [1731]. XPath-like [1731]. XScale^{T M} [1104]. XSnippet [1666]. XTREM [1104]. XVF [866].

year [1202, 2167]. Years [2485, 1633, 1436, 1486, 1539, 1144]. Yi [2212]. yield [2244]. Ynot [2083].

Z [1123]. Z0 [1422]. ZebraNet [664, 1586]. Zephyr [586]. Zero [20, 21]. Zero-Instruction [20, 21]. zip [641]. zip-like [641]. zips [1152]. Zobel [396]. zone [1914].

References

Clinger:1990:HRF

[1] William D. Clinger. How to read floating point numbers accurately. ACM SIGPLAN Notices, 25(6):92– 101, June 1990. CODEN SIN-ODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/93542/p92-clinger/

. See also output algorithms in [?, 2, ?, ?, 1041].

Steele:1990:HPF

[2] Guy L. Steele, Jr. and Jon L. White. How to print floating-point numbers accurately. ACM SIGPLAN Notices, 25(6):112–126, June 1990. CODEN SINODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/93542/p112-steele/

. See also input algorithm in [1, 1040], and a faster output algorithm in [?] and [?], IBM S/360 algorithms in [?] for both IEEE 754 and S/360 formats, and a twenty-year retrospective in [1041]. In electronic mail dated Wed, 27 Jun 1990 11:55:36 EDT, Guy Steele reported that an intrepid pre-SIGPLAN 90 conference implementation of what is stated in the paper revealed 3 mistakes:

1. Table 5 (page 124): insert k <-- 0 after assertion, and also delete k <-- 0 from Table 6.

2. Table 9 (page 125):

     for  -1:USER!("");
     substitute -1:USER!("0");

and delete the comment.

3. Table 10 (page 125):

     for  fill(-k, "0")
     substitute fill(-k-1, "0")

Sirer:2000:UPG

[3] Emin Gün Sirer and Brian N. Bershad. Using production grammars in software testing. ACM SIGPLAN Notices, 35 (1):1–13, January 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nakatani:2000:JDE

[4] Lloyd H. Nakatani, Mark A. Ardis, Robert G. Olsen, and Paul M. Pontrelli. Jargons for domain engineering. ACM SIGPLAN Notices, 35(1): 15–24, January 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reichwein:2000:SSI

[5] James Reichwein, Gregg Rothermel, and Margaret Burnett. Slicing spreadsheets: An integrated methodology for spreadsheet testing and debugging. ACM SIGPLAN Notices, 35(1):25–38, January 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Guyer:2000:ALO

[6] Samuel Z. Guyer and Calvin Lin. An annotation language for optimizing software libraries. ACM SIGPLAN Notices, 35(1):39–52, January 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Menon:2000:CSL

[7] Vijay Menon and Keshav Pingali. A case for source-level transformations in MATLAB. ACM SIGPLAN Notices, 35(1):53–65, January 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Parson:2000:UJR

[8] Dale E. Parson. Using Java reflection to automate extension language parsing. ACM SIGPLAN Notices, 35(1): 67–80, January 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sheard:2000:DIU

[9] Tim Sheard, Zine el-abidine Benaissa, and Emir Pasalic. DSL implementation using staging and monads. ACM SIG-PLAN Notices, 35(1):81–94, January 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Peterson:2000:MR

[10] John Peterson and Greg Hager. Monadic robotics. ACM SIGPLAN Notices, 35(1):95–108, January 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Leijen:2000:DSE

[11] Daan Leijen and Erik Meijer. Domainspecific embedded compilers. ACM SIGPLAN Notices, 35(1):109–122, January 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jennings:2000:VVE

[12] James Jennings and Eric Beuscher. Verischemelog: Verilog embedded in Scheme. ACM SIGPLAN Notices, 35 (1):123–134, January 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fernandez:2000:DSD

[13] Mary Fernández, Dan Suciu, and Igor Tatarinov. Declarative specification of data-intensive Web sites. ACM SIG-PLAN Notices, 35(1):135–148, January 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Li:2000:CSL

[14] Du Li and Richard R. Muntz. A collaboration specification language. ACM SIGPLAN Notices, 35(1):149–162, January 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bonachea:2000:HLP

[15] Dan Bonachea, Kathleen Fisher, Anne Rogers, and Frederick Smith. Hancock: A language for processing very large-scale data. ACM SIGPLAN Notices, 35(1):163–176, January 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ryder:2000:CCP

[16] Barbara G. Ryder. Conference corner: PLDI ’99: Programming language design and implementation (report). ACM SIGPLAN Notices, 35(2):9, February 2000. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/r/Ryder:
Barbara_G=.html.

Ryder:2000:PPL

[17] Barbara Ryder. PLDI ’99: Programming language design and implementation. ACM SIGPLAN Notices, 35(2): 9–10, 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clark:2000:PPPa

[18] Chis Clark. Practical parsing patterns: Uniform abstract syntax trees. ACM SIGPLAN Notices, 35(2):11–16, February 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//ftp.informatik.rwth-aachen.de/

                                        dblp/db/indices/a-tree/c/Clark:
                                        Chis.html.

Clark:2000:UAS

[19] Chris Clark. Uniform abstract syntax trees. ACM SIGPLAN Notices, 35(2): 11–16, 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:2000:FUR

[20] Paul Frenger. Forth: The ultimate RISC: A zero-instruction computer. ACM SIGPLAN Notices, 35(2):17–23, February 2000. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.

                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/f/Frenger:Paul.
                                        html.

Frenger:2000:URZ

[21] Paul Frenger. The Ultimate RISC: A zero-instruction computer. ACM SIGPLAN Notices, 35(2):17–24, 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reinholtz:2000:JWF

[22] Kirk Reinholtz. Java will be faster than C++. ACM SIGPLAN Notices, 35(2): 25–28, 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reinholtz:2000:TCJ

[23] Kirk Reinholtz. Technical correspondence: Java will be faster than C++. ACM SIGPLAN Notices, 35(2):25–28, February 2000. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/r/Reinholtz:Kirk.
html.

Davis:2000:OOA

[24] Matthew S. Davis. An object oriented approach to constructing recursive descent parsers. ACM SIG-PLAN Notices, 35(2):29–35, February 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/d/Davis:
Matthew_S=.html.

Milicev:2000:SPS

[25] Dragan Milicev and Zoran Jovanovic. Sources of parallelism in software pipelining loops with conditional branches. ACM SIGPLAN Notices, 35(2):36–45, 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Milicev:2000:SSP

[26] Dragan Milicev and Zoran Jovanovic. Sources in software pipelining loops with conditions branches. ACM SIG-PLAN Notices, 35(2):36–45, February 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/j/Jovanovic:Zoran.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
m/Milicev:Dragan.html.

Simonis:2000:ABO

[27] Volker Simonis. Adapters and binders — overcoming problems in the design and implementation of the C++- STL. ACM SIGPLAN Notices, 35 (2):46–53, February 2000. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/s/Simonis:Volker.
html.

Xie:2000:CST

[28] Gaoyan Xie, Yongsen Xu, Yu Li, and Qian Li. Codebugger: a software tool for cooperative debugging. ACM SIG-PLAN Notices, 35(2):54–60, February 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/l/Li:Qian.html;
 http://ftp.informatik.rwth-aachen.
de/dblp/db/indices/a-tree/l/Li:
Yu.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/x/Xie:Gaooyan.html;   http:
//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/x/Xu:YongSen.
html.

Swen:2000:OOP

[29] Bing Swen. Object-oriented programming with induction. ACM SIG-PLAN Notices, 35(2):61–67, February 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/s/Swen:Bing.html.

Anonymous:2000:IIA

[30] Anonymous. Important information for ACM SIGPLAN Notices authors. ACM SIGPLAN Notices, 35(3):1, March 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Soffa:2000:ECN

[31] Mary Lou Soffa. Executive Committee news: 1999 SIGPLAN awards. ACM SIGPLAN Notices, 35(3):2–3, March 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergmann:2000:CCCa

[32] Seth Bergmann. Conference corner: Calendar. ACM SIGPLAN Notices, 35 (3):4–15, March 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pontelli:2000:CCC

[33] Enrico Pontelli and Vitor Santos Costa. Conference corner: Conference report: PADL ’00: Workshop on Practical Aspects of Declarative Languages. ACM SIGPLAN Notices, 35(3):16–17, March 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chalk:2000:CCC

[34] Peter Chalk. Conference corner: Conference report: JICC4: Java in the Computing Curricula. ACM SIG-PLAN Notices, 35(3):18–19, March 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chalk:2000:JJC

[35] Peter Chalk. JICC4: Java in the computing curricula. ACM SIGPLAN Notices, 35(3):18–19, March 2000. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2000:CPI

[36] Anonymous. Call for papers: 2000 International Symposium on Memory Management (ISMM), Minneapolis, Minnesota, October 15–16, 2000. ACM SIGPLAN Notices, 35(3):20, March 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Benson:2000:JRS

[37] Brent W. Benson, Jr. Java reflections: Servlets: Web-based applications you can live with. ACM SIG-PLAN Notices, 35(3):21–24, March 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Benson:2000:WBA

[38] Brent W. Benson, Jr. Web-based applications you can live with. ACM SIGPLAN Notices, 35(3):21–24, March 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:2000:FRL

[39] Paul Frenger. Forth report: Learning Forth with Modular Forth. ACM SIGPLAN Notices, 35(3):25–30, March 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:2000:LFM

[40] Paul Frenger. Learning Forth with Modular Forth. ACM SIGPLAN Notices, 35(3):25–30, March 2000. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergin:2000:HPL

[41] Tim Bergin. History of programming languages and software engineering: A Web-based tool. ACM SIG-PLAN Notices, 35(3):31, March 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Padua:2000:CIA

[42] David Padua and Ron Cytron. Compilers and interpreters archive. ACM SIG-PLAN Notices, 35(3):32, March 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergmann:2000:IA

[43] A. Michael Bergmann. Information for authors. ACM SIGPLAN Notices, 35(3):33, March 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Li:2000:SPN

[44] Yu Li, Gaoyan Xie, Yongsen Xu, and Yi Yang. The security problem of nested classes. ACM SIG-PLAN Notices, 35(3):34–38, March 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yu:2000:SPN

[45] Li Yu, Xie Gaoyan, Xu Yongsen, and Yang Yi. The security problem of nested classes. ACM SIG-PLAN Notices, 35(3):34–38, March 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Heering:2000:SPL

[46] Jan Heering and Paul Klint. Semantics of programming languages: A tool-oriented approach. ACM SIG-PLAN Notices, 35(3):39–48, March 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Naik:2000:EMD

[47] Mayur Naik and Rajeev Kumar. Efficient message dispatch in objectoriented systems. ACM SIGPLAN Notices, 35(3):49–58, March 2000. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kokol:2000:RCN

[48] Peter Kokol and Vili Podgorelec. Ranking the complexity of NIAM conceptual schemas by alpha metric. ACM SIGPLAN Notices, 35(3):59–64, March 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Li:2000:INN

[49] Tao Li and XiaoJie Liu. An intelligent Neural Network Programming System (NNPS). ACM SIGPLAN Notices, 35 (3):65–72, March 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ramsey:2000:IPC

[50] Norman Ramsey and Kevin Scott. The 1999 ICFP programming contest. ACM SIGPLAN Notices, 35(3):73–83, March 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berman:2000:LEW

[51] A. Michael Berman. Letter from the editor; welcome ever smiles / and farewell goes out sighing. ACM SIG-PLAN Notices, 35(4):1, April 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berman:2000:IIA

[52] A. Michael Berman. Important information for ACM SIGPLAN Notices authors. ACM SIGPLAN Notices, 35(4): 2, April 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergmann:2000:CCCb

[53] Seth Bergmann. Conference corner: Calendar. ACM SIGPLAN Notices, 35 (4):3–16, April 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2000:CPA

[54] Anonymous. Call for papers: The 28th Annual ACM-SIGACT Symposium Principles of Programming Languages, London, January 17–19, 2001 (POPL’01). ACM SIGPLAN Notices, 35(4):17, April 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clark:2000:NLS

[55] Chris Clark. Newlines and lexer states. ACM SIGPLAN Notices, 35(4):18–24, April 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clark:2000:PPPb

[56] Chris Clark. Practical parsing patterns: Newlines and lexer states. ACM SIGPLAN Notices, 35(4):18–24, April 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Norris:2000:IA

[57] Cindy Norris. Information for authors. ACM SIGPLAN Notices, 35 (4):25, April 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bruce:2000:CDF

[58] Kim B. Bruce. Curriculum 2001 draft found lacking in programming languages. ACM SIGPLAN Notices, 35 (4):26–28, April 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2000:PKU

[59] Anonymous. Proposed knowledge units for programming languages for Curriculum 2001. ACM SIGPLAN Notices, 35(4):29–43, April 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Group:2000:PKU

[60] Programming Language Knowledge Focus Group. Proposed knowledge units for programming languages for Curriculum 2001. ACM SIGPLAN Notices, 35(4):29–43, April 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Manjunath:2000:SHJ

[61] Geetha Manjunath and Venkatesh Krishnan. A Small hybrid JIT for embedded systems. ACM SIGPLAN Notices, 35(4):44–50, April 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sakharov:2000:HSM

[62] Alexander Sakharov. A hybrid state machine notation for component specification. ACM SIGPLAN Notices, 35 (4):51–56, April 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bala:2000:DTDa

[63] Vasanth Bala, Evelyn Duesterwald, and Sanjeev Banerjia. Dynamo: a transparent dynamic optimization system. ACM SIGPLAN Notices, 35(5):1–12, May 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/pldi/
349299/p1-bala/p1-bala.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
pldi/349299/p1-bala/.

Cierniak:2000:PJJ

[64] Michał Cierniak, Guei-Yuan Lueh, and James M. Stichnoth. Practicing JUDO: Java under dynamic optimizations. ACM SIGPLAN Notices, 35(5):13–26, May 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/pldi/
349299/p13-cierniak/p13-cierniak.
pdf;   http://www.acm.org/pubs/

citations/proceedings/pldi/349299/
p13-cierniak/.

Lucco:2000:SSD

[65] Steven Lucco. Split-stream dictionary program compression. ACM SIGPLAN Notices, 35(5):27–34, May 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://

delivery.acm.org/10.1145/350000/
349307/p27-lucco.pdf;   http://
www.acm.org/pubs/articles/proceedings/
pldi/349299/p27-lucco/p27-lucco.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pldi/349299/
p27-lucco/.

Das:2000:UBP

[66] Manuvir Das. Unification-based pointer analysis with directional assignments. ACM SIGPLAN Notices, 35(5):35–46, May 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/pldi/
349299/p35-das/p35-das.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
pldi/349299/p35-das/.

Rountev:2000:LVS

[67] Atanas Rountev and Satish Chandra. Off-line variable substitution for scaling points-to analysis. ACM SIGPLAN Notices, 35(5): 47–56, May 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/pldi/
349299/p47-rountev/p47-rountev.
pdf;   http://www.acm.org/pubs/

citations/proceedings/pldi/349299/
p47-rountev/.

Cheng:2000:MIP

[68] Ben-Chung Cheng and Wen mei W. Hwu. Modular interprocedural pointer analysis using access paths: design, implementation, and evaluation. ACM SIGPLAN Notices, 35(5):57–69, May 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/pldi/
349299/p57-cheng/p57-cheng.pdf;
http://www.acm.org/pubs/citations/
proceedings/pldi/349299/p57-cheng/

.

Xu:2000:SCM

[69] Zhichen Xu, Barton P. Miller, and Thomas Reps. Safety checking of machine code. ACM SIGPLAN Notices, 35(5):70–82, May 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/pldi/
349299/p70-xu/p70-xu.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
pldi/349299/p70-xu/.

Necula:2000:TVO

[70] George C. Necula. Translation validation for an optimizing compiler. ACM SIGPLAN Notices, 35(5):83–94, May 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/pldi/
349299/p83-necula/p83-necula.pdf;
http://www.acm.org/pubs/citations/

proceedings/pldi/349299/p83-necula/

.

Colby:2000:CCJ

[71] Christopher Colby, Peter Lee, George C. Necula, Fred Blau, Mark Plesko, and Kenneth Cline. A certifying compiler for Java. ACM SIGPLAN Notices, 35(5):95–107, May 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/pldi/
349299/p95-colby/p95-colby.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pldi/349299/p95-colby/

.

Stephenson:2000:BAA

[72] Mark Stephenson, Jonathan Babb, and Saman Amarasinghe. Bidwidth analysis with application to silicon compilation. ACM SIGPLAN Notices, 35(5):108–120, May 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/pldi/
349299/p108-stephenson/p108-stephenson.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pldi/349299/
p108-stephenson/.

Wilken:2000:OIS

[73] Kent Wilken, Jack Liu, and Mark Heffernan. Optimal instruction scheduling using integer programming. ACM SIGPLAN Notices, 35 (5):121–133, May 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/pldi/
349299/p121-wilken/p121-wilken.

pdf;   http://www.acm.org/pubs/
citations/proceedings/pldi/349299/
p121-wilken/.

Zalamea:2000:ISC

[74] Javier Zalamea, Josep Llosa, Eduard Ayguadé, and Mateo Valero. Improved spill code generation for software pipelined loops. ACM SIGPLAN Notices, 35(5):134–144, May 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/articles/proceedings/pldi/
349299/p134-zalamea/p134-zalamea.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pldi/349299/
p134-zalamea/.

Larsen:2000:ESL

[75] Samuel Larsen and Saman Amarasinghe. Exploiting superword level parallelism with multimedia instruction sets. ACM SIGPLAN Notices, 35(5):145–156, May 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/pldi/
349299/p145-larsen/p145-larsen.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pldi/349299/
p145-larsen/.

Lin:2000:CAI

[76] Yuan Lin and David Padua. Compiler analysis of irregular memory accesses. ACM SIGPLAN Notices, 35 (5):157–168, May 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/pldi/

                                        349299/p157-lin/p157-lin.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/pldi/349299/p157-lin/

.

Yi:2000:TLR

[77] Qing Yi, Vikram Adve, and Ken Kennedy. Transforming loops to recursion for multi-level memory hierarchies. ACM SIGPLAN Notices, 35(5):169–181, May 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/articles/proceedings/pldi/
                                        349299/p169-yi/p169-yi.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        pldi/349299/p169-yi/.

Rugina:2000:SBA

[78] Radu Rugina and Martin Rinard. Symbolic bounds analysis of pointers, array indices, and accessed memory regions. ACM SIGPLAN Notices, 35 (5):182–195, May 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/articles/proceedings/pldi/
                                        349299/p182-rugina/p182-rugina.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pldi/349299/
                                        p182-rugina/.

Sreedhar:2000:FIO

[79] Vugranam C. Sreedhar, Michael Burke, and Jong-Deok Choi. A framework for interprocedural optimization in the presence of dynamic class loading. ACM SIGPLAN Notices, 35 (5):196–207, May 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/pldi/
349299/p196-sreedhar/p196-sreedhar.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pldi/349299/
p196-sreedhar/.

Ruf:2000:ESR

[80] Erik Ruf. Effective synchronization removal for Java. ACM SIGPLAN Notices, 35(5):208–218, May 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/articles/proceedings/pldi/
349299/p208-ruf/p208-ruf.pdf;
http://www.acm.org/pubs/citations/
proceedings/pldi/349299/p208-ruf/

.

Flanagan:2000:TBR

[81] Cormac Flanagan and Stephen N. Freund. Type-based race detection for Java. ACM SIGPLAN Notices, 35(5):219–232, May 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/pldi/
349299/p219-flanagan/p219-flanagan.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pldi/349299/
p219-flanagan/.

Ramalingam:2000:LDD

[82] G. Ramalingam. On loops, dominators, and dominance frontier. ACM SIGPLAN Notices, 35(5):233– 241, May 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/pldi/

                                        349299/p233-ramalingam/p233-ramalingam.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pldi/349299/
                                        p233-ramalingam/.

Wan:2000:FRP

[83] Zhanyong Wan and Paul Hudak. Functional reactive programming from first principles. ACM SIGPLAN Notices, 35(5):242–252, May 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/articles/proceedings/pldi/
                                        349299/p242-wan/p242-wan.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/pldi/349299/p242-wan/

.

Fahndrich:2000:SCS

[84] Manuel Fähndrich, Jakob Rehof, and Manuvir Das. Scalable contextsensitive flow analysis using instantiation constraints. ACM SIGPLAN Notices, 35(5):253–263, May 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

                                        org/pubs/articles/proceedings/pldi/
                                        349299/p253-fahndrich/p253-fahndrich.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pldi/349299/
                                        p253-fahndrich/.

Cannarozzi:2000:CGC

[85] Dante J. Cannarozzi, Michael P. Plezbert, and Ron K. Cytron. Contaminated garbage collection. ACM SIGPLAN Notices, 35(5):264– 273, May 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.acm.org/

pubs/articles/proceedings/pldi/
349299/p264-cannarozzi/p264-cannarozzi.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pldi/349299/
p264-cannarozzi/.

Domani:2000:GFG

[86] Tamar Domani, Elliot K. Kolodner, and Erez Petrank. A generational on-the-fly garbage collector for Java. ACM SIGPLAN Notices, 35(5):274–284, May 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/pldi/
349299/p274-domani/p274-domani.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pldi/349299/
p274-domani/.

Ramsey:2000:SIL

[87] Norman Ramsey and Simon Peyton Jones. A single intermediate language that supports multiple implementations of exceptions. ACM SIGPLAN Notices, 35(5):285– 298, May 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/pldi/
349299/p285-ramsey/p285-ramsey.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pldi/349299/
p285-ramsey/.

Boothe:2000:EAB

[88] Bob Boothe. Efficient algorithms for bidirectional debugging. ACM SIG-PLAN Notices, 35(5):299–310, May 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.

                                        acm.org/pubs/articles/proceedings/
                                        pldi/349299/p299-boothe/p299-boothe.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pldi/349299/
                                        p299-boothe/.

Heydon:2000:CFC

[89] Allan Heydon, Roy Levin, and Yuan Yu. Caching function calls using precise dependencies. ACM SIGPLAN Notices, 35(5):311–320, May 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

                                        org/pubs/articles/proceedings/pldi/
                                        349299/p311-heydon/p311-heydon.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pldi/349299/
                                        p311-heydon/.

Bodik:2000:AEA

[90] Rastislav Bodík, Rajiv Gupta, and Vivek Sarkar. ABCD: eliminating array bounds checks on demand. ACM SIGPLAN Notices, 35 (5):321–333, May 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/articles/proceedings/pldi/
                                        349299/p321-bodik/p321-bodik.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pldi/349299/p321-bodik/

.

Ghemawat:2000:FAG

[91] Sanjay Ghemawat, Keith H. Randall, and Daniel J. Scales. Field analysis: getting useful and low-cost interprocedural information. ACM SIGPLAN Notices, 35(5):334–344, May 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/articles/proceedings/pldi/
349299/p334-ghemawat/p334-ghemawat.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pldi/349299/
p334-ghemawat/.

Dolby:2000:AOI

[92] Julian Dolby and Andrew Chien. An automatic object inlining optimization and its evaluation. ACM SIGPLAN Notices, 35(5):345–357, May 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/articles/proceedings/pldi/
349299/p345-dolby/p345-dolby.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pldi/349299/p345-dolby/

.

Cytron:2000:LC

[93] Ron Cytron. Letter from the Chair. ACM SIGPLAN Notices, 35(6):1, June 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergmann:2000:CCCc

[94] Seth Bergmann. Conference corner: Calendar. ACM SIGPLAN Notices, 35(6):2–14, June 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clark:2000:PPPc

[95] Chris Clark. Practical parsing patterns: Off-and-on tokens. ACM SIG-PLAN Notices, 35(6):15–20, June 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clark:2000:T

[96] Chris Clark. Off-and-on tokens. ACM SIGPLAN Notices, 35(6):15–20, June 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:2000:FRG

[97] Paul Frenger. Forth report: The GOTO machine. ACM SIGPLAN Notices, 35(6):21–24, June 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:2000:GM

[98] Paul Frenger. GOTO machine. ACM SIGPLAN Notices, 35(6):21–24, June 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fenwick:2000:IAa

[99] Jan Fenwick and Cindy Norris. Information for authors. ACM SIG-PLAN Notices, 35(6):25, June 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vanDeursen:2000:DSL

[100] Arie van Deursen, Paul Klint, and Joost Visser. Domain-specific languages: An annotated bibliography. ACM SIGPLAN Notices, 35(6):26–36, June 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Srikant:2000:EPA

[101] Y. N. Srikant and D. V. Ravindra. Effective parametrization of architectural registers for register allocation algorithms. ACM SIGPLAN Notices, 35 (6):37–46, June 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schupp:2000:LLC

[102] Sibylle Schupp. Lazy lists in C++. ACM SIGPLAN Notices, 35(6):47–54, June 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smith:2000:OCF

[103] Michael D. Smith. Overcoming the challenges to feedback-directed optimization (keynote talk). ACM SIG-PLAN Notices, 35(7):1–11, July 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Engler:2000:DTA

[104] Dawson R. Engler and Wilson C. Hsieh. Derive: a tool that automatically reverse-engineers instruction encodings. ACM SIGPLAN Notices, 35 (7):12–22, July 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Johansson:2000:PGO

[105] Erik Johansson and Sven olof Nyström. Profile-guided optimization across process boundaries. ACM SIGPLAN Notices, 35(7):23–31, July 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Voss:2000:FRD

[106] Michael J. Voss and Rudolf Eigenmann. A framework for remote dynamic program optimization. ACM SIGPLAN Notices, 35(7):32–40, July 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ung:2000:MAD

[107] David Ung and Cristina Cifuentes. Machine-adaptable dynamic binary translation. ACM SIGPLAN Notices, 35(7):41–51, July 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Arnold:2000:CSS

[108] Matthew Arnold, Stephen Fink, Vivek Sarkar, and Peter F. Sweeney. A comparative study of static and profilebased heuristics for inlining. ACM SIG-PLAN Notices, 35(7):52–64, July 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mehofer:2000:PDF

[109] Eduard Mehofer and Bernhard Scholz. Probabilistic data flow system with two-edge profiling. ACM SIGPLAN Notices, 35(7):65–72, July 2000. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chambers:2000:D

[110] Craig Chambers. DyC. ACM SIG-PLAN Notices, 35(7):73–74, July 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chambers:2000:DPS

[111] Craig Chambers. DyC (panel session). ACM SIGPLAN Notices, 35(7):73–74, July 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bala:2000:DTDb

[112] Vasanth Bala. Dynamo: A transparent, dynamic, native binary optimizer. ACM SIGPLAN Notices, 35(7): 75, July 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cytron:2000:DPS

[113] Ron Cytron and Vasanth Bala. Dynamo (panel session): a transparent, dynamic, native binary optimizer. ACM SIGPLAN Notices, 35(7): 75, July 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cytron:2000:TPS

[114] Ron Cytron and Renaud Marlet. Tempo, a program specializer for C (panel session). ACM SIGPLAN Notices, 35(7):76–77, July 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marlet:2000:TPS

[115] Renaud Marlet. Tempo, a program specializer for C. ACM SIGPLAN Notices, 35(7):76–77, July 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cytron:2000:DCJ

[116] Ron Cytron and Vivek Sarkar. Dynamic compilation in Jalapeño (panel session). ACM SIGPLAN Notices, 35(7):78, July 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sarkar:2000:DCJ

[117] Vivek Sarkar. Dynamic compilation in Jalapeño. ACM SIGPLAN Notices, 35(7):78, July 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chung:2000:SDP

[118] Yoo C. Chung and Michael J. Voss. Summary of the Dynamo ’00 panel discussion. ACM SIGPLAN Notices, 35 (7):79–81, July 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cytron:2000:SDP

[119] Ron Cytron, Yoo C. Chung, and Michael J. Voss. Summary of the Dynamo ’00 panel discussion (panel session). ACM SIGPLAN Notices, 35 (7):79–81, July 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cytron:2000:ECN

[120] Ron Cytron. Executive Committee news: SIGPLAN awards. ACM SIGPLAN Notices, 35(8):1–2, August 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergmann:2000:CCCd

[121] Seth Bergmann. Conference corner: Calendar. ACM SIGPLAN Notices, 35 (8):3–14, August 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:2000:FFB

[122] Paul Frenger. Forth and the FreeBSD bootloader. ACM SIGPLAN Notices, 35(8):15–17, August 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:2000:FRF

[123] Paul Frenger. Forth report: Forth and the FreeBSD bootloader. ACM SIGPLAN Notices, 35(8):15–17, August 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fenwick:2000:IAb

[124] Jan Fenwick and Cindy Norris. Information for authors. ACM SIG-PLAN Notices, 35(8):18, August 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fenwick:2000:IPAa

[125] Jan Fenwick and Cindy Norris. Information for past authors. ACM SIG-PLAN Notices, 35(8):19, August 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fereira:2000:UPL

[126] Ulisses Fereira. uu for programming languages. ACM SIGPLAN Notices, 35 (8):20–30, August 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ferreira:2000:UPL

[127] Ulisses Ferreira. uu for programming languages. ACM SIGPLAN Notices, 35 (8):20–30, August 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Juric:2000:JDO

[128] Matjaz B. Juric, Ivan Rozman, and Simon Nash. Java 2 distributed object middleware performance analysis and optimization. ACM SIG-PLAN Notices, 35(8):31–40, August 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wong:2000:FGK

[129] Limsoon Wong. The functional guts of the Kleisli query system. ACM SIGPLAN Notices, 35(9):1– 10, September 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/fp/351240/
p1-wong/p1-wong.pdf;   http://www.
acm.org/pubs/citations/proceedings/
fp/351240/p1-wong/.

Hosoya:2000:RET

[130] Haruo Hosoya, Jérôme Vouillon, and Benjamin C. Pierce. Regular expression types for XML. ACM SIGPLAN Notices, 35(9):11– 22, September 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/fp/351240/
p11-hosoya/p11-hosoya.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
fp/351240/p11-hosoya/.

Queinnec:2000:IBE

[131] Christian Queinnec. The influence of browsers on evaluators or, continuations to program Web servers. ACM SIGPLAN Notices, 35(9):23– 33, September 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/fp/351240/
p23-queinnec/p23-queinnec.pdf;
http://www.acm.org/pubs/citations/
proceedings/fp/351240/p23-queinnec/

.

Skalka:2000:SES

[132] Christian Skalka and Scott Smith. Static enforcement of security with types. ACM SIGPLAN Notices, 35 (9):34–45, September 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/fp/351240/
p34-skalka/p34-skalka.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
fp/351240/p34-skalka/.

Pottier:2000:IFI

[133] François Pottier and Sylvain Conchon. Information flow inference for free. ACM SIGPLAN Notices, 35 (9):46–57, September 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/fp/351240/
p46-pottier/p46-pottier.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
fp/351240/p46-pottier/.

Weirich:2000:TSC

[134] Stephanie Weirich. Type-safe cast:

(functional pearl). ACM SIGPLAN Notices, 35(9):58–67, September 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/articles/proceedings/fp/
351240/p58-weirich/p58-weirich.
pdf;   http://www.acm.org/pubs/
citations/proceedings/fp/351240/
p58-weirich/.

Crary:2000:TCI

[135] Karl Crary. Typed compilation of inclusive subtyping. ACM SIGPLAN Notices, 35(9):68–81, September 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/articles/proceedings/fp/
351240/p68-crary/p68-crary.pdf;
 http://www.acm.org/pubs/citations/
proceedings/fp/351240/p68-crary/

.

Trifonov:2000:FRI

[136] Valery Trifonov, Bratin Saha, and Zhong Shao. Fully reflexive intensional type analysis. ACM SIGPLAN Notices, 35(9):82–93, September 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/fp/351240/
p82-trifonov/p82-trifonov.pdf;
http://www.acm.org/pubs/citations/
proceedings/fp/351240/p82-trifonov/

.

Chakravarty:2000:MTN

[137] Manuel M. T. Chakravarty and Gabriele Keller. More types for nested data parallel programming. ACM SIGPLAN Notices, 35(9):94– 105, September 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/fp/351240/
p94-chakravarty/p94-chakravarty.
pdf;   http://www.acm.org/pubs/
citations/proceedings/fp/351240/
p94-chakravarty/.

Sage:2000:FDG

[138] Meurig Sage. FranTk — a declarative GUI language for Haskell. ACM SIGPLAN Notices, 35(9):106– 117, September 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/fp/351240/
p106-sage/p106-sage.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
fp/351240/p106-sage/.

McNamara:2000:FPC

[139] Brian McNamara and Yannis Smaragdakis. Functional programming in C++. ACM SIGPLAN Notices, 35 (9):118–129, September 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/fp/351240/
p118-mcnamara/p118-mcnamara.pdf;
http://www.acm.org/pubs/citations/
proceedings/fp/351240/p118-mcnamara/

.

Harper:2000:AMS

[140] Robert Harper and Benjamin C. Pierce. Advanced module systems (invited talk): a guide for the perplexed. ACM SIGPLAN Notices, 35 (9):130, September 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/fp/351240/
p130-harper/p130-harper.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
fp/351240/p130-harper/.

Okasaki:2000:BFN

[141] Chris Okasaki. Breadth-first numbering: lessons from a small exercise in algorithm design. ACM SIGPLAN Notices, 35(9):131–136, September 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/fp/351240/
p131-okasaki/p131-okasaki.pdf;
http://www.acm.org/pubs/citations/
proceedings/fp/351240/p131-okasaki/

.

Sasano:2000:MIP

[142] Isao Sasano, Zhenjiang Hu, Masato Takeichi, and Mizuhito Ogawa. Make it practical: a generic linear-time algorithm for solving maximum-weightsum problems. ACM SIGPLAN Notices, 35 (9):137–149, September 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/fp/351240/
p137-sasano/p137-sasano.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
fp/351240/p137-sasano/.

Faxen:2000:CES

[143] Karl-Filip Faxén. Cheap eagerness: speculative evaluation in a lazy functional language. ACM SIGPLAN Notices, 35(9):150–161, September 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/fp/351240/
p150-faxen/p150-faxen.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
fp/351240/p150-faxen/.

Baker-Finch:2000:OSP

[144] Clem Baker-Finch, David J. King, and Phil Trinder. An operational semantics for parallel lazy evaluation. ACM SIGPLAN Notices, 35 (9):162–173, September 2000. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/citations/proceedings/
fp/351240/p162-baker-finch/.

Erkok:2000:RMB

[145] Levent Erkök and John Launchbury. Recursive monadic bindings. ACM SIGPLAN Notices, 35(9):174– 185, September 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/fp/351240/
p174-erkok/p174-erkok.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
fp/351240/p174-erkok/.

Hinze:2000:DBM

[146] Ralf Hinze. Deriving backtracking monad transformers. ACM SIGPLAN Notices, 35(9):186–197, September 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/fp/351240/
p186-hinze/p186-hinze.pdf;   http:

                                        //www.acm.org/pubs/citations/proceedings/
                                        fp/351240/p186-hinze/.

Davies:2000:ITC

[147] Rowan Davies and Frank Pfenning. Intersection types and computational effects. ACM SIGPLAN Notices, 35(9): 198–208, September 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/articles/proceedings/fp/351240/
                                        p198-davies/p198-davies.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        fp/351240/p198-davies/.

Damian:2000:SAP

[148] Daniel Damian and Olivier Danvy. Syntactic accidents in program analysis: on the impact of the CPS transformation. ACM SIGPLAN Notices, 35(9):209–220, September 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

                                        org/pubs/citations/proceedings/
                                        fp/351240/p209-damian/.

Gapeyev:2000:RSR

[149] Vladimir Gapeyev, Michael Y. Levin, and Benjamin C. Pierce. Recursive subtyping revealed: (functional pearl). ACM SIGPLAN Notices, 35 (9):221–231, September 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/articles/proceedings/fp/351240/
                                        p221-gapeyev/p221-gapeyev.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/fp/351240/p221-gapeyev/

.

Curien:2000:DC

[150] Pierre-Louis Curien and Hugo Herbelin. The duality of computation. ACM SIGPLAN Notices, 35(9): 233–243, September 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/fp/351240/
p233-curien/p233-curien.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
fp/351240/p233-curien/.

Seger:2000:CFP

[151] Carl Seger. Combining functional programming and hardware verification (invited talk). ACM SIGPLAN Notices, 35(9):244, September 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/articles/proceedings/fp/
351240/p244-seger/p244-seger.pdf;
http://www.acm.org/pubs/citations/
proceedings/fp/351240/p244-seger/

.

Serrano:2000:UMA

[152] Manuel Serrano and Hans-J. Boehm. Understanding memory allocation of scheme programs. ACM SIGPLAN Notices, 35(9):245–256, September 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/articles/proceedings/fp/
351240/p245-serrano/p245-serrano.
pdf;   http://www.acm.org/pubs/
citations/proceedings/fp/351240/
p245-serrano/.

Cheadle:2000:NSH

[153] A. M. Cheadle, A. J. Field, S. Marlow, S. L. Peyton Jones, and R. L. While. Non-stop Haskell. ACM SIGPLAN Notices, 35(9):257–267, September 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/fp/351240/
p257-cheadle/p257-cheadle.pdf;
http://www.acm.org/pubs/citations/
proceedings/fp/351240/p257-cheadle/

.

Claessen:2000:QLT

[154] Koen Claessen and John Hughes. QuickCheck: a lightweight tool for random testing of Haskell programs. ACM SIGPLAN Notices, 35(9):268– 279, September 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/fp/351240/
p268-claessen/p268-claessen.pdf;
 http://www.acm.org/pubs/citations/
proceedings/fp/351240/p268-claessen/

.

Jones:2000:CCA

[155] Simon Peyton Jones, Jean-Marc Eber, and Julian Seward. Composing contracts: an adventure in financial engineering (functional pearl). ACM SIGPLAN Notices, 35(9):280– 292, September 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/fp/351240/
p280-jones/p280-jones.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
fp/351240/p280-jones/.

Maessen:2000:IJM

[156] Jan-Willem Maessen and Xiaowei Shen. Improving the Java memory model using CRF. ACM SIGPLAN Notices, 35(10):1–12, October 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/citations/proceedings/
oops/353171/p1-maessen/.

Noble:2000:ECO

[157] James Noble, David Holmes, and John Potter. Exclusion for composite objects. ACM SIGPLAN Notices, 35 (10):13–28, October 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/citations/proceedings/oops/
353171/p13-noble/.

Buhr:2000:OOR

[158] Peter A. Buhr, Ashif S. Harji, Philipp E. Lim, and Jiongxiong Chen. Object-oriented real-time concurrency. ACM SIGPLAN Notices, 35 (10):29–46, October 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/citations/proceedings/oops/
353171/p29-buhr/.

Arnold:2000:AOJ

[159] Matthew Arnold, Stephen Fink, David Grove, Michael Hind, and Peter F. Sweeney. Adaptive optimization in the Jalapeño JVM. ACM SIGPLAN Notices, 35(10):47–65, October 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

                                        org/pubs/citations/proceedings/
                                        oops/353171/p47-arnold/.

Serrano:2000:QQS

[160] Mauricio Serrano, Rajesh Bordawekar, Sam Midkiff, and Manish Gupta. Quicksilver: a quasi-static compiler for Java. ACM SIGPLAN Notices, 35 (10):66–82, October 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/citations/proceedings/oops/
                                        353171/p66-serrano/.

Zaks:2000:SCJ

[161] Ayal Zaks, Vitaly Feldman, and Nava Aizikowitz. Sealed calls in Java packages. ACM SIGPLAN Notices, 35 (10):83–92, October 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/citations/proceedings/oops/
                                        353171/p83-zaks/.

Lee:2000:RWO

[162] Ji Y. Lee, Hye J. Kim, and Kyo C. Kang. A real world object modeling method for creating simulation environment of real-time systems. ACM SIGPLAN Notices, 35 (10):93–104, October 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/citations/proceedings/oops/
                                        353171/p93-lee/.

Bastide:2000:FSC

[163] Rémi Bastide, Philippe Palanque, Ousmane Sy, and David Navarre. Formal specification of CORBA services: experience and lessons learned. ACM SIGPLAN Notices, 35(10):105– 117, October 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/citations/proceedings/oops/
353171/p105-bastide/.

Kiernan:2000:MOQ

[164] Jerry Kiernan and Michael J. Carey. Middleware object query processing with deferred updates and autonomous sources. ACM SIGPLAN Notices, 35 (10):118–129, October 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/citations/proceedings/oops/
353171/p118-kiernan/.

Clifton:2000:MMO

[165] Curtis Clifton, Gary T. Leavens, Craig Chambers, and Todd Millstein. MultiJava: modular open classes and symmetric multiple dispatch for Java. ACM SIGPLAN Notices, 35 (10):130–145, October 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/citations/proceedings/oops/
353171/p130-clifton/.

Viroli:2000:PPJ

[166] Mirko Viroli and Antonio Natali. Parametric polymorphism in Java: an approach to translation based on reflective features. ACM SIGPLAN Notices, 35(10):146–165, October 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

                                        org/pubs/citations/proceedings/
                                        oops/353171/p146-viroli/.

Demeyer:2000:FRC

[167] Serge Demeyer, Stéphane Ducasse, and Oscar Nierstrasz. Finding refactorings via change metrics. ACM SIGPLAN Notices, 35(10):166–177, October 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

                                        org/pubs/citations/proceedings/
                                        oops/353171/p166-demeyer/.

Harrison:2000:MUD

[168] William Harrison, Charles Barton, and Mukund Raghavachari. Mapping UML designs to Java. ACM SIGPLAN Notices, 35(10):178–187, October 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

                                        org/pubs/citations/proceedings/
                                        oops/353171/p178-harrison/.

Herrmann:2000:PCS

[169] Stephan Herrmann and Mira Mezini. PIROL: a case study for multidimensional separation of concerns in software engineering environments. ACM SIGPLAN Notices, 35(10):188– 207, October 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/citations/proceedings/oops/
                                        353171/p188-herrmann/.

Ruby:2000:SCC

[170] Clyde Ruby and Gary T. Leavens. Safely creating correct subclasses without seeing superclass code.

ACM SIGPLAN Notices, 35(10):208– 228, October 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/citations/proceedings/oops/
353171/p208-ruby/.

Lujan:2000:OOO

[171] Mikel Luján, T. L. Freeman, and John R. Gurd. OoLALA: an object oriented analysis and design of numerical linear algebra. ACM SIGPLAN Notices, 35(10):229–252, October 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/citations/proceedings/
oops/353171/p229-lujan/.

Ortigosa:2000:TAO

[172] Alvaro Ortigosa, Marcelo Campo, and Roberto Moriyón. Towards agentoriented assistance for framework instantiation. ACM SIGPLAN Notices, 35(10):253–263, October 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/citations/proceedings/
oops/353171/p253-ortigosa/.

Sundaresan:2000:PVM

[173] Vijay Sundaresan, Laurie Hendren, Chrislain Razafimahefa, Raja Vallée-Rai, Patrick Lam, Etienne Gagnon, and Charles Godin. Practical virtual method call resolution for Java. ACM SIGPLAN Notices, 35(10):264– 280, October 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/

                                        pubs/citations/proceedings/oops/
                                        353171/p264-sundaresan/.

Tip:2000:SPB

[174] Frank Tip and Jens Palsberg. Scalable propagation-based call graph construction algorithms. ACM SIGPLAN Notices, 35(10):281–293, October 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

                                        org/pubs/citations/proceedings/
                                        oops/353171/p281-tip/.

Ishizaki:2000:SDT

[175] Kazuaki Ishizaki, Motohiro Kawahito, Toshiaki Yasue, Hideaki Komatsu, and Toshio Nakatani. A study of devirtualization techniques for a Java^{T M} Just-In-Time compiler. ACM SIGPLAN Notices, 35(10):294–310, October 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

                                        org/pubs/citations/proceedings/
                                        oops/353171/p294-ishizaki/.

Glew:2000:ECO

[176] Neal Glew. An efficient class and object encoding. ACM SIGPLAN Notices, 35 (10):311–324, October 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/citations/proceedings/oops/
                                        353171/p311-glew/.

Qian:2000:FSJ

[177] Zhenyu Qian, Allen Goldberg, and Alessandro Coglio. A formal specification of Java^{T M} class loading. ACM SIGPLAN Notices, 35(10):325– 336, October 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/citations/proceedings/oops/
353171/p325-qian/.

Rayside:2000:AUO

[178] Derek Rayside and Gerard T. Campbell. An Aristotelian understanding of object-oriented programming. ACM SIGPLAN Notices, 35(10):337– 353, October 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/citations/proceedings/oops/
353171/p337-rayside/.

Czajkowski:2000:AIJ

[179] Grzegorz Czajkowski. Application isolation in the Java^{T M} Virtual Machine. ACM SIGPLAN Notices, 35 (10):354–366, October 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/citations/proceedings/oops/
353171/p354-czajkowski/.

Bryce:2000:ASO

[180] Ciarán Bryce and Chrislain Razafimahefa. An approach to safe object sharing. ACM SIGPLAN Notices, 35 (10):367–381, October 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/citations/proceedings/oops/
353171/p367-bryce/.

Bacon:2000:GDJ

[181] David F. Bacon, Robert E. Strom, and Ashis Tarafdar. Guava: a dialect of Java without data races. ACM SIGPLAN Notices, 35(10):382– 400, October 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/citations/proceedings/oops/
                                        353171/p382-bacon/.

Nagle:2000:DCS

[182] David Nagle, Steven Schlosser, John Linwood Griffin, and Gregory Ganger. Designing computer systems with MEMSbased storage. ACM SIGPLAN Notices, 35(11):1–12, November 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schlosser:2000:DCS

[183] Steven W. Schlosser, John Linwood Griffin, David F. Nagle, and Gregory R. Ganger. Designing computer systems with MEMS-based storage. ACM SIG-PLAN Notices, 35(11):1–12, November 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gharachorloo:2000:ADA

[184] Kourosh Gharachorloo, Madhu Sharma, Simon Steely, and Stephen Van Doren. Architecture and design of AlphaServer GS320. ACM SIGPLAN Notices, 35 (11):13–24, November 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Martin:2000:TSA

[185] Milo M. K. Martin, Daniel J. Sorin, Anastassia Ailamaki, Alaa R. Alameldeen, Ross M. Dickson, Carl J.

Mauer, Kevin E. Moore, Manoj Plakal, Mark D. Hill, and David A. Wood. Timestamp snooping: An approach for extending SMPs. ACM SIGPLAN Notices, 35(11):25–36, November 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nanda:2000:MPR

[186] Ashwini Nanda, Kwok-Ken Mak, Krishnan Sugavanam, Ramendra K. Sahoo, Vijayaraghavan Soundararajan, and T. Basil Smith. MemorIES: A programmable, real-time hardware emulation tool for multiprocessor server design. ACM SIGPLAN Notices, 35(11): 37–48, November 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gibson:2000:FVS

[187] Jeff Gibson, Robert Kunz, David Ofelt, Mark Horowitz, John Hennessy, and Mark Heinrich. FLASH vs. (simulated) FLASH: closing the simulation loop. ACM SIGPLAN Notices, 35(11): 49–58, November 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Heinrich:2000:SAA

[188] Mark Heinrich, Jeff Givson, Robert Kunz, David Ofelt, Mark Horowitz, and John Hennessy. FLASH vs. (simulated) FLASH: Closing the simulation loop. ACM SIGPLAN Notices, 35(11): 49–58, November 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chou:2000:UML

[189] Andy Chou, Benjamin Chelf, Dawson Engler, and Mark Heinrich. Using meta-level compilation to check FLASH protocol code. ACM SIG-PLAN Notices, 35(11):59–70, November 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bhoedjang:2000:EDA

[190] Raoul A. F. Bhoedjang, Kees Verstoep, Tim Rühl, Henri E. Bal, and Rutger F. H. Hofman. Evaluating design alternatives for reliable communication on high-speed networks. ACM SIGPLAN Notices, 35(11):71–81, November 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mattson:2000:CS

[191] Peter Mattson, William J. Dally, Scott W. Rixner, Ujval J. Kapasi, and John D. Owens. Communication scheduling. ACM SIGPLAN Notices, 35(11):82–92, November 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hill:2000:SAD

[192] Jason Hill, Robert Szewczyk, Alec Woo, Seth Hollar, David Culler, and Kristofer Pister. System architecture directions for networked sensors. ACM SIGPLAN Notices, 35(11):93– 104, November 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lebeck:2000:PAP

[193] Alvin R. Lebeck, Xiaobo Fan, Heng Zeng, and Carla Ellis. Power aware page allocation. ACM SIGPLAN Notices, 35(11):105–116, November 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berger:2000:HSM

[194] Emery D. Berger, Kathryn S. McKinley, Robert D. Blumofe, and Paul R. Wilson. Hoard: A scalable memory allocator for multithreaded applications. ACM SIGPLAN Notices, 35(11):117– 128, November 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Flautner:2000:TLP

[195] Krisztián Flautner, Rich Uhlig, Steve Reinhardt, and Trevor Mudge. Thread level parallelism and interactive performance of desktop applications. ACM SIGPLAN Notices, 35(11):129–138, November 2000. CODEN SINODQ. ISBN 1-58113-317-0. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://

delivery.acm.org/10.1145/360000/
357001/p129-flautner.pdf.

Kawahito:2000:ENP

[196] Motohiro Kawahito, Hideaki Komatsu, and Toshio Nakatani. Effective null pointer check elimination utilizing hardware trap. ACM SIGPLAN Notices, 35(11):139–149, November 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:2000:FVL

[197] Youtao Zhang, Jun Yang, and Rajiv Gupta. Frequent value locality and value-centric data cache design. ACM SIGPLAN Notices, 35(11):150– 159, November 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Burrows:2000:EFV

[198] Michael Burrows, Ulfar Erlingson, Shun-Tak Leung, Mark Vandevoorde, Carl Waldspurger, Kip Walker, and Bill Weihl. Efficient and flexible value sampling. ACM SIGPLAN Notices, 35(11):160–167, November 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lie:2000:ASC

[199] David Lie, Chandramohan Thekkath, Mark Mitchell, Patrick Lincoln, Dan Boneh, John Mitchell, and Mark Horowitz. Architectural support for copy and tamper resistant software. ACM SIGPLAN Notices, 35(11):168– 177, November 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Austin:2000:ASF

[200] Todd Austin, Jerome Burke, and John McDonald. Architectural support for fast symmetric-key cryptography. ACM SIGPLAN Notices, 35(11):178– 189, November 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Burke:2000:ASF

[201] Jerome Burke, John McDonald, and Todd Austin. Architectural support for fast symmetric-key cryptography. ACM SIGPLAN Notices, 35(11):178– 189, November 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kubiatowicz:2000:OAG

[202] John Kubiatowicz, David Bindel, Yan Chen, Steven Czerwinski, Patrick Eaton, Dennis Geels, Ramakrishan Gummadi, Sean Rhea, Hakim Weatherspoon, Westley Weimer, Chris Wells, and Ben Zhao. OceanStore: An architecture for global-scale persistent storage. ACM SIGPLAN Notices, 35(11): 190–201, November 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Duesterwald:2000:SPH

[203] Evelyn Duesterwald and Vasanth Bala. Software profiling for hot path prediction: Less is more. ACM SIGPLAN Notices, 35(11):202–211, November 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zahir:2000:CCD

[204] Rumi Zahir, Jonathan Ross, Dale Morris, and Drew Hess. OS and compiler considerations in the design of the IA-64 architecture. ACM SIGPLAN Notices, 35(11):212–221, November 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/devresource.hp.com/devresource/
Docs/TechPapers/IA64/IA64oscompilercon.
pdf.

Connors:2000:HSD

[205] Daniel A. Connors, Hillery C. Hunter, Ben-Chung Cheng, and Wen-Mei W. Hwu. Hardware support for dynamic activation of compiler-directed computation reuse. ACM SIGPLAN Notices, 35(11):222–233, November 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Snavely:2000:SJS

[206] Allan Snavely and Dean M. Tullsen. Symbiotic jobscheduling for a simultaneous multithreading processor. ACM SIGPLAN Notices, 35(11):234–244, November 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Redstone:2000:AOS

[207] Joshua A. Redstone, Susan J. Eggers, and Henry M. Levy. An analysis of operating system behavior on a simultaneous multithreaded architecture. ACM SIGPLAN Notices, 35(11): 245–256, November 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sundaramoorthy:2000:SPI

[208] Karthik Sundaramoorthy, Zachary Purser, and Eric Rotenberg. Slipstream processors: Improving both performance and fault tolerance. ACM SIGPLAN Notices, 35(11):257–268, November 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2000:AI

[209] Anonymous. Author index. ACM SIGPLAN Notices, 35(11):269–271, November 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergmann:2000:CCCe

[210] Seth Bergmann. Conference corner: Calendar. ACM SIGPLAN Notices, 35 (12):1–12, December 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2000:CCL

[211] Anonymous. Conference corner: LCTES ’01 call for papers. ACM SIG-PLAN Notices, 35(12):13, December 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Benson:2000:JR

[212] Brent W. Benson, Jr. Java reflections. ACM SIGPLAN Notices, 35(12): 14–18, December 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Benson:2000:JRJ

[213] Brent W. Benson, Jr. Java reflections: Java server pages: Servlets, inside out. ACM SIGPLAN Notices, 35(12):14– 18, December 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jakeman:2000:FR

[214] Chris Jakeman. Forth report. ACM SIGPLAN Notices, 35(12):19–21, December 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jakeman:2000:FU

[215] Chris Jakeman. Forth in the UK. ACM SIGPLAN Notices, 35(12):19–21, December 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fenwick:2000:IAc

[216] Jay Fenwick and Cindy Norris. Information for authors. ACM SIG-PLAN Notices, 35(12):22, December 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fenwick:2000:IPAb

[217] Jay Fenwick and Cindy Norris. Information for past authors. ACM SIG-PLAN Notices, 35(12):23, December 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Corchuelo:2000:MCA

[218] Rafael Corchuelo, José A. Pérez, and Miguel Toro. A multiparty coordination aspect language. ACM SIGPLAN Notices, 35(12):24–32, December 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kuhl:2000:OOL

[219] Bernd Kühl and Axel-Tobias Schreiner. An object-oriented LL(1) parser generator. ACM SIGPLAN Notices, 35(12): 33–40, December 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gutierrez:2000:IQC

[220] David Basanta Gutiérrez, Cándida Luengo Díez, Raul Izquierdo Castanedo, Jose Emilio Labra Gayo, and Juan Manuel Cueva Lovelle. Improving the quality of compiler construction with object-oriented techniques. ACM SIG-PLAN Notices, 35(12):41–50, December 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Knutson:2000:MDS

[221] Charles D. Knutson, Timothy A. Budd, and Hugh Vidos. Multiparadigm design of a simple relational database. ACM SIGPLAN Notices, 35(12):51–61, December 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hirzel:2001:TAG

[222] Martin Hirzel and Amer Diwan. On the type accuracy of garbage collection. ACM SIGPLAN Notices, 36(1): 1–11, January 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shaham:2001:EGJ

[223] Ran Shaham, Elliot K. Kolodner, and Mooly Sagiv. On the effectiveness of GC in Java. ACM SIGPLAN Notices, 36(1):12–17, January 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Steensgaard:2001:TSH

[224] Bjarne Steensgaard. Thread-specific heaps for multi-threaded programs. ACM SIGPLAN Notices, 36(1):18–24, January 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Makholm:2001:RBM

[225] Henning Makholm. A region-based memory manager for Prolog. ACM SIGPLAN Notices, 36(1):25–34, January 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chilimbi:2001:DTF

[226] Trishul Chilimbi, Richard Jones, and Benjamin Zorn. Designing a trace format for heap allocation events. ACM SIGPLAN Notices, 36(1):35–49, January 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tarditi:2001:CGC

[227] David Tarditi. Compact garbage collection tables. ACM SIGPLAN Notices, 36(1):50–58, January 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boehm:2001:RGC

[228] Hans-J. Boehm. Reducing garbage collector cache misses. ACM SIG-PLAN Notices, 36(1):59–64, January 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chung:2001:MAL

[229] Yoo C. Chung and Soo-Mook Moon. Memory allocation with lazy fits. ACM SIGPLAN Notices, 36(1):65–70, January 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rodriguez-Rivera:2001:CGC

[230] Gustavo Rodriguez-Rivera, Mike Spertus, and Charles Fiterman. Conservative garbage collection for general memory allocators. ACM SIGPLAN Notices, 36(1):71–79, January 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Heil:2001:CGC

[231] Timothy H. Heil and James E. Smith. Concurrent garbage collection using hardware-assisted profiling. ACM SIG-PLAN Notices, 36(1):80–93, January 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Plakal:2001:CGC

[232] Manoj Plakal and Charles N. Fischer. Concurrent garbage collection using program slices on multithreaded processors. ACM SIGPLAN Notices, 36 (1):94–100, January 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hudson:2001:CRG

[233] Richard L. Hudson, J. Eliot B. Moss, Sreenivas Subramoney, and Weldon Washburn. Cycles to recycle: Garbage collection on the IA-64. ACM SIG-PLAN Notices, 36(1):101–110, January 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fitzgerald:2001:CPD

[234] Robert Fitzgerald and David Tarditi. The case for profile-directed selection of garbage collectors. ACM SIGPLAN Notices, 36(1):111–120, January 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Agesen:2001:EOS

[235] Ole Agesen and Alex Garthwaite. Efficient object sampling via weak references. ACM SIGPLAN Notices, 36(1): 121–126, January 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Harris:2001:DAP

[236] Timothy L. Harris. Dynamic adaptive pre-tenuring. ACM SIGPLAN Notices, 36(1):127–136, January 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stefanovic:2001:MOL

[237] Darko Stefanović, Kathryn S. McKinley, and J. Eliot B. Moss. On models for object lifetime distributions. ACM SIGPLAN Notices, 36(1):137–142, January 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Printezis:2001:GMC

[238] Tony Printezis and David Detlefs. A generational mostly-concurrent garbage collector. ACM SIGPLAN Notices, 36(1):143–154, January 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Domani:2001:IFG

[239] Tamar Domani, Elliot K. Kolodner, Ethan Lewis, Eliot E. Salant, Katherine Barabash, Itai Lahan, Yossi Levanoni, Erez Petrank, and Igor Yanorer. Implementing an on-the-fly garbage collector for Java. ACM SIGPLAN Notices, 36(1):155–166, January 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dickman:2001:DTR

[240] Peter Dickman. Diffusion tree restructuring for indirect reference counting. ACM SIGPLAN Notices, 36(1): 167–177, January 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:2001:EF

[241] Paul Frenger. Forth: Extreme Forth. ACM SIGPLAN Notices, 36(2):20– 23, February 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kumar:2001:NNC

[242] Rajeev Kumar. A neural net compiler system for hierarchical organization. ACM SIGPLAN Notices, 36(2): 26–36, February 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sunik:2001:DOL

[243] Boris Sunik. Definition of open language. ACM SIGPLAN Notices, 36(2): 37–46, February 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:2001:SGI

[244] Yukong Zhang, Young-Jun Kwon, and Hyuk Jae Lee. A systematic generation of initial register-reuse chains for dependence minimization. ACM SIG-PLAN Notices, 36(2):47–54, February 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mitchell:2001:PLM

[245] John C. Mitchell. Programming language methods in computer security. ACM SIGPLAN Notices, 36(3):1–3, March 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/plan/
360204/p1-mitchell/p1-mitchell.
pdf;   http://www.acm.org/pubs/
citations/proceedings/plan/360204/
p1-mitchell/.

Sangiorgi:2001:EIA

[246] Davide Sangiorgi. Extensionality and intensionality of the ambient logics. ACM SIGPLAN Notices, 36(3):4–13, March 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ishtiaq:2001:BAL

[247] Samin S. Ishtiaq and Peter W. O’Hearn. BI as an assertion language for mutable data structures. ACM SIGPLAN Notices, 36(3):14–26, March 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yahav:2001:VSP

[248] Eran Yahav. Verifying safety properties of concurrent Java programs using 3valued logic. ACM SIGPLAN Notices, 36(3):27–40, March 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/plan/
360204/p27-yahav/p27-yahav.pdf;
 http://www.acm.org/pubs/citations/
proceedings/plan/360204/p27-yahav/

.

Odersky:2001:CLT

[249] Martin Odersky, Christoph Zenger, and Matthias Zenger. Colored local type inference. ACM SIGPLAN Notices, 36(3):41–53, March 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/articles/proceedings/plan/
360204/p41-odersky/p41-odersky.
pdf;   http://www.acm.org/pubs/
citations/proceedings/plan/360204/
p41-odersky/.

Rehof:2001:TBF

[250] Jakob Rehof and Manuel Fähndrich. Type-base flow analysis: from polymorphic subtyping to CFL-reachability. ACM SIGPLAN Notices, 36(3):54– 66, March 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/plan/
360204/p54-rehof/p54-rehof.pdf;
 http://www.acm.org/pubs/citations/
proceedings/plan/360204/p54-rehof/

.

Hosoya:2001:REP

[251] Haruo Hosoya and Benjamin Pierce. Regular expression pattern matching for XML. ACM SIGPLAN Notices, 36(3):67–80, March 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/articles/proceedings/plan/
                                        360204/p67-hosoya/p67-hosoya.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/plan/360204/p67-hosoya/

.

Lee:2001:SCP

[252] Chin Soon Lee, Neil D. Jones, and Amir M. Ben-Amram. The sizechange principle for program termination. ACM SIGPLAN Notices, 36(3):81–92, March 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/articles/proceedings/plan/
                                        360204/p81-lee/p81-lee.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        plan/360204/p81-lee/.

Monniaux:2001:AMC

[253] David Monniaux. An abstract Monte-Carlo method for the analysis of probabilistic programs. ACM SIGPLAN Notices, 36(3):93–101, March 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

                                        org/pubs/articles/proceedings/plan/
                                        360204/p93-monniaux/p93-monniaux.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/plan/360204/
                                        p93-monniaux/.

Ganzinger:2001:EDM

[254] Harald Ganzinger. Efficient deductive methods for program analysis. ACM SIGPLAN Notices, 36 (3):102–103, March 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/plan/
360204/p102-ganzinger/p102-ganzinger.
pdf;   http://www.acm.org/pubs/
citations/proceedings/plan/360204/
p102-ganzinger/.

Abadi:2001:MVN

[255] Martín Abadi and Cédric Fournet. Mobile values, new names, and secure communication. ACM SIGPLAN Notices, 36(3):104–115, March 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/articles/proceedings/plan/
360204/p104-abadi/p104-abadi.pdf;
http://www.acm.org/pubs/citations/
proceedings/plan/360204/p104-abadi/

.

Unyapoth:2001:NPC

[256] Asis Unyapoth and Peter Sewell. Nomadic pict: correct communication infrastructure for mobile computation. ACM SIGPLAN Notices, 36 (3):116–127, March 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/plan/
360204/p116-unyapoth/p116-unyapoth.
pdf;   http://www.acm.org/pubs/
citations/proceedings/plan/360204/
p116-unyapoth/.

Igarashi:2001:GTS

[257] Atsushi Igarashi and Naoki Kobayashi. A generic type system for the Picalculus. ACM SIGPLAN Notices, 36 (3):128–141, March 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/plan/
360204/p128-igarashi/p128-igarashi.
pdf;   http://www.acm.org/pubs/
citations/proceedings/plan/360204/
p128-igarashi/.

Necula:2001:OBC

[258] George C. Necula and S. P. Rahul. Oracle-based checking of untrusted software. ACM SIGPLAN Notices, 36(3):142–154, March 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/plan/
360204/p142-necula/p142-necula.
pdf;   http://www.acm.org/pubs/
citations/proceedings/plan/360204/
p142-necula/.

Calcagno:2001:SOS

[259] Cristiano Calcagno. Stratified operational semantics for safety and correctness of the region calculus. ACM SIGPLAN Notices, 36(3):155– 165, March 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/plan/
360204/p155-calcagno/p155-calcagno.
pdf;   http://www.acm.org/pubs/
citations/proceedings/plan/360204/
p155-calcagno/.

Wang:2001:TPG

[260] Daniel C. Wang and Andrew W. Appel. Type-preserving garbage collectors. ACM SIGPLAN Notices, 36 (3):166–178, March 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/plan/
360204/p166-wang/p166-wang.pdf;
 http://www.acm.org/pubs/citations/
proceedings/plan/360204/p166-wang/

.

Kandemir:2001:CTI

[261] Mahmut Taylan Kandemir. A compiler technique for improving whole-program locality. ACM SIGPLAN Notices, 36 (3):179–192, March 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/plan/
360204/p179-kandemir/p179-kandemir.
pdf;   http://www.acm.org/pubs/
citations/proceedings/plan/360204/
p179-kandemir/.

Flanagan:2001:AEE

[262] Cormac Flanagan and James B. Saxe. Avoiding exponential explosion: generating compact verification conditions. ACM SIGPLAN Notices, 36 (3):193–205, March 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/plan/
360204/p193-flanagan/p193-flanagan.
pdf;   http://www.acm.org/pubs/
citations/proceedings/plan/360204/
p193-flanagan/.

Bhargavan:2001:WPM

[263] Karthikeyan Bhargavan, Satish Chandra, Peter J. McCann, and Carl A. Gunter. What packets may come: automata for network monitoring. ACM SIGPLAN Notices, 36(3):206– 219, March 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/articles/proceedings/plan/
                                        360204/p206-bhargavan/p206-bhargavan.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/plan/360204/
                                        p206-bhargavan/.

Milner:2001:CF

[264] Robin Milner. Computational flux. ACM SIGPLAN Notices, 36(3):220– 221, March 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/articles/proceedings/plan/
                                        360204/p220-milner/p220-milner.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/plan/360204/
                                        p220-milner/.

Bugliesi:2001:SSA

[265] Michele Bugliesi and Giuseppe Castagna. Secure safe ambients. ACM SIGPLAN Notices, 36(3):222–235, March 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

                                        org/pubs/articles/proceedings/plan/
                                        360204/p222-bugliesi/p222-bugliesi.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/plan/360204/
                                        p222-bugliesi/.

Sewell:2001:MAT

[266] Peter Sewell. Modules, abstract types, and distributed versioning. ACM SIGPLAN Notices, 36(3):236– 247, March 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/plan/
360204/p236-sewell/p236-sewell.
pdf;   http://www.acm.org/pubs/
citations/proceedings/plan/360204/
p236-sewell/.

Gordon:2001:TML

[267] Andrew D. Gordon and Don Syme. Typing a multi-language intermediate code. ACM SIGPLAN Notices, 36 (3):248–260, March 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/plan/
360204/p248-gordon/p248-gordon.
pdf;   http://www.acm.org/pubs/
citations/proceedings/plan/360204/
p248-gordon/.

Shields:2001:TIR

[268] Mark Shields and Erik Meijer. Typeindexed rows. ACM SIGPLAN Notices, 36(3):261–275, March 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/plan/
360204/p261-shields/p261-shields.
pdf;   http://www.acm.org/pubs/
citations/proceedings/plan/360204/
p261-shields/.

Gil:2001:SAT

[269] Joseph (Yossi) Gil. Subtyping arithmetical types. ACM SIGPLAN Notices, 36(3):276–289, March 2001. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

                                        org/pubs/articles/proceedings/plan/
                                        360204/p276-gil/p276-gil.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/plan/360204/p276-gil/

.

Vouillon:2001:CSB

[270] Jérôme Vouillon. Combining subsumption and binary methods: an object calculus with views. ACM SIGPLAN Notices, 36(3):290–303, March 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

                                        org/pubs/articles/proceedings/plan/
                                        360204/p290-vouillon/p290-vouillon.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/plan/360204/
                                        p290-vouillon/.

Ball:2001:MPS

[271] Thomas Ball and Hans Boehm. A modest proposal for SIGPLAN in the 21st Century. ACM SIGPLAN Notices, 36(4):1–2, April 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ball:2001:S

[272] Thomas Ball and Hans Boehm. SIG-PLANet. ACM SIGPLAN Notices, 36 (4):1–2, April 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

                                        URL http://www.acm.org/sigplan/
                                        sigplanet/.

Cytron:2001:CCD

[273] Ron Cytron and Matthias Felleisen. Computing curricula 2001 draft: SIG-PLAN response to the Ironman Draft of the computing curricula 2001. ACM SIGPLAN Notices, 36(4):3–4, April 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fenwick:2001:CCD

[274] Jay Fenwick, Cindy Norris, Ron Cytron, and Matthias Felleisen. Computing Curricula 2001 draft. ACM SIGPLAN Notices, 36(4):3–4, April 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergmann:2001:CCC

[275] Seth Bergmann. Conference corner: Calendar. ACM SIGPLAN Notices, 36 (4):5–20, April 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:2001:CEF

[276] Paul Frenger. Close encounters of the Forth kind. ACM SIGPLAN Notices, 36(4):21–24, April 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McNamara:2001:FPC

[277] Brian McNamara and Yannis Smaragdakis. Functional programming in C++ using the FC++ library. ACM SIGPLAN Notices, 36(4):25–30, April 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McNamara:2001:FPI

[278] Brian McNamara and Yannis Smaragdakis. Functional programming in C++ using the FC++ library. ACM SIGPLAN Notices, 36(4):25–30, April 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fenwick:2001:IAa

[279] Jay Fenwick and Cindy Norris. Information for authors. ACM SIG-PLAN Notices, 36(4):31, April 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fenwick:2001:IPAa

[280] Jay Fenwick and Cindy Norris. Information for past authors. ACM SIG-PLAN Notices, 36(4):32, April 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:2001:SOO

[281] Zhenqiang Chen and Baowen Xu. Slicing object-oriented Java programs. ACM SIGPLAN Notices, 36(4):33–40, April 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:2001:SCJ

[282] Zhenqiang Chen and Baowen Xu. Slicing concurrent Java programs. ACM SIGPLAN Notices, 36(4):41–47, April 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shaofeng:2001:SIT

[283] Wang Shaofeng, Yang Fuquing, and Sun Jiaguang. Some issues on the translator of Smalltalk to C++. ACM SIGPLAN Notices, 36(4):48–55, April 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wang:2001:SIT

[284] Shaofeng Wang, Fuqing Yang, and Jiaguang Sun. Some issues on the translator of Smalltalk to C++. ACM SIGPLAN Notices, 36(4):48–55, April 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kennedy:2001:DIG

[285] Andrew Kennedy and Don Syme. Design and implementation of generics for the .NET common language runtime. ACM SIGPLAN Notices, 36(5):1–12, May 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hicks:2001:DSU

[286] Michael Hicks, Jonathan T. Moore, and Scott Nettles. Dynamic software updating. ACM SIGPLAN Notices, 36 (5):13–23, May 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Heintze:2001:DDP

[287] Nevin Heintze and Olivier Tardieu. Demand-driven pointer analysis. ACM SIGPLAN Notices, 36(5):24–34, May 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vivien:2001:IPE

[288] Frédéric Vivien and Martin Rinard. Incrementalized pointer and escape analysis. ACM SIGPLAN Notices, 36 (5):35–46, May 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vivient:2001:IPE

[289] Frederic Vivient and Martin Rinard. Incrementalized pointer and escape analysis. ACM SIGPLAN Notices, 36 (5):35–46, May 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ghiya:2001:IPA

[290] Rakesh Ghiya, Daniel Lavery, and David Sehr. On the importance of points-to analysis and other memory disambiguation methods for C programs. ACM SIGPLAN Notices, 36 (5):47–58, May 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

DeLine:2001:EHL

[291] Robert DeLine and Manuel Fähndrich. Enforcing high-level protocols in lowlevel software. ACM SIGPLAN Notices, 36(5):59–69, May 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gay:2001:LSR

[292] David Gay and Alex Aiken. Language support for regions. ACM SIG-PLAN Notices, 36(5):70–80, May 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Monnier:2001:PS

[293] Stefan Monnier, Bratin Saha, and Zhong Shao. Principled scavenging. ACM SIGPLAN Notices, 36(5):81–91, May 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bacon:2001:JCB

[294] David F. Bacon, Clement R. Attanasio, Han B. Lee, V. T. Rajan, and Stephen Smith. Java without the coffee breaks: a nonintrusive multiprocessor garbage collector. ACM SIGPLAN Notices, 36 (5):92–103, May 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shaham:2001:HPS

[295] Ran Shaham, Elliot K. Kolodner, and Mooly Sagiv. Heap profiling for spaceefficient Java. ACM SIGPLAN Notices, 36(5):104–113, May 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berger:2001:CHP

[296] Emery D. Berger, Benjamin G. Zorn, and Kathryn S. McKinley. Composing high-performance memory allocators. ACM SIGPLAN Notices, 36(5): 114–124, May 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cheng:2001:PRT

[297] Perry Cheng and Guy E. Blelloch. A parallel, real-time garbage collector. ACM SIGPLAN Notices, 36(5): 125–136, May 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Amme:2001:STS

[298] Wolfram Amme, Niall Dalton, Jeffery von Ronne, and Michael Franz. SafeTSA: A type safe and referentially secure mobile-code representation based on static single assignment form. ACM SIGPLAN Notices, 36(5): 137–147, May 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Evans:2001:BCP

[299] William S. Evans and Christopher W. Fraser. Bytecode compression via profiled grammar rewriting. ACM SIG-PLAN Notices, 36(5):148–155, May 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Krintz:2001:UAR

[300] Chandra Krintz and Brad Calder. Using annotations to reduce dynamic optimization time. ACM SIGPLAN Notices, 36(5):156–167, May 2001. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Arnold:2001:FRC

[301] Matthew Arnold and Barbara G. Ryder. A framework for reducing the cost of instrumented code. ACM SIG-PLAN Notices, 36(5):168–179, May 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:2001:TWP

[302] Youtao Zhang and Rajiv Gupta. Timestamped whole program path representation and its applications.

ACM SIGPLAN Notices, 36(5):180– 190, May 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chilimbi:2001:ERA

[303] Trishul M. Chilimbi. Efficient representations and abstractions for quantifying and exploiting data reference locality. ACM SIGPLAN Notices, 36 (5):191–202, May 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ball:2001:APA

[304] Thomas Ball, Rupak Majumdar, Todd Millstein, and Sriram K. Rajamani. Automatic predicate abstraction of C programs. ACM SIGPLAN Notices, 36 (5):203–213, May 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Aggarwal:2001:RFA

[305] Aneesh Aggarwal and Keith H. Randall. Related field analysis. ACM SIG-PLAN Notices, 36(5):214–220, May 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Moller:2001:PAL

[306] Anders Møller and Michael I. Schwartzbach. The pointer assertion logic engine. ACM SIGPLAN Notices, 36(5):221– 231, May 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Thies:2001:UFS

[307] William Thies, Frédéric Vivien, Jeffrey Sheldon, and Saman Amarasinghe. A unified framework for schedule and storage optimization. ACM SIG-PLAN Notices, 36(5):232–242, May 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Appel:2001:OSC

[308] Andrew W. Appel and Lal George. Optimal spilling for CISC machines with few registers. ACM SIGPLAN Notices, 36(5):243–253, May 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Heintze:2001:UFA

[309] Nevin Heintze and Olivier Tardieu. Ultra-fast aliasing analysis using CLA: A million lines of C code in a second. ACM SIGPLAN Notices, 36(5): 254–263, May 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hanson:2001:DV

[310] David R. Hanson and Todd A. Proebsting. Dynamic variables. ACM SIG-PLAN Notices, 36(5):264–273, May 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marlow:2001:AEH

[311] Simon Marlow, Simon Peyton Jones, Andrew Moran, and John Reppy. Asynchronous exceptions in Haskell. ACM SIGPLAN Notices, 36(5):274– 285, May 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chatterjee:2001:EAC

[312] Siddhartha Chatterjee, Erin Parker, Philip J. Hanlon, and Alvin R. Lebeck. Exact analysis of the cache behavior of nested loops. ACM SIG-PLAN Notices, 36(5):286–297, May 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xiong:2001:SLC

[313] Jianxin Xiong, Jeremy Johnson, Robert Johnson, and David Padua. SPL: A language and compiler for DSP algorithms. ACM SIGPLAN Notices, 36(5):298–308, May 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kumar:2001:ELP

[314] Sanjeev Kumar, Yitzhak Mandelbaum, Xiang Yu, and Kai Li. ESP: A language for programmable devices. ACM SIG-PLAN Notices, 36(5):309–320, May 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schnarr:2001:FLC

[315] Eric C. Schnarr, Mark D. Hill, and James R. Larus. Facile: A language and compiler for high-performance processor simulators. ACM SIGPLAN Notices, 36(5):321–331, May 2001. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2001:AIa

[316] Anonymous. Author index. ACM SIG-PLAN Notices, 36(5):333, May 2001.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Knobe:2001:ECN

[317] Kathleen Knobe, Wilson Hsieh, and Uli Kremer. Executive committee news: SIGPLAN’s Professional Activities Committee. ACM SIGPLAN Notices, 36(6):1–2, June 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2001:CCCa

[318] Anonymous. Conference corner: Calendar. ACM SIGPLAN Notices, 36 (6):3–18, June 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Thiemann:2001:CPP

[319] Pieter Thiemann. Call for papers — PEPM’02: 2002 ACM Sigplan Workshop on Partial Evaluation and Semantics-Based Program Manipulation, Portland, Oregon, USA, January 14–15, 2002, Preceding POPL’02. ACM SIGPLAN Notices, 36(6):19, June 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergin:2001:CPU

[320] Joe Bergin. Curricular patterns: The undergraduate language course: What to do? ACM SIGPLAN Notices, 36 (6):20–22, June 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergin:2001:ULC

[321] Joe Bergin. The undergraduate language course: what to do? ACM SIGPLAN Notices, 36(6):20–22, June 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:2001:FD

[322] Paul Frenger. Is Forth dead? ACM SIGPLAN Notices, 36(6):23–25, June 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:2001:FRF

[323] Paul Frenger. Forth report: Is Forth dead? ACM SIGPLAN Notices, 36 (6):23–25, June 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fenwick:2001:IAb

[324] Jay Fenwick and Cindy Norris. Information for authors. ACM SIG-PLAN Notices, 36(6):26, June 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fenwick:2001:IPAb

[325] Jay Fenwick and Cindy Norris. Information for past authors. ACM SIG-PLAN Notices, 36(6):27, June 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Li:2001:DSN

[326] Tao Li and XiaoJie Liu. The denotational semantics of NIPL. ACM SIG-PLAN Notices, 36(6):28–37, June 2001.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gofen:2001:PDO

[327] Alexander Gofen. From Pascal to Delphi to Object Pascal-2000. ACM SIG-PLAN Notices, 36(6):38–49, June 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yu:2001:CTI

[328] Sheng Yu. Class-is-type is inadequate for object reuse. ACM SIG-PLAN Notices, 36(6):50–59, June 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Foster:2001:PCO

[329] Ian Foster. Parallel computing in 2010: opportunities and challenges in a networked world. ACM SIGPLAN Notices, 36(7):1, July 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/ppopp/
379539/p1-foster/p1-foster.pdf;
 http://www.acm.org/pubs/citations/
proceedings/ppopp/379539/p1-foster/

.

Kim:2001:RIA

[330] Seon Wook Kim, Chong liang Ooi, Rudolf Eigenmann, Babak Falsafi, and T. N. Vijaykumar. Reference idempotency analysis: a framework for optimizing speculative execution. ACM SIGPLAN Notices, 36(7):2–11, July 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/ppopp/
379539/p2-kim/p2-kim.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
ppopp/379539/p2-kim/.

Salcianu:2001:PEA

[331] Alexandru Salcianu and Martin Rinard. Pointer and escape analysis for multithreaded programs. ACM SIGPLAN Notices, 36(7):12– 23, July 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/ppopp/
379539/p12-salcianu/p12-salcianu.
pdf;   http://www.acm.org/pubs/
citations/proceedings/ppopp/379539/
p12-salcianu/.

Wise:2001:LSM

[332] David S. Wise, Jeremy D. Frens, Yuhong Gu, and Gregory A. Alexander. Language support for Mortonorder matrices. ACM SIGPLAN Notices, 36(7):24–33, July 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/ppopp/
379539/p24-wise/p24-wise.pdf;
http://www.acm.org/pubs/citations/
proceedings/ppopp/379539/p24-wise/

.

vanNieuwpoort:2001:ELB

[333] Rob V. van Nieuwpoort, Thilo Kielmann, and Henri E. Bal. Efficient load balancing for widearea divide-and-conquer applications. ACM SIGPLAN Notices, 36(7):34– 43, July 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/articles/proceedings/ppopp/
                                        379539/p34-van_nieuwpoort/p34-van_
                                        nieuwpoort.pdf;   http://www.acm.
                                        org/pubs/citations/proceedings/
                                        ppopp/379539/p34-van_nieuwpoort/

.

Scott:2001:SQB

[334] Michael L. Scott and William N. Scherer. Scalable queue-based spin locks with timeout. ACM SIGPLAN Notices, 36(7):44–52, July 2001. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

                                        org/pubs/articles/proceedings/ppopp/
                                        379539/p44-scott/p44-scott.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/ppopp/379539/p44-scott/

.

Lu:2001:CER

[335] Honghui Lu, Alan L. Cox, and Willy Zwaenepoel. Contention elimination by replication of sequential sections in distributed shared memory programs. ACM SIGPLAN Notices, 36(7):53–61, July 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/articles/proceedings/ppopp/
                                        379539/p53-lu/p53-lu.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        ppopp/379539/p53-lu/.

Morris:2001:ADR

[336] Donald G. Morris and David K. Lowenthal. Accurate data redistribution cost estimation in software distributed shared memory systems. ACM SIGPLAN Notices, 36(7):62–71, July 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/ppopp/
379539/p62-morris/p62-morris.pdf;
http://www.acm.org/pubs/citations/
proceedings/ppopp/379539/p62-morris/

.

Rencuzogullari:2001:DAA

[337] Umit Rencuzogullari and Sandhya Dwardadas. Dynamic adaptation to available resources for parallel computing in an autonomous network of workstations. ACM SIGPLAN Notices, 36(7):72–81, July 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/ppopp/
379539/p72-rencuzogullari/p72-rencuzogullari.
pdf;   http://www.acm.org/pubs/
citations/proceedings/ppopp/379539/
p72-rencuzogullari/.

Chien:2001:PPC

[338] Andrew A. Chien. Parallel programming challenges for Internet-scale computing (entropia). ACM SIGPLAN Notices, 36(7):82, July 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/ppopp/
379539/p82-chien/p82-chien.pdf;
http://www.acm.org/pubs/citations/
proceedings/ppopp/379539/p82-chien/

.

Veldema:2001:SLG

[339] R. Veldema, R. F. H. Hofman, R. A. F.

Bhoedjang, C. J. H. Jacobs, and H. E. Bal. Source-level global optimizations for fine-grain distributed shared memory systems. ACM SIGPLAN Notices, 36(7):83–92, July 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/ppopp/
379539/p83-veldema/p83-veldema.
pdf;   http://www.acm.org/pubs/
citations/proceedings/ppopp/379539/
p83-veldema/.

Voss:2001:HLA

[340] Michael J. Voss and Rudolf Eigemann. High-level adaptive program optimization with ADAPT. ACM SIGPLAN Notices, 36(7):93– 102, July 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/ppopp/
379539/p93-voss/p93-voss.pdf;
http://www.acm.org/pubs/citations/
proceedings/ppopp/379539/p93-voss/

.

Lim:2001:BAC

[341] Amy W. Lim, Shih-Wei Liao, and Monica S. Lam. Blocking and array contraction across arbitrarily nested loops using affine partitioning. ACM SIGPLAN Notices, 36 (7):103–112, July 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/ppopp/
379539/p103-lim/p103-lim.pdf;
http://www.acm.org/pubs/citations/
proceedings/ppopp/379539/p103-lim/

.

Carrera:2001:EVP

[342] Enrique V. Carrera and Ricardo Bianchini. Efficiency vs. portability in cluster-based network servers. ACM SIGPLAN Notices, 36(7):113– 122, July 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/ppopp/
379539/p113-carrera/p113-carrera.
pdf;   http://www.acm.org/pubs/
citations/proceedings/ppopp/379539/
p113-carrera/.

Vetter:2001:SSA

[343] Jeffrey S. Vetter and Michael O. McCracken. Statistical scalability analysis of communication operations in distributed applications. ACM SIGPLAN Notices, 36(7):123– 132, July 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/ppopp/
379539/p123-vetter/p123-vetter.
pdf;   http://www.acm.org/pubs/
citations/proceedings/ppopp/379539/
p123-vetter/.

Ino:2001:LPC

[344] Fumihiko Ino, Noriyuki Fujimoto, and Kenichi Hagihara. LogGPS: a parallel computational model for synchronization analysis. ACM SIGPLAN Notices, 36(7):133–142, July 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/articles/proceedings/ppopp/
379539/p133-ino/p133-ino.pdf;

                                        http://www.acm.org/pubs/citations/
                                        proceedings/ppopp/379539/p133-ino/

.

Kuck:2001:PPD

[345] David Kuck. Peer to peer and distributed computing. ACM SIGPLAN Notices, 36(7):143, July 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/articles/proceedings/ppopp/
                                        379539/p143-kuck/p143-kuck.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/ppopp/379539/p143-kuck/

.

Lee:2001:AIL

[346] Sheayun Lee, Andreas Ermedahl, Sang Lyul Min, and Naehyuck Chang. An Accurate Instruction-Level Energy Consumption Model for Embedded RISC Processors. ACM SIG-PLAN Notices, 36(8):1–10, August 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kim:2001:HRT

[347] Minyoung Kim and Soonhoi Ha. Hybrid Run-time Power Management Technique for Real-time Embedded System with Voltage Scalable Processor. ACM SIGPLAN Notices, 36(8): 11–19, August 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kang:2001:PAD

[348] Dong-In Kang, Stephen Crago, and Jinwoo Suh. Power-aware Design Synthesis Techniques for Distributed Real-Time Systems. ACM SIGPLAN Notices, 36(8):20–28, August 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

DeSutter:2001:CGC

[349] Bjorn De Sutter, Bruno De Bus, Koen De Bosschere, and Saumya Debray. Combining global code and data compaction. ACM SIGPLAN Notices, 36 (8):29–38, August 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Park:2001:RAB

[350] Jinpyo Park, Je-Hyung Lee, and Soo-Mook Moon. Register allocation for banked register file. ACM SIGPLAN Notices, 36(8):39–47, August 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Huang:2001:LTA

[351] Xianglong Huang, Steve Carr, and Philip Sweany. Loop Transformations for Architectures with Partitioned Register Banks. ACM SIG-PLAN Notices, 36(8):48–55, August 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cadot:2001:ECL

[352] Sidney Cadot, Frits Kuijlman, Koen Langendoen, Kees van Reeuwijk, and Henk Sips. ENSEMBLE: A communication layer for embedded multiprocessor systems. ACM SIGPLAN Notices, 36(8):56–63, August 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Henzinger:2001:ECS

[353] Thomas A. Henzinger, Benjamin Horowitz, and Christoph Meyer Kirsch. Embedded Control Systems Development with Giotto. ACM SIGPLAN Notices, 36(8):64–72, August 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Palopoli:2001:TSF

[354] Luigi Palopoli, Giuseppe Lipari, Luca Abeni, Marco Di Natale, Paolo Ancilotti, and Fabio Conticelli. A tool for simulation and fast prototyping of embedded control systems. ACM SIG-PLAN Notices, 36(8):73–81, August 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bakshi:2001:MMB

[355] A. Bakshi, V. K. Prasanna, and A. Ledeczi. MILAN: A model based integrated simulation framework for design of embedded systems. ACM SIG-PLAN Notices, 36(8):82–87, August 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vivancos:2001:PTA

[356] Emilio Vivancos, Christopher Healy, Frank Mueller, and David Whalley. Parametric Timing Analysis. ACM SIGPLAN Notices, 36(8):88–93, August 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ziegenbein:2001:IBA

[357] Dirk Ziegenbein, Fabian Wolf, Kai Richter, Marek Jersak, and Rolf Ernst.

Interval-based Analysis of Software Processes. ACM SIGPLAN Notices, 36 (8):94–101, August 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Unnikrishnan:2001:AAL

[358] Leena Unnikrishnan, Scott D. Stoller, and Yanhong A. Liu. Automatic Accurate Live Memory Analysis for Garbage-Collected Languages. ACM SIGPLAN Notices, 36(8):102–111, August 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Theiling:2001:GDT

[359] Henrik Theiling. Generating Decision Trees for Decoding Binaries. ACM SIGPLAN Notices, 36(8):112–120, August 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Merillion:2001:DHE

[360] Fabrice Mérillion and Gilles Muller. Dealing with Hardware in Embedded Software: A General Framework Based on the Devil Language. ACM SIG-PLAN Notices, 36(8):121–127, August 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Merillon:2001:DHE

[361] Fabrice Merillon and Gilles Muller. Dealing with hardware in embedded software: A general framework based on the Devil language. ACM SIG-PLAN Notices, 36(8):121–127, August 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kadayif:2001:MCA

[362] Ismail Kadayif, Mahmut Taylan Kandemir, Narayanan Vijaykrishnan, Mary Jane Irwin, and Jagannathan Ramanujam. Morphable Cache Architectures: Potential Benefits. ACM SIG-PLAN Notices, 36(8):128–137, August 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Granston:2001:SPI

[363] Elana Granston, Eric Stotzer, and Joe Zbiciak. Software pipelining irregular loops on the TMS320C6000 VLIW DSP architecture. ACM SIGPLAN Notices, 36(8):138–144, August 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kastner:2001:IBI

[364] Daniel Kästner and Sebastian Winkel. ILP-based instruction scheduling for IA-64. ACM SIGPLAN Notices, 36(8): 145–154, August 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wagner:2001:CCD

[365] Jens Wagner and Rainer Leupers. C compiler design for an industrial network processor. ACM SIGPLAN Notices, 36(8):155–164, August 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kessler:2001:DPA

[366] Christoph Keßler and Andrzej Bednarski. A Dynamic Programming Approach to Optimal Integrated Code Generation. ACM SIGPLAN Notices, 36(8):165–174, August 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yellin:2001:SMC

[367] Daniel M. Yellin. Stuck in the Middle: Challenges and Trends in Optimizing Middleware. ACM SIG-PLAN Notices, 36(8):175–180, August 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.

cs.wisc.edu/~bodik/om2001/danny-
abstract.html.

Hogstedt:2001:OCI

[368] Karin Högstedt, Doug Kimelman, and Nan Wang. Optimizing Component Interaction. ACM SIGPLAN Notices, 36 (8):181, August 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wegman:2001:OCI

[369] Mark Wegman. Optimizing component interaction. ACM SIGPLAN Notices, 36(8):181, August 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.cs.wisc.
edu/~bodik/om2001/mark-abstract.
html.

Larus:2001:UCS

[370] James R. Larus and Michael Parkes. Using cohort scheduling to enhance server performance (extended abstract). ACM SIGPLAN Notices, 36 (8):182–187, August 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.cs.wisc.
                                        edu/~bodik/om2001/mark-abstract.
                                        html.

Narain:2001:MBA

[371] Sanjai Narain, Ravichander Vaidyanathan, Stanley Moyer, William Stephens, Kirthika Parmeswaran, and Abdul Rahim Shareef. Middleware for building adaptive systems via configuration. ACM SIGPLAN Notices, 36(8): 188–195, August 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gore:2001:DOS

[372] Pradeep Gore, Ron Cytron, Douglas Schmidt, and Carlos O’Ryan. Designing and optimizing a scalable CORBA notification service. ACM SIGPLAN Notices, 36(8):196–204, August 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Othman:2001:IDA

[373] Ossama Othman and Douglas C. Schmidt. Issues in the Design of Adaptive Middleware Load Balancing. ACM SIGPLAN Notices, 36(8):205–213, August 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pyarali:2001:EOT

[374] Irfan Pyarali, Marina Spivak, Ron Cytron, and Douglas C. Schmidt. Evaluating and optimizing thread pool strategies for real-time CORBA. ACM SIGPLAN Notices, 36(8):214–222, August 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brunsch:2001:DES

[375] Darrell Brunsch, Carlos O’Ryan, and Douglas C. Schmidt. Designing an efficient and scalable server-side asynchrony model for CORBA. ACM SIG-PLAN Notices, 36(8):223–229, August 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Krishnamurthy:2001:IQE

[376] Yamuna Krishnamurthy, Vishal Kachroo, David A. Karr, Craig Rodrigues, Joseph P. Loyall, Richard Schantz, and Douglas C. Schmidt. Integration of QoS-enabled distributed object computing middleware for developing next-generation distributed applications. ACM SIGPLAN Notices, 36(8): 230–237, August 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Adve:2001:LCS

[377] Vikram Adve, Vinh Vi Lam, and Brian Ensink. Language and compiler support for adaptive distributed applications. ACM SIGPLAN Notices, 36(8): 238–246, August 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lee:2001:ESC

[378] Edward A. Lee. Embedded software from concurrent component models. ACM SIGPLAN Notices, 36 (8):??, August 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://redwood.snu.
ac.kr/lctes2001/edward-abstract.

html. Invited paper; not in conference proceedings.

Palem:2001:WCW

[379] Krishna V. Palem. What is a compiler when the architecture is not hard? ACM SIGPLAN Notices, 36(8):??, August 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Listed in conference program, but did not appear in conference proceedings.

Boehm:2001:LNE

[380] Hans Boehm. Letter from the newly elected Chair. ACM SIGPLAN Notices, 36(9):1–2, September 2001. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Whalley:2001:LLS

[381] Dave Whalley. Letter from the LCTES Steering Committee. ACM SIGPLAN Notices, 36(9):3, September 2001. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2001:C

[382] Anonymous. Calendar. ACM SIG-PLAN Notices, 36(9):4–16, September 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2001:P

[383] Anonymous. In parallel. ACM SIG-PLAN Notices, 36(9):17, September 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2001:LS

[384] Anonymous. LCTES’02 and SCOPES’02. ACM SIGPLAN Notices, 36(9):18, September 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Page:2001:FPW

[385] Rex Page. Functional programming, and where you can put it. ACM SIG-PLAN Notices, 36(9):19–24, September 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2001:AOC

[386] Anonymous. ASTs for optimizing compilers. ACM SIGPLAN Notices, 36(9): 25–30, September 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clark:2001:AOC

[387] Chris Clark. ASTs for optimizing compilers. ACM SIGPLAN Notices, 36(9): 25–30, September 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Christopher:2001:TPO

[388] Thomas W. Christopher. Test programs to observe parallelism. ACM SIGPLAN Notices, 36(9):31– 36, September 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2001:IA

[389] Anonymous. Information for authors. ACM SIGPLAN Notices, 36(9): 37, September 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2001:IPA

[390] Anonymous. Information for past authors. ACM SIGPLAN Notices, 36(9): 38, September 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhu:2001:DSP

[391] Ming-Yuan Zhu. Denotational semantics of programming languages and compiler generation in PowerEpsilon. ACM SIGPLAN Notices, 36(9):39– 53, September 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhu:2001:FSD

[392] Ming-Yuan Zhu. Formal specifications of debuggers. ACM SIGPLAN Notices, 36(9):54–63, September 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kienle:2001:USR

[393] Holger M. Kienle. Using smgn for rapid prototyping of small domainspecific languages. ACM SIGPLAN Notices, 36(9):64–73, September 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Moore:2001:SSH

[394] David L. Moore. smgn: Some historical remarks. ACM SIGPLAN Notices, 36(9):74–76, September 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gassanenko:2001:ACL

[395] Michael L. Gassanenko. About code layout for Lisp. ACM SIGPLAN Notices, 36(9):77–80, September 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lu:2001:DZA

[396] Lunjin Lu. On Dart-Zobel algorithm for testing regular type inclusion. ACM SIGPLAN Notices, 36(9): 81–85, September 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shaofeng:2001:FDW

[397] Wang Shaofeng and Sun Jiaguang. A framework design of workflow management system with Java RMI. ACM SIGPLAN Notices, 36(9):86– 93, September 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wang:2001:FDW

[398] Shaofeng Wang and Jiaguang Sun. A framework design for workflow management system with Java RMI. ACM SIGPLAN Notices, 36(9):86– 93, September 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2001:FP

[399] Anonymous. Functional programming. ACM SIGPLAN Notices, 36(9): ??, September 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2001:PPP

[400] Anonymous. Practical parsing patterns. ACM SIGPLAN Notices, 36(9): ??, September 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2001:TCa

[401] Anonymous. Technical correspondence. ACM SIGPLAN Notices, 36 (9):??, September 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pitts:2001:FAR

[402] Andrew M. Pitts. A fresh approach to representing syntax with static binders in functional programming. ACM SIGPLAN Notices, 36(10):1, October 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://cristal.inria.
fr/ICFP2001/Abstracts/pitts.html.

Invited talk; abstract only.

Fluet:2001:CUD

[403] Matthew Fluet and Stephen Weeks. Contification using dominators. ACM SIGPLAN Notices, 36(10):2–13, October 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://

cristal.inria.fr/ICFP2001/Abstracts/
48.html.

Dimock:2001:FCT

[404] Allyn Dimock, Ian Westmacott, Robert Muller, Franklyn Turbak, and J. B. Wells. Functioning without closure: Type-safe customized function representations for Standard ML. ACM SIGPLAN Notices, 36 (10):14–25, October 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://cristal.inria.
fr/ICFP2001/Abstracts/55.html.

LeFessant:2001:OPM

[405] Fabrice Le Fessant and Luc Maranget. Optimizing pattern matching. ACM SIGPLAN Notices, 36(10):26–37, October 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://cristal.inria.
fr/ICFP2001/Abstracts/8.html.

Neubauer:2001:ELD

[406] Matthias Neubauer and Michael Sperber. Down with Emacs Lisp: Dynamic scope analysis. ACM SIGPLAN Notices, 36(10):38–49, October 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://cristal.

inria.fr/ICFP2001/Abstracts/50.
html.

Russo:2001:RSS

[407] Claudio V. Russo. Recursive structures for Standard ML. ACM SIGPLAN Notices, 36(10):50–61, October 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://cristal.

inria.fr/ICFP2001/Abstracts/28.
html.

Duggan:2001:TBH

[408] Dominic Duggan. Type-based hot swapping of running modules (extended abstract). ACM SIGPLAN Notices, 36(10):62–73, October 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://cristal.

                                        inria.fr/ICFP2001/Abstracts/21.
                                        html.

Ganz:2001:MMS

[409] Steven E. Ganz, Amr Sabry, and Walid Taha. Macros as multi-stage computations: Type-safe, generative, binding macros in MacroML. ACM SIGPLAN Notices, 36(10):74–85, October 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://cristal.

                                        inria.fr/ICFP2001/Abstracts/42.
                                        html.

Sheard:2001:GUT

[410] Tim Sheard. Generic unification via two-level types and parameterized modules. ACM SIGPLAN Notices, 36 (10):86–97, October 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://cristal.inria.
                                        fr/ICFP2001/Abstracts/13.html.

Chuang:2001:GVS

[411] Tyng-Ruey Chuang. Generic validation of structural content with parametric modules. ACM SIGPLAN Notices, 36 (10):98–109, October 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://cristal.inria.
                                        fr/ICFP2001/Abstracts/22.html.

Hinze:2001:SIT

[412] Ralf Hinze. A simple implementation technique for priority search queues. ACM SIGPLAN Notices, 36 (10):110–121, October 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://cristal.inria.
fr/ICFP2001/Abstracts/23.html.

Sperber:2001:DSL

[413] Michael Sperber. Developing a stage lighting system from scratch. ACM SIGPLAN Notices, 36(10):122– 133, October 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://cristal.inria.
fr/ICFP2001/Abstracts/32.html.

Anand:2001:CPP

[414] Saswat Anand, Wei-Ngan Chin, and Siau-Cheng Khoo. Charting patterns on price history. ACM SIGPLAN Notices, 36(10):134–145, October 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://cristal.

inria.fr/ICFP2001/Abstracts/64.
html.

Wan:2001:RTF

[415] Zhanyong Wan, Walid Taha, and Paul Hudak. Real-time FRP. ACM SIGPLAN Notices, 36(10):146– 156, October 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://cristal.inria.
fr/ICFP2001/Abstracts/43.html.

Russell:2001:EHH

[416] George Russell. Events in Haskell, and how to implement them. ACM SIGPLAN Notices, 36(10):157–168, October 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://cristal.inria.
fr/ICFP2001/Abstracts/24.html.

Hongwei:2001:DTA

[417] Hongwei Xi and Robert Harper. A dependently typed assembly language. ACM SIGPLAN Notices, 36 (10):169–180, October 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://cristal.inria.
                                        fr/ICFP2001/Abstracts/27.html.

Xi:2001:DTA

[418] Hongwei Xi and Robert Harper. A dependently typed assembly language. ACM SIGPLAN Notices, 36(10):169– 180, October 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Walker:2001:RLT

[419] David Walker and Kevin Watkins. On regions and linear types (extended abstract). ACM SIGPLAN Notices, 36 (10):181–192, October 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://cristal.inria.
                                        fr/ICFP2001/Abstracts/36.html.

Chitil:2001:CET

[420] Olaf Chitil. Compositional explanation of types and algorithmic debugging of type errors. ACM SIGPLAN Notices, 36(10):193–204, October 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://cristal.

                                        inria.fr/ICFP2001/Abstracts/40.
                                        html.

Chakravarty:2001:FAF

[421] Manuel M. T. Chakravarty and Gabriele Keller. Functional array fusion. ACM SIGPLAN Notices, 36 (10):205–216, October 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://cristal.inria.
fr/ICFP2001/Abstracts/38.html.

Nanevski:2001:AGS

[422] Aleksandar Nanevski, Guy Blelloch, and Robert Harper. Automatic generation of staged geometric predicates. ACM SIGPLAN Notices, 36 (10):217–228, October 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://cristal.inria.
fr/ICFP2001/Abstracts/14.html.

Paterson:2001:NNA

[423] Ross Paterson. A new notation for arrows. ACM SIGPLAN Notices, 36 (10):229–240, October 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://cristal.inria.
fr/ICFP2001/Abstracts/49.html.

Zenger:2001:EAD

[424] Matthias Zenger and Martin Odersky. Extensible algebraic datatypes with defaults. ACM SIGPLAN Notices, 36 (10):241–252, October 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://cristal.inria.
fr/ICFP2001/Abstracts/34.html.

Grobauer:2001:CRD

[425] Bernd Grobauer. Cost recurrences for DML programs. ACM SIGPLAN Notices, 36(10):253–264, October 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://cristal.

                                        inria.fr/ICFP2001/Abstracts/45.
                                        html.

Gustavsson:2001:PLC

[426] Jörgen Gustavsson and David Sands. Possibilities and limitations of callby-need space improvement. ACM SIGPLAN Notices, 36(10):265–276, October 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://cristal.inria.
                                        fr/ICFP2001/Abstracts/60.html.

Anonymous:2001:AIc

[427] Anonymous. Author index. ACM SIGPLAN Notices, 36(10):277, October 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2001:IT

[428] Anonymous. Invited talk 1. ACM SIGPLAN Notices, 36(10):??, October 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2001:Sa

[429] Anonymous. Session 1. ACM SIG-PLAN Notices, 36(10):??, October 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2001:Sb

[430] Anonymous. Session 2. ACM SIG-PLAN Notices, 36(10):??, October 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2001:Sc

[431] Anonymous. Session 3. ACM SIG-PLAN Notices, 36(10):??, October 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2001:Sd

[432] Anonymous. Session 4. ACM SIG-PLAN Notices, 36(10):??, October 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2001:Se

[433] Anonymous. Session 5. ACM SIG-PLAN Notices, 36(10):??, October 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2001:Sf

[434] Anonymous. Session 6. ACM SIG-PLAN Notices, 36(10):??, October 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2001:Sg

[435] Anonymous. Session 7. ACM SIG-PLAN Notices, 36(10):??, October 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Augustsson:2001:BHD

[436] Lennart Augustsson. Bluespec — a hardware description language. ACM SIGPLAN Notices, 36(10):??, October 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Invited talk; no abstract or article in volume.

Frigo:2001:FFT

[437] Matteo Frigo. The Fastest Fourier Transform in the West. ACM SIG-PLAN Notices, 36(10):??, October 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://

cristal.inria.fr/ICFP2001/Abstracts/

frigo.html. Invited talk; no abstract or article in volume.

Findler:2001:CSO

[438] Robert Bruce Findler and Matthias Felleisen. Contract soundness for object-oriented languages. ACM SIG-PLAN Notices, 36(11):1–15, November 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOPSLA’01).

Ancona:2001:CCJ

[439] D. Ancona, G. Lagorio, and E. Zucca. A core calculus for Java exceptions. ACM SIGPLAN Notices, 36(11):16–30, November 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOPSLA’01).

Bachrach:2001:JSE

[440] Jonthan Bachrach and Keith Playford. The Java Syntactic Extender (JSE). ACM SIGPLAN Notices, 36(11):31–42, November 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOPSLA’01).

Rountev:2001:PAJ

[441] Atanas Rountev, Ana Milanova, and Barbara G. Ryder. Points-to analysis for Java using annotated constraints. ACM SIGPLAN Notices, 36(11):43–55, November 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOPSLA’01).

Boyapati:2001:PTS

[442] Chandrasekhar Boyapati and Martin Rinard. A parameterized type system for race-free Java programs. ACM SIG-PLAN Notices, 36(11):56–69, November 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOPSLA’01).

vonPraun:2001:ORD

[443] Christoph von Praun and Thomas R. Gross. Object race detection. ACM SIGPLAN Notices, 36(11):70–82, November 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOPSLA’01).

Ogasawara:2001:SEH

[444] Takeshi Ogasawara, Hideaki Komatsu, and Toshio Nakatani. A study of exception handling and its dynamic optimization in Java. ACM SIGPLAN Notices, 36(11):83–95, November 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOP-SLA’01).

Zibin:2001:EST

[445] Yoav Zibin and Joseph Yossi Gil. Efficient subtyping tests with PQencoding. ACM SIGPLAN Notices, 36(11):96–107, November 2001. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOP-SLA’01).

Alpern:2001:EIJ

[446] Bowen Alpern, Anthony Cocchi, Stephen Fink, and David Grove. Efficient implementation of Java interfaces: Invokeinterface considered harmless. ACM SIGPLAN Notices, 36(11):108– 124, November 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIG-PLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOPSLA’01).

Czajkowski:2001:MCV

[447] Grzegorz Czajkowski and Laurent Daynés. Multitasking without compromise: a virtual machine evolution. ACM SIGPLAN Notices, 36(11):125– 138, November 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIG-PLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOPSLA’01).

Binder:2001:PRC

[448] Walter Binder, Jane G. Hulaas, and Alex Villazón. Portable resource control in Java. ACM SIGPLAN Notices, 36(11):139–155, November 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOP-SLA’01).

Nakamura:2001:ICC

[449] Hiroaki Nakamura. Incremental computation of complex object queries. ACM SIGPLAN Notices, 36(11):156– 165, November 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIG-PLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOPSLA’01).

Whaley:2001:PMC

[450] John Whaley. Partial method compilation using dynamic profile information. ACM SIGPLAN Notices, 36 (11):166–179, November 2001. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOP-SLA’01).

Suganuma:2001:DOF

[451] Toshio Suganuma, Toshiaki Yasue, Motohiro Kawahito, Hideaki Komatsu, and Toshio Nakatani. A dynamic optimization framework for a Java justin-time compiler. ACM SIGPLAN Notices, 36(11):180–195, November 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOP-SLA’01).

Pechtchanski:2001:DOI

[452] Igor Pechtchanski and Vivek Sarkar. Dynamic optimistic interprocedural analysis: a framework and an application. ACM SIGPLAN Notices, 36 (11):195–210, November 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOP-SLA’01).

McDirmid:2001:JNA

[453] Sean McDirmid, Matthew Flatt, and Wilson C. Hsieh. Jiazzi: new-age components for old-fashioned Java. ACM SIGPLAN Notices, 36(11):211–222, November 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOPSLA’01).

Duggan:2001:MMB

[454] Dominic Duggan and Ching-Ching Techaubol. Modular mixin-based inheritance for application frameworks. ACM SIGPLAN Notices, 36(11):223– 240, November 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIG-PLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOPSLA’01).

Grothoff:2001:EOC

[455] Christian Grothoff, Jens Palsberg, and Jan Vitek. Encapsulating objects with confined types. ACM SIGPLAN Notices, 36(11):241–253, November 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOP-SLA’01).

Eugster:2001:OE

[456] Patrick Th. Eugster, Rachid Guerraoui, and Christian Heide Damm. On objects and events. ACM SIGPLAN Notices, 36(11):254–269, November 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOPSLA’01).

Visser:2001:VCT

[457] Joost Visser. Visitor combination and traversal control. ACM SIGPLAN Notices, 36(11):270–282, November 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOP-SLA’01).

Ostermann:2001:OOC

[458] Klaus Ostermann and Mira Mezini. Object-oriented composition untangled. ACM SIGPLAN Notices, 36(11): 283–299, November 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOP-SLA’01).

Lanza:2001:CCB

[459] Michele Lanza and Stéphane Ducasse. A categorization of classes based on the visualization of their internal structure: the class blueprint. ACM SIGPLAN Notices, 36(11):300–311, November 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOPSLA’01).

Harrold:2001:RTS

[460] Mary Jean Harrold, James A. Jones, Tongyu Li, Donglin Liang, and Ashish Gujarathi. Regression test selection for Java software. ACM SIGPLAN Notices, 36(11):312–326, November 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOP-SLA’01).

Riehle:2001:AUV

[461] Dirk Riehle, Steven Fraleigh, Dirk Bucka-Lassen, and Nosa Omorogbe. The architecture of a UML virtual machine. ACM SIGPLAN Notices, 36 (11):327–341, November 2001. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOP-SLA’01).

Blackburn:2001:PJ

[462] Stephen M. Blackburn, Sharad Singhai, Matthew Hertz, Kathryn S. McKinely, and J. Eliot B. Moss. Pretenuring for Java. ACM SIGPLAN Notices, 36(11):342–352, November 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOP-SLA’01).

Brecht:2001:CGC

[463] Tim Brecht, Eshrat Arjomandi, Chang Li, and Hang Pham. Controlling garbage collection and heap growth to reduce the execution time of Java applications. ACM SIGPLAN Notices, 36(11):353–366, November 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOP-SLA’01).

Levanoni:2001:FRC

[464] Yossi Levanoni and Erez Petrank. An on-the-fly reference counting garbage collector for Java. ACM SIGPLAN Notices, 36(11):367–380, November 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOP-SLA’01).

Vlissides:2001:IIS

[465] John Vlissides and Doug Lea. Introduction to this issue’s special section on Intriguing Technology from OOPSLA. ACM SIGPLAN Notices, 36(12):1, December 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2001:CCCb

[466] Anonymous. Conference corner: Calendar. ACM SIGPLAN Notices, 36(12): 2–11, December 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2001:CCIa

[467] Anonymous. Conference corner: International Corner on Software Engineering (ICSE). ACM SIGPLAN Notices, 36(12):12, December 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2001:CCF

[468] Anonymous. Conference corner: Foundations of Software Engineering (FSE). ACM SIGPLAN Notices, 36(12):13, December 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2001:CCP

[469] Anonymous. Conference corner: Program Analysis for System Tools and Engineering (PASTE). ACM SIG-PLAN Notices, 36(12):14, December 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2001:CCIb

[470] Anonymous. Conference corner: International Symposium on Memory Management (ISMM). ACM SIGPLAN Notices, 36(12):15, December 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:2001:FM

[471] Paul Frenger. Forth mindstorms. ACM SIGPLAN Notices, 36(12):16–19, December 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Varela:2001:PDR

[472] Carlos Varela and Gul Agha. Programming dynamically reconfigurable open systems with SALSA. ACM SIGPLAN Notices, 36(12):20–34, December 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nolte:2001:TPH

[473] Jörg Nolte, Yutaka Ishikawa, and Mitsuhisa Sato. TACO — prototyping high-level object-oriented programming constructs by means of template based programming techniques. ACM SIGPLAN Notices, 36(12):35–49, December 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yoder:2001:ADA

[474] Joseph W. Yoder, Federico Balaguer, and Ralph Johnson. Architecture and design of adaptive object-models. ACM SIGPLAN Notices, 36(12):50–60, December 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pawson:2001:NOT

[475] Richard Pawson and Robert Matthews. Naked objects: a technique for designing more expressive systems. ACM SIGPLAN Notices, 36(12):61–67, December 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fenwick:2001:IAc

[476] Jay Fenwick and Cindy Norris. Information for authors. ACM SIG-PLAN Notices, 36(12):68, December 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fenwick:2001:IPAc

[477] Jay Fenwick and Cindy Norris. Information for past authors. ACM SIG-PLAN Notices, 36(12):69, December 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xu:2001:DAR

[478] Baowen Xu and Zhenqiang Chen. Dependence analysis for recursive Java programs. ACM SIGPLAN Notices, 36 (12):70–76, December 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dolog:2001:RCA

[479] Peter Dolog, Valentino Vranić, and Mária Bieliková. Representing change by aspect. ACM SIGPLAN Notices, 36(12):77–83, December 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Avdicausevic:2001:AED

[480] Enis Avdičaušević, Mitja Lenič, Marjan Mernik, and Viljem Žumer. AspectCOOL: An experiment in design and implementation of aspect-oriented language. ACM SIGPLAN Notices, 36 (12):84–94, December 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2001:F

[481] Anonymous. Forth. ACM SIGPLAN Notices, 36(12):??, December 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2001:ITO

[482] Anonymous. Intriguing technology from OOPSLA. ACM SIGPLAN Notices, 36(12):??, December 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2001:TCb

[483] Anonymous. Technical correspondence. ACM SIGPLAN Notices, 36 (12):??, December 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ball:2002:SPD

[484] Thomas Ball and Sriram K. Rajamani. The Slam project: debugging system software via static analysis. ACM SIGPLAN Notices, 37(1):1–3, January 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ammons:2002:MS

[485] Glenn Ammons, Rastislav Bodík, and James R. Larus. Mining specifications. ACM SIGPLAN Notices, 37(1):4–16, January 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kuncak:2002:RA

[486] Viktor Kuncak, Patrick Lam, and Martin Rinard. Role analysis. ACM SIG-PLAN Notices, 37(1):17–32, January 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Abadi:2002:ASP

[487] Martín Abadi and Bruno Blanchet. Analyzing security protocols with secrecy types and logic programs. ACM SIG-PLAN Notices, 37(1):33–44, January 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chaki:2002:TMM

[488] Sagar Chaki, Sriram K. Rajamani, and Jakob Rehof. Types as models: model checking message-passing programs. ACM SIGPLAN Notices, 37(1): 45–57, January 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Henzinger:2002:LA

[489] Thomas A. Henzinger, Ranjit Jhala, Rupak Majumdar, and Grégoire Sutre. Lazy abstraction. ACM SIGPLAN Notices, 37(1):58–70, January 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Merro:2002:BCS

[490] Massimo Merro and Matthew Hennessy. Bisimulation congruences in safe ambients. ACM SIGPLAN Notices, 37 (1):71–80, January 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Honda:2002:UTS

[491] Kohei Honda and Nobuko Yoshida. A uniform type structure for secure information flow. ACM SIGPLAN Notices, 37(1):81–92, January 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boehm:2002:BSU

[492] Hans-J. Boehm. Bounding space usage of conservative garbage collectors. ACM SIGPLAN Notices, 37(1):93–100, January 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Petrank:2002:HCC

[493] Erez Petrank and Dror Rawitz. The hardness of cache conscious data placement. ACM SIGPLAN Notices, 37(1): 101–112, January 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Philipose:2002:TAC

[494] Matthai Philipose, Craig Chambers, and Susan J. Eggers. Towards automatic construction of staged compilers. ACM SIGPLAN Notices, 37(1): 113–125, January 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Felleisen:2002:PCB

[495] Matthias Felleisen. From POPL to the classroom and back. ACM SIG-PLAN Notices, 37(1):126–127, January 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Necula:2002:CTS

[496] George C. Necula, Scott McPeak, and Westley Weimer. CCured: type-safe retrofitting of legacy code. ACM SIG-PLAN Notices, 37(1):128–139, January 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rubin:2002:EPA

[497] Shai Rubin, Rastislav Bodík, and Trishul Chilimbi. An efficient profileanalysis framework for data-layout optimizations. ACM SIGPLAN Notices, 37(1):140–153, January 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ramsey:2002:SLC

[498] Norman Ramsey and Avi Pfeffer. Stochastic lambda calculus and monads of probability distributions. ACM SIGPLAN Notices, 37(1):154–165, January 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Banerjee:2002:RIC

[499] Anindya Banerjee and David A. Naumann. Representation independence, confinement and access control [extended abstract]. ACM SIGPLAN Notices, 37(1):166–177, January 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cousot:2002:SDP

[500] Patrick Cousot and Radhia Cousot. Systematic design of program transformation frameworks by abstract interpretation. ACM SIGPLAN Notices, 37(1):178–190, January 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Flanagan:2002:PAS

[501] Cormac Flanagan and Shaz Qadeer. Predicate abstraction for software verification. ACM SIGPLAN Notices, 37 (1):191–202, January 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Su:2002:FOT

[502] Zhendong Su, Alexander Aiken, Joachim Niehren, Tim Priesnitz, and Ralf Treinen. The first-order theory of subtyping constraints. ACM SIG-PLAN Notices, 37(1):203–216, January 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shao:2002:TSC

[503] Zhong Shao, Bratin Saha, Valery Trifonov, and Nikolaos Papaspyrou. A type system for certified binaries. ACM SIGPLAN Notices, 37(1):217–232, January 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Neubauer:2002:FLO

[504] Matthias Neubauer, Peter Thiemann, Martin Gasbichler, and Michael Sperber. Functional logic overloading. ACM SIGPLAN Notices, 37(1):233–244, January 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Manohar:2002:SFD

[505] Rajit Manohar. Scalable formal design methods for asynchronous VLSI. ACM SIGPLAN Notices, 37(1):245–246, January 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Acar:2002:AFP

[506] Umut A. Acar, Guy E. Blelloch, and Robert Harper. Adaptive functional programming. ACM SIGPLAN Notices, 37(1):247–259, January 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hofmann:2002:SNS

[507] Martin Hofmann. The strength of non-size increasing computation. ACM SIGPLAN Notices, 37(1):260–269, January 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lerner:2002:CDA

[508] Sorin Lerner, David Grove, and Craig Chambers. Composing dataflow analyses and transformations. ACM SIG-PLAN Notices, 37(1):270–282, January 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lacey:2002:PCC

[509] David Lacey, Neil D. Jones, Eric Van Wyk, and Carl Christian Frederiksen. Proving correctness of compiler optimizations by temporal logic. ACM SIGPLAN Notices, 37(1):283–294, January 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shuf:2002:EPT

[510] Yefim Shuf, Manish Gupta, Rajesh Bordawekar, and Jaswinder Pal Singh. Exploiting prolific types for memory management and optimizations. ACM SIGPLAN Notices, 37(1):295–306, January 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fournet:2002:SIT

[511] Cédric Fournet and Andrew D. Gordon. Stack inspection: theory and variants. ACM SIGPLAN Notices, 37(1): 307–318, January 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pottier:2002:IFI

[512] François Pottier and Vincent Simonet. Information flow inference for ML. ACM SIGPLAN Notices, 37(1):319– 330, January 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Igarashi:2002:RUA

[513] Atsushi Igarashi and Naoki Kobayashi. Resource usage analysis. ACM SIG-PLAN Notices, 37(1):331–342, January 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Noble:2002:FRJ

[514] Julian V. Noble. Forth report: jump tables and state machines. ACM SIG-PLAN Notices, 37(2):14–19, February 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Christopher:2002:PFP

[515] Thomas W. Christopher. In parallel: Facilities for parallelism: what should be in a parallel library. ACM SIG-PLAN Notices, 37(2):20–22, February 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Creak:2002:HTWa

[516] Alan Creak. How things were: Programming lessons from days gone by: backward ran sentences until reeled the mind. ACM SIGPLAN Notices, 37(2): 23–26, February 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Singh:2002:ADC

[517] Prabhat K. Singh and Arun Lakhotia. Analysis and detection of computer viruses and worms: an annotated bibliography. ACM SIGPLAN Notices, 37(2):29–35, February 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rashid:2002:WAP

[518] Awais Rashid. Weaving aspects in a persistent environment. ACM SIG-PLAN Notices, 37(2):36–44, February 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kuhl:2002:OLA

[519] Bernd Kühl and Axel-Tobias Schreiner. Objects for lexical analysis. ACM SIG-PLAN Notices, 37(2):45–52, February 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Siegfried:2002:SLH

[520] Robert M. Siegfried. A scripting language to help the blind to program visually. ACM SIGPLAN Notices, 37(2): 53–56, February 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nedjah:2002:ECD

[521] Nadia Nedjah and Luiza de Macedo Mourelle. Efficient concise deterministic patternmatching automata for ambiguous patterns. ACM SIGPLAN Notices, 37(2): 57–67, February 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zongyan:2002:BVE

[522] Qiu Zongyan. Bit-vector encoding of n-queen problem. ACM SIGPLAN Notices, 37(2):68–70, February 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chambers:2002:SC

[523] Craig Chambers. Staged compilation (invited talk). ACM SIGPLAN Notices, 37(3):1–8, March 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2002 ACM SIGPLAN Workshop Partial Evaluation and Semantics-Based Program Manipulation (PEPM’02).

Meur:2002:TBG

[524] Anne-Françoise Le Meur, Julia L. Lawall, and Charles Consel. Towards bridging the gap between programming languages and partial evaluation. ACM SIGPLAN Notices, 37(3):9–18, March 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2002 ACM SIGPLAN Workshop Partial Evaluation and Semantics-Based Program Manipulation (PEPM’02).

Asai:2002:OPE

[525] Kenichi Asai. Online partial evaluation for shift and reset. ACM SIGPLAN Notices, 37(3):19–30, March 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2002 ACM SIGPLAN Workshop Partial Evaluation and Semantics-Based Program Manipulation (PEPM’02).

Brabrand:2002:GLM

[526] Claus Brabrand and Michael I. Schwartzbach. Growing languages with metamorphic syntax macros. ACM SIGPLAN Notices, 37(3):31–40, March 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2002 ACM SIGPLAN Workshop Partial Evaluation and Semantics-Based Program Manipulation (PEPM’02).

Secher:2002:CID

[527] Jens Peter Secher and Morten Heine Sørensen. From checking to inference via driving and dag grammars. ACM SIGPLAN Notices, 37(3):41–51, March 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2002 ACM SIGPLAN Workshop Partial Evaluation and Semantics-Based Program Manipulation (PEPM’02).

Vidal:2002:CAN

[528] Germán Vidal. Cost-augmented narrowing-driven specialization. ACM SIGPLAN Notices, 37(3):52–62, March 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2002 ACM SIGPLAN Workshop Partial Evaluation and Semantics-Based Program Manipulation (PEPM’02).

Lu:2002:PDA

[529] Lunjin Lu. Path dependent analysis of logic programs. ACM SIG-PLAN Notices, 37(3):63–74, March 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2002 ACM SIGPLAN Workshop Partial Evaluation and Semantics-Based Program Manipulation (PEPM’02).

Gomez:2002:ATB

[530] Gustavo Gómez and Yanhong A. Liu. Automatic time-bound analysis for a higher-order language. ACM SIG-PLAN Notices, 37(3):75–86, March 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2002 ACM SIGPLAN Workshop Partial Evaluation and Semantics-Based Program Manipulation (PEPM’02).

Giacobazzi:2002:CPS

[531] Roberto Giacobazzi and Isabella Mastroeni. Compositionality in the puzzle of semantics. ACM SIGPLAN Notices, 37(3):87–97, March 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2002 ACM SIGPLAN Workshop Partial Evaluation and Semantics-Based Program Manipulation (PEPM’02).

Bischof:2002:IAD

[532] Christian H. Bischof, Paul D. Hovland, and Boyana Norris. Implementation of automatic differentiation tools (invited talk). ACM SIGPLAN Notices, 37(3): 98–107, March 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2002 ACM SIG-PLAN Workshop Partial Evaluation and Semantics-Based Program Manipulation (PEPM’02).

Liu:2002:POU

[533] Yanhong A. Liu and Scott D. Stoller. Program optimization using indexed and recursive data structures. ACM SIGPLAN Notices, 37(3):108–118, March 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2002 ACM SIG-PLAN Workshop Partial Evaluation and Semantics-Based Program Manipulation (PEPM’02).

Ramakrishnan:2002:MII

[534] Naren Ramakrishnan, Robert Capra, and Manuel A. Pérez-Quiñones. Mixedinitiative interaction = mixed computation. ACM SIGPLAN Notices, 37(3): 119–130, March 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2002 ACM SIG-PLAN Workshop Partial Evaluation and Semantics-Based Program Manipulation (PEPM’02).

Barthe:2002:CTI

[535] Gilles Barthe and Tarmo Uustalu. CPS translating inductive and coinductive types. ACM SIGPLAN Notices, 37(3): 131–142, March 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2002 ACM SIG-PLAN Workshop Partial Evaluation and Semantics-Based Program Manipulation (PEPM’02).

Ramsey:2002:LSP

[536] Normal Ramsey. L^{A}TEX support for proceedings. ACM SIGPLAN Notices, 37(4):1–3, April 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cytron:2002:SAA

[537] Ron Cytron. SIGPLAN Awards: The 2001 Award Winners. ACM SIG-PLAN Notices, 37(4):4, April 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fenwick:2002:CCC

[538] Jay Fenwick. Conference corner: Calendar. ACM SIGPLAN Notices, 37 (4):5–14, April 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2002:CPF

[539] Anonymous. Call for participation: First International IFIP/ACM Working Conference on Component Deployment, June 20–21, 2002, Berlin Germany. ACM SIGPLAN Notices, 37 (4):15, April 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2002:CPI

[540] Anonymous. Call for participation: ISSTA 2002: International Symposium on Software Testing and Analysis. ACM SIGPLAN Notices, 37(4):16, April 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:2002:FRS

[541] Paul Frenger. Forth report: Strap-on Forth. ACM SIGPLAN Notices, 37 (4):17–20, April 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Creak:2002:HTWb

[542] Alan Creak. How things were: Programming lessons from days gone by: Cobol, stewed for students. ACM SIGPLAN Notices, 37(4):21–26, April 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Christopher:2002:PCR

[543] Thomas W. Christopher. In parallel: Communicating reactive objects: Message-driven parallelism. ACM SIGPLAN Notices, 37(4):27–28, April 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fenwick:2002:IAa

[544] Jay Fenwick and Cindy Norris. Information for authors. ACM SIG-PLAN Notices, 37(4):29, April 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fenwick:2002:IPAa

[545] Jay Fenwick and Cindy Norris. Information for past authors. ACM SIG-PLAN Notices, 37(4):30, April 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Weisert:2002:IYO

[546] Conrad Weisert. If your only tool is a hammer, everything looks like a nail: Pseudo object-oriented programming considered harmful. ACM SIG-PLAN Notices, 37(4):31, April 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zee:2002:WBRa

[547] Karen Zee and Martin Rinard. Write barrier removal by static analysis. ACM SIGPLAN Notices, 37(4):32–41, April 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:2002:EGR

[548] Yumin Zhang, Xiaobo (Sharon) Hu, and Danny Z. Chen. Efficient global register allocation for minimizing energy consumption. ACM SIGPLAN Notices, 37(4):42–53, April 2002. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Spinellis:2002:MPC

[549] Diomidis Spinellis. A modest proposal for curing the public field phobia. ACM SIGPLAN Notices, 37(4):54–56, April 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ledru:2002:SPJ

[550] Pascal Ledru. Smart proxies for Jini services. ACM SIGPLAN Notices, 37 (4):57–61, April 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Acebal:2002:GDP

[551] César F. Acebal, Raúl Izquierdo Castanedo, and Juan M. Cueva Lovelle. Good design principles in a compiler university course. ACM SIGPLAN Notices, 37(4):62–73, April 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wells:2002:HMA

[552] Paul Damian Wells. The HSSM macroarchitecture, Virtual Machine and H languages. ACM SIGPLAN Notices, 37 (4):74–82, April 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Foster:2002:FST

[553] Jeffrey S. Foster, Tachio Terauchi, and Alex Aiken. Flow-sensitive type qualifiers. ACM SIGPLAN Notices, 37(5):1– 12, May 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fahndrich:2002:AFP

[554] Manuel Fahndrich and Robert DeLine. Adoption and focus: practical linear types for imperative programming. ACM SIGPLAN Notices, 37(5):13–24, May 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Budimlic:2002:FCC

[555] Zoran Budimlic, Keith D. Cooper, Timothy J. Harvey, Ken Kennedy, Timothy S. Oberg, and Steven W. Reeves. Fast copy coalescing and live-range identification. ACM SIG-PLAN Notices, 37(5):25–32, May 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Koseki:2002:PDG

[556] Akira Koseki, Hideaki Komatsu, and Toshio Nakatani. Preference-directed graph coloring. ACM SIGPLAN Notices, 37(5):33–44, May 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gargi:2002:SAP

[557] Karthik Gargi. A sparse algorithm for predicated global value numbering. ACM SIGPLAN Notices, 37(5):45–56, May 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Das:2002:EPS

[558] Manuvir Das, Sorin Lerner, and Mark Seigle. ESP: path-sensitive program verification in polynomial time. ACM SIGPLAN Notices, 37(5):57–68, May 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hallem:2002:SLB

[559] Seth Hallem, Benjamin Chelf, Yichen Xie, and Dawson Engler. A system and language for building systemspecific, static analyses. ACM SIG-PLAN Notices, 37(5):69–82, May 2002.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ramalingam:2002:DSP

[560] G. Ramalingam, Alex Warshavsky, John Field, Deepak Goyal, and Mooly Sagiv. Deriving specialized program analyses for certifying componentclient conformance. ACM SIGPLAN Notices, 37(5):83–94, May 2002. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Debray:2002:PGC

[561] Saumya Debray and William Evans. Profile-guided code compression. ACM SIGPLAN Notices, 37(5):95–105, May 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rajagopalan:2002:PDO

[562] Mohan Rajagopalan, Saumya K. Debray, Matti A. Hiltunen, and Richard D. Schlichting. Profile-directed optimization of event-based programs. ACM SIGPLAN Notices, 37(5):106– 116, May 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Liao:2002:PPB

[563] Steve S. W. Liao, Perry H. Wang, Hong Wang, Gerolf Hoflehner, Daniel Lavery, and John P. Shen. Post-pass binary adaptation for software-based speculative precomputation. ACM SIG-PLAN Notices, 37(5):117–128, May 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ossia:2002:PIC

[564] Yoav Ossia, Ori Ben-Yitzhak, Irit Goft, Elliot K. Kolodner, Victor Leikehman, and Avi Owshanko. A parallel, incremental and concurrent GC for servers. ACM SIGPLAN Notices, 37 (5):129–140, May 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hallenberg:2002:CRI

[565] Niels Hallenberg, Martin Elsman, and Mads Tofte. Combining region inference and garbage collection. ACM SIG-PLAN Notices, 37(5):141–152, May 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Blackburn:2002:BGA

[566] Stephen M. Blackburn, Richard Jones, Kathryn S. McKinley, and J. Eliot B. Moss. Beltway: getting around garbage collection gridlock. ACM SIGPLAN Notices, 37(5):153–164, May 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

So:2002:CAF

[567] Byoungro So, Mary W. Hall, and Pedro C. Diniz. A compiler approach to fast hardware design space exploration in FPGA-based systems. ACM SIG-PLAN Notices, 37(5):165–176, May 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cociorva:2002:STT

[568] Daniel Cociorva, Gerald Baumgartner, Chi-Chung Lam, P. Sadayappan, J. Ramanujam, Marcel Nooijen, David E. Bernholdt, and Robert Harrison. Space-time trade-off optimization for a class of electronic structure calculations. ACM SIGPLAN Notices, 37 (5):177–186, May 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kawahito:2002:ESE

[569] Motohiro Kawahito, Hideaki Komatsu, and Toshio Nakatani. Effective sign extension elimination. ACM SIG-PLAN Notices, 37(5):187–198, May 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chilimbi:2002:DHD

[570] Trishul M. Chilimbi and Martin Hirzel. Dynamic hot data stream prefetching for general-purpose programs. ACM SIGPLAN Notices, 37(5):199– 209, May 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wu:2002:EDR

[571] Youfeng Wu. Efficient discovery of regular stride patterns in irregular programs and its use in compiler prefetching. ACM SIGPLAN Notices, 37(5): 210–221, May 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Burtscher:2002:SLC

[572] Martin Burtscher, Amer Diwan, and Matthias Hauswirth. Static load classification for improving the value predictability of data-cache misses. ACM SIGPLAN Notices, 37(5):222– 233, May 2002. CODEN SINODQ.

ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Flanagan:2002:ESC

[573] Cormac Flanagan, K. Rustan M. Leino, Mark Lillibridge, Greg Nelson, James B. Saxe, and Raymie Stata. Extended static checking for Java. ACM SIGPLAN Notices, 37(5): 234–245, May 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Leino:2002:UDG

[574] K. Rustan M. Leino, Arnd Poetzsch-Heffter, and Yunhong Zhou. Using data groups to specify and check side effects. ACM SIGPLAN Notices, 37 (5):246–257, May 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Choi:2002:EPD

[575] Jong-Deok Choi, Keunwoo Lee, Alexey Loginov, Robert O’Callahan, Vivek Sarkar, and Manu Sridharan. Efficient and precise datarace detection for multithreaded object-oriented programs. ACM SIGPLAN Notices, 37 (5):258–269, May 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:2002:MMD

[576] Jason Baker and Wilson C. Hsieh. Maya: multiple-dispatch syntax extension in Java. ACM SIGPLAN Notices, 37(5):270–281, May 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Grossman:2002:RBM

[577] Dan Grossman, Greg Morrisett, Trevor Jim, Michael Hicks, Yanling Wang, and James Cheney. Region-based memory management in cyclone. ACM SIG-PLAN Notices, 37(5):282–293, May 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Almasi:2002:MCM

[578] George Almási and David Padua. Ma-JIC: compiling MATLAB for speed and responsiveness. ACM SIGPLAN Notices, 37(5):294–303, May 2002. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Joshi:2002:DGD

[579] Rajeev Joshi, Greg Nelson, and Keith Randall. Denali: a goal-directed superoptimizer. ACM SIGPLAN Notices, 37 (5):304–314, May 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Henzinger:2002:EMP

[580] Thomas A. Henzinger and Christoph M. Kirsch. The embedded machine: predictable, portable real-time code. ACM SIGPLAN Notices, 37(5):315– 326, May 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:2002:FRF

[581] Paul Frenger. Forth report: Forth’s DOOM. ACM SIGPLAN Notices, 37 (6):14–17, June 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergin:2002:UOO

[582] Joseph Bergin and Russel Winder. Understanding object-oriented programming. ACM SIGPLAN Notices, 37 (6):18–25, June 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fenwick:2002:IAb

[583] Jay Fenwick and Cindy Norris. Information for authors. ACM SIG-PLAN Notices, 37(6):26, June 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fenwick:2002:IPAb

[584] Jay Fenwick and Cindy Norris. Information for past authors. ACM SIG-PLAN Notices, 37(6):27, June 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:2002:YAM

[585] Haiming Chen and Yunmei Dong. Yet another meta-language for programming language processing. ACM SIG-PLAN Notices, 37(6):28–37, June 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Guilan:2002:RCC

[586] Dai Guilan, Tian Jinlan, Zhang Suqin, Jiang Weidu, and Dai Jun. Retargetable cross compilation techniques: comparison and analysis of GCC and Zephyr. ACM SIGPLAN Notices, 37 (6):38–44, June 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Guilan:2002:SCT

[587] Dai Guilan, Zhang Suqing, Tian Jinlan, and Jiang Weidu. A study of compiler techniques for multiple targets in compiler infrastructures. ACM SIGPLAN Notices, 37(6):45–51, June 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kiselyov:2002:SS

[588] Oleg Kiselyov. SXML specification. ACM SIGPLAN Notices, 37(6):52–58, June 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cockshott:2002:VPR

[589] Paul Cockshott. Vector Pascal reference manual. ACM SIGPLAN Notices, 37(6):59–81, June 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lee:2002:AOI

[590] Ji-Hyun Lee, Cheol-Jung Yoo, and Ok-Bae Chang. Analysis of object interaction during the enterprise JavaBeans lifecycle using formal specification technique. ACM SIGPLAN Notices, 37 (6):82–92, June 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Magarshack:2002:SCN

[591] Philippe Magarshack. Systems-on-chip needs for embedded software development: an industrial perspective. ACM SIGPLAN Notices, 37(7):1, July 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Saputra:2002:ECC

[592] H. Saputra, M. Kandemir, N. Vijaykrishnan, M. J. Irwin, J. S. Hu, C-H. Hsu, and U. Kremer. Energy-conscious compilation based on voltage scaling. ACM SIGPLAN Notices, 37(7):2–11, July 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Oh:2002:FRD

[593] Hyunok Oh and Soonhoi Ha. Fractional rate dataflow model and efficient code synthesis for multimedia applications. ACM SIGPLAN Notices, 37(7):12–17, July 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mohanty:2002:RDS

[594] S. Mohanty, V. K. Prasanna, S. Neema, and J. Davis. Rapid design space exploration of heterogeneous embedded systems using symbolic search and multi-granular simulation. ACM SIG-PLAN Notices, 37(7):18–27, July 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Palem:2002:DSO

[595] Krishna V. Palem, Rodric M. Rabbah, Vincent J. Mooney III, Pinar Korkmaz, and Kiran Puttaswamy. Design space optimization of embedded memory systems via data remapping. ACM SIGPLAN Notices, 37(7):28–37, July 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hunleth:2002:FFM

[596] Frank Hunleth and Ron K. Cytron. Footprint and feature management using aspect-oriented programming techniques. ACM SIGPLAN Notices, 37 (7):38–45, July 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kastner:2002:GCF

[597] Daniel Kästner and Stephan Wilhelm. Generic control flow reconstruction from assembly code. ACM SIG-PLAN Notices, 37(7):46–55, July 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Krishnaswamy:2002:PGS

[598] Arvind Krishnaswamy and Rajiv Gupta. Profile guided selection of ARM and thumb instructions. ACM SIG-PLAN Notices, 37(7):56–64, July 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wakabayashi:2002:SAI

[599] Takayuki Wakabayashi and Hiroaki Takada. Standardization approach of ITRON debugging interface specification and evaluation of its adaptability. ACM SIGPLAN Notices, 37(7):65–74, July 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jhumka:2002:SDG

[600] Arshad Jhumka, Martin Hiller, Vilgot Claesson, and Neeraj Suri. On systematic design of globally consistent executable assertions in embedded software. ACM SIGPLAN Notices, 37 (7):75–84, July 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Feng:2002:AFV

[601] Xiushan Feng and Alan J. Hu. Automatic formal verification for scheduled VLIW code. ACM SIGPLAN Notices, 37(7):85–92, July 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fettweis:2002:DWD

[602] Gerhard Fettweis. DSPs: why don’t they just go away! ACM SIG-PLAN Notices, 37(7):93, July 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lorenz:2002:EAC

[603] Markus Lorenz, Lars Wehmeyer, and Thorsten Dräger. Energy aware compilation for DSPs with SIMD instructions. ACM SIGPLAN Notices, 37 (7):94–101, July 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kessler:2002:OIC

[604] Christoph Kessler and Andrzej Bednarski. Optimal integrated code generation for clustered VLIW architectures. ACM SIGPLAN Notices, 37 (7):102–111, July 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Qian:2002:LFC

[605] Yi Qian, Steve Carr, and Philip Sweany. Loop fusion for clustered VLIW architectures. ACM SIGPLAN Notices, 37(7):112–119, July 2002. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Naik:2002:CCS

[606] Mayur Naik and Jens Palsberg. Compiling with code-size constraints. ACM SIGPLAN Notices, 37(7):120–129, July 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cho:2002:ERM

[607] Jeonghun Cho, Yunheung Paek, and David Whalley. Efficient register and memory assignment for non-orthogonal architectures via graph coloring and MST algorithms. ACM SIGPLAN Notices, 37(7):130–138, July 2002. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Scholz:2002:RAI

[608] Bernhard Scholz and Erik Eckstein. Register allocation for irregular architectures. ACM SIGPLAN Notices, 37 (7):139–148, July 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Barthelmann:2002:ITR

[609] Volker Barthelmann. Inter-task register-allocation for static operating systems. ACM SIGPLAN Notices, 37 (7):149–154, July 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhao:2002:VSI

[610] Wankang Zhao, Baosheng Cai, David Whalley, Mark W. Bailey, Robert van Engelen, Xin Yuan, Jason D. Hiser, Jack W. Davidson, Kyle Gallivan, and Douglas L. Jones. VISTA: a system for interactive code improvement. ACM SIGPLAN Notices, 37(7):155–164, July 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hu:2002:CDC

[611] J. S. Hu, M. Kandemir, N. Vijaykrishnan, M. J. Irwin, H. Saputra, and W. Zhang. Compiler-directed cache polymorphism. ACM SIGPLAN Notices, 37(7):165–174, July 2002. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Milner:2002:QPF

[612] Christopher W. Milner and Jack W. Davidson. Quick piping: a fast, highlevel model for describing processor pipelines. ACM SIGPLAN Notices, 37 (7):175–184, July 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wahlen:2002:ASC

[613] Oliver Wahlen, Tilman Glökler, Achim Nohl, Andreas Hoffmann, Rainer Leupers, and Heinrich Meyr. Application specific compiler/architecture codesign: a case study. ACM SIG-PLAN Notices, 37(7):185–193, July 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Palm:2002:WUC

[614] Jeffrey Palm, Han Lee, Amer Diwan, and J. Eliot B. Moss. When to use a compilation service? ACM SIG-PLAN Notices, 37(7):194–203, July 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Newsome:2002:PCD

[615] Matt Newsome and Des Watson. Proxy compilation of dynamically loaded Java classes with MoJo. ACM SIGPLAN Notices, 37(7):204–212, July 2002. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dudani:2002:ECF

[616] Ajay Dudani, Frank Mueller, and Yifan Zhu. Energy-conserving feedback EDF scheduling for embedded systems with real-time constraints. ACM SIG-PLAN Notices, 37(7):213–222, July 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kim:2002:PPT

[617] Saehwa Kim, Seongsoo Hong, and Tae-Hyung Kim. Perfecting preemption threshold scheduling for objectoriented real-time system design: from the perspective of real-time synchronization. ACM SIGPLAN Notices, 37 (7):223–232, July 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:2002:FRQ

[618] Paul Frenger. Forth report: Quartus Forth for the Palm Pilot. ACM SIGPLAN Notices, 37(8):6–8, August 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clark:2002:PPP

[619] Chris Clark. Practical parsing patterns: Conflicts. ACM SIGPLAN Notices, 37(8):9–14, August 2002. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Christopher:2002:PHS

[620] Thomas W. Christopher. In parallel: HPC is a strange land. ACM SIG-PLAN Notices, 37(8):15–16, August 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fenwick:2002:IAc

[621] Jay Fenwick and Cindy Norris. Information for authors. ACM SIG-PLAN Notices, 37(8):17, August 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fenwick:2002:IPAc

[622] Jay Fenwick and Cindy Norris. Information for past authors. ACM SIG-PLAN Notices, 37(8):18, August 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hughes:2002:HMT

[623] Elliott Hughes. How many trivial getter methods does Java have? ACM SIGPLAN Notices, 37(8):19–24, August 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jeon:2002:EBT

[624] Taewoong Jeon, Hyon Woo Seung, and Sungyoung Lee. Embedding builtin tests in hot spots of an objectoriented framework. ACM SIGPLAN Notices, 37(8):25–34, August 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ortin:2002:IRC

[625] Francisco Ortín and Juan Manuel Cueva. Implementing a real computationalenvironment jump in order to develop a runtime-adaptable reflective platform. ACM SIGPLAN Notices, 37(8):35–44, August 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:2002:OMD

[626] Zhenqiang Chen, Baowen Xu, and Jianjun Zhao. An overview of methods for dependence analysis of concurrent programs. ACM SIGPLAN Notices, 37 (8):45–52, August 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dhamdhere:2002:PPR

[627] Dhananjay M. Dhamdhere. Epath PRE: partial redundancy elimination made easy. ACM SIGPLAN Notices, 37(8):53–65, August 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Doh:2002:DIL

[628] Kyung Goo Doh and Seung Cheol Shin. Detection of information leak by data flow analysis. ACM SIG-PLAN Notices, 37(8):66–71, August 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Myers:2002:TMN

[629] Brad A. Myers. Towards more natural functional programming languages. ACM SIGPLAN Notices, 37 (9):1, September 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hinze:2002:BOS

[630] Ralf Hinze. Bootstrapping one-sided flexible arrays. ACM SIGPLAN Notices, 37(9):2–13, September 2002. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Voigtlander:2002:CRM

[631] Janis Voigtländer. Concatenate, reverse and map vanish for free. ACM SIGPLAN Notices, 37(9):14– 25, September 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Carlsson:2002:MIC

[632] Magnus Carlsson. Monads for incremental computing. ACM SIGPLAN Notices, 37(9):26–35, September 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ford:2002:PPS

[633] Bryan Ford. Packrat parsing: simple, powerful, lazy, linear time, functional pearl. ACM SIGPLAN Notices, 37(9):

36–47, September 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Findler:2002:CHO

[634] Robert Bruce Findler and Matthias Felleisen. Contracts for higher-order functions. ACM SIGPLAN Notices, 37 (9):48–59, September 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ohori:2002:ICE

[635] Atsushi Ohori and Kiyoshi Yamatodani. An interoperable calculus for external object access. ACM SIGPLAN Notices, 37(9):60–71, September 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Flatt:2002:CCM

[636] Matthew Flatt. Composable and compilable macros: you want it when? ACM SIGPLAN Notices, 37(9):72– 83, September 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dube:2002:DDA

[637] Danny Dubé and Marc Feeley. A demand-driven adaptive type analysis. ACM SIGPLAN Notices, 37(9): 84–97, September 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Glynn:2002:EAN

[638] Kevin Glynn, Peter J. Stuckey, Martin Sulzmann, and Harald Søndergaard. Exception analysis for non-strict languages. ACM SIGPLAN Notices, 37 (9):98–109, September 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Millstein:2002:MTH

[639] Todd Millstein, Colin Bleckner, and Craig Chambers. Modular typechecking for hierarchically extensible datatypes and functions. ACM SIG-PLAN Notices, 37(9):110–122, September 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Moore:2002:FFM

[640] J. Strother Moore. Functional formal methods. ACM SIGPLAN Notices, 37 (9):123, September 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Svenningsson:2002:SFA

[641] Josef Svenningsson. Shortcut fusion for accumulating parameters & ziplike functions. ACM SIGPLAN Notices, 37(9):124–132, September 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Luth:2002:CMU

[642] Christoph Lüth and Neil Ghani. Composing monads using coproducts. ACM SIGPLAN Notices, 37(9):133– 144, September 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hanna:2002:IVF

[643] Keith Hanna. Interactive visual functional programming. ACM SIG-PLAN Notices, 37(9):145–156, September 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baars:2002:TDT

[644] Arthur I. Baars and S. Doaitse Swierstra. Typing dynamic typing. ACM SIGPLAN Notices, 37(9):157– 166, September 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stuckey:2002:TO

[645] Peter J. Stuckey and Martin Sulzmann. A theory of overloading. ACM SIG-PLAN Notices, 37(9):167–178, September 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Neubauer:2002:TCM

[646] Matthias Neubauer and Peter Thiemann. Type classes with more higherorder polymorphism. ACM SIG-PLAN Notices, 37(9):179–190, September 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Crary:2002:EST

[647] Karl Crary and Joseph C. Vanderwaart. An expressive, scalable type theory for certified code. ACM SIG-PLAN Notices, 37(9):191–205, September 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nanevski:2002:MPN

[648] Aleksandar Nanevski. Meta-programming with names and necessity. ACM SIG-PLAN Notices, 37(9):206–217, September 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pasalic:2002:TSI

[649] Emir Pašalić, Walid Taha, and Tim Sheard. Tagless staged interpreters for typed languages. ACM SIGPLAN Notices, 37(9):218–229, September 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Danvy:2002:TBA

[650] Olivier Danvy and Mayer Goldberg. There and back again. ACM SIG-PLAN Notices, 37(9):230–234, September 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gregoire:2002:CIS

[651] Benjamin Grégoire and Xavier Leroy. A compiled implementation of strong reduction. ACM SIGPLAN Notices, 37 (9):235–246, September 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hansen:2002:ESR

[652] Lars T. Hansen and William D. Clinger. An experimental study of renewal-older-first garbage collection. ACM SIGPLAN Notices, 37(9):247– 258, September 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Serpette:2002:CSJ

[653] Bernard Paul Serpette and Manuel Serrano. Compiling scheme to JVM bytecode: a performance study. ACM SIG-PLAN Notices, 37(9):259–270, September 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gasbichler:2002:FSC

[654] Martin Gasbichler and Michael Sperber. Final shift for call/cc: direct implementation of shift and reset. ACM SIGPLAN Notices, 37(9):271– 282, September 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jones:2002:PGT

[655] Neil D. Jones and Arne J. Glenstrup. Program generation, termination, and binding-time analysis. ACM SIGPLAN Notices, 37(9):283, September 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Estrin:2002:SNR

[656] Deborah Estrin. Sensor network research: Emerging challenges for architecture, systems, and languages. ACM SIGPLAN Notices, 37(10):1, October 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rajwar:2002:TLF

[657] Ravi Rajwar and James R. Goodman. Transactional lock-free execution of lock-based programs. ACM SIG-PLAN Notices, 37(10):5–17, October 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Martinez:2002:SSA

[658] José F. Martínez and Josep Torrellas. Speculative synchronization: applying thread-level speculation to explicitly parallel applications. ACM SIGPLAN Notices, 37(10):18–29, October 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lepak:2002:TSS

[659] Kevin M. Lepak and Mikko H. Lipasti. Temporally silent stores. ACM SIG-PLAN Notices, 37(10):30–41, October 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sherwood:2002:ACL

[660] Timothy Sherwood, Erez Perelman, Greg Hamerly, and Brad Calder. Automatically characterizing large scale program behavior. ACM SIGPLAN Notices, 37(10):45–57, October 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ogata:2002:BFO

[661] Kazunori Ogata, Hideaki Komatsu, and Toshio Nakatani. Bytecode fetch optimization for a Java interpreter. ACM SIGPLAN Notices, 37(10):58– 67, October 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Li:2002:UIO

[662] Tao Li, Lizy Kurian John, Anand Sivasubramaniam, N. Vijaykrishnan, and Juan Rubio. Understanding and improving operating system effects in control flow prediction. ACM SIGPLAN Notices, 37(10):68–80, October 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Levis:2002:MTV

[663] Philip Levis and David Culler. Maté: a tiny virtual machine for sensor networks. ACM SIGPLAN Notices, 37 (10):85–95, October 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Juang:2002:EEC

[664] Philo Juang, Hidekazu Oki, Yong Wang, Margaret Martonosi, Li Shiuan Peh, and Daniel Rubenstein. Energyefficient computing for wildlife tracking: design tradeoffs and early experiences with ZebraNet. ACM SIGPLAN Notices, 37(10):96–107, October 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kirovski:2002:ETS

[665] Darko Kirovski, Milenko Drinić, and Miodrag Potkonjak. Enabling trusted software integrity. ACM SIGPLAN Notices, 37(10):108–120, October 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zeng:2002:EME

[666] Heng Zeng, Carla S. Ellis, Alvin R. Lebeck, and Amin Vahdat. ECOSystem: managing energy as a first class operating system resource. ACM SIG-PLAN Notices, 37(10):123–132, October 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ashok:2002:CMC

[667] Raksit Ashok, Saurabh Chheda, and Csaba Andras Moritz. Cool-Mem: combining statically speculative memory accessing with selective address translation for energy efficiency. ACM SIGPLAN Notices, 37(10):133–143, October 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sasanka:2002:JLG

[668] Ruchira Sasanka, Christopher J. Hughes, and Sarita V. Adve. Joint local and global hardware adaptations for energy. ACM SIGPLAN Notices, 37(10):144–155, October 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kim:2002:DEC

[669] Dongkeun Kim and Donald Yeung. Design and evaluation of compiler algorithms for pre-execution. ACM SIG-PLAN Notices, 37(10):159–170, October 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhai:2002:COS

[670] Antonia Zhai, Christopher B. Colohan, J. Gregory Steffan, and Todd C. Mowry. Compiler optimization of scalar value communication between speculative threads. ACM SIGPLAN Notices, 37(10):171–183, October 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Oplinger:2002:ESR

[671] Jeffrey Oplinger and Monica S. Lam. Enhancing software reliability with speculative threads. ACM SIGPLAN Notices, 37(10):184–196, October 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Butts:2002:DDI

[672] J. Adam Butts and Guri Sohi. Dynamic dead-instruction detection and elimination. ACM SIGPLAN Notices, 37(10): 199–210, October 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kim:2002:ANU

[673] Changkyu Kim, Doug Burger, and Stephen W. Keckler. An adaptive, nonuniform cache structure for wire-delay dominated on-chip caches. ACM SIG-PLAN Notices, 37(10):211–222, October 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mukherjee:2002:CSA

[674] Shubhendu S. Mukherjee, Federico Silla, Peter Bannon, Joel Emer, Steve Lang, and David Webb. A comparative study of arbitration algorithms for the Alpha 21364 pipelined router. ACM SIGPLAN Notices, 37(10):223– 234, October 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kim:2002:IWS

[675] Hyong youb Kim, Vijay S. Pai, and Scott Rixner. Increasing Web server throughput with network interface data caching. ACM SIGPLAN Notices, 37 (10):239–250, October 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kohler:2002:PLO

[676] Eddie Kohler, Robert Morris, and Benjie Chen. Programming language optimizations for modular router configurations. ACM SIGPLAN Notices, 37(10): 251–263, October 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sivathanu:2002:ERA

[677] Muthian Sivathanu, Andrea C. Arpaci-Dusseau, and Remzi H. Arpaci-Dusseau. Evolving RPC for active storage. ACM SIGPLAN Notices, 37(10): 264–276, October 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cooksey:2002:SCD

[678] Robert Cooksey, Stephan Jourdan, and Dirk Grunwald. A stateless, contentdirected data prefetching mechanism. ACM SIGPLAN Notices, 37(10):279– 290, October 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gordon:2002:SCC

[679] Michael I. Gordon, William Thies, Michal Karczmarek, Jasper Lin, Ali S. Meli, Andrew A. Lamb, Chris Leger, Jeremy Wong, Henry Hoffmann, David Maze, and Saman Amarasinghe. A stream compiler for communicationexposed architectures. ACM SIGPLAN Notices, 37(10):291–303, October 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Witchel:2002:MMP

[680] Emmett Witchel, Josh Cates, and Krste Asanović. Mondrian memory protection. ACM SIGPLAN Notices, 37(10):304–316, October 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berger:2002:RCM

[681] Emery D. Berger, Benjamin G. Zorn, and Kathryn S. McKinley. Reconsidering custom memory allocation. ACM SIGPLAN Notices, 37(11):1– 12, November 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shuf:2002:CPL

[682] Yefim Shuf, Manish Gupta, Hubertus Franke, Andrew Appel, and Jaswinder Pal Singh. Creating and preserving locality of Java applications at allocation and garbage collection times. ACM SIGPLAN Notices, 37(11):13– 25, November 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bjornsson:2002:BHP

[683] Magnus E. Bjornsson and Liuba Shrira. BuddyCache: high-performance object storage for collaborative strongconsistency applications in a WAN. ACM SIGPLAN Notices, 37(11):26– 39, November 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Manolescu:2002:WEC

[684] Dragos A. Manolescu. Workflow enactment with continuation and future objects. ACM SIGPLAN Notices, 37(11): 40–51, November 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mezini:2002:IIC

[685] Mira Mezini and Klaus Ostermann. Integrating independent components with on-demand remodularization. ACM SIGPLAN Notices, 37(11):52– 67, November 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Killijian:2002:PSC

[686] Marc-Olivier Killijian, Juan-Carlos Ruiz, and Jean-Charles Fabre. Portable serialization of CORBA objects: a reflective approach. ACM SIGPLAN Notices, 37(11):68–82, November 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marelly:2002:MIS

[687] Rami Marelly, David Harel, and Hillel Kugler. Multiple instances and symbolic variables in executable sequence charts. ACM SIGPLAN Notices, 37(11):83–100, November 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Michel:2002:CBA

[688] Laurent Michel and Pascal Van Hentenryck. A constraint-based architecture for local search. ACM SIGPLAN Notices, 37(11):101–110, November 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Arnold:2002:OFD

[689] Matthew Arnold, Michael Hind, and Barbara G. Ryder. Online feedbackdirected optimization of Java. ACM SIGPLAN Notices, 37(11):111–129, November 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kawachiya:2002:LRJ

[690] Kiyokuni Kawachiya, Akira Koseki, and Tamiya Onodera. Lock reservation: Java locks can mostly do without atomic operations. ACM SIGPLAN Notices, 37(11):130–141, November 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zibin:2002:FAC

[691] Yoav Zibin and Joseph Yossi Gil. Fast algorithm for creating space efficient dispatching tables with application to multi-dispatching. ACM SIGPLAN Notices, 37(11):142–160, November 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hannemann:2002:DPI

[692] Jan Hannemann and Gregor Kiczales. Design pattern implementation in Java and aspectJ. ACM SIGPLAN Notices, 37(11):161–173, November 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Soares:2002:IDP

[693] Sergio Soares, Eduardo Laureano, and Paulo Borba. Implementing distribution and persistence aspects with aspectJ. ACM SIGPLAN Notices, 37 (11):174–190, November 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zee:2002:WBRb

[694] Karen Zee and Martin Rinard. Write barrier removal by static analysis. ACM SIGPLAN Notices, 37(11):191– 210, November 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boyapati:2002:OTS

[695] Chandrasekhar Boyapati, Robert Lee, and Martin Rinard. Ownership types for safe programming: preventing data races and deadlocks. ACM SIGPLAN Notices, 37(11):211–230, November 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Khurshid:2002:AAL

[696] Sarfraz Khurshid, Darko Marinov, and Daniel Jackson. An analyzable annotation language. ACM SIGPLAN Notices, 37(11):231–245, November 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cecchet:2002:PSE

[697] Emmanuel Cecchet, Julie Marguerite, and Willy Zwaenepoel. Performance and scalability of EJB applications.

ACM SIGPLAN Notices, 37(11):246– 261, November 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Palacz:2002:IEC

[698] Krzysztof Palacz, Jan Vitek, Grzegorz Czajkowski, and Laurent Daynas. Incommunicado: efficient communication for isolates. ACM SIGPLAN Notices, 37(11):262–274, November 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

DeSutter:2002:SML

[699] Bjorn De Sutter, Bruno De Bus, and Koen De Bosschere. Sifting out the mud: low level C++ code reuse. ACM SIGPLAN Notices, 37(11):275– 291, November 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clarke:2002:OED

[700] Dave Clarke and Sophia Drossopoulou. Ownership, encapsulation and the disjointness of type and effect. ACM SIGPLAN Notices, 37(11):292–310, November 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Aldrich:2002:AAP

[701] Jonathan Aldrich, Valentin Kostadinov, and Craig Chambers. Alias annotations for program understanding. ACM SIGPLAN Notices, 37(11):311– 330, November 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pucella:2002:TFC

[702] Riccardo Pucella. Towards a formalization for COM. Part I: The primitive calculus. ACM SIGPLAN Notices, 37(11):331–342, November 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Printezis:2002:GAH

[703] Tony Printezis and Richard Jones. GCspy: an adaptable heap visualisation framework. ACM SIGPLAN Notices, 37(11):343–358, November 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Koved:2002:ARA

[704] Larry Koved, Marco Pistoia, and Aaron Kershenbaum. Access rights analysis for Java. ACM SIGPLAN Notices, 37(11):359–372, November 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dmitriev:2002:LSM

[705] Mikhail Dmitriev. Language-specific make technology for the Java programming language. ACM SIGPLAN Notices, 37(11):373–385, November 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:2002:FRG

[706] Paul Frenger. Forth report: Going APE.FORTH to Mars. ACM SIG-PLAN Notices, 37(12):9–13, December 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Creak:2002:HTWc

[707] Alan Creak. How things were: Edsger W. Dijkstra. ACM SIGPLAN Notices, 37(12):14–16, December 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Norris:2002:ISI

[708] Cindy Norris. Introduction to special issue. ACM SIGPLAN Notices, 37 (12):17, December 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Arts:2002:TAE

[709] Thomas Arts and Lars-Åke Fredlund. Trace analysis of Erlang programs. ACM SIGPLAN Notices, 37(12):18– 24, December 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wiger:2002:WCP

[710] Ulf Wiger, Gösta Ask, and Kent Boortz. World-class product certification using Erlang. ACM SIGPLAN Notices, 37(12):25–34, December 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Claessen:2002:ELA

[711] Koen Claessen and Gordon J. Pace. An embedded language approach to teaching hardware compilation. ACM SIG-PLAN Notices, 37(12):35–46, December 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Claessen:2002:TMC

[712] Koen Claessen and John Hughes. Testing monadic code with QuickCheck. ACM SIGPLAN Notices, 37(12):47– 59, December 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sheard:2002:TMP

[713] Tim Sheard and Simon Peyton Jones. Template meta-programming for Haskell. ACM SIGPLAN Notices, 37(12):60–75, December 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Giavitto:2002:PMR

[714] Jean-Louis Giavitto, Olivier Michel, and Julien Cohen. Pattern-matching and rewriting rules for group indexed data structures. ACM SIGPLAN Notices, 37(12):76–87, December 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Erwig:2002:RBL

[715] Martin Erwig and Delin Ren. A rulebased language for programming software updates. ACM SIGPLAN Notices, 37(12):88–97, December 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Groote:2002:CAM

[716] Jan Friso Groote and Bert Lisser. Computer assisted manipulation of algebraic process specifications. ACM SIG-PLAN Notices, 37(12):98–107, December 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hornos:2002:FMC

[717] Miguel J. Hornos and Manuel I. Capel. On-the-fly model checking from interval logic specifications. ACM SIG-PLAN Notices, 37(12):108–119, December 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Simeon:2003:EX

[718] Jérôme Siméon and Philip Wadler. The essence of XML. ACM SIG-PLAN Notices, 38(1):1–13, January 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Acar:2003:SM

[719] Umut A. A. Acar, Guy E. Blelloch, and Robert Harper. Selective memoization. ACM SIGPLAN Notices, 38 (1):14–25, January 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Taha:2003:EC

[720] Walid Taha and Michael Florentin Nielsen. Environment classifiers. ACM SIGPLAN Notices, 38(1):26–37, January 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jensen:2003:BT

[721] Ole Høgh Jensen and Robin Milner. Bigraphs and transitions. ACM SIG-PLAN Notices, 38(1):38–49, January 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schmitt:2003:CHO

[722] Alan Schmitt and Jean-Bernard Stefani. The m-calculus: a higher-order distributed process calculus. ACM SIGPLAN Notices, 38(1):50–61, January 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bouajjani:2003:GAS

[723] Ahmed Bouajjani, Javier Esparza, and Tayssir Touili. A generic approach to the static analysis of concurrent programs with procedures. ACM SIG-PLAN Notices, 38(1):62–73, January 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gulwani:2003:DAE

[724] Sumit Gulwani and George C. Necula. Discovering affine equalities using random interpretation. ACM SIG-PLAN Notices, 38(1):74–84, January 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tallam:2003:BAG

[725] Sriraman Tallam and Rajiv Gupta. Bitwidth aware global register allocation. ACM SIGPLAN Notices, 38(1): 85–96, January 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ball:2003:SCL

[726] Thomas Ball, Mayur Naik, and Sriram K. Rajamani. From symptom to cause: localizing errors in counterexample traces. ACM SIGPLAN Notices, 38(1):97–105, January 2003.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Carter:2003:FCR

[727] Larry Carter, Jeanne Ferrante, and Clark Thomborson. Folklore confirmed: reducible flow graphs are exponentially larger. ACM SIGPLAN Notices, 38(1):106–114, January 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chakaravarthy:2003:NRC

[728] Venkatesan T. Chakaravarthy. New results on the computability and complexity of points–to analysis. ACM SIGPLAN Notices, 38(1):115–125, January 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zibin:2003:IAD

[729] Yoav Zibin and Joseph (Yossi) Gil. Incremental algorithms for dispatching in dynamically typed languages. ACM SIGPLAN Notices, 38(1):126–138, January 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Thielecke:2003:CET

[730] Hayo Thielecke. From control effects to typed continuation passing. ACM SIG-PLAN Notices, 38(1):139–149, January 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:2003:CSC

[731] Gang Chen. Coercive subtyping for the calculus of constructions. ACM SIG-PLAN Notices, 38(1):150–159, January 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zibin:2003:EAI

[732] Yoav Zibin, Joseph (Yossi) Gil, and Jeffrey Considine. Efficient algorithms for isomorphisms of simple types. ACM SIGPLAN Notices, 38(1):160–171, January 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Petersen:2003:TTM

[733] Leaf Petersen, Robert Harper, Karl Crary, and Frank Pfenning. A type theory for memory allocation and data layout. ACM SIGPLAN Notices, 38(1): 172–184, January 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hofmann:2003:SPH

[734] Martin Hofmann and Steffen Jost. Static prediction of heap space usage for first-order functional programs. ACM SIGPLAN Notices, 38(1):185– 197, January 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Crary:2003:TFT

[735] Karl Crary. Toward a foundational typed assembly language. ACM SIG-PLAN Notices, 38(1):198–212, January 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boyapati:2003:OTO

[736] Chandrasekhar Boyapati, Barbara Liskov, and Liuba Shrira. Ownership types for object encapsulation. ACM SIGPLAN Notices, 38(1):213–223, January 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xi:2003:GRD

[737] Hongwei Xi, Chiyan Chen, and Gang Chen. Guarded recursive datatype constructors. ACM SIGPLAN Notices, 38(1):224–235, January 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dreyer:2003:TSH

[738] Derek Dreyer, Karl Crary, and Robert Harper. A type system for higherorder modules. ACM SIGPLAN Notices, 38(1):236–249, January 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Barthe:2003:PPT

[739] Gilles Barthe, Horatiu Cirstea, Claude Kirchner, and Luigi Liquori. Pure patterns type systems. ACM SIGPLAN Notices, 38(1):250–261, January 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boehm:2003:DFS

[740] Hans-J. Boehm. Destructors, finalizers, and synchronization. ACM SIG-PLAN Notices, 38(1):262–272, January 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gheorghioiu:2003:ICA

[741] Ovidiu Gheorghioiu, Alexandru Salcianu, and Martin Rinard. Interprocedural compatibility analysis for static object preallocation. ACM SIGPLAN Notices, 38(1):273–284, January 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bacon:2003:RTG

[742] David F. Bacon, Perry Cheng, and V. T. Rajan. A real-time garbage collector with low overhead and consistent utilization. ACM SIGPLAN Notices, 38(1):285–298, January 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ding:2003:CDR

[743] Chen Ding and Yutao Zhong. Compilerdirected run-time monitoring of program data access. ACM SIGPLAN Notices, 38(2s):1–12, February 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Guilan:2003:AIR

[744] Dai Guilan, Tian Jinlan, Zhang Suqing, and Jiang Weidu. An abstract intermediate representation in compilation systems. ACM SIGPLAN Notices, 38(2): 12–18, February 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lattner:2003:APA

[745] Chris Lattner and Vikram Adve. Automatic pool allocation for disjoint data structures. ACM SIGPLAN Notices, 38(2s):13–24, February 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hamilton:2003:LIC

[746] Jennifer Hamilton. Language integration in the common language runtime. ACM SIGPLAN Notices, 38(2): 19–28, February 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stefanovic:2003:OFG

[747] Darko Stefanović, Matthew Hertz, Stephen M. Blackburn, Kathryn S. McKinley, and J. Eliot B. Moss. Olderfirst garbage collection in practice: evaluation in a Java Virtual Machine. ACM SIGPLAN Notices, 38(2s):25– 36, February 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jackson:2003:EFO

[748] Quinn Tyler Jackson. Efficient formalism-only parsing of XML/ HTML using the §-calculus. ACM SIG-PLAN Notices, 38(2):29–35, February 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schilling:2003:SHM

[749] Jonathan L. Schilling. The simplest heuristics may be the best in Java JIT compilers. ACM SIGPLAN Notices, 38(2):36–46, February 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Almasi:2003:CSD

[750] George Almási, Călin Caşcaval, and David A. Padua. Calculating stack distances efficiently. ACM SIGPLAN Notices, 38(2s):37–43, February 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nethercote:2003:CBL

[751] Nicholas Nethercote and Alan Mycroft. The cache behaviour of large lazy functional programs on stock hardware. ACM SIGPLAN Notices, 38(2s): 44–55, February 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sinclair:2003:IMT

[752] Robert Sinclair. Interactive mathematics textbooks. ACM SIGPLAN Notices, 38(2):47–56, February 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lopes:2003:SPC

[753] Ricardo Lopes, Luís Fernando Castro, and Vítor Santos Costa. From simulation to practice: cache performance study of a Prolog system. ACM SIG-PLAN Notices, 38(2s):56–64, February 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Queinnec:2003:IBI

[754] Christian Queinnec. Inverting back the inversion of control or, continuations versus page-centric programming. ACM SIGPLAN Notices, 38(2):

57–64, February 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ykman-Couvreur:2003:MOA

[755] Ch. Ykman-Couvreur, J. Lambrecht, A. van der Togt, and F. Catthoor. Multi-objective abstract data type refinement for mapping tables in telecom network applications. ACM SIG-PLAN Notices, 38(2s):65–75, February 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schulz:2003:PNH

[756] Martin Schulz, Jie Tao, Jürgen Jeitner, and Wolfgang Karl. A proposal for a new hardware cache monitoring architecture. ACM SIGPLAN Notices, 38(2s):76–85, February 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mahapatra:2003:PAA

[757] Nihar R. Mahapatra, Jiangjiang Liu, and Krishnan Sundaresan. The performance advantage of applying compression to the memory system. ACM SIG-PLAN Notices, 38(2s):86–96, February 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cooper:2003:ESA

[758] Keith D. Cooper and Li Xu. An efficient static analysis algorithm to detect redundant memory operations. ACM SIGPLAN Notices, 38(2s):97– 107, February 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Robertz:2003:APM

[759] Sven G. Robertz. Applying priorities to memory allocation. ACM SIG-PLAN Notices, 38(2s):108–118, February 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Endo:2003:RPT

[760] Toshio Endo and Kenjiro Taura. Reducing pause time of conservative collectors. ACM SIGPLAN Notices, 38 (2s):119–131, February 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Deters:2003:ADS

[761] Morgan Deters and Ron K. Cytron. Automated discovery of scoped memory regions for real-time Java. ACM SIGPLAN Notices, 38(2s):132–142, February 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hirzel:2003:UCH

[762] Martin Hirzel, Johannes Henkel, Amer Diwan, and Michael Hind. Understanding the connectivity of heap objects. ACM SIGPLAN Notices, 38(2s):143– 156, February 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Printezis:2003:VTG

[763] Tony Printezis and Alex Garthwaite. Visualising the train garbage collector. ACM SIGPLAN Notices, 38(2s):157– 170, February 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shaham:2003:EIH

[764] Ran Shaham, Elliot K. Kolodner, and Mooly Sagiv. Estimating the impact of heap liveness information on space consumption in Java. ACM SIGPLAN Notices, 38(2s):171–182, February 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Domani:2003:TLH

[765] Tamar Domani, Gal Goldshtein, Elliot K. Kolodner, Ethan Lewis, Erez Petrank, and Dafna Sheinwald. Thread-local heaps for Java. ACM SIGPLAN Notices, 38(2s):183–194, February 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Johansson:2003:HAC

[766] Erik Johansson, Konstantinos Sagonas, and Jesper Wilhelmsson. Heap architectures for concurrent languages using message passing. ACM SIGPLAN Notices, 38(2s):195–206, February 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ben-Yitzhak:2003:API

[767] Ori Ben-Yitzhak, Irit Goft, Elliot K. Kolodner, Kean Kuiper, and Victor Leikehman. An algorithm for parallel incremental compaction. ACM SIG-PLAN Notices, 38(2s):207–212, February 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vardhan:2003:UPO

[768] Abhay Vardhan and Gul Agha. Using passive object garbage collection algorithms for garbage collection of active objects. ACM SIGPLAN Notices, 38(2s):213–220, February 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kaplan:2003:ACD

[769] Scott F. Kaplan, Lyle A. McGeoch, and Megan F. Cole. Adaptive caching for demand prepaging. ACM SIG-PLAN Notices, 38(2s):221–232, February 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Qian:2003:ARB

[770] Feng Qian and Laurie Hendren. An adaptive, region-based allocator for Java. ACM SIGPLAN Notices, 38(2s): 233–244, February 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kumar:2003:DMM

[771] Sanjeev Kumar and Kai Li. Dynamic memory management for programmable devices. ACM SIG-PLAN Notices, 38(2s):245–255, February 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Henderson:2003:AGC

[772] Fergus Henderson. Accurate garbage collection in an uncooperative environment. ACM SIGPLAN Notices, 38(2s): 256–262, February 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Aggarwal:2003:SCV

[773] Aneesh Aggarwal. Software caching vs. prefetching. ACM SIGPLAN Notices, 38(2s):263–268, February 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dice:2003:MLF

[774] Dave Dice and Alex Garthwaite. Mostly lock-free malloc. ACM SIG-PLAN Notices, 38(2s):269–280, February 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Blackburn:2003:PWB

[775] Stephen M. Blackburn and Kathryn S. McKinley. In or out?: putting write barriers in their place. ACM SIG-PLAN Notices, 38(2s):281–290, February 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Flanagan:2003:TA

[776] Cormac Flanagan and Shaz Qadeer. Types for atomicity. ACM SIGPLAN Notices, 38(3):1–12, March 2003. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Grossman:2003:TSM

[777] Dan Grossman. Type-safe multithreading in cyclone. ACM SIG-PLAN Notices, 38(3):13–25, March 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lammel:2003:SYB

[778] Ralf Lämmel and Simon Peyton Jones. Scrap your boilerplate: a practical design pattern for generic programming. ACM SIGPLAN Notices, 38(3):26–37, March 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Park:2003:CPL

[779] Sungwoo Park. A calculus for probabilistic languages. ACM SIGPLAN Notices, 38(3):38–49, March 2003. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kobayashi:2003:TRE

[780] Naoki Kobayashi. Time regions and effects for resource usage analysis. ACM SIGPLAN Notices, 38(3):50–61, March 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Calcagno:2003:DVS

[781] Cristiano Calcagno, Luca Cardelli, and Andrew D. Gordon. Deciding validity in a spatial logic for trees. ACM SIGPLAN Notices, 38(3):62–73, March 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ahmed:2003:LAS

[782] Amal Ahmed and David Walker. The logical approach to stack typing. ACM SIGPLAN Notices, 38(3):74–85, March 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Amtoft:2003:IAT

[783] Torben Amtoft and Robert Muller. Inferring annotated types for interprocedural register allocation with constructor flattening. ACM SIGPLAN Notices, 38(3):86–97, March 2003. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vanderwaart:2003:TCR

[784] Joseph C. Vanderwaart, Derek Dreyer, Leaf Petersen, Karl Crary, Robert Harper, and Perry Cheng. Typed compilation of recursive datatypes. ACM SIGPLAN Notices, 38(3):98–108, March 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vanderwaart:2003:TIG

[785] Joseph C. Vanderwaart and Karl Crary. A typed interface for garbage collection. ACM SIGPLAN Notices, 38(3):109–122, March 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Elsman:2003:GCS

[786] Martin Elsman. Garbage collection safety for region-based memory management. ACM SIGPLAN Notices, 38(3):123–134, March 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Creak:2003:EFO

[787] Alan Creak. Everything is Fortran, in its own way. ACM SIGPLAN Notices, 38(4):7–12, April 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:2003:EFW

[788] Paul Frenger. Evaluating Forth in the Windows environment. ACM SIGPLAN Notices, 38(4):13–15, April 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Das:2003:FIVa

[789] Dibyendu Das. Function inlining versus function cloning. ACM SIGPLAN Notices, 38(4):18–24, April 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

deOliveiraGuimaraes:2003:EBC

[790] José de Oliveira Guimarães. Experiences in building a compiler for an object-oriented language. ACM SIGPLAN Notices, 38(4):25–33, April 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kwon:2003:AJP

[791] Jagun Kwon, Andy Wellings, and Steve King. Assessment of the Java programming language for use in high integrity systems. ACM SIGPLAN Notices, 38 (4):34–46, April 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Luo:2003:FSD

[792] Lei Luo, Ming-Yuan Zhu, and Qing-Li Zhang. A formal semantic definition of DEVIL. ACM SIGPLAN Notices, 38 (4):47–56, April 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhuge:2003:KKG

[793] Hai Zhuge and Jie Liu. KGOL: a Knowledge Grid operating language. ACM SIGPLAN Notices, 38(4):57–66, April 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gay:2003:NLH

[794] David Gay, Philip Levis, Robert von Behren, Matt Welsh, Eric Brewer, and David Culler. The nesC language: A holistic approach to networked embedded systems. ACM SIGPLAN Notices, 38(5):1–11, May 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lamb:2003:LAO

[795] Andrew A. Lamb, William Thies, and Saman Amarasinghe. Linear analysis and optimization of stream programs. ACM SIGPLAN Notices, 38(5):12–25, May 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

George:2003:TIN

[796] Lal George and Matthias Blume. Taming the IXP network processor. ACM SIGPLAN Notices, 38(5):26–37, May 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hsu:2003:DIE

[797] Chung-Hsing Hsu and Ulrich Kremer. The design, implementation, and evaluation of a compiler algorithm for CPU energy reduction. ACM SIG-PLAN Notices, 38(5):38–48, May 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xie:2003:CTD

[798] Fen Xie, Margaret Martonosi, and Sharad Malik. Compile-time dynamic voltage scaling settings: opportunities and limits. ACM SIGPLAN Notices, 38(5):49–62, May 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yotov:2003:CEM

[799] Kamen Yotov, Xiaoming Li, Gang Ren, Michael Cibulskis, Gerald De-Jong, Maria Garzaran, David Padua, Keshav Pingali, Paul Stodghill, and Peng Wu. A comparison of empirical and model-driven optimization. ACM SIGPLAN Notices, 38(5):63–76, May 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stephenson:2003:MOI

[800] Mark Stephenson, Saman Amarasinghe, Martin Martin, and Una-May O’Reilly. Meta optimization: improving compiler heuristics with machine learning. ACM SIGPLAN Notices, 38 (5):77–90, May 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Strout:2003:CTC

[801] Michelle Mills Strout, Larry Carter, and Jeanne Ferrante. Compile-time composition of run-time data and iteration reorderings. ACM SIGPLAN Notices, 38(5):91–102, May 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berndl:2003:PAU

[802] Marc Berndl, Ondrej Lhoták, Feng Qian, Laurie Hendren, and Navindra Umanee. Points-to analysis using BDDs. ACM SIGPLAN Notices, 38 (5):103–114, May 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vonPraun:2003:SCA

[803] Christoph von Praun and Thomas R. Gross. Static conflict analysis for multi-threaded object-oriented programs. ACM SIGPLAN Notices, 38 (5):115–128, May 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Aiken:2003:CIL

[804] Alex Aiken, Jeffrey S. Foster, John Kodumal, and Tachio Terauchi. Checking and inferring local non-aliasing. ACM SIGPLAN Notices, 38(5):129– 140, May 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Liblit:2003:BIR

[805] Ben Liblit, Alex Aiken, Alice X. Zheng, and Michael I. Jordan. Bug isolation via remote program sampling. ACM SIGPLAN Notices, 38 (5):141–154, May 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dor:2003:CTR

[806] Nurit Dor, Michael Rodeh, and Mooly Sagiv. CSSV: towards a realistic tool for statically detecting all buffer overflows in C. ACM SIGPLAN Notices, 38 (5):155–167, May 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Heine:2003:PFS

[807] David L. Heine and Monica S. Lam. A practical flow-sensitive and contextsensitive C and C++ memory leak detector. ACM SIGPLAN Notices, 38 (5):168–181, May 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ammons:2003:DTS

[808] Glenn Ammons, David Mandelin, Rastislav Bodík, and James R. Larus. Debugging temporal specifications with concept analysis. ACM SIG-PLAN Notices, 38(5):182–195, May 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Blanchet:2003:SAL

[809] Bruno Blanchet, Patrick Cousot, Radhia Cousot, Jérome Feret, Laurent Mauborgne, Antoine Miné, David Monniaux, and Xavier Rival. A static analyzer for large safety-critical software. ACM SIGPLAN Notices, 38(5): 196–207, May 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:2003:PST

[810] Juan Chen, Dinghao Wu, Andrew W. Appel, and Hai Fang. A provably sound TAL for back-end optimization. ACM SIGPLAN Notices, 38(5): 208–219, May 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lerner:2003:APC

[811] Sorin Lerner, Todd Millstein, and Craig Chambers. Automatically proving the correctness of compiler optimizations. ACM SIGPLAN Notices, 38 (5):220–231, May 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Condit:2003:CRW

[812] Jeremy Condit, Matthew Harren, Scott McPeak, George C. Necula, and Westley Weimer. CCured in the real world. ACM SIGPLAN Notices, 38 (5):232–244, May 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ding:2003:PWP

[813] Chen Ding and Yutao Zhong. Predicting whole-program locality through reuse distance analysis. ACM SIG-PLAN Notices, 38(5):245–257, May 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Joisha:2003:SAS

[814] Pramod G. Joisha and Prithviraj Banerjee. Static array storage optimization in MATLAB. ACM SIG-PLAN Notices, 38(5):258–268, May 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Inagaki:2003:SPD

[815] Tatsushi Inagaki, Tamiya Onodera, Hideaki Komatsu, and Toshio Nakatani.

Stride prefetching by dynamically inspecting objects. ACM SIGPLAN Notices, 38(5):269–277, May 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ertl:2003:OIB

[816] M. Anton Ertl and David Gregg. Optimizing indirect branch prediction accuracy in virtual machine interpreters. ACM SIGPLAN Notices, 38(5):278– 288, May 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lin:2003:CFS

[817] Jin Lin, Tong Chen, Wei-Chung Hsu, Pen-Chung Yew, Roy Dz-Ching Ju, Tin-Fook Ngai, and Sun Chan. A compiler framework for speculative analysis and optimizations. ACM SIG-PLAN Notices, 38(5):289–299, May 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chu:2003:RBH

[818] Michael Chu, Kevin Fan, and Scott Mahlke. Region-based hierarchical operation partitioning for multicluster processors. ACM SIGPLAN Notices, 38(5):300–311, May 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Suganuma:2003:RBC

[819] Toshio Suganuma, Toshiaki Yasue, and Toshio Nakatani. A region-based compilation technique for a Java just-in-time compiler. ACM SIGPLAN Notices, 38(5):312–323, May 2003. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boyapati:2003:OTS

[820] Chandrasekhar Boyapati, Alexandru Salcianu, William Beebee, Jr., and Martin Rinard. Ownership types for safe region-based memory management in real-time Java. ACM SIG-PLAN Notices, 38(5):324–337, May 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Flanagan:2003:TES

[821] Cormac Flanagan and Shaz Qadeer. A type and effect system for atomicity. ACM SIGPLAN Notices, 38(5): 338–349, May 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Christopher:2003:SPS

[822] Thomas W. Christopher. A simple parallel system. ACM SIGPLAN Notices, 38(6):6–8, June 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergin:2003:SDC

[823] Joe Bergin, Achla Agarwal, and Krishna Agarwal. Some deficiencies of C++ in teaching CS1 and CS2. ACM SIGPLAN Notices, 38(6):9–13, June 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ryder:2003:IDE

[824] Barbara G. Ryder and Mary Lou Soffa. Influences on the design of exception handling: ACM SIGSOFT project on the impact of software engineering research on programming language design. ACM SIGPLAN Notices, 38 (6):16–22, June 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Das:2003:FIVb

[825] Dibyendu Das. Function inlining versus function cloning. ACM SIG-PLAN Notices, 38(6):23–29, June 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ChunYan:2003:TEL

[826] Yu ChunYan, Wu Minghui, Liu Nairuo, Zhuang Yueting, and Pan Yunhe. Translating EXPRESS language model into C language model. ACM SIG-PLAN Notices, 38(6):30–39, June 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ortin:2003:RNA

[827] Francisco Ortin, Juan Manuel Cueva, and Ana Belen Martinez. The reflective nitrO abstract machine. ACM SIGPLAN Notices, 38(6):40–49, June 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sunik:2003:POC

[828] Boris Sunik. The paradigm of open C++. ACM SIGPLAN Notices, 38 (6):50–59, June 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:2003:AGD

[829] Qing-Li Zhang, Ming-Yuan Zhu, and Shuo-Ying Chen. Automatic generation of device drivers. ACM SIG-PLAN Notices, 38(6):60–69, June 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhao:2003:PIO

[830] Min Zhao, Bruce Childers, and Mary Lou Soffa. Predicting the impact of optimizations for embedded systems. ACM SIGPLAN Notices, 38(7):1–11, July 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kulkarni:2003:FEO

[831] Prasad Kulkarni, Wankang Zhao, Hwashin Moon, Kyunghwan Cho, David Whalley, Jack Davidson, Mark Bailey, Yunheung Paek, and Kyle Gallivan. Finding effective optimization phase sequences. ACM SIGPLAN Notices, 38(7):12–23, July 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vanbroekhoven:2003:ACP

[832] Peter Vanbroekhoven, Gerda Janssens, Maurice Bruynooghe, Henk Corporaal, and Francky Catthoor. Advanced copy propagation for arrays. ACM SIG-PLAN Notices, 38(7):24–33, July 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shaylor:2003:JVM

[833] Nik Shaylor, Douglas N. Simon, and William R. Bush. A Java Virtual Machine architecture for very small devices. ACM SIGPLAN Notices, 38 (7):34–41, July 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schultz:2003:CJL

[834] Ulrik Pagh Schultz, Kim Burgaard, Flemming Gram Christensen, and Jørgen Lindskov Knudsen. Compiling Java for low-end embedded systems. ACM SIGPLAN Notices, 38(7):42–50, July 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Corsaro:2003:EMR

[835] Angelo Corsaro and Ron K. Cytron. Efficient memory-reference checks for real-time Java. ACM SIGPLAN Notices, 38(7):51–58, July 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ananian:2003:DSO

[836] C. Scott Ananian and Martin Rinard. Data size optimizations for Java programs. ACM SIGPLAN Notices, 38 (7):59–68, July 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dhurjati:2003:MSR

[837] Dinakar Dhurjati, Sumant Kowshik, Vikram Adve, and Chris Lattner. Memory safety without runtime checks or garbage collection. ACM SIG-PLAN Notices, 38(7):69–80, July 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bacon:2003:CFS

[838] David F. Bacon, Perry Cheng, and V. T. Rajan. Controlling fragmentation and space consumption in the metronome, a real-time garbage collector for Java. ACM SIGPLAN Notices, 38(7):81–92, July 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Robertz:2003:TTG

[839] Sven Gestegard Robertz and Roger Henriksson. Time-triggered garbage collection: robust and adaptive realtime GC scheduling for embedded systems. ACM SIGPLAN Notices, 38 (7):93–102, July 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Karczmarek:2003:PSS

[840] Michal Karczmarek, William Thies, and Saman Amarasinghe. Phased scheduling of stream programs. ACM SIGPLAN Notices, 38(7):103–112, July 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pop:2003:SDF

[841] Paul Pop, Petru Eles, and Zebo Peng. Schedulability-driven frame packing for multi-cluster distributed embedded systems. ACM SIGPLAN Notices, 38 (7):113–122, July 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ait-Ameur:2003:RAA

[842] Y. Ait-Ameur, G. Bel, F. Boniol, S. Pairault, and V. Wiels. Robustness analysis of avionics embedded systems. ACM SIGPLAN Notices, 38(7):

123–132, July 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

deNiz:2003:TWS

[843] Dionisio de Niz and Raj Rajkumar. Time weaver: a software-through-models framework for embedded realtime systems. ACM SIGPLAN Notices, 38(7):133–143, July 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pratap:2003:TLA

[844] Ravi Pratap M., Ron K. Cytron, David Sharp, and Edward Pla. Transport layer abstraction in event channels for embedded systems. ACM SIG-PLAN Notices, 38(7):144–152, July 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Caspi:2003:SSL

[845] Paul Caspi, Adrian Curic, Aude Maignan, Christos Sofronis, Stavros Tripakis, and Peter Niebert. From Simulink to SCADE/lustre to TTA: a layered approach for distributed embedded applications. ACM SIG-PLAN Notices, 38(7):153–162, July 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wasowski:2003:EPS

[846] Andrzej Wasowski. On efficient program synthesis from statecharts. ACM SIGPLAN Notices, 38(7):163–170, July 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Alur:2003:GES

[847] Rajeev Alur, Franjo Ivancic, Jesung Kim, Insup Lee, and Oleg Sokolsky. Generating embedded software from hierarchical hybrid models. ACM SIG-PLAN Notices, 38(7):171–182, July 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lee:2003:AML

[848] Jong eun Lee, Kiyoung Choi, and Nikil D. Dutt. An algorithm for mapping loops onto coarse-grained reconfigurable architectures. ACM SIG-PLAN Notices, 38(7):183–188, July 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Suresh:2003:PTH

[849] Dinesh C. Suresh, Walid A. Najjar, Frank Vahid, Jason R. Villarreal, and Greg Stitt. Profiling tools for hardware/software partitioning of embedded applications. ACM SIG-PLAN Notices, 38(7):189–198, July 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Guo:2003:GCA

[850] Yuanqing Guo, Gerard J. M. Smit, Hajo Broersma, and Paul M. Heysters. A graph covering algorithm for a coarse grain reconfigurable system. ACM SIGPLAN Notices, 38(7):199–208, July 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:2003:TRW

[851] Tao Zhang, Santosh Pande, and Antonio Valverde. Tamper-resistant whole program partitioning. ACM SIG-PLAN Notices, 38(7):209–219, July 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhuang:2003:SAO

[852] Xiaotong Zhuang, ChokSheak Lau, and Santosh Pande. Storage assignment optimizations through variable coalescence for embedded processors. ACM SIGPLAN Notices, 38(7):220–231, July 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Corliss:2003:DID

[853] Marc L. Corliss, E. Christopher Lewis, and Amir Roth. A DISE implementation of dynamic code decompression. ACM SIGPLAN Notices, 38(7): 232–243, July 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

DeSutter:2003:SEC

[854] Bjorn De Sutter, Hans Vandierendonck, Bruno De Bus, and Koen De Bosschere. On the side-effects of code abstraction. ACM SIGPLAN Notices, 38(7):244–253, July 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Krishnaswamy:2003:EPB

[855] Arvind Krishnaswamy. Enhancing the performance of 16-bit code using augmenting instructions. ACM SIGPLAN Notices, 38(7):254–264, July 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yu:2003:EBT

[856] Yang Yu and Viktor K. Prasanna. Energy-balanced task allocation for collaborative processing in networked embedded systems. ACM SIGPLAN Notices, 38(7):265–274, July 2003. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kim:2003:AIL

[857] H. S. Kim, N. Vijaykrishnan, M. Kandemir, and M. J. Irwin. Adapting instruction level parallelism for optimizing leakage in VLIW architectures. ACM SIGPLAN Notices, 38 (7):275–283, July 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

AbouGhazaleh:2003:EMR

[858] Nevine AbouGhazaleh, Bruce Childers, Daniel Mosse, Rami Melhem, and Matthew Craven. Energy management for real-time embedded applications with compiler support. ACM SIGPLAN Notices, 38(7):284–293, July 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Christopher:2003:DSP

[859] Thomas W. Christopher. Details on a simple parallel system. ACM SIG-PLAN Notices, 38(8):12–14, August 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:2003:JF

[860] Paul Frenger. The JOY of Forth. ACM SIGPLAN Notices, 38(8):15–17, August 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pronk:2003:FDP

[861] C. Pronk and M. Schönhacker. Formal definition of programming language standards. ACM SIGPLAN Notices, 38 (8):20–21, August 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xu:2003:MEJ

[862] Baomin Xu, Weimin Lian, and Qiang Gao. Migration of enterprise JavaBeans with ProActive Interposition Objects. ACM SIGPLAN Notices, 38(8):22–28, August 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Costanza:2003:DSF

[863] Pascal Costanza. Dynamically scoped functions as the essence of AOP. ACM SIGPLAN Notices, 38(8):29–36, August 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Iliasov:2003:TBP

[864] Alex Iliasov. Templates-based portable just-in-time compiler. ACM SIG-PLAN Notices, 38(8):37–43, August 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hericko:2003:OSA

[865] Marjan Hericko, Matjaz B. Juric, Ivan Rozman, Simon Beloglavec, and Ales Zivkovic. Object serialization analysis and comparison in Java and .NET. ACM SIGPLAN Notices, 38(8):44–54, August 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stephens:2003:XCI

[866] Kurt Stephens. XVF: C++ introspection by extensible visitation. ACM SIGPLAN Notices, 38(8):55–59, August 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kumar:2003:VRO

[867] K. V. Seshu Kumar. Value reuse optimization: reuse of evaluated math library function calls through compiler generated cache. ACM SIG-PLAN Notices, 38(8):60–66, August 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Knight:2003:CIA

[868] Thomas Knight, Jr. Conservation of information: applications in functional, reversible, and quantum computing. ACM SIGPLAN Notices, 38 (9):1, September 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Heeren:2003:STI

[869] Bastiaan Heeren, Jurriaan Hage, and S. Doaitse Swierstra. Scripting the type inference process. ACM SIG-PLAN Notices, 38(9):3–13, September 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Neubauer:2003:DST

[870] Matthias Neubauer and Peter Thiemann. Discriminative sum types locate the source of type errors. ACM SIG-PLAN Notices, 38(9):15–26, September 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

LeBotlan:2003:MRM

[871] Didier Le Botlan and Didier Rémy. ML^{F}: raising ML to the power of system F. ACM SIGPLAN Notices, 38(9): 27–38, September 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Simonet:2003:EHX

[872] Vincent Simonet. An extension of HM(X) with bounded existential and universal data-types. ACM SIGPLAN Notices, 38(9):39–50, September 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Benzaken:2003:CXC

[873] Véronique Benzaken, Giuseppe Castagna, and Alain Frisch. CDuce: an XML-centric general-purpose language. ACM SIGPLAN Notices, 38(9): 51–63, September 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Levin:2003:CRP

[874] Michael Y. Levin. Compiling regular patterns. ACM SIGPLAN Notices, 38(9):65–77, September 2003.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Page:2003:SDM

[875] Rex L. Page. Software is discrete mathematics. ACM SIGPLAN Notices, 38(9):79–86, September 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Leifer:2003:GAS

[876] James J. Leifer, Gilles Peskine, Peter Sewell, and Keith Wansbrough. Global abstraction-safe marshalling with hash types. ACM SIGPLAN Notices, 38(9): 87–98, September 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bierman:2003:DRM

[877] Gavin Bierman, Michael Hicks, Peter Sewell, Gareth Stoyle, and Keith Wansbrough. Dynamic rebinding for marshalling and update, with destructtime λ. ACM SIGPLAN Notices, 38(9): 99–110, September 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ogawa:2003:IFP

[878] Mizuhito Ogawa, Zhenjiang Hu, and Isao Sasano. Iterative-free program analysis. ACM SIGPLAN Notices, 38 (9):111–123, September 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mairson:2003:HSD

[879] Harry G. Mairson. From Hilbert space to Dilbert space: context semantics as a language for games and flow analysis. ACM SIGPLAN Notices, 38(9): 125, September 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Walker:2003:TA

[880] David Walker, Steve Zdancewic, and Jay Ligatti. A theory of aspects. ACM SIGPLAN Notices, 38(9):127– 139, September 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Loh:2003:DSG

[881] Andres Löh, Dave Clarke, and Johan Jeuring. Dependency-style generic Haskell. ACM SIGPLAN Notices, 38 (9):141–152, September 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nilsson:2003:FAD

[882] Henrik Nilsson. Functional automatic differentiation with Dirac impulses. ACM SIGPLAN Notices, 38(9):153– 164, September 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jones:2003:UCA

[883] Simon Peyton Jones, Alan Blackwell, and Margaret Burnett. A usercentred approach to functions in Excel. ACM SIGPLAN Notices, 38(9):165– 176, September 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kameyama:2003:SCA

[884] Yukiyoshi Kameyama and Masahito Hasegawa. A sound and complete axiomatization of delimited continuations. ACM SIGPLAN Notices, 38(9): 177–188, September 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wadler:2003:CVD

[885] Philip Wadler. Call-by-value is dual to call-by-name. ACM SIGPLAN Notices, 38(9):189–201, September 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Beffara:2003:DNF

[886] Emmanuel Beffara and Vincent Danos. Disjunctive normal forms and local exceptions. ACM SIGPLAN Notices, 38(9):203–211, September 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mandelbaum:2003:ETT

[887] Yitzhak Mandelbaum, David Walker, and Robert Harper. An effective theory of type refinements. ACM SIG-PLAN Notices, 38(9):213–225, September 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Higuchi:2003:STS

[888] Tomoyuki Higuchi and Atsushi Ohori. A static type system for JVM access control. ACM SIGPLAN Notices, 38(9):227–237, September 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hughes:2003:PPS

[889] R. John M. Hughes and S. Doaitse Swierstra. Polish parsers, step by step. ACM SIGPLAN Notices, 38(9):239– 248, September 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Washburn:2003:BGB

[890] Geoffrey Washburn and Stephanie Weirich. Boxes go bananas: encoding higher-order abstract syntax with parametric polymorphism. ACM SIG-PLAN Notices, 38(9):249–262, September 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shinwell:2003:FPB

[891] Mark R. Shinwell, Andrew M. Pitts, and Murdoch J. Gabbay. FreshML: programming with binders made simple. ACM SIGPLAN Notices, 38(9): 263–274, September 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:2003:MPT

[892] Chiyan Chen and Hongwei Xi. Metaprogramming through typeful code representation. ACM SIGPLAN Notices, 38(9):275–286, September 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ennals:2003:OEA

[893] Robert Ennals and Simon Peyton Jones. Optimistic evaluation: an adaptive evaluation strategy for nonstrict programs. ACM SIGPLAN Notices, 38(9):287–298, September 2003.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wand:2003:UAE

[894] Mitchell Wand. Understanding aspects: extended abstract. ACM SIG-PLAN Notices, 38(9):299–300, September 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kazar:2003:HPS

[895] Baris Kazar. High performance spatial data mining for very large data-sets (citation only). ACM SIGPLAN Notices, 38(10):1, October 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Puppin:2003:ACS

[896] Diego Puppin. Adapting convergent scheduling using machine learning (citation only). ACM SIGPLAN Notices, 38(10):1, October 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Prabhu:2003:UTL

[897] Manohar K. Prabhu and Kunle Olukotun. Using thread-level speculation to simplify manual parallelization. ACM SIGPLAN Notices, 38(10):1–12, October 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Coarfa:2003:ECA

[898] Cristian Coarfa and Yuri Dotsenko. An emerging co-array Fortran compiler (citation only). ACM SIGPLAN Notices, 38(10):2, October 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Squyres:2003:CAL

[899] Jeffrey M. Squyres. A component architecture for LAM/MPI (citation only). ACM SIGPLAN Notices, 38(10):2, October 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lee:2003:PPA

[900] Jaejin Lee. The Pensieve Project: automatic implementation of programming language consistency models (citation only). ACM SIGPLAN Notices, 38 (10):3, October 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nienaltowski:2003:SIC

[901] Piotr Nienaltowski. SCOOP It up! (citation only). ACM SIGPLAN Notices, 38(10):3, October 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cintra:2003:TER

[902] Marcelo Cintra and Diego R. Llanos. Toward efficient and robust software speculative parallelization on multiprocessors. ACM SIGPLAN Notices, 38 (10):13–24, October 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:2003:CSS

[903] Peng-Sheng Chen, Ming-Yu Hung, Yuan-Shin Hwang, Roy Dz-Ching Ju, and Jenq Kuen Lee. Compiler support for speculative multithreading architecture with probabilistic points-to analysis. ACM SIGPLAN Notices, 38(10): 25–36, October 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McDowell:2003:ISS

[904] Luke K. McDowell, Susan J. Eggers, and Steven D. Gribble. Improving server software support for simultaneous multithreaded processors. ACM SIGPLAN Notices, 38(10):37–48, October 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fraguela:2003:PFP

[905] Basilio B. Fraguela, Jose Renau, Paul Feautrier, David Padua, and Josep Torrellas. Programming the FlexRAM parallel intelligent memory system. ACM SIGPLAN Notices, 38(10):49–60, October 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kim:2003:ETL

[906] Hyong youb Kim, Vijay S. Pai, and Scott Rixner. Exploiting task-level concurrency in a programmable network interface. ACM SIGPLAN Notices, 38 (10):61–72, October 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McCurdy:2003:UCC

[907] Collin McCurdy and Charles Fischer. User-controllable coherence for high performance shared memory multiprocessors. ACM SIGPLAN Notices, 38 (10):73–82, October 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bronevetsky:2003:AAL

[908] Greg Bronevetsky, Daniel Marques, Keshav Pingali, and Paul Stodghill. Automated application-level checkpointing of MPI programs. ACM SIG-PLAN Notices, 38(10):84–94, October 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Karwande:2003:CMC

[909] Amit Karwande, Xin Yuan, and David K. Lowenthal. CC–MPI: a compiled communication capable MPI prototype for Ethernet switched clusters. ACM SIGPLAN Notices, 38(10):95– 106, October 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Liu:2003:IMS

[910] Ting Liu and Margaret Martonosi. Impala: a middleware system for managing autonomic, parallel sensor systems. ACM SIGPLAN Notices, 38(10): 107–118, October 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chu:2003:ODA

[911] Lingkun Chu, Hong Tang, Tao Yang, and Kai Shen. Optimizing data aggregation for cluster-based Internet services. ACM SIGPLAN Notices, 38(10): 119–130, October 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:2003:EHL

[912] DeQing Chen, Chunqiang Tang, Brandon Sanders, Sandhya Dwarkadas, and Michael L. Scott. Exploiting high-level coherence information to optimize distributed shared state. ACM SIGPLAN Notices, 38(10):131–142, October 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frens:2003:FMO

[913] Jeremy D. Frens and David S. Wise. Factorization with Morton-ordered quadtree matrices for memory re-use and parallelism. ACM SIGPLAN Notices, 38(10):144–154, October 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frens:2003:QFM

[914] Jeremy D. Frens and David S. Wise. QR factorization with Morton-ordered quadtree matrices for memory re-use and parallelism. ACM SIGPLAN Notices, 38(10):144–154, October 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Deitz:2003:DIP

[915] Steven J. Deitz, Bradford L. Chamberlain, Sung-Eun Choi, and Lawrence Snyder. The design and implementation of a parallel array operator for the arbitrary remapping of data. ACM SIGPLAN Notices, 38(10):155– 166, October 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

OCallahan:2003:HDD

[916] Robert O’Callahan and Jong-Deok Choi. Hybrid dynamic data race detection. ACM SIGPLAN Notices, 38(10): 167–178, October 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pozniansky:2003:EFD

[917] Eli Pozniansky and Assaf Schuster. Efficient on-the-fly data race detection in multithreaded C++ programs. ACM SIGPLAN Notices, 38(10):179– 190, October 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Goel:2003:CRP

[918] Ankit Goel, Abhik Roychoudhury, and Tulika Mitra. Compactly representing parallel program executions. ACM SIGPLAN Notices, 38(10):191– 202, October 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tan:2003:UGD

[919] Kai Tan, Duane Szafron, Jonathan Schaeffer, John Anvik, and Steve Mac-Donald. Using generative design patterns to generate parallel code for a distributed memory environment. ACM SIGPLAN Notices, 38(10):203– 215, October 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Taura:2003:PPP

[920] Kenjiro Taura, Kenji Kaneda, Toshio Endo, and Akinori Yonezawa. Phoenix: a parallel programming model for accommodating dynamically joining/ leaving resources. ACM SIGPLAN Notices, 38(10):216–229, October 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Saunders:2003:AAP

[921] Steven Saunders and Lawrence Rauchwerger. ARMI: an adaptive, platform independent communication library. ACM SIGPLAN Notices, 38(10): 230–240, October 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hatcliff:2003:SPE

[922] John Hatcliff, William Deng, Matthew B. Dwyer, Georg Jung, Venkatesh Ranganath, and Robby. Slicing and partial evaluation of CORBA component model designs for avionics system. ACM SIGPLAN Notices, 38(10): 241–242, October 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ager:2003:FPE

[923] Mads Sig Ager, Olivier Danvy, and Henning Korsholm Rohde. Fast partial evaluation of pattern matching in strings. ACM SIGPLAN Notices, 38(10):243–249, October 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gluck:2003:TII

[924] Robert Glück, Youhei Kawada, and Takuya Hashimoto. Transforming interpreters into inverse interpreters by partial evaluation. ACM SIGPLAN Notices, 38(10):250–259, October 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:2003:ITP

[925] Chiyan Chen and Hongwei Xi. Implementing typeful program transformations. ACM SIGPLAN Notices, 38(10): 260–268, October 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Puebla:2003:ASA

[926] Germán Puebla and Manuel Hermenegildo. Abstract specialization and its applications. ACM SIGPLAN Notices, 38(10): 269–283, October 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Allen:2003:SJP

[927] Matthew Allen and Susan Horwitz. Slicing Java programs that throw and catch exceptions. ACM SIGPLAN Notices, 38(10):284–294, October 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bozga:2003:SSA

[928] Marius Bozga, Radu Iosif, and Yassine Laknech. Storeless semantics and alias logic. ACM SIGPLAN Notices, 38(10): 295–305, October 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Attardi:2003:CCF

[929] Giuseppe Attardi, Antonio Cisternino, and Andrew Kennedy. CodeBricks: code fragments as building blocks. ACM SIGPLAN Notices, 38(10):306– 314, October 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chin:2003:EST

[930] Wei-Ngan Chin, Siau-Cheng Khoo, and Dana N. Xu. Extending sized type with collection analysis. ACM SIG-PLAN Notices, 38(10):315–324, October 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Liu:2003:OAF

[931] Yanhong A. Liu and Scott D. Stoller. Optimizing Ackermann’s function by incrementalization. ACM SIGPLAN Notices, 38(10):325–331, October 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lafferty:2003:LIA

[932] Donal Lafferty and Vinny Cahill. Language-independent aspect-oriented programming. ACM SIGPLAN Notices, 38(11):1–12, November 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tip:2003:RGU

[933] Frank Tip, Adam Kiezun, and Dirk Bäumer. Refactoring for generalization using type constraints. ACM SIG-PLAN Notices, 38(11):13–26, November 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tanter:2003:PBR

[934] Éric Tanter, Jacques Noyé, Denis Caromel, and Pierre Cointe. Partial behavioral reflection: spatial and temporal selection of reification. ACM SIG-PLAN Notices, 38(11):27–46, November 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Black:2003:ATS

[935] Andrew P. Black, Nathanael Schärli, and Stéphane Ducasse. Applying traits to the Smalltalk collection classes. ACM SIGPLAN Notices, 38(11):47– 64, November 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mougin:2003:OIA

[936] Philippe Mougin and Stéphane Ducasse. OOPAL: integrating array programming in object-oriented programming. ACM SIGPLAN Notices, 38(11):65– 77, November 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Demsky:2003:ADR

[937] Brian Demsky and Martin Rinard. Automatic detection and repair of errors in data structures. ACM SIGPLAN Notices, 38(11):78–95, November 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Allen:2003:FCA

[938] Eric Allen, Jonathan Bannet, and Robert Cartwright. A first-class approach to genericity. ACM SIG-PLAN Notices, 38(11):96–114, November 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Garcia:2003:CSL

[939] Ronald Garcia, Jaakko Jarvi, Andrew Lumsdaine, Jeremy Siek, and Jeremiah Willcock. A comparative study of language support for generic programming. ACM SIGPLAN Notices, 38(11):115–134, November 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhao:2003:LCF

[940] Tian Zhao, Jens Palsber, and Jan Vite. Lightweight confinement for featherweight Java. ACM SIGPLAN Notices, 38(11):135–148, November 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dufour:2003:DMJ

[941] Bruno Dufour, Karel Driesen, Laurie Hendren, and Clark Verbrugge. Dynamic metrics for Java. ACM SIGPLAN Notices, 38(11):149–168, November 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Eeckhout:2003:HJP

[942] Lieven Eeckhout, Andy Georges, and Koen De Bosschere. How Java programs interact with virtual machines at the microarchitectural level. ACM SIGPLAN Notices, 38(11):169– 186, November 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ishizaki:2003:ECP

[943] Kazuaki Ishizaki, Mikio Takeuchi, Kiyokuni Kawachiya, Toshio Suganuma, Osamu Gohda, Tatsushi Inagaki, Akira Koseki, Kazunori Ogata, Motohiro Kawahito, Toshiaki Yasue, Takeshi Ogasawara, Tamiya Onodera, Hideaki Komatsu, and Toshio Nakatani. Effectiveness of crossplatform optimizations for a Java justin-time compiler. ACM SIGPLAN Notices, 38(11):187–204, November 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lee:2003:HOO

[944] Keunwoo Lee, Anthony LaMarca, and Craig Chambers. HydroJ: objectoriented pattern matching for evolvable distributed systems. ACM SIGPLAN Notices, 38(11):205–223, November 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Millstein:2003:RMB

[945] Todd Millstein, Mark Reay, and Craig Chambers. Relaxed MultiJava: balancing extensibility and modular typechecking. ACM SIGPLAN Notices, 38 (11):224–240, November 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Corwin:2003:MRM

[946] John Corwin, David F. Bacon, David Grove, and Chet Murthy. MJ: a rational module system for Java and its applications. ACM SIGPLAN Notices, 38(11):241–254, November 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Barabash:2003:MCG

[947] Katherine Barabash, Yoav Ossia, and Erez Petrank. Mostly concurrent garbage collection revisited. ACM SIGPLAN Notices, 38(11):255–268, November 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Azatchi:2003:FMS

[948] Hezi Azatchi, Yossi Levanoni, Harel Paz, and Erez Petrank. An on-the-fly mark and sweep garbage collector based on sliding views. ACM SIGPLAN Notices, 38(11):269–281, November 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:2003:HCM

[949] G. Chen, M. Kandemir, N. Vijaykrishnan, M. J. Irwin, B. Mathiske, and M. Wolczko. Heap compression for memory-constrained Java environments. ACM SIGPLAN Notices, 38(11):282–301, November 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fahndrich:2003:DCN

[950] Manuel Fähndrich and K. Rustan M. Leino. Declaring and checking non-null types in an object-oriented language. ACM SIGPLAN Notices, 38(11):302– 312, November 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marinov:2003:OEP

[951] Darko Marinov and Robert O’Callahan. Object equality profiling. ACM SIGPLAN Notices, 38(11):313–325, November 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sachindran:2003:MCF

[952] Narendran Sachindran, J. Eliot, and B. Moss. Mark-copy: fast copying GC with less space overhead. ACM SIGPLAN Notices, 38(11):326– 343, November 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Blackburn:2003:URC

[953] Stephen M. Blackburn and Kathryn S. McKinley. Ulterior reference counting: fast garbage collection without a long wait. ACM SIGPLAN Notices, 38(11):344–358, November 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hirzel:2003:CBG

[954] Martin Hirzel, Amer Diwan, and Matthew Hertz. Connectivity-based garbage collection. ACM SIGPLAN Notices, 38(11):359–373, November 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clarke:2003:SWB

[955] Dave Clarke, Michael Richmond, and James Noble. Saving the world from bad beans: deployment-time confinement checking. ACM SIGPLAN Notices, 38(11):374–387, November 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Harris:2003:LSL

[956] Tim Harris and Keir Fraser. Language support for lightweight transactions. ACM SIGPLAN Notices, 38(11):388– 402, November 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boyapati:2003:LMU

[957] Chandrasekhar Boyapati, Barbara Liskov, Liuba Shrira, Chuang-Hue Moh, and Steven Richman. Lazy modular upgrades in persistent object stores. ACM SIGPLAN Notices, 38 (11):403–417, November 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Burke:2003:ECO

[958] Michael G. Burke. Executive Committee October 2003 meeting report. ACM SIGPLAN Notices, 38(12):1, December 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fenwick:2003:LE

[959] Jay Fenwick and Cindy Norris. Letters from the Editors. ACM SIG-PLAN Notices, 38(12):2, December 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wadler:2003:RIP

[960] Philip Wadler, Krzysztof R. Apt, and Matthias Felleisen. Reminiscences on influential papers [definitional interpreters for higher-order programming languages; structured programming; call-by-name, call-by-value, and the λcalculus]. ACM SIGPLAN Notices, 38 (12):2, December 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Creak:2003:PLD

[961] Alan Creak. Programming lessons from days gone by: extreme design. ACM SIGPLAN Notices, 38(12):11–16, December 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fenwick:2003:IA

[962] Jay Fenwick and Cindy Norris. Information for authors. ACM SIG-PLAN Notices, 38(12):17, December 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Griffith:2003:LES

[963] Walter L. (Larry) Griffith. Letter to the Editor: Some deficiencies of C++ in teaching CS1 and CS2. ACM SIG-PLAN Notices, 38(12):18–19, December 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Avvenuti:2003:JBV

[964] Marco Avvenuti, Cinzia Bernardeschi, and Nicoletta De Francesco. Java bytecode verification for secure information flow. ACM SIGPLAN Notices, 38(12): 20–27, December 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Aycock:2003:ACC

[965] John Aycock. The ART of compiler construction projects. ACM SIGPLAN Notices, 38(12):28–32, December 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lopes:2003:BAT

[966] Cristina Videira Lopes, Paul Dourish, David H. Lorenz, and Karl Lieberherr. Beyond AOP: toward naturalistic programming. ACM SIGPLAN Notices, 38(12):34–43, December 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kamin:2003:RRT

[967] Sam Kamin. Routine run-time code generation. ACM SIGPLAN Notices, 38(12):44–56, December 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rinard:2003:AOC

[968] Martin Rinard. Acceptability-oriented computing. ACM SIGPLAN Notices, 38(12):57–75, December 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Biddle:2003:NNJ

[969] Robert Biddle, Angela Martin, and James Noble. No name: just notes on software reuse. ACM SIGPLAN Notices, 38(12):76–96, December 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rival:2004:STF

[970] Xavier Rival. Symbolic transfer function-based approaches to certified compilation. ACM SIGPLAN Notices, 39(1):1–13, January 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Benton:2004:SRC

[971] Nick Benton. Simple relational correctness proofs for static analyses and program transformations. ACM SIG-PLAN Notices, 39(1):14–25, January 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sittampalam:2004:IET

[972] Ganesh Sittampalam, Oege de Moor, and Ken Friis Larsen. Incremental execution of transformation specifications. ACM SIGPLAN Notices, 39(1):26–38, January 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yu:2004:FGN

[973] Dachuan Yu, Andrew Kennedy, and Don Syme. Formalization of generics for the .NET common language runtime. ACM SIGPLAN Notices, 39(1): 39–51, January 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vouillon:2004:STF

[974] Jerome Vouillon and Paul-André Melliès. Semantic types: a fresh look at the ideal model for types. ACM SIG-PLAN Notices, 39(1):52–63, January 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Balat:2004:ENT

[975] Vincent Balat, Roberto Di Cosmo, and Marcelo Fiore. Extensional normalisation and type-directed partial evaluation for typed lambda calculus with sums. ACM SIGPLAN Notices, 39(1): 64–76, January 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fiore:2004:IGR

[976] Marcelo Fiore. Isomorphisms of generic recursive polynomial types. ACM SIG-PLAN Notices, 39(1):77–88, January 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pottier:2004:PTD

[977] François Pottier and Nadji Gauthier. Polymorphic typed defunctionalization. ACM SIGPLAN Notices, 39(1): 89–98, January 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Johann:2004:FTP

[978] Patricia Johann and Janis Voigtländer. Free theorems in the presence of seq. ACM SIGPLAN Notices, 39(1):99–110, January 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ford:2004:PEG

[979] Bryan Ford. Parsing expression grammars: a recognition-based syntactic foundation. ACM SIGPLAN Notices, 39(1):111–122, January 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Caromel:2004:ADO

[980] Denis Caromel, Ludovic Henrio, and Bernard Paul Serpette. Asynchronous and deterministic objects. ACM SIG-PLAN Notices, 39(1):123–134, January 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zilio:2004:LYC

[981] Silvano Dal Zilio, Denis Lugiez, and Charles Meyssonnier. A logic you can count on. ACM SIGPLAN Notices, 39(1):135–146, January 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yoshida:2004:CDT

[982] Nobuko Yoshida. Channel dependent types for higher-order mobile processes. ACM SIGPLAN Notices, 39(1):147– 160, January 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sumii:2004:BDS

[983] Eijiro Sumii and Benjamin C. Pierce. A bisimulation for dynamic sealing. ACM SIGPLAN Notices, 39(1):161–172, January 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cousot:2004:AIB

[984] Patrick Cousot and Radhia Cousot. An abstract interpretation-based framework for software watermarking. ACM SIGPLAN Notices, 39(1):173–185, January 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Giacobazzi:2004:ANI

[985] Roberto Giacobazzi and Isabella Mastroeni. Abstract non-interference: parameterizing non-interference by abstract interpretation. ACM SIGPLAN Notices, 39(1):186–197, January 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bhargavan:2004:SWS

[986] Karthikeyan Bhargavan, Cédric Fournet, and Andrew D. Gordon. A semantics for Web services authentication. ACM SIGPLAN Notices, 39(1): 198–209, January 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boehm:2004:SCL

[987] Hans-J. Boehm. The space cost of lazy reference counting. ACM SIG-PLAN Notices, 39(1):210–219, January 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Birkedal:2004:LRA

[988] Lars Birkedal, Noah Torp-Smith, and John C. Reynolds. Local reasoning about a copying garbage collector. ACM SIGPLAN Notices, 39(1): 220–231, January 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Henzinger:2004:AP

[989] Thomas A. Henzinger, Ranjit Jhala, Rupak Majumdar, and Kenneth L. McMillan. Abstractions from proofs. ACM SIGPLAN Notices, 39(1):232– 244, January 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Qadeer:2004:SPC

[990] Shaz Qadeer, Sriram K. Rajamani, and Jakob Rehof. Summarizing procedures in concurrent programs. ACM SIG-PLAN Notices, 39(1):245–255, January 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Flanagan:2004:ADA

[991] Cormac Flanagan and Stephen N. Freund. Atomizer: a dynamic atomicity checker for multithreaded programs. ACM SIGPLAN Notices, 39(1): 256–267, January 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

OHearn:2004:SIH

[992] Peter W. O’Hearn, Hongseok Yang, and John C. Reynolds. Separation and information hiding. ACM SIGPLAN Notices, 39(1):268–280, January 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dunfield:2004:TT

[993] Joshua Dunfield and Frank Pfenning. Tridirectional typechecking. ACM SIGPLAN Notices, 39(1):281–292, January 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dreyer:2004:TSW

[994] Derek Dreyer. A type system for wellfounded recursion. ACM SIGPLAN Notices, 39(1):293–305, January 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ancona:2004:PTJ

[995] Davide Ancona and Elena Zucca. Principal typings for Java-like languages. ACM SIGPLAN Notices, 39(1):306– 317, January 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sankaranarayanan:2004:NLL

[996] Sriram Sankaranarayanan, Henny B. Sipma, and Zohar Manna. Nonlinear loop invariant generation using Gröbner bases. ACM SIGPLAN Notices, 39(1):318–329, January 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Muller-Olm:2004:PIA

[997] Markus Müller-Olm and Helmut Seidl. Precise interprocedural analysis through linear algebra. ACM SIG-PLAN Notices, 39(1):330–341, January 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gulwani:2004:GVN

[998] Sumit Gulwani and George C. Necula. Global value numbering using random interpretation. ACM SIGPLAN Notices, 39(1):342–352, January 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:2004:DT

[999] Paul Frenger. Dutch treat. ACM SIG-PLAN Notices, 39(2):7–10, February 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Singh:2004:ALO

[1000] Awadhesh Kumar Singh and Anup Kumar Bandyopadhyay. Adding the leadsto operator to Dijkstra’s calculus. ACM SIGPLAN Notices, 39(2):12–17, February 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lu:2004:CTO

[1001] Hongmin Lu, Yuming Zhou, Jiangtao Lu, and Baowen Xu. A compile-time optimization framework for Ada rendezvous. ACM SIGPLAN Notices, 39 (2):18–25, February 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Song:2004:WCW

[1002] Litong Song and Krishna Kavi. What can we gain by unfolding loops? ACM SIGPLAN Notices, 39(2):26–33, February 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Liu:2004:PAM

[1003] Yuan Liu and Baowen Xu. Process algebra model of Ada protected objects. ACM SIGPLAN Notices, 39(2): 34–39, February 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:2004:DEM

[1004] Paul Frenger. Deus Ex Macintosh. ACM SIGPLAN Notices, 39(3):7–11, March 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:2004:SSD

[1005] Yingzhou Zhang and Baowen Xu. A survey of semantic description frameworks for programming languages. ACM SIGPLAN Notices, 39(3):14–30, March 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yavner:2004:BPK

[1006] Jonathan Yavner. Back-propagation of knowledge from syntax tree to C source code. ACM SIGPLAN Notices, 39 (3):31–37, March 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kumar:2004:WWC

[1007] K. V. Seshu Kumar. When and what to compile/optimize in a virtual machine? ACM SIGPLAN Notices, 39 (3):38–45, March 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lam:2004:GTC

[1008] Patrick Lam, Viktor Kuncak, and Martin Rinard. Generalized typestate checking using set interfaces and pluggable analyses. ACM SIGPLAN Notices, 39(3):46–55, March 2004. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2004:FRa

[1009] Anonymous. Forth report. ACM SIG-PLAN Notices, 39(3):??, March 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cattell:2004:CGM

[1010] R. G. G. Cattell, Joseph M. Newcomer, and Bruce W. Leverett. Code generation in a machine-independent compiler. ACM SIGPLAN Notices, 39 (4):1–13, April 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pennello:2004:ECL

[1011] Thomas J. Pennello and Frank DeRemer. Efficient computation of LALR(1) look-ahead sets. ACM SIGPLAN Notices, 39(4):14–27, April 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Minimair:2004:MMR

[1012] Manfred Minimair. MR: Macaulay Resultant package for Maple. ACM SIGPLAN Notices, 39(4):26–29, April 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Allen:2004:ASO

[1013] Frances E. Allen and Janet Fabri. Automatic storage optimization. ACM SIGPLAN Notices, 39(4):28–37, April 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Auslander:2004:OPC

[1014] Marc Auslander and Martin Hopkins. An overview of the PL.8 compiler. ACM SIGPLAN Notices, 39(4):38–48, April 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Graham:2004:GCG

[1015] Susan L. Graham, Peter B. Kessler, and Marshall K. McKusick. gprof: a call graph execution profiler. ACM SIGPLAN Notices, 39(4):49–57, April 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Markstein:2004:ORC

[1016] Victoria Markstein, Peter Markstein, and John Cocke. Optimization of range checking. ACM SIGPLAN Notices, 39 (4):58–65, April 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chaitin:2004:RAS

[1017] Gregory Chaitin. Register allocation and spilling via graph coloring. ACM SIGPLAN Notices, 39(4):66–74, April 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Allen:2004:ALI

[1018] Randy Allen and Ken Kennedy. Automatic loop interchange. ACM SIGPLAN Notices, 39(4):75–90, April 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chow:2004:RAP

[1019] Fred Chow and John Hennessy. Register allocation by priority-based coloring. ACM SIGPLAN Notices, 39 (4):91–103, April 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Davidson:2004:AGP

[1020] Jack W. Davidson and Christopher W. Fraser. Automatic generation of peephole optimizations. ACM SIGPLAN Notices, 39(4):104–111, April 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fisher:2004:PPS

[1021] Joseph A. Fisher, John R. Ellis, John C. Ruttenberg, and Alexandru Nicolau. Parallel processing: a smart compiler and a dumb machine. ACM SIGPLAN Notices, 39(4): 112–124, April 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Johnsson:2004:ECL

[1022] Thomas Johnsson. Efficient compilation of lazy evaluation. ACM SIG-PLAN Notices, 39(4):125–138, April 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Burke:2004:IDA

[1023] Michael G. Burke and Ron K. Cytron. Interprocedural dependence analysis and parallelization. ACM SIGPLAN Notices, 39(4):139–154, April 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Callahan:2004:ICP

[1024] David Callahan, Keith D. Cooper, Ken Kennedy, and Linda Torczon. Interprocedural constant propagation. ACM SIGPLAN Notices, 39(4):155– 166, April 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Muchnick:2004:EIS

[1025] Steven S. Muchnick and Phillip B. Gibbons. Efficient instruction scheduling for a pipelined architecture. ACM SIG-PLAN Notices, 39(4):167–174, April 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kranz:2004:OOC

[1026] David Kranz, Richard Kelsey, Jonathan Rees, Paul Hudak, James Philbin, and Norman Adams. Orbit: an optimizing compiler for Scheme. ACM SIG-PLAN Notices, 39(4):175–191, April 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wall:2004:GRA

[1027] David W. Wall. Global register allocation at link time. ACM SIG-PLAN Notices, 39(4):192–204, April 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Appel:2004:RTC

[1028] Andrew W. Appel. Real-time concurrent collection on stock multiprocessors. ACM SIGPLAN Notices, 39(4): 205–216, April 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cooper:2004:ISE

[1029] Keith D. Cooper and Ken Kennedy. Interprocedural side-effect analysis in linear time. ACM SIGPLAN Notices, 39 (4):217–228, April 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Horwitz:2004:ISU

[1030] Susan Horwitz, Thomas Reps, and David Binkley. Interprocedural slicing using dependence graphs. ACM SIG-PLAN Notices, 39(4):229–243, April 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lam:2004:SPE

[1031] Monica S. Lam. Software pipelining: an effective scheduling technique for VLIW machines. ACM SIGPLAN Notices, 39(4):244–256, April 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shivers:2004:HOC

[1032] Olin Shivers. Higher-order controlflow analysis in retrospect: lessons learned, lessons abandoned. ACM SIG-PLAN Notices, 39(4):257–269, April 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wall:2004:RWV

[1033] David W. Wall. Register windows vs. register allocation. ACM SIG-PLAN Notices, 39(4):270–282, April 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Briggs:2004:CHR

[1034] Preston Briggs, Keith D. Cooper, Ken Kennedy, and Linda Torczon. Coloring heuristics for register allocation. ACM SIGPLAN Notices, 39(4): 283–294, April 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chambers:2004:RCO

[1035] Craig Chambers and David Ungar. A retrospective on: “Customization: Optimizing Compiler Technology for Self, a Dynamically-Typed Object-Oriented Programming Language”. ACM SIG-PLAN Notices, 39(4):295–312, April 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schonberg:2004:FDA

[1036] Edith Schonberg. On-the-fly detection of access anomalies. ACM SIG-PLAN Notices, 39(4):313–327, April 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Callahan:2004:IRA

[1037] David Callahan, Steve Carr, and Ken Kennedy. Improving register allocation for subscripted variables. ACM SIG-PLAN Notices, 39(4):328–342, April 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chase:2004:APS

[1038] David Chase, Mark Wegman, and F. Ken Zadeck. Analysis of pointers and structures. ACM SIGPLAN Notices, 39(4):343–359, April 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clinger:2004:HRF

[1039] William D. Clinger. How to read floating point numbers accurately. ACM SIGPLAN Notices, 39(4):360– 371, April 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clinger:2004:RHR

[1040] William D. Clinger. Retrospective: How to read floating point numbers accurately. ACM SIGPLAN Notices, 39 (4):360–371, April 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Best of PLDI 1979–1999. Reprint of, and retrospective on, [1].

Steele:2004:RHP

[1041] Guy L. Steele Jr. and Jon L. White. Retrospective: How to print floatingpoint numbers accurately. ACM SIG-PLAN Notices, 39(4):372–389, April 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Best of PLDI 1979– 1999. Reprint of, and retrospective on, [2].

Kessler:2004:FBD

[1042] Peter B. Kessler. Fast breakpoints: design and implementation. ACM SIG-PLAN Notices, 39(4):390–397, April 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pettis:2004:PGC

[1043] Karl Pettis, Robert C. Hansen, and Jack W. Davidson. Profile guided code positioning. ACM SIGPLAN Notices, 39(4):398–411, April 2004.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cartwright:2004:ST

[1044] Robert Cartwright and Mike Fagan.

Soft typing. ACM SIGPLAN Notices, 39(4):412–428, April 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wall:2004:PPB

[1045] David W. Wall. Predicting program behavior using real or estimated profiles. ACM SIGPLAN Notices, 39(4): 429–441, April 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lam:2004:DLO

[1046] Monica S. Lam and Michael E. Wolf.

A data locality optimizing algorithm. ACM SIGPLAN Notices, 39(4):442– 459, April 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Knoop:2004:LCM

[1047] Jens Knoop, Oliver Rüthing, and Bernhard Steffen. Lazy code motion. ACM SIGPLAN Notices, 39(4): 460–472, April 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Landi:2004:SAA

[1048] William Landi and Barbara G. Ryder.

A safe approximate algorithm for interprocedural pointer aliasing. ACM SIG-PLAN Notices, 39(4):473–489, April 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boehm:2004:SEC

[1049] Hans J. Boehm. Space efficient conservative garbage collection. ACM SIG-PLAN Notices, 39(4):490–501, April 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Flanagan:2004:ECC

[1050] Cormac Flanagan, Amr Sabry, Bruce F. Duba, and Matthias Felleisen. The essence of compiling with continuations. ACM SIGPLAN Notices, 39(4): 502–514, April 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kerns:2004:BSI

[1051] Daniel R. Kerns and Susan J. Eggers. Balanced scheduling: instruction scheduling when memory latency is uncertain. ACM SIGPLAN Notices, 39 (4):515–527, April 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Srivastava:2004:ASB

[1052] Amitabh Srivastava and Alan Eustace. ATOM: a system for building customized program analysis tools. ACM SIGPLAN Notices, 39(4):528– 539, April 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lee:2004:OMR

[1053] Peter Lee and Mark Leone. Optimizing ML with run-time code generation. ACM SIGPLAN Notices, 39(4):

540–553, April 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tarditi:2004:TTD

[1054] David Tarditi, Greg Morrisett, Perry Cheng, Chris Stone, Robert Harper, and Peter Lee. TIL: a type-directed, optimizing compiler for ML. ACM SIGPLAN Notices, 39(4):554–567, April 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ammons:2004:IDF

[1055] Glenn Ammons and James R. Larus.

Improving data-flow analysis with path profiles. ACM SIGPLAN Notices, 39 (4):568–582, April 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bacon:2004:TLF

[1056] David F. Bacon, Ravi Konuru, Chet Murthy, and Mauricio J. Serrano. Thin locks: featherweight synchronization for Java. ACM SIGPLAN Notices, 39 (4):583–595, April 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bodik:2004:CRR

[1057] Rastislav Bodík, Rajiv Gupta, and Mary Lou Soffa. Complete removal of redundant expressions. ACM SIG-PLAN Notices, 39(4):596–611, April 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Necula:2004:DIC

[1058] George C. Necula and Peter Lee. The design and implementation of a certifying compiler. ACM SIGPLAN Notices, 39(4):612–625, April 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Blelloch:2004:BTS

[1059] Guy E. Blelloch and Perry Cheng. On bounding time and space for multiprocessor garbage collection. ACM SIG-PLAN Notices, 39(4):626–641, April 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frigo:2004:FFT

[1060] Matteo Frigo. A fast Fourier transform compiler. ACM SIGPLAN Notices, 39 (4):642–655, April 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Grant:2004:RES

[1061] Brian Grant, Matthai Philipose, Markus Mock, Craig Chambers, and Susan J. Eggers. A retrospective on: “an evaluation of staged run-time optimizations in DyC’’. ACM SIGPLAN Notices, 39(4):656–669, April 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shan:2004:STA

[1062] Chung chieh Shan. Sexy types in action. ACM SIGPLAN Notices, 39 (5):15–22, May 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ghosh:2004:GJC

[1063] Debasish Ghosh. Generics in Java and C++: a comparative model. ACM SIGPLAN Notices, 39(5):40–47, May 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wells:2004:UIR

[1064] Paul Damian Wells. A universal intermediate representation for massively parallel software development. ACM SIGPLAN Notices, 39(5):48–57, May 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Juric:2004:JRR

[1065] Matjaz B. Juric, Bostjan Kezmah, Marjan Hericko, Ivan Rozman, and Ivan Vezocnik. Java RMI, RMI tunneling and Web services comparison and performance analysis. ACM SIG-PLAN Notices, 39(5):58–65, May 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kantorovitz:2004:LAT

[1066] Isaiah Pinchas Kantorovitz. Lexical analysis tool. ACM SIGPLAN Notices, 39(5):66–74, May 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Henzinger:2004:RCC

[1067] Thomas A. Henzinger, Ranjit Jhala, and Rupak Majumdar. Race checking by context inference. ACM SIG-PLAN Notices, 39(6):1–13, May 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Qadeer:2004:KKI

[1068] Shaz Qadeer and Dinghao Wu. KISS: keep it simple and sequential. ACM SIGPLAN Notices, 39(6):14–24, May 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yahav:2004:VSP

[1069] Eran Yahav and G. Ramalingam. Verifying safety properties using separation and heterogeneous abstractions. ACM SIGPLAN Notices, 39(6):25–34, May 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Michael:2004:SLF

[1070] Maged M. Michael. Scalable lockfree dynamic memory allocation. ACM SIGPLAN Notices, 39(6):35–46, May 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Flatt:2004:KSS

[1071] Matthew Flatt and Robert Bruce Findler. Kill-safe synchronization abstractions. ACM SIGPLAN Notices, 39 (6):47–58, May 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Johnson:2004:MCP

[1072] Troy A. Johnson, Rudolf Eigenmann, and T. N. Vijaykumar. Min-cut program decomposition for thread-level speculation. ACM SIGPLAN Notices, 39(6):59–70, May 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Du:2004:CDC

[1073] Zhao-Hui Du, Chu-Cheow Lim, Xiao-Feng Li, Chen Yang, Qingyu Zhao, and Tin-Fook Ngai. A cost-driven compilation framework for speculative parallelization of sequential programs. ACM SIGPLAN Notices, 39(6):71–81, May 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Eichenberger:2004:VSA

[1074] Alexandre E. Eichenberger, Peng Wu, and Kevin O’Brien. Vectorization for SIMD architectures with alignment constraints. ACM SIGPLAN Notices, 39(6):82–93, May 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:2004:CED

[1075] Xiangyu Zhang and Rajiv Gupta.

Cost effective dynamic program slicing. ACM SIGPLAN Notices, 39(6):94–106, May 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Collberg:2004:DPB

[1076] C. Collberg, E. Carter, S. Debray, A. Huntwork, J. Kececioglu, C. Linn, and M. Stepp. Dynamic path-based software watermarking. ACM SIG-PLAN Notices, 39(6):107–118, May 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wang:2004:PAA

[1077] Cheng Wang and Zhiyuan Li. Parametric analysis for adaptive computation offloading. ACM SIGPLAN Notices, 39 (6):119–130, May 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Whaley:2004:CBC

[1078] John Whaley and Monica S. Lam. Cloning-based context-sensitive pointer alias analysis using binary decision diagrams. ACM SIGPLAN Notices, 39 (6):131–144, May 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhu:2004:SPA

[1079] Jianwen Zhu and Silvian Calman. Symbolic pointer analysis revisited. ACM SIGPLAN Notices, 39(6):145– 157, May 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lhotak:2004:JBB

[1080] Ondřej Lhoták and Laurie Hendren. Jedd: a BDD-based relational extension of Java. ACM SIGPLAN Notices, 39(6):158–169, May 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kulkarni:2004:FSE

[1081] Prasad Kulkarni, Stephen Hines, Jason Hiser, David Whalley, Jack Davidson, and Douglas Jones. Fast searches for effective optimization phase sequences. ACM SIGPLAN Notices, 39 (6):171–182, May 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cavazos:2004:IHD

[1082] John Cavazos, J. Eliot, and B. Moss. Inducing heuristics to decide whether to schedule. ACM SIGPLAN Notices, 39(6):183–194, May 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vachharajani:2004:LSS

[1083] Manish Vachharajani, Neil Vachharajani, and David I. August. The liberty structural specification language: a high-level modeling language for component reuse. ACM SIGPLAN Notices, 39(6):195–206, May 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kodumal:2004:SCC

[1084] John Kodumal and Alex Aiken. The set constraint/CFL reachability connection in practice. ACM SIGPLAN Notices, 39(6):207–218, May 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Liu:2004:PRP

[1085] Yanhong A. Liu, Tom Rothamel, Fuxiang Yu, Scott D. Stoller, and Nanjun Hu. Parametric regular path queries. ACM SIGPLAN Notices, 39 (6):219–230, May 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Venet:2004:PES

[1086] Arnaud Venet and Guillaume Brat.

Precise and efficient static array bound checking for large embedded C programs. ACM SIGPLAN Notices, 39 (6):231–242, May 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chin:2004:RIO

[1087] Wei-Ngan Chin, Florin Craciun, Shengchao Qin, and Martin Rinard. Region inference for an object-oriented language. ACM SIGPLAN Notices, 39 (6):243–254, May 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhong:2004:ARS

[1088] Yutao Zhong, Maksim Orlovich, Xipeng Shen, and Chen Ding. Array regrouping and structure splitting using whole-program reference affinity. ACM SIGPLAN Notices, 39(6): 255–266, May 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Adl-Tabatabai:2004:PIB

[1089] Ali-Reza Adl-Tabatabai, Richard L. Hudson, Mauricio J. Serrano, and Sreenivas Subramoney. Prefetch inection based on hardware monitoring and object metadata. ACM SIG-PLAN Notices, 39(6):267–276, May 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smith:2004:GAG

[1090] Michael D. Smith, Norman Ramsey, and Glenn Holloway. A generalized algorithm for graph-coloring register allocation. ACM SIGPLAN Notices, 39 (6):277–288, May 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhuang:2004:BRA

[1091] Xiaotong Zhuang and Santosh Pande. Balancing register allocation across threads for a multithreaded network processor. ACM SIGPLAN Notices, 39 (6):289–300, May 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Antochi:2004:GGB

[1092] Iosif Antochi, Ben Juurlink, Stamatis Vassiliadis, and Petri Liuha. Graal-Bench: a 3D graphics benchmark suite for mobile phones. ACM SIG-PLAN Notices, 39(7):1–9, July 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Martin:2004:MSE

[1093] Thomas Martin, Mark Jones, Joshua Edmison, Tanwir Sheikh, and Zahi Nakad. Modeling and simulating electronic textile applications. ACM SIG-PLAN Notices, 39(7):10–19, July 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Willmann:2004:SLB

[1094] Paul Willmann, Michael Brogioli, and Vijay S. Pai. Spinach: a liberty-based simulator for programmable network interface architectures. ACM SIG-PLAN Notices, 39(7):20–29, July 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Conway:2004:NDS

[1095] Christopher L. Conway and Stephen A.

Edwards. NDL: a domain-specific language for device drivers. ACM SIG-PLAN Notices, 39(7):30–36, July 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kumar:2004:AST

[1096] Nagendra J. Kumar, Siddhartha Shivshankar, and Alexander G. Dean. Asynchronous software thread integration for efficient software. ACM SIG-PLAN Notices, 39(7):37–46, July 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Qin:2004:FCM

[1097] Wei Qin, Subramanian Rajagopalan, and Sharad Malik. A formal concurrency model based architecture description language for synthesis of software development tools. ACM SIG-PLAN Notices, 39(7):47–56, July 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jejurikar:2004:PSF

[1098] Ravindra Jejurikar and Rajesh Gupta. Procrastination scheduling in fixed priority real-time systems. ACM SIG-PLAN Notices, 39(7):57–66, July 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhuang:2004:PEP

[1099] Xiaotong Zhuang and Santosh Pande. Power-efficient prefetching via bitdifferential offset assignment on embedded processors. ACM SIGPLAN Notices, 39(7):67–77, July 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pokam:2004:SSM

[1100] Gilles Pokam, Olivier Rochecouste, André Seznec, and François Bodin. Speculative software management of datapath-width for energy optimization. ACM SIGPLAN Notices, 39 (7):78–87, July 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Im:2004:DVS

[1101] Chaeseok Im and Soonhoi Ha. Dynamic voltage scaling for real-time multi-task scheduling using buffers. ACM SIGPLAN Notices, 39(7):88–94, July 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Li:2004:TBB

[1102] Lian Li and Jingling Xue. A tracebased binary compilation framework for energy-aware computing. ACM SIGPLAN Notices, 39(7):95–106, July 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lapalme:2004:ENN

[1103] J. Lapalme, E. M. Aboulhamid, G. Nicolescu, L. Charest, F. R. Boyer, J. P. David, and G. Bois. ESys.Net: a new solution for embedded systems modeling and simulation. ACM SIG-PLAN Notices, 39(7):107–114, July 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Contreras:2004:XPS

[1104] Gilberto Contreras, Margaret Martonosi, Jinzhan Peng, Roy Ju, and Guei-Yuan Lueh. XTREM: a power simulator for the Intel XScale^{T M}; core. ACM SIGPLAN Notices, 39(7):115–125, July 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kastens:2004:FDI

[1105] Uwe Kastens, Dinh Khoi Le, Adrian Slowik, and Michael Thies. Feedback driven instruction-set extension. ACM SIGPLAN Notices, 39(7):126–135, July 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Patil:2004:CSI

[1106] Kaustubh Patil, Kiran Seth, and Frank Mueller. Compositional static instruction cache simulation. ACM SIG-PLAN Notices, 39(7):136–145, July 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Starner:2004:MCI

[1107] Johan Stärner and Lars Asplund. Measuring the cache interference cost in preemptive real-time systems. ACM SIGPLAN Notices, 39(7):146–154, July 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:2004:ACU

[1108] Lingli Zhang and Chandra Krintz. Adaptive code unloading for resourceconstrained JVMs. ACM SIGPLAN Notices, 39(7):155–164, July 2004. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Li:2004:ACF

[1109] Peng Li and Steve Zdancewic. Advanced control flow in Java card programming. ACM SIGPLAN Notices, 39(7):165–174, July 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zeng:2004:GFC

[1110] Jia Zeng, Cristian Soviani, and Stephen A. Edwards. Generating fast code from concurrent program dependence graphs. ACM SIGPLAN Notices, 39(7):175–181, July 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hiser:2004:EEM

[1111] Jason D. Hiser and Jack W. Davidson.

EMBARC: an efficient memory bank assignment algorithm for retargetable compilers. ACM SIGPLAN Notices, 39 (7):182–191, July 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhuang:2004:HMR

[1112] Xiaotong Zhuang, Tao Zhang, and Santosh Pande. Hardware-managed register allocation for embedded processors. ACM SIGPLAN Notices, 39(7): 192–201, July 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Daveau:2004:RRA

[1113] Jean-Marc Daveau, Thomas Thery, Thierry Lepley, and Miguel Santana. A retargetable register allocation framework for embedded processors. ACM SIGPLAN Notices, 39(7):202–210, July 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

DeBus:2004:LTO

[1114] Bruno De Bus, Bjorn De Sutter, Ludo Van Put, Dominique Chanet, and Koen De Bosschere. Link-time optimization of ARM binaries. ACM SIG-PLAN Notices, 39(7):211–220, July 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Scholz:2004:OST

[1115] Bernhard Scholz, Nigel Horspool, and Jens Knoop. Optimizing for space and time usage with speculative partial redundancy elimination. ACM SIGPLAN Notices, 39(7):221–230, July 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Almagor:2004:FEC

[1116] L. Almagor, Keith D. Cooper, Alexander Grosul, Timothy J. Harvey, Steven W. Reeves, Devika Subramanian, Linda Torczon, and Todd Waterman. Finding effective compilation sequences. ACM SIGPLAN Notices, 39 (7):231–239, July 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Saputra:2004:CPR

[1117] H. Saputra, G. Chen, R. Brooks, N. Vijaykrishnan, M. Kandemir, and M. J. Irwin. Code protection for resourceconstrained embedded devices. ACM SIGPLAN Notices, 39(7):240–248, July 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Guo:2004:IDR

[1118] Zhi Guo, Betul Buyukkurt, and Walid Najjar. Input data reuse in compiling window operations onto reconfigurable hardware. ACM SIGPLAN Notices, 39 (7):249–256, July 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wasowski:2004:FSE

[1119] Andrzej Wasowski. Flattening statecharts without explosions. ACM SIG-PLAN Notices, 39(7):257–266, July 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:2004:EF

[1120] Paul Frenger. Embed with Forth. ACM SIGPLAN Notices, 39(8):8–11, August 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Creak:2004:PNA

[1121] Alan Creak. Parsing by numbers and asparagus. ACM SIGPLAN Notices, 39 (8):12–19, August 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reid:2004:OF

[1122] John Reid. An overview of Fortran 2003. ACM SIGPLAN Notices, 39(8): 31–38, August 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wu:2004:SZS

[1123] Fangjun Wu and Tong Yi. Slicing Z specifications. ACM SIGPLAN Notices, 39(8):39–48, August 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bronnikov:2004:PAP

[1124] Dmitri Bronnikov. A practical adoption of partial redundancy elimination. ACM SIGPLAN Notices, 39(8):49–53, August 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

deOliveiraGuimaraes:2004:CST

[1125] José de Oliveira Guimarães. Closures for statically-typed object-oriented languages. ACM SIGPLAN Notices, 39 (8):54–60, August 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Babu:2004:MDM

[1126] Chitra Babu and D. Janakiram. Method driven model: a unified model for an object composition language. ACM SIGPLAN Notices, 39(8):61–71, August 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sandya:2004:JJL

[1127] S. M. Sandya. Jazzing up JVMs with off-line profile data: does it pay? ACM SIGPLAN Notices, 39(8):72–80, August 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ramsey:2004:CCI

[1128] Norman Ramsey and Simon L. Peyton Jones. The C–compiler infrastructure. ACM SIGPLAN Notices, 39 (9):1, September 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Launchbury:2004:GHA

[1129] John Launchbury. Galois: high assurance software. ACM SIGPLAN Notices, 39(9):3, September 2004. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marlow:2004:MFC

[1130] Simon Marlow and Simon Peyton Jones. Making a fast curry: push/ enter vs. eval/apply for higher-order languages. ACM SIGPLAN Notices, 39(9):4–15, September 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Herman:2004:ISA

[1131] David Herman and Philippe Meunier.

Improving the static analysis of embedded languages via partial evaluation. ACM SIGPLAN Notices, 39(9): 16–27, September 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Christiansen:2004:SDW

[1132] Jan Christiansen and Frank Huch.

Searching for deadlocks while debugging Concurrent Haskell programs. ACM SIGPLAN Notices, 39(9):28– 39, September 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ariola:2004:TTF

[1133] Zena M. Ariola, Hugo Herbelin, and Amr Sabry. A type-theoretic foundation of continuations and prompts. ACM SIGPLAN Notices, 39(9):40– 53, September 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wand:2004:RMB

[1134] Mitchell Wand and Dale Vaillancourt. Relating models of backtracking. ACM SIGPLAN Notices, 39(9): 54–65, September 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Graham:2004:DMW

[1135] Paul Graham. Don’t make the wrong mistakes: programming as debugging. ACM SIGPLAN Notices, 39(9): 66, September 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Broberg:2004:REP

[1136] Niklas Broberg, Andreas Farre, and Josef Svenningsson. Regular expression patterns. ACM SIGPLAN Notices, 39(9):67–78, September 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shivers:2004:MRF

[1137] Olin Shivers and David Fisher. Multireturn function call. ACM SIGPLAN Notices, 39(9):79–89, September 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tolmach:2004:IFL

[1138] Andrew Tolmach, Sergio Antoy, and Marius Nita. Implementing functional logic languages using multiple threads and stores. ACM SIGPLAN Notices, 39(9):90–102, September 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fluet:2004:MR

[1139] Matthew Fluet and Greg Morrisett.

Monadic regions. ACM SIGPLAN Notices, 39(9):103–114, September 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tse:2004:TDP

[1140] Stephen Tse and Steve Zdancewic.

Translating dependency into parametricity. ACM SIGPLAN Notices, 39 (9):115–125, September 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Colazzo:2004:TPC

[1141] Dario Colazzo, Giorgio Ghelli, Paolo Manghi, and Carlo Sartiani. Types for path correctness of XML queries. ACM SIGPLAN Notices, 39(9):126– 137, September 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Neergaard:2004:TPI

[1142] Peter Møller Neergaard and Harry G.

Mairson. Types, potency, and idempotency: why nonlinearity and amnesia make a type system work. ACM SIG-PLAN Notices, 39(9):138–149, September 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gauthier:2004:NMF

[1143] Nadji Gauthier and François Pottier. Numbering matters: first-order canonical forms for second-order recursive types. ACM SIGPLAN Notices, 39 (9):150–161, September 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wiger:2004:YIF

[1144] Ulf Wiger. 20 years of industrial functional programming. ACM SIGPLAN Notices, 39(9):162, September 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Honda:2004:PLP

[1145] Kohei Honda. From process logic to program logic. ACM SIGPLAN Notices, 39(9):163–174, September 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yu:2004:VSP

[1146] Dachuan Yu and Zhong Shao. Verification of safety properties for concurrent assembly code. ACM SIGPLAN Notices, 39(9):175–188, September 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Blume:2004:SCM

[1147] Matthias Blume and David McAllester. A sound (and complete) model of contracts. ACM SIGPLAN Notices, 39 (9):189–200, September 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sarkar:2004:NIC

[1148] Dipanwita Sarkar, Oscar Waddell, and R. Kent Dybvig. A nanopass infrastructure for compiler education. ACM SIGPLAN Notices, 39(9):201– 212, September 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Forsberg:2004:FM

[1149] Markus Forsberg and Aarne Ranta.

Functional morphology. ACM SIG-PLAN Notices, 39(9):213–223, September 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Findler:2004:SFP

[1150] Robert Bruce Findler and Matthew Flatt. Slideshow: functional presentations. ACM SIGPLAN Notices, 39 (9):224–235, September 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hinze:2004:GM

[1151] Ralf Hinze. Generics for the masses.

ACM SIGPLAN Notices, 39(9):236– 243, September 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lammel:2004:SMB

[1152] Ralf Lämmel and Simon Peyton Jones.

Scrap more boilerplate: reflection, zips, and generalised casts. ACM SIG-PLAN Notices, 39(9):244–255, September 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Perera:2004:RRB

[1153] Roly Perera. Refactoring: to the Rubicon… and beyond! ACM SIG-PLAN Notices, 39(10):2–3, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Freeman:2004:JSR

[1154] Steve Freeman, Tim Mackinnon, Nat Pryce, and Joe Walnes. jMock: supporting responsibility-based design with mock objects. ACM SIGPLAN Notices, 39(10):4–5, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Aldrich:2004:MISa

[1155] Jonathan Aldrich, David Garlan, Bradley Schmerl, and Tony Tseng. Modeling and implementing software architecture with acme and archJava. ACM SIGPLAN Notices, 39(10):6–7, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fuhrer:2004:ARE

[1156] Robert Fuhrer, Frank Tip, and Adam Kieżun. Advanced refactorings in eclipse. ACM SIGPLAN Notices, 39 (10):8, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McCormick:2004:JFY

[1157] Edward McCormick and Kris De Volder. JQuery: finding your way through tangled code. ACM SIG-PLAN Notices, 39(10):9–10, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kaundinya:2004:MEDa

[1158] Murali Kaundinya and Ali Syed. Modeling event driven applications with a specification language (MEDASL). ACM SIGPLAN Notices, 39(10):11– 12, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sazawal:2004:DSP

[1159] Vibha Sazawal and David Notkin. Design snippets: partial design representations extracted from source code. ACM SIGPLAN Notices, 39(10):13– 14, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fowler:2004:MPR

[1160] Matthew Fowler and Brahm van Niekerk. Meta-programming for the real world. ACM SIGPLAN Notices, 39 (10):15, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ostlund:2004:JOA

[1161] Helena Åberg Östlund. JRA: offline analysis of runtime behaviour. ACM SIGPLAN Notices, 39(10):16–17, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Spinczyk:2004:MBS

[1162] Olaf Spinczyk and Danilo Beuche.

Modeling and building software product lines with eclipse. ACM SIGPLAN Notices, 39(10):18–19, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:2004:PRA

[1163] Charles Zhang and Hans-Arno Jacobsen. PRISM is research in aSpect mining. ACM SIGPLAN Notices, 39(10): 20–21, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bhattacharrya:2004:SDE

[1164] Arnab Bhattacharrya and Robert Fuhrer. Smell detection for eclipse. ACM SIGPLAN Notices, 39(10):22, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lippert:2004:AAEa

[1165] Martin Lippert. AJEER: an aspectJenabled eclipse runtime. ACM SIG-PLAN Notices, 39(10):23–24, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Akers:2004:PTR

[1166] Robert L. Akers, Ira D. Baxter, and Michael Mehlich. Program transformations for re-engineering C++ components [OOPSLA/GPCE]. ACM SIG-PLAN Notices, 39(10):25–26, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gray:2004:CSG

[1167] Jeff Gray, Jing Zhang, Suman Roychoudhury, and Ira Baxter. C-SAW and genAWeave: a two-level aspect weaving toolsuite. ACM SIGPLAN Notices, 39(10):27–28, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tarr:2004:CME

[1168] Peri Tarr, William Chung, William Harrison, Vincent Kruskal, Harold Ossher, Stanley M. Sutton, Jr., Andrew Clement, Matthew Chapman, Helen Hawkins, and Sian January. The concern manipulation environment [OOPSLA/GPCE]. ACM SIGPLAN Notices, 39(10):29–30, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hauswirth:2004:PEU

[1169] Matthias Hauswirth, Peter F. Sweeney, and Amer Diwan. Performance explorer: understanding Java application behavior. ACM SIGPLAN Notices, 39 (10):31–32, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gordon:2004:C

[1170] Donald Gordon, James Noble, and Robert Biddle. ConstrainedJava. ACM SIGPLAN Notices, 39(10):33–34, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nguyen:2004:OOS

[1171] Tien N. Nguyen, Ethan V. Munson, and John T. Boyland. Objectoriented, structural software configuration management. ACM SIGPLAN Notices, 39(10):35–36, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Parigot:2004:TDD

[1172] Didier Parigot. Towards domaindriven development: the smartTools software factory. ACM SIGPLAN Notices, 39(10):37–38, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tolvanen:2004:MDS

[1173] Juha-Pekka Tolvanen. MetaEdit+: domain-specific modeling for full code generation demonstrated [GPCE]. ACM SIGPLAN Notices, 39(10):39– 40, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Guerrero:2004:IDM

[1174] Miguel Guerrero, Edward Pizzi, Robert Rosenbaum, Kedar Swadi, and Walid Taha. Implementing DSLs in metaO-Caml. ACM SIGPLAN Notices, 39 (10):41–42, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Eichberg:2004:XCA

[1175] Michael Eichberg and Thorsten Schäfer. XIRC: cross-artifact information retrieval [GPCE]. ACM SIGPLAN Notices, 39(10):43–44, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cao:2004:MDA

[1176] Lan Cao. Modeling dynamics of agile software development. ACM SIG-PLAN Notices, 39(10):46–47, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gestwicki:2004:IVO

[1177] Paul V. Gestwicki. Interactive visualization of object-oriented programs. ACM SIGPLAN Notices, 39(10):48– 49, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Potanin:2004:GOP

[1178] Alex Potanin, James Noble, and Robert Biddle. Generic ownership: practical ownership control in programming languages. ACM SIGPLAN Notices, 39(10):50–51, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Roychoudhury:2004:LIAa

[1179] Suman Roychoudhury. A languageindependent approach to software maintenance using grammar adapters. ACM SIGPLAN Notices, 39(10):52– 53, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Siek:2004:MG

[1180] Jeremy Siek. Modular generics. ACM SIGPLAN Notices, 39(10):54–55, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nienaltowski:2004:EDR

[1181] Piotr Nienaltowski. Efficient data race and deadlock prevention in concurrent object-oriented programs. ACM SIG-PLAN Notices, 39(10):56–57, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Trofin:2004:FRRa

[1182] Mircea Trofin. A framework for removing redundant context management services in Enterprise JavaBeans application servers. ACM SIGPLAN Notices, 39(10):58–59, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Silaghi:2004:RDAa

[1183] Raul Silaghi. Refining designs along middleware-specific concerndimensions at different MDA-levels of abstraction. ACM SIGPLAN Notices, 39(10):60–62, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schmolitzky:2004:OFI

[1184] Axel Schmolitzky. “objects first, interfaces next” or interfaces before inheritance. ACM SIGPLAN Notices, 39(10):64–67, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ventura:2004:AWC

[1185] Phil Ventura, Christopher Egert, and Adrienne Decker. Ancestor worship in CS1: on the primacy of arrays. ACM SIGPLAN Notices, 39(10):68–72, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Henriksen:2004:GCO

[1186] Poul Henriksen and Michael Kölling.

greenfoot: combining object visualisation with interaction. ACM SIGPLAN Notices, 39(10):73–82, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Malan:2004:ECD

[1187] Katherine Malan and Ken Halland. Examples that can do harm in learning programming. ACM SIGPLAN Notices, 39(10):83–87, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fenstermacher:2004:IMI

[1188] Kurt D. Fenstermacher. If I had a model, I’d model in the mornin’. ACM SIGPLAN Notices, 39(10):88–95, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bruce:2004:EDP

[1189] Kim B. Bruce and Andrea Danyluk.

Event-driven programming facilitates learning standard programming concepts. ACM SIGPLAN Notices, 39(10): 96–100, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cheng:2004:AFS

[1190] Eric Cheng, Dung Zung Nguyen, Mathias Ricken, and Stephen Wong. Abstract factories and the shape calculator. ACM SIGPLAN Notices, 39(10): 101–102, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wick:2004:UGL

[1191] Michael R. Wick. Using the game of life to introduce freshman students to the power and elegance of design patterns. ACM SIGPLAN Notices, 39(10): 103–105, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nevison:2004:CAP

[1192] Christopher H. Nevison. From concrete to abstract: the power of generalization. ACM SIGPLAN Notices, 39(10): 106–108, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cheng:2004:MBS

[1193] Eric Cheng, Dung Zung Nguyen, Mathias Ricken, and Stephen Wong. Marine biology simulation. ACM SIGPLAN Notices, 39(10):109–110, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Noble:2004:NNPa

[1194] James Noble and Robert Biddle. Notes on notes on postmodern programming: radio edit. ACM SIGPLAN Notices, 39(10):112–115, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sheard:2004:LFa

[1195] Tim Sheard. Languages of the future. ACM SIGPLAN Notices, 39(10): 116–119, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Singh:2004:PPPa

[1196] Munindar P. Singh, Amit K. Chopra, Nirmit V. Desai, and Ashok U. Mallya. Protocols for processes: programming in the large for open systems (extended abstract). ACM SIGPLAN Notices, 39(10):120–123, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Edwards:2004:ECPa

[1197] Jonathan Edwards. Example centric programming. ACM SIGPLAN Notices, 39(10):124, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marick:2004:MWOa

[1198] Brian Marick. Methodology work is ontology work. ACM SIGPLAN Notices, 39(10):125, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

West:2004:LLAa

[1199] David West. Looking for love: (in all the wrong places). ACM SIGPLAN Notices, 39(10):126–127, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhao:2004:GOPa

[1200] Yinliang Zhao. Granule-oriented programming (extended abstract). ACM SIGPLAN Notices, 39(10):128–131, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hovemeyer:2004:FBEa

[1201] David Hovemeyer and William Pugh. Finding bugs is easy. ACM SIGPLAN Notices, 39(10):132–136, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Miller:2004:MDA

[1202] Granville Miller, Scott Ambler, Steve Cook, Stephen Mellor, Karl Frank, and Jon Kern. Model driven architecture: the realities, a year later. ACM SIG-PLAN Notices, 39(10):138–140, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Haungs:2004:SDA

[1203] Jim Haungs, Martin Fowler, Ralph Johnson, Steve McConnell, and Richard Gabriel. Software development: arts & crafts or math & science? ACM SIGPLAN Notices, 39(10): 141–142, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fowler:2004:GJV

[1204] Martin Fowler, Don Box, Anders Hejlsberg, Alan Knight, Rob High, and John Crupi. The great J2EE vs. Microsoft .NET shootout. ACM SIGPLAN Notices, 39(10):143–144, October 2004.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fraser:2004:COG

[1205] Steven Fraser, Lougie Anderson, Ron Crocker, Richard Gabriel, Martin Fowler, Ricardo Lopez, and Dave Thomas. Challenges in outsourcing and global development: how will your job change? ACM SIGPLAN Notices, 39(10):145–147, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fraser:2004:RCS

[1206] Steven Fraser, Angela Martin, Robert Biddle, David Hussman, Granville Miller, Mary Poppendieck, Linda Rising, and Mark Striebeck. The role of the customer in software development: the XP customer — fad or fashion? ACM SIGPLAN Notices, 39(10): 148–150, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Josuttis:2004:VUI

[1207] Nicolai M. Josuttis, Jutta Eckstein, Linda Rising, Lise B. Hvatum, Mary Lynn Manns, and Rebecca Wirfs-Brock. The view: the ultimate IT chat. ACM SIGPLAN Notices, 39(10): 151–152, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Eliassen:2004:QBR

[1208] Frank Eliassen, Richard Staehli, Gordon Blair, and Jan Øyvind Aagedal. QuA: building with reusable QoSaware components. ACM SIGPLAN Notices, 39(10):154–155, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Aldrich:2004:MISb

[1209] Jonathan Aldrich, David Garlan, Bradley Schmerl, and Tony Tseng. Modeling and implementing software architecture with acme and archJava. ACM SIGPLAN Notices, 39(10):156– 157, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Carleton:2004:IDP

[1210] Lawrence Carleton. An invitation to the dance of progress. ACM SIGPLAN Notices, 39(10):158–159, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Niculescu:2004:OOO

[1211] Virginia Niculescu. OOLACA: an object oriented library for abstract and computational algebra. ACM SIG-PLAN Notices, 39(10):160–161, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sobering:2004:PCV

[1212] Geoff Sobering, Levi Cook, and Steve Anderson. Pseudo-classes: very simple and lightweight mockobject-like classes for unit-testing. ACM SIGPLAN Notices, 39(10):162–163, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Paquet:2004:TFG

[1213] Joey Paquet, Aihua Wu, and Peter Grogono. Towards a framework for the general intensional programming compiler in the GIPSY. ACM SIGPLAN Notices, 39(10):164–165, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kulesza:2004:AOG

[1214] Uirá Kulesza, Alessandro Garcia, and Carlos Lucena. An aspect-oriented generative approach. ACM SIGPLAN Notices, 39(10):166–167, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kaundinya:2004:MEDb

[1215] Murali Kaundinya and Ali Syed. Modeling event driven applications with a specification language (MEDASL). ACM SIGPLAN Notices, 39(10):168– 169, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hussain:2004:IQC

[1216] Tauqeer Hussain, Shafay Shamail, and Mian M. Awais. Improving quality in conceptual modeling. ACM SIGPLAN Notices, 39(10):170–171, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Buytaert:2004:BAJ

[1217] Dries Buytaert, Andy Georges, Lieven Eeckhout, and Koen De Bosschere. Bottleneck analysis in Java applications using hardware performance monitors. ACM SIGPLAN Notices, 39(10):

172–173, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Abrahamsson:2004:MDA

[1218] Pekka Abrahamsson, Antti Hanhineva, Hanna Hulkko, Tuomas Ihme, Juho Jäälinoja, Mikko Korkala, Juha Koskela, Pekka Kyllönen, and Outi Salo. Mobile-D: an agile approach for mobile application development. ACM SIGPLAN Notices, 39(10):174– 175, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Szep:2004:DDS

[1219] A. Szep, R. Smeikal, M. Jandl, and K. M. Goeschka. Dependable distributed systems. ACM SIGPLAN Notices, 39(10):176–177, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Khaled:2004:VLO

[1220] Rilla Khaled, Anna Maria Luxton, James Noble, Leo Ferres, Judy Brown, and Robert Biddle. Visualisation for learning OOP, using AOP and eclipse. ACM SIGPLAN Notices, 39(10):178– 179, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lippert:2004:AAEb

[1221] Martin Lippert. AJEER: an aspectJenabled eclipse runtime. ACM SIG-PLAN Notices, 39(10):180–181, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vogler:2004:OMI

[1222] Gabriel Vögler, Thomas Flor, Hajo Eichler, and Matthias Kasprowicz. An open model infrastructure for automotive software. ACM SIGPLAN Notices, 39(10):182–183, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Phattarasukol:2004:DPI

[1223] Somsak Phattarasukol and Daisy Sang.

Design pattern integrated tool. ACM SIGPLAN Notices, 39(10):184–185, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stephenson:2004:AMO

[1224] Ben Stephenson and Wade Holst. Advancements in multicode optimization. ACM SIGPLAN Notices, 39(10):186– 187, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lancia:2004:JPA

[1225] Maurizio Lancia, Paola Garzenini, Roberto Puccinelli, and Alessio Marchetti. J2EE for the public administration: a success story. ACM SIG-PLAN Notices, 39(10):188–189, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Holst:2004:MEU

[1226] Wade Holst. Meta: extending and unifying languages. ACM SIGPLAN Notices, 39(10):190–191, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Amir:2004:UPS

[1227] Rafik Amir and Amir Zeid. A UML profile for service oriented architectures. ACM SIGPLAN Notices, 39(10): 192–193, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chan:2004:ATA

[1228] Hoi Chan and Trieu C. Chieu. An approach and tools to automate externalization of application logic. ACM SIG-PLAN Notices, 39(10):194–195, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Skotiniotis:2004:CAC

[1229] Therapon Skotiniotis and David H. Lorenz. Cona: aspects for contracts and contracts for aspects. ACM SIG-PLAN Notices, 39(10):196–197, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Freeman:2004:CAT

[1230] Russ Freeman and Phil Webb. <CTRL> + <ALT> + <TOOL PARADIGM SHIFT>? ACM SIG-PLAN Notices, 39(10):198–199, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Washizaki:2004:SSJ

[1231] Hironori Washizaki and Yoshiaki Fukazawa. A search system for Java programs by using extracted JavaBeans components. ACM SIGPLAN Notices, 39(10):200–201, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cole:2004:DRA

[1232] Leonardo Cole and Paulo Borba. Deriving refactorings for aspectJ. ACM SIG-PLAN Notices, 39(10):202–203, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pugh:2004:RRW

[1233] Bill Pugh and Jaime Spacco. RUBiS revisited: why J2EE benchmarking is hard. ACM SIGPLAN Notices, 39(10): 204–205, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ueda:2004:MMA

[1234] Leo Kazuhiro Ueda and Fabio Kon.

Mobile musical agents: the andante project. ACM SIGPLAN Notices, 39 (10):206–207, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bettini:2004:CCM

[1235] Lorenzo Bettini, Viviana Bono, and Silvia Likavec. A core calculus of mixins and incomplete objects. ACM SIG-PLAN Notices, 39(10):208–209, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cabral:2004:RCI

[1236] Bruno Cabral, Paulo Marques, and Luís Silva. RAIL: code instrumentation for .NET. ACM SIGPLAN Notices, 39(10):210–211, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boshernitsan:2004:IIS

[1237] Marat Boshernitsan and Susan L. Graham. iXj: interactive source-to-source transformations for Java. ACM SIG-PLAN Notices, 39(10):212–213, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rheinheimer:2004:WAB

[1238] Leticia R. Rheinheimer, Junior M. Martins, and Sérgio Crespo C. S. Pinto. WSAgent: an agent based on Web services to promote interoperability between heterogeneous systems in the health domain. ACM SIGPLAN Notices, 39(10):214–215, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kojarski:2004:AFC

[1239] Sergei Kojarski and David H. Lorenz. AOP as a first class reflective mechanism. ACM SIGPLAN Notices, 39(10): 216–217, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gokhale:2004:CAC

[1240] Aniruddha Gokhale, Krishnakumar Balasubramanian, and Tao Lu. CoS-MIC: addressing crosscutting deployment and configuration concerns of distributed real-time and embedded systems. ACM SIGPLAN Notices, 39(10): 218–219, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hamza:2004:SSO

[1241] Haitham S. Hamza. SODA: a stabilityoriented domain analysis method. ACM SIGPLAN Notices, 39(10):220– 221, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ivkovic:2004:MSP

[1242] Igor Ivkovic and Kostas Kontogiannis. Model synchronization as a problem of maximizing model dependencies. ACM SIGPLAN Notices, 39(10): 222–223, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Holland:2004:RCD

[1243] Simon Holland. Reflective composition: the declarative composition of roles to unify objects, roles, and aspects. ACM SIGPLAN Notices, 39(10): 224–225, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gestwicki:2004:JJI

[1244] Paul V. Gestwicki and Bharat Jayaraman. JIVE: Java interactive visualization environment. ACM SIGPLAN Notices, 39(10):226–228, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McBride:2004:SAE

[1245] Matthew R. McBride. The software architect: essence, intuition, and guiding principles. ACM SIGPLAN Notices, 39(10):230–235, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Freeman:2004:MRO

[1246] Steve Freeman, Tim Mackinnon, Nat Pryce, and Joe Walnes. Mock roles, objects. ACM SIGPLAN Notices, 39(10): 236–246, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berenbach:2004:CUT

[1247] Brian A. Berenbach. Comparison of UML and text based requirements engineering. ACM SIGPLAN Notices, 39(10):247–252, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Blotner:2004:PPP

[1248] Joseph A. Blotner. PIP: a product planning strategy for the whole family or… how we became the brady bunch. ACM SIGPLAN Notices, 39(10):253– 259, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Namioka:2004:EI

[1249] Aki Namioka and Cary Bran. eXtreme ISO ?!? ACM SIGPLAN Notices, 39(10):260–263, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Antion:2004:WRM

[1250] Daniel Antion. Why reuse matters: ANI’s digital archive system. ACM SIGPLAN Notices, 39(10):264– 267, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dvorak:2004:HRT

[1251] Daniel L. Dvorak and William K.

Reinholtz. Hard real-time: C++ versus RTSJ. ACM SIGPLAN Notices, 39(10):268–274, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Murphy-Hill:2004:TEL

[1252] Emerson R. Murphy-Hill and Andrew P. Black. Traits: experience with a language feature. ACM SIGPLAN Notices, 39(10):275–282, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zimmermann:2004:SGW

[1253] Olaf Zimmermann, Sven Milinski, Michael Craes, and Frank Oellermann. Second generation Web servicesoriented architecture in production in the finance industry. ACM SIGPLAN Notices, 39(10):283–289, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kainz:2004:MSD

[1254] Joachim F. Kainz. Migrating to simpler distributed applications. ACM SIG-PLAN Notices, 39(10):290–293, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reimer:2004:VSP

[1255] Darrell Reimer, Edith Schonberg, Kavitha Srinivas, Harini Srinivasan, Julian Dolby, Aaron Kershenbaum, and Larry Koved. Validating structural properties of nested objects. ACM SIG-PLAN Notices, 39(10):294–304, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bodden:2004:LLR

[1256] Eric Bodden. A lightweight LTL runtime verification tool for Java. ACM SIGPLAN Notices, 39(10):306– 307, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Salimi:2004:NGF

[1257] Hadi Salimi. NGMF: a generic framework for constructing graph-based systems. ACM SIGPLAN Notices, 39(10): 308–309, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:2004:SSE

[1258] Jing Zhang. Supporting software evolution through model-driven program transformation. ACM SIGPLAN Notices, 39(10):310–311, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Trofin:2004:FRRb

[1259] Mircea Trofin. A framework for removing redundant context management services in enterprise JavaBeans application servers. ACM SIGPLAN Notices, 39(10):312–313, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Roychoudhury:2004:LIAb

[1260] Suman Roychoudhury. A languageindependent approach to software maintenance using grammar adapters. ACM SIGPLAN Notices, 39(10):314– 315, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Parsons:2004:FDA

[1261] Trevor Parsons. A framework for detecting, assessing and visualizing performance antipatterns in component based systems. ACM SIGPLAN Notices, 39(10):316–317, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Silaghi:2004:RDAb

[1262] Raul Silaghi. Refining designs along middleware-specific concerndimensions at different MDA-levels of abstraction. ACM SIGPLAN Notices, 39(10):318–319, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Krishna:2004:EDO

[1263] Arvind S. Krishna. Enhancing distributed object middleware qualities. ACM SIGPLAN Notices, 39(10):320– 321, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xanthos:2004:IRC

[1264] Spiros Xanthos. Identification of reusable components within an objectoriented software system using algebraic graph theory. ACM SIGPLAN Notices, 39(10):322–323, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xie:2004:AIC

[1265] Tao Xie. Automatic identification of common and special object-oriented unit tests. ACM SIGPLAN Notices, 39(10):324–325, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Keschenau:2004:REU

[1266] Martin Keschenau. Reverse engineering of UML specifications from Java programs. ACM SIGPLAN Notices, 39(10):326–327, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Layman:2004:EII

[1267] Lucas Layman. Empirical investigation of the impact of extreme programming practices on software projects. ACM SIGPLAN Notices, 39(10):328– 329, October 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hammond:2004:PTC

[1268] Lance Hammond, Brian D. Carlstrom, Vicky Wong, Ben Hertzberg, Mike Chen, Christos Kozyrakis, and Kunle Olukotun. Programming with transactional coherence and consistency (TCC). ACM SIGPLAN Notices, 39 (11):1–13, November 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Budiu:2004:SC

[1269] Mihai Budiu, Girish Venkataramani, Tiberiu Chelcea, and Seth Copen Goldstein. Spatial computation. ACM SIG-PLAN Notices, 39(11):14–26, November 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ekanayake:2004:ULP

[1270] Virantha Ekanayake, Clinton Kelly IV, and Rajit Manohar. An ultra lowpower processor for sensor networks. ACM SIGPLAN Notices, 39(11):27– 36, November 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lumb:2004:DSD

[1271] Christopher R. Lumb and Richard Golding. D-SPTF: decentralized request distribution in brick-based storage systems. ACM SIGPLAN Notices, 39(11):37–47, November 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Saito:2004:FBD

[1272] Yasushi Saito, Svend Frølund, Alistair Veitch, Arif Merchant, and Susan Spence. FAB: building distributed enterprise disk arrays from commodity components. ACM SIGPLAN Notices, 39(11):48–58, November 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Denehy:2004:DSA

[1273] Timothy E. Denehy, John Bent, Florentina I. Popovici, Andrea C.

Arpaci-Dusseau, and Remzi H. Arpaci-Dusseau. Deconstructing storage arrays. ACM SIGPLAN Notices, 39(11): 59–71, November 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhuang:2004:HIE

[1274] Xiaotong Zhuang, Tao Zhang, and Santosh Pande. HIDE: an infrastructure for efficiently protecting information leakage on the address bus. ACM SIG-PLAN Notices, 39(11):72–84, November 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Suh:2004:SPE

[1275] G. Edward Suh, Jae W. Lee, David Zhang, and Srinivas Devadas. Secure program execution via dynamic information flow tracking. ACM SIGPLAN Notices, 39(11):85–96, November 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Huh:2004:CDM

[1276] Jaehyuk Huh, Jichuan Chang, Doug Burger, and Gurindar S. Sohi. Coherence decoupling: making use of incoherence. ACM SIGPLAN Notices, 39(11):97–106, November 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Srinivasan:2004:CFP

[1277] Srikanth T. Srinivasan, Ravi Rajwar, Haitham Akkary, Amit Gandhi, and Mike Upton. Continual flow pipelines. ACM SIGPLAN Notices, 39(11):107– 119, November 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Desikan:2004:SSR

[1278] Rajagopalan Desikan, Simha Sethumadhavan, Doug Burger, and Stephen W. Keckler. Scalable selective re-execution for EDGE architectures. ACM SIGPLAN Notices, 39(11):120–132, November 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Regehr:2004:HSA

[1279] John Regehr and Alastair Reid.

HOIST: a system for automatically deriving static analyzers for embedded systems. ACM SIGPLAN Notices, 39(11):133–143, November 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wang:2004:HTV

[1280] Perry H. Wang, Jamison D. Collins, Hong Wang, Dongkeun Kim, Bill Greene, Kai-Ming Chan, Aamir B. Yunus, Terry Sych, Stephen F. Moore, and John P. Shen. Helper threads via virtual multithreading on an experimental Itanium-2 processor-based platform. ACM SIGPLAN Notices, 39 (11):144–155, November 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hauswirth:2004:LOM

[1281] Matthias Hauswirth and Trishul M.

Chilimbi. Low-overhead memory leak detection using adaptive statistical profiling. ACM SIGPLAN Notices, 39 (11):156–164, November 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shen:2004:LPP

[1282] Xipeng Shen, Yutao Zhong, and Chen Ding. Locality phase prediction. ACM SIGPLAN Notices, 39(11):165– 176, November 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhou:2004:DTP

[1283] Pin Zhou, Vivek Pandey, Jagadeesan Sundaresan, Anand Raghuraman, Yuanyuan Zhou, and Sanjeev Kumar. Dynamic tracking of page miss ratio curve for memory management. ACM SIGPLAN Notices, 39(11):177– 188, November 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rabbah:2004:COP

[1284] Rodric M. Rabbah, Hariharan Sandanagobalane, Mongkol Ekpanyapong, and Weng-Fai Wong. Compiler orchestrated prefetching via speculation and predication. ACM SIGPLAN Notices, 39(11):189–198, November 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cher:2004:SPM

[1285] Chen-Yong Cher, Antony L. Hosking, and T. N. Vijaykumar. Software prefetching for mark-sweep garbage collection: hardware analysis and software redesign. ACM SIGPLAN Notices, 39(11):199–210, November 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lowell:2004:DVM

[1286] David E. Lowell, Yasushi Saito, and Eileen J. Samberg. Devirtualizable virtual machines enabling general, single-node, online maintenance. ACM SIGPLAN Notices, 39(11):211– 223, November 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smolens:2004:FBS

[1287] Jared C. Smolens, Brian T. Gold, Jangwoo Kim, Babak Falsafi, James C. Hoe, and Andreas G. Nowatzyk. Fingerprinting: bounding soft-error detection latency and bandwidth. ACM SIGPLAN Notices, 39(11):224–234, November 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bronevetsky:2004:ALC

[1288] Greg Bronevetsky, Daniel Marques, Keshav Pingali, Peter Szwed, and Martin Schulz. Application-level checkpointing for shared memory programs. ACM SIGPLAN Notices, 39(11):235– 247, November 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wu:2004:FOM

[1289] Qiang Wu, Philo Juang, Margaret Martonosi, and Douglas W. Clark. Formal online methods for voltage/ frequency control in multiple clock domain microprocessors. ACM SIGPLAN Notices, 39(11):248–259, November 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gomaa:2004:HRL

[1290] Mohamed Gomaa, Michael D. Powell, and T. N. Vijaykumar. Heat-and-run: leveraging SMT and CMP to manage power density through the operating system. ACM SIGPLAN Notices, 39(11):260–270, November 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Li:2004:PDE

[1291] Xiaodong Li, Zhenmin Li, Francis David, Pin Zhou, Yuanyuan Zhou, Sarita Adve, and Sanjeev Kumar. Performance directed energy management for main memory and disks. ACM SIGPLAN Notices, 39(11):271– 283, November 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2004:A

[1292] Anonymous. Architecture. ACM SIG-PLAN Notices, 39(11):??, November 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2004:MSA

[1293] Anonymous. Memory system analysis and optimization. ACM SIG-PLAN Notices, 39(11):??, November 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2004:NMA

[1294] Anonymous. New models and architectures. ACM SIGPLAN Notices, 39 (11):??, November 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2004:Pa

[1295] Anonymous. Potpourri. ACM SIG-PLAN Notices, 39(11):??, November 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2004:Pb

[1296] Anonymous. Power. ACM SIG-PLAN Notices, 39(11):??, November 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2004:R

[1297] Anonymous. Reliability. ACM SIG-PLAN Notices, 39(11):??, November 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2004:Sa

[1298] Anonymous. Storage. ACM SIG-PLAN Notices, 39(11):??, November 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2004:Sb

[1299] Anonymous. Security. ACM SIG-PLAN Notices, 39(11):??, November 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:2004:FAR

[1300] Paul Frenger. Forth and AI revisited: BRAIN.FORTH. ACM SIG-PLAN Notices, 39(12):11–16, December 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cabana:2004:PPJ

[1301] Brian Cabana, Suad Alagić, and Jeff Faulkner. Parametric polymorphism for Java: is there any hope in sight? ACM SIGPLAN Notices, 39(12):22– 31, December 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hughes:2004:CSS

[1302] Elliott Hughes. Checking spelling in source code. ACM SIGPLAN Notices, 39(12):32–38, December 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Noble:2004:NNPb

[1303] James Noble and Robert Biddle. Notes on notes on postmodern programming. ACM SIGPLAN Notices, 39(12):40– 56, December 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

West:2004:LLAb

[1304] David West. Looking for love (in all the wrong places). ACM SIGPLAN Notices, 39(12):57–63, December 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marick:2004:MWOb

[1305] Brian Marick. Methodology work is ontology work. ACM SIGPLAN Notices, 39(12):64–72, December 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Singh:2004:PPPb

[1306] Munindar P. Singh, Amit K. Chopra, Nirmit Desai, and Ashok U. Mallya. Protocols for processes: programming in the large for open systems. ACM SIGPLAN Notices, 39(12):73–83, December 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Edwards:2004:ECPb

[1307] Jonathan Edwards. Example centric programming. ACM SIGPLAN Notices, 39(12):84–91, December 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hovemeyer:2004:FBEb

[1308] David Hovemeyer and William Pugh.

Finding bugs is easy. ACM SIG-PLAN Notices, 39(12):92–106, December 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhao:2004:GOPb

[1309] Yinliang Zhao. Granule-oriented programming. ACM SIGPLAN Notices, 39(12):107–118, December 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sheard:2004:LFb

[1310] Tim Sheard. Languages of the future.

ACM SIGPLAN Notices, 39(12):119– 132, December 2004. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2004:FRb

[1311] Anonymous. Forth report. ACM SIG-PLAN Notices, 39(12):??, December 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2004:OO

[1312] Anonymous. OOPSLA onward! ACM SIGPLAN Notices, 39(12):??, December 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chakravarty:2005:ATC

[1313] Manuel M. T. Chakravarty, Gabriele Keller, Simon Peyton Jones, and Simon Marlow. Associated types with class. ACM SIGPLAN Notices, 40(1): 1–13, January 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cobbe:2005:EAR

[1314] Richard Cobbe and Matthias Felleisen. Environmental acquisition revisited. ACM SIGPLAN Notices, 40(1):14–25, January 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ancona:2005:PBC

[1315] Davide Ancona, Ferruccio Damiani, Sophia Drossopoulou, and Elena Zucca. Polymorphic bytecode: compositional compilation for Java-like languages. ACM SIGPLAN Notices, 40 (1):26–37, January 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:2005:STI

[1316] Juan Chen and David Tarditi. A simple typed intermediate language for object-oriented languages. ACM SIG-PLAN Notices, 40(1):38–49, January 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hosoya:2005:PPX

[1317] Haruo Hosoya, Alain Frisch, and Giuseppe Castagna. Parametric polymorphism for XML. ACM SIGPLAN Notices, 40(1):50–62, January 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sumii:2005:BTA

[1318] Eijiro Sumii and Benjamin C. Pierce.

A bisimulation for type abstraction and recursion. ACM SIGPLAN Notices, 40 (1):63–74, January 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Goguen:2005:SAE

[1319] Healfdene Goguen. A syntactic approach to eta equality in type theory. ACM SIGPLAN Notices, 40(1):75–84, January 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ghica:2005:SGQ

[1320] Dan R. Ghica. Slot games: a quantitative model of computation. ACM SIG-PLAN Notices, 40(1):85–97, January 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Alur:2005:SIS

[1321] Rajeev Alur, Pavol Černý, P. Madhusudan, and Wonhong Nam. Synthesis of interface specifications for Java classes. ACM SIGPLAN Notices, 40 (1):98–109, January 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Flanagan:2005:DPO

[1322] Cormac Flanagan and Patrice Godefroid. Dynamic partial-order reduction for model checking software. ACM SIGPLAN Notices, 40(1):110–121, January 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Grumberg:2005:PGU

[1323] Orna Grumberg, Flavio Lerda, Ofer Strichman, and Michael Theobald. Proof-guided underapproximationwidening for multi-process systems. ACM SIGPLAN Notices, 40(1):122– 131, January 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Podelski:2005:TPA

[1324] Andreas Podelski and Andrey Rybalchenko. Transition predicate abstraction and fair termination. ACM SIGPLAN Notices, 40(1):132–144, January 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gay:2005:CQP

[1325] Simon J. Gay and Rajagopal Nagarajan. Communicating quantum processes. ACM SIGPLAN Notices, 40(1):

145–157, January 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Li:2005:DPR

[1326] Peng Li and Steve Zdancewic. Downgrading policies and relaxed noninterference. ACM SIGPLAN Notices, 40 (1):158–170, January 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Park:2005:PLB

[1327] Sungwoo Park, Frank Pfenning, and Sebastian Thrun. A probabilistic language based upon sampling functions. ACM SIGPLAN Notices, 40(1):171– 182, January 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stoyle:2005:MMS

[1328] Gareth Stoyle, Michael Hicks, Gavin Bierman, Peter Sewell, and Iulian Neamtiu. Mutatis mutandis: safe and predictable dynamic software updating. ACM SIGPLAN Notices, 40(1): 183–194, January 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Field:2005:TPM

[1329] John Field and Carlos A. Varela.

Transactors: a programming model for maintaining globally consistent distributed state in unreliable environments. ACM SIGPLAN Notices, 40(1): 195–208, January 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bruni:2005:TFC

[1330] Roberto Bruni, Hernán Melgratti, and Ugo Montanari. Theoretical foundations for compensations in flow composition languages. ACM SIGPLAN Notices, 40(1):209–220, January 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Neubauer:2005:SPM

[1331] Matthias Neubauer and Peter Thiemann. From sequential programs to multi-tier applications by program transformation. ACM SIGPLAN Notices, 40(1):221–232, January 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Foster:2005:CBD

[1332] J. Nathan Foster, Michael B. Greenwald, Jonathan T. Moore, Benjamin C. Pierce, and Alan Schmitt. Combinators for bi-directional tree transformations: a linguistic approach to the view update problem. ACM SIGPLAN Notices, 40(1):233–246, January 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Parkinson:2005:SLA

[1333] Matthew Parkinson and Gavin Bierman. Separation logic and abstraction. ACM SIGPLAN Notices, 40(1): 247–258, January 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bornat:2005:PAS

[1334] Richard Bornat, Cristiano Calcagno, Peter O’Hearn, and Matthew Parkinson. Permission accounting in separation logic. ACM SIGPLAN Notices, 40(1):259–270, January 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Calcagno:2005:CLT

[1335] Cristiano Calcagno, Philippa Gardner, and Uri Zarfaty. Context logic and tree update. ACM SIGPLAN Notices, 40(1):271–282, January 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boyland:2005:CEU

[1336] John Tang Boyland and William Retert. Connecting effects and uniqueness with adoption. ACM SIGPLAN Notices, 40(1):283–295, January 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rinetzky:2005:SPL

[1337] Noam Rinetzky, Jörg Bauer, Thomas Reps, Mooly Sagiv, and Reinhard Wilhelm. A semantics for procedure local heaps and its abstractions. ACM SIG-PLAN Notices, 40(1):296–309, January 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hackett:2005:RBS

[1338] Brian Hackett and Radu Rugina.

Region-based shape analysis with tracked locations. ACM SIGPLAN Notices, 40(1):310–323, January 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gulwani:2005:PIA

[1339] Sumit Gulwani and George C. Necula. Precise interprocedural analysis using random interpretation. ACM SIGPLAN Notices, 40(1):324–337, January 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gopan:2005:FNA

[1340] Denis Gopan, Thomas Reps, and Mooly Sagiv. A framework for numeric analysis of array operations. ACM SIG-PLAN Notices, 40(1):338–350, January 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xie:2005:SED

[1341] Yichen Xie and Alex Aiken. Scalable error detection using boolean satisfiability. ACM SIGPLAN Notices, 40(1): 351–363, January 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lerner:2005:ASP

[1342] Sorin Lerner, Todd Millstein, Erika Rice, and Craig Chambers. Automated soundness proofs for dataflow analyses and transformations via local rules. ACM SIGPLAN Notices, 40(1): 364–377, January 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Manson:2005:JMM

[1343] Jeremy Manson, William Pugh, and Sarita V. Adve. The Java memory model. ACM SIGPLAN Notices, 40(1): 378–391, January 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:2005:TIP

[1344] Paul Frenger. Tina: an improbable 3-pin microcontroller. ACM SIG-PLAN Notices, 40(2):5–10, February 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Badjonski:2005:AJA

[1345] Mihal Badjonski, Mirjana Ivanovic, and Zoran Budimac. Adaptable Java Agents (AJA): a tool for programming of multi-agent systems. ACM SIG-PLAN Notices, 40(2):17–26, February 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jiang:2005:ERO

[1346] Shujuan Jiang and Baowen Xu. An efficient and reliable object-oriented exception handling mechanism. ACM SIGPLAN Notices, 40(2):27–32, February 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shaw:2005:LDO

[1347] Sachin Shaw and Pawan Kumar. Loopdead optimization. ACM SIGPLAN Notices, 40(2):33–40, February 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zerzelidis:2005:RRT

[1348] A. Zerzelidis and A. J. Wellings. Requirements for a real-time .NET framework. ACM SIGPLAN Notices, 40(2): 41–50, February 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2005:F

[1349] Anonymous. Forth. ACM SIGPLAN Notices, 40(2):??, February 2005. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2005:TC

[1350] Anonymous. Technical correspondence. ACM SIGPLAN Notices, 40 (2):??, February 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Proulx:2005:CPD

[1351] Viera K. Proulx and Tanya Cashorali. Calculator problem and the design recipe. ACM SIGPLAN Notices, 40 (3):4–11, March 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mueller:2005:ACM

[1352] Conrad Mueller. Axiomatic computational model. ACM SIGPLAN Notices, 40(3):29–39, March 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xian:2005:MMR

[1353] Yuqiang Xian and Guangze Xiong. Minimizing memory requirement of real-time systems with concurrent garbage collector. ACM SIGPLAN Notices, 40(3):40–48, March 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2005:CP

[1354] Anonymous. Circular patterns. ACM SIGPLAN Notices, 40(3):??, March 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:2005:TFY

[1355] Paul Frenger. Tired by Forth?: you should be! ACM SIGPLAN Notices, 40(4):5–8, April 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Saitou:2005:DAI

[1356] Tetsuo Saitou, Mitsugu Suzuki, and Tan Watanabe. Dominance analysis of irreducible CFGs by reduction. ACM SIGPLAN Notices, 40(4):10–19, April 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Beloglavec:2005:ALM

[1357] Simon Beloglavec, Marjan Heričko, Matjaž B. Jurič, and Ivan Rozman. Analysis of the limitations of multiple client handling in a Java server environment. ACM SIGPLAN Notices, 40 (4):20–28, April 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Crepinsek:2005:EGPa

[1358] Matej Črepinšek, Marjan Mernik, and Viljem Žumer. Extracting grammar from programs: brute force approach. ACM SIGPLAN Notices, 40(4):29–38, April 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Crepinsek:2005:EGPb

[1359] Matej Črepinšek, Marjan Mernik, Faizan Javed, Barrett R. Bryant, and Alan Sprague. Extracting grammar from programs: evolutionary approach. ACM SIGPLAN Notices, 40(4):39–46, April 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vanderSpek:2005:SER

[1360] P. van der Spek, N. Plat, and C. Pronk. Syntax error repair for a Java-based parser generator. ACM SIGPLAN Notices, 40(4):47–50, April 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bulic:2005:EWF

[1361] Patricio Bulić and Veselko Guštin. An efficient way to filter out data dependences with a sufficiently large distance between memory references. ACM SIGPLAN Notices, 40(4):51–60, April 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gluck:2005:RAP

[1362] Robert Glück and Masahiko Kawabe. Revisiting an automatic program inverter for Lisp. ACM SIGPLAN Notices, 40(5):8–17, May 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kumar:2005:RMS

[1363] Rajeev Kumar, Vikram Agrawal, and Anil Mangolia. Realization of multimethods in single dispatch object oriented languages. ACM SIGPLAN Notices, 40(5):18–27, May 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sunik:2005:SL

[1364] Boris Sunik. The specification language T. ACM SIGPLAN Notices, 40(5):28– 38, May 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Derk:2005:TSM

[1365] M. D. Derk. Towards a simpler method of operational semantics for language definition. ACM SIGPLAN Notices, 40 (5):39–44, May 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Barbosa:2005:RPT

[1366] Manuel Barbosa, Alcino Cunha, and Jorge Sousa Pinto. Recursion patterns and time-analysis. ACM SIG-PLAN Notices, 40(5):45–54, May 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xu:2005:SVD

[1367] Min Xu, Rastislav Bodík, and Mark D.

Hill. A serializability violation detector for shared-memory server programs. ACM SIGPLAN Notices, 40 (6):1–14, June 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Liblit:2005:SSB

[1368] Ben Liblit, Mayur Naik, Alice X.

Zheng, Alex Aiken, and Michael I. Jordan. Scalable statistical bug isolation.

ACM SIGPLAN Notices, 40(6):15–26, June 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Elmas:2005:VVC

[1369] Tayfun Elmas, Serdar Tasiran, and Shaz Qadeer. VYRD: VerifYing concurrent programs by runtime Refinement-violation Detection. ACM SIGPLAN Notices, 40(6):27–37, June 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jhala:2005:PS

[1370] Ranjit Jhala and Rupak Majumdar. Path slicing. ACM SIGPLAN Notices, 40(6):38–47, June 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mandelin:2005:JMH

[1371] David Mandelin, Lin Xu, Rastislav Bodík, and Doug Kimelman. Jungloid mining: helping to navigate the API jungle. ACM SIGPLAN Notices, 40 (6):48–61, June 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Furr:2005:CTS

[1372] Michael Furr and Jeffrey S. Foster. Checking type safety of foreign function calls. ACM SIGPLAN Notices, 40 (6):62–72, June 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Siek:2005:ELS

[1373] Jeremy Siek and Andrew Lumsdaine. Essential language support for generic programming. ACM SIGPLAN Notices, 40(6):73–84, June 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chin:2005:STQ

[1374] Brian Chin, Shane Markstrum, and Todd Millstein. Semantic type qualifiers. ACM SIGPLAN Notices, 40 (6):85–95, June 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Krishnaswami:2005:PBO

[1375] Neel Krishnaswami and Jonathan Aldrich. Permission-based ownership: encapsulating state in higherorder typed languages. ACM SIG-PLAN Notices, 40(6):96–106, June 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jimenez:2005:CPI

[1376] Daniel A. Jiménez. Code placement for improving dynamic branch prediction accuracy. ACM SIGPLAN Notices, 40 (6):107–116, June 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Avgustinov:2005:OA

[1377] Pavel Avgustinov, Aske Simon Christensen, Laurie Hendren, Sascha Kuzins, Jennifer Lhoták, Ondřej Lhoták, Oege de Moor, Damien Sereni, Ganesh Sittampalam, and Julian Tibble. Optimising aspectJ. ACM SIG-PLAN Notices, 40(6):117–128, June 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lattner:2005:APA

[1378] Chris Lattner and Vikram Adve. Automatic pool allocation: improving performance by controlling data structure layout in the heap. ACM SIG-PLAN Notices, 40(6):129–142, June 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hertz:2005:GCP

[1379] Matthew Hertz, Yi Feng, and Emery D. Berger. Garbage collection without paging. ACM SIGPLAN Notices, 40 (6):143–153, June 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rong:2005:RAS

[1380] Hongbo Rong, Alban Douillet, and Guang R. Gao. Register allocation for software pipelined multi-dimensional loops. ACM SIGPLAN Notices, 40 (6):154–167, June 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhuang:2005:DRA

[1381] Xiaotong Zhuang and Santosh Pande. Differential register allocation. ACM SIGPLAN Notices, 40(6):168–179, June 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Aleta:2005:DFS

[1382] Alex Aletà, Josep M. Codina, Antonio Gonzàlez, and David Kaeli. Demystifying on-the-fly spill code. ACM SIG-PLAN Notices, 40(6):180–189, June 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Luk:2005:PBC

[1383] Chi-Keung Luk, Robert Cohn, Robert Muth, Harish Patil, Artur Klauser, Geoff Lowney, Steven Wallace, Vijay Janapa Reddi, and Kim Hazelwood. Pin: building customized program analysis tools with dynamic instrumentation. ACM SIGPLAN Notices, 40 (6):190–200, June 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ayers:2005:TFF

[1384] Andrew Ayers, Richard Schooler, Chris Metcalf, Anant Agarwal, Junghwan Rhee, and Emmett Witchel. Trace-Back: first fault diagnosis by reconstruction of distributed control flow. ACM SIGPLAN Notices, 40(6):201– 212, June 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Godefroid:2005:DDA

[1385] Patrice Godefroid, Nils Klarlund, and Koushik Sen. DART: directed automated random testing. ACM SIG-PLAN Notices, 40(6):213–223, June 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:2005:SAH

[1386] Michael K. Chen, Xiao Feng Li, Ruiqi Lian, Jason H. Lin, Lixia Liu, Tao Liu, and Roy Ju. Shangri-La: achieving high performance from compiled network applications while enabling ease of programming. ACM SIG-PLAN Notices, 40(6):224–236, June 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dai:2005:APP

[1387] Jinquan Dai, Bo Huang, Long Li, and Luddy Harrison. Automatically partitioning packet processing applications for pipelined architectures. ACM SIG-PLAN Notices, 40(6):237–248, June 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ni:2005:PAH

[1388] Yang Ni, Ulrich Kremer, Adrian Stere, and Liviu Iftode. Programming adhoc networks of mobile and resourceconstrained devices. ACM SIGPLAN Notices, 40(6):249–260, June 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boehm:2005:TCI

[1389] Hans-J. Boehm. Threads cannot be implemented as a library. ACM SIG-PLAN Notices, 40(6):261–268, June 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Quinones:2005:MCI

[1390] Carlos García Quiñones, Carlos Madriles, Jesús Sánchez, Pedro Marcuello, Antonio González, and Dean M. Tullsen. Mitosis compiler: an infrastructure for speculative threading based on pre-computation slices. ACM SIGPLAN Notices, 40(6):269– 279, June 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Solar-Lezama:2005:PSB

[1391] Armando Solar-Lezama, Rodric Rabbah, Rastislav Bodík, and Kemal Ebcioğlu. Programming by sketching for bit-streaming programs. ACM SIG-PLAN Notices, 40(6):281–294, June 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fisher:2005:PDS

[1392] Kathleen Fisher and Robert Gruber.

PADS: a domain-specific language for processing ad hoc data. ACM SIG-PLAN Notices, 40(6):295–304, June 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bauer:2005:CSP

[1393] Lujo Bauer, Jay Ligatti, and David Walker. Composing security policies with polymer. ACM SIGPLAN Notices, 40(6):305–314, June 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Franchetti:2005:FLM

[1394] Franz Franchetti, Yevgen Voronenko, and Markus Püschel. Formal loop merging for signal transforms. ACM SIGPLAN Notices, 40(6):315–326, June 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xu:2005:EEP

[1395] Ruibin Xu, Dakai Zhu, Cosmin Rusu, Rami Melhem, and Daniel Mossé. Energy-efficient policies for embedded clusters. ACM SIGPLAN Notices, 40 (7):1–10, July 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:2005:EAM

[1396] Kun Zhang and Santosh Pande. Efficient application migration under compiler guidance. ACM SIGPLAN Notices, 40(7):10–20, July 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Henzinger:2005:CCG

[1397] Thomas A. Henzinger, Christoph M. Kirsch, and Slobodan Matic. Composable code generation for distributed giotto. ACM SIGPLAN Notices, 40 (7):21–30, July 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Farcas:2005:TDR

[1398] Emilia Farcas, Claudiu Farcas, Wolfgang Pree, and Josef Templ. Transparent distribution of real-time components based on logical execution time. ACM SIGPLAN Notices, 40(7):31–39, July 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gay:2005:SDP

[1399] David Gay, Phil Levis, and David Culler. Software design patterns for TinyOS. ACM SIGPLAN Notices, 40 (7):40–49, July 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Regehr:2005:PIO

[1400] John Regehr and Usit Duongsaa. Preventing interrupt overload. ACM SIG-PLAN Notices, 40(7):50–58, July 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Titzer:2005:NPI

[1401] Ben L. Titzer and Jens Palsberg. Nonintrusive precision instrumentation of microcontroller software. ACM SIG-PLAN Notices, 40(7):59–68, July 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cooper:2005:AAC

[1402] Keith D. Cooper, Alexander Grosul, Timothy J. Harvey, Steven Reeves, Devika Subramanian, Linda Torczon, and Todd Waterman. ACME: adaptive compilation made efficient. ACM SIG-PLAN Notices, 40(7):69–77, July 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Franke:2005:PSL

[1403] Björn Franke, Michael O’Boyle, John Thomson, and Grigori Fursin. Probabilistic source-level optimisation of embedded programs. ACM SIGPLAN Notices, 40(7):78–86, July 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Oi:2005:DLV

[1404] Hitoshi Oi. On the design of the local variable cache in a hardware translation-based Java Virtual Machine. ACM SIGPLAN Notices, 40 (7):87–94, July 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chanet:2005:SWC

[1405] Dominique Chanet, Bjorn De Sutter, Bruno De Bus, Ludo Van Put, and Koen De Bosschere. System-wide compaction and specialization of the Linux kernel. ACM SIGPLAN Notices, 40 (7):95–104, July 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brisk:2005:DCT

[1406] Philip Brisk, Jamie Macbeth, Ani Nahapetian, and Majid Sarrafzadeh. A dictionary construction technique for code compression systems with echo instructions. ACM SIGPLAN Notices, 40 (7):105–114, July 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sermulins:2005:CAO

[1407] Janis Sermulins, William Thies, Rodric Rabbah, and Saman Amarasinghe. Cache aware optimization of stream programs. ACM SIGPLAN Notices, 40 (7):115–126, July 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sassone:2005:SSS

[1408] Peter G. Sassone, D. Scott Wills, and Gabriel H. Loh. Static strands: safely collapsing dependence chains for increasing embedded power efficiency. ACM SIGPLAN Notices, 40 (7):127–136, July 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

So:2005:CSP

[1409] Won So and Alexander G. Dean. Complementing software pipelining with software thread integration. ACM SIGPLAN Notices, 40(7):137–146, July 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kudriavtsev:2005:GPS

[1410] Alexei Kudriavtsev and Peter Kogge.

Generation of permutations for SIMD processors. ACM SIGPLAN Notices, 40(7):147–156, July 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Staschulat:2005:SPC

[1411] Jan Staschulat and Rolf Ernst. Scalable precision cache analysis for preemptive scheduling. ACM SIGPLAN Notices, 40(7):157–165, July 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xu:2005:SBC

[1412] Rong Xu and Zhiyuan Li. A samplebased cache mapping scheme. ACM SIGPLAN Notices, 40(7):166–174, July 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tan:2005:WAU

[1413] Yudong Tan and Vincent J. Mooney III. WCRT analysis for a uniprocessor with a unified prioritized cache. ACM SIGPLAN Notices, 40(7):175–182, July 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bacon:2005:SGR

[1414] David F. Bacon, Perry Cheng, David Grove, and Martin T. Vechev. Syncopation: generational real-time garbage collection in the metronome. ACM SIGPLAN Notices, 40(7):183–192, July 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mann:2005:SDA

[1415] Tobias Mann, Morgan Deters, Rob LeGrand, and Ron K. Cytron. Static determination of allocation rates to support real-time garbage collection. ACM SIGPLAN Notices, 40(7):193– 202, July 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhu:2005:FES

[1416] Yifan Zhu and Frank Mueller. Feedback EDF scheduling exploiting hardware-assisted asynchronous dynamic voltage scaling. ACM SIG-PLAN Notices, 40(7):203–212, July 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kandemir:2005:CME

[1417] Mahmut Kandemir, Guangyu Chen, and Ismail Kadayif. Compiling for memory emergency. ACM SIGPLAN Notices, 40(7):213–221, July 2005. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Defoe:2005:UBD

[1418] Delvin C. Defoe, Sharath R. Cholleti, and Ron K. Cytron. Upper bound for defragmenting buddy heaps. ACM SIGPLAN Notices, 40(7):222–229, July 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Griffin:2005:EEG

[1419] Paul Griffin, Witawas Srisa-an, and J. Morris Chang. An energy efficient garbage collector for Java embedded devices. ACM SIGPLAN Notices, 40 (7):230–238, July 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:2005:FSF

[1420] Paul Frenger. Forth sorts fruit down under. ACM SIGPLAN Notices, 40 (8):7–10, August 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Furlong:2005:HSA

[1421] Jeff Furlong and Atanas Radenski.

Handling the subclassing anomaly with Object Teams. ACM SIGPLAN Notices, 40(8):12–18, August 2005. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Greiner:2005:ZTE

[1422] Saso Greiner, Damijan Rebernak, Janez Brest, and Viljem Zumer. Z0 — a tiny experimental language. ACM SIGPLAN Notices, 40(8):19–28, August 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Draheim:2005:GPC

[1423] Dirk Draheim, Christof Lutteroth, and Gerald Weber. Generative programming for C#. ACM SIGPLAN Notices, 40(8):29–33, August 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jiang:2005:AAT

[1424] Shujuan Jiang, Yongping Zhang, Dashun Yan, and Yuanpeng Jiang. An approach to automatic testing exception handling. ACM SIGPLAN Notices, 40(8):34–39, August 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pottier:2005:MTI

[1425] François Pottier. From ML type inference to stratified type inference. ACM SIGPLAN Notices, 40(9):1, September 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shivers:2005:ALS

[1426] Olin Shivers. The anatomy of a loop: a story of scope and control. ACM SIG-PLAN Notices, 40(9):2–14, September 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sewell:2005:AHL

[1427] Peter Sewell, James J. Leifer, Keith Wansbrough, Francesco Zappa Nardelli, Mair Allen-Williams, Pierre Habouzit, and Viktor Vafeiadis. Acute: high-level programming language design for distributed computation. ACM SIGPLAN Notices, 40(9):15–26, September 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ramsey:2005:ELS

[1428] Norman Ramsey, Kathleen Fisher, and Paul Govereau. An expressive language of signatures. ACM SIGPLAN Notices, 40(9):27–40, September 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dreyer:2005:RTG

[1429] Derek Dreyer. Recursive type generativity. ACM SIGPLAN Notices, 40(9): 41–53, September 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nilsson:2005:DOF

[1430] Henrik Nilsson. Dynamic optimization for functional reactive programming using generalized algebraic data types. ACM SIGPLAN Notices, 40(9): 54–65, September 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:2005:CPT

[1431] Chiyan Chen and Hongwei Xi. Combining programming with theorem proving. ACM SIGPLAN Notices, 40(9): 66–77, September 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ahmed:2005:SIM

[1432] Amal Ahmed, Matthew Fluet, and Greg Morrisett. A step-indexed model of substructural state. ACM SIGPLAN Notices, 40(9):78–91, September 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ringenburg:2005:AFC

[1433] Michael F. Ringenburg and Dan Grossman. AtomCaml: first-class atomicity via rollback. ACM SIGPLAN Notices, 40(9):92–104, September 2005.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Terauchi:2005:WSE

[1434] Tachio Terauchi and Alex Aiken. Witnessing side-effects. ACM SIG-PLAN Notices, 40(9):105–115, September 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hallgren:2005:PAO

[1435] Thomas Hallgren, Mark P. Jones, Rebekah Leslie, and Andrew Tolmach. A principled approach to operating system construction in Haskell. ACM SIG-PLAN Notices, 40(9):116–128, September 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Eich:2005:JTY

[1436] Brendan Eich. JavaScript at ten years. ACM SIGPLAN Notices, 40(9): 129, September 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Remy:2005:SPT

[1437] Didier Rémy. Simple, partial typeinference for System F based on typecontainment. ACM SIGPLAN Notices, 40(9):130–143, September 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Leijen:2005:QTM

[1438] Daan Leijen and Andres Löh. Qualified types for MLF. ACM SIGPLAN Notices, 40(9):144–155, September 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Makholm:2005:TIP

[1439] Henning Makholm and J. B. Wells.

Type inference, principal typings, and let-polymorphism for first-class mixin modules. ACM SIGPLAN Notices, 40 (9):156–167, September 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Diatchki:2005:HLV

[1440] Iavor S. Diatchki, Mark P. Jones, and Rebekah Leslie. High-level views on low-level representations. ACM SIG-PLAN Notices, 40(9):168–179, September 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cheney:2005:SYN

[1441] James Cheney. Scrap your nameplate: (functional pearl). ACM SIG-PLAN Notices, 40(9):180–191, September 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kiselyov:2005:BIT

[1442] Oleg Kiselyov, Chung chieh Shan, Daniel P. Friedman, and Amr Sabry. Backtracking, interleaving, and terminating monad transformers: (functional pearl). ACM SIGPLAN Notices, 40(9):192–203, September 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lammel:2005:SYB

[1443] Ralf Lämmel and Simon Peyton Jones. Scrap your boilerplate with class: extensible generic functions. ACM SIG-PLAN Notices, 40(9):204–215, September 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pettyjohn:2005:CGS

[1444] Greg Pettyjohn, John Clements, Joe Marshall, Shriram Krishnamurthi, and Matthias Felleisen. Continuations from generalized stack inspection. ACM SIGPLAN Notices, 40(9):216– 227, September 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ramos:2005:FND

[1445] J. Guadalupe Ramos, Josep Silva, and Germán Vidal. Fast narrowingdriven partial evaluation for inductively sequential programs. ACM SIG-PLAN Notices, 40(9):228–239, September 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Harper:2005:MMT

[1446] Robert Harper. Mechanizing the metatheory of programming languages. ACM SIGPLAN Notices, 40(9):240, September 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chakravarty:2005:ATS

[1447] Manuel M. T. Chakravarty, Gabriele Keller, and Simon Peyton Jones. Associated type synonyms. ACM SIG-PLAN Notices, 40(9):241–253, September 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Feng:2005:MVC

[1448] Xinyu Feng and Zhong Shao. Modular verification of concurrent assembly code with dynamic thread creation and termination. ACM SIGPLAN Notices, 40(9):254–267, September 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Westbrook:2005:LBA

[1449] Edwin Westbrook, Aaron Stump, and Ian Wehrman. A language-based approach to functionally correct imperative programming. ACM SIG-PLAN Notices, 40(9):268–279, September 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berger:2005:LAA

[1450] Martin Berger, Kohei Honda, and Nobuko Yoshida. A logical analysis of aliasing in imperative higherorder functions. ACM SIGPLAN Notices, 40(9):280–293, September 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ghani:2005:MAG

[1451] Neil Ghani, Patricia Johann, Tarmo Uustalu, and Varmo Vene. Monadic augment and generalised short cut fusion. ACM SIGPLAN Notices, 40(9): 294–305, September 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dantas:2005:PPA

[1452] Daniel S. Dantas, David Walker, Geoffrey Washburn, and Stephanie Weirich. PolyAML: a polymorphic aspect-oriented functional programming language. ACM SIGPLAN Notices, 40(9):306–319, September 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Masuhara:2005:ACA

[1453] Hidehiko Masuhara, Hideaki Tatsuzawa, and Akinori Yonezawa. Aspectual Caml: an aspect-oriented functional language. ACM SIGPLAN Notices, 40(9):320–330, September 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jarvi:2005:ATC

[1454] Jaakko Järvi, Jeremiah Willcock, and Andrew Lumsdaine. Associated types and constraint propagation for mainstream object-oriented generics. ACM SIGPLAN Notices, 40(10):1–19, October 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kennedy:2005:GAD

[1455] Andrew Kennedy and Claudio V. Russo. Generalized algebraic data types and object-oriented programming. ACM SIGPLAN Notices, 40(10): 21–40, October 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Odersky:2005:SCA

[1456] Martin Odersky and Matthias Zenger. Scalable component abstractions. ACM SIGPLAN Notices, 40(10):41–57, October 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sridharan:2005:DDP

[1457] Manu Sridharan, Denis Gopan, Lexin Shan, and Rastislav Bodík. Demanddriven points-to analysis for Java. ACM SIGPLAN Notices, 40(10):59– 76, October 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nanda:2005:DOT

[1458] Mangala Gowri Nanda, Christian Grothoff, and Satish Chandra. Deriving object typestates in the presence of inter-object references. ACM SIG-PLAN Notices, 40(10):77–96, October 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gil:2005:MPJ

[1459] Joseph (Yossi) Gil and Itay Maman.

Micro patterns in Java code. ACM SIGPLAN Notices, 40(10):97–116, October 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Diaz-Pace:2005:AAS

[1460] J. Andrés Díaz-Pace and Marcelo R.

Campo. ArchMatE: from architectural styles to object-oriented models through exploratory tool support. ACM SIGPLAN Notices, 40(10):117– 132, October 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zdun:2005:MAP

[1461] Uwe Zdun and Paris Avgeriou. Modeling architectural patterns using architectural primitives. ACM SIGPLAN Notices, 40(10):133–146, October 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Oancea:2005:PPS

[1462] Cosmin E. Oancea and Stephen M. Watt. Parametric polymorphism for software component architectures. ACM SIGPLAN Notices, 40(10):147– 166, October 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sangal:2005:UDM

[1463] Neeraj Sangal, Ev Jordan, Vineet Sinha, and Daniel Jackson. Using dependency models to manage complex software architecture. ACM SIGPLAN Notices, 40(10):167–176, October 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergel:2005:CJC

[1464] Alexandre Bergel, Stéphane Ducasse, and Oscar Nierstrasz. Classbox/J: controlling the scope of change in Java. ACM SIGPLAN Notices, 40(10):177– 189, October 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Liu:2005:IBP

[1465] Yu David Liu and Scott F. Smith. Interaction-based programming with classages. ACM SIGPLAN Notices, 40 (10):191–209, October 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tschantz:2005:JAR

[1466] Matthew S. Tschantz and Michael D.

Ernst. Javari: adding reference immutability to Java. ACM SIGPLAN Notices, 40(10):211–230, October 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gray:2005:FGI

[1467] Kathryn E. Gray, Robert Bruce Findler, and Matthew Flatt. Fine-grained interoperability through mirrors and contracts. ACM SIGPLAN Notices, 40 (10):231–245, October 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kojarski:2005:PAD

[1468] Sergei Kojarski and David H. Lorenz.

Pluggable AOP: designing aspect mechanisms for third-party composition. ACM SIGPLAN Notices, 40(10): 247–263, October 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Balaban:2005:RSC

[1469] Ittai Balaban, Frank Tip, and Robert Fuhrer. Refactoring support for class library migration. ACM SIGPLAN Notices, 40(10):265–279, October 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hauswirth:2005:AVP

[1470] Matthias Hauswirth, Amer Diwan, Peter F. Sweeney, and Michael C. Mozer. Automating vertical profiling. ACM SIGPLAN Notices, 40(10): 281–296, October 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Arnold:2005:IVM

[1471] Matthew Arnold, Adam Welc, and V. T. Rajan. Improving virtual machine performance using a cross-run profile repository. ACM SIGPLAN Notices, 40(10):297–311, October 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hertz:2005:QPG

[1472] Matthew Hertz and Emery D. Berger. Quantifying the performance of garbage collection vs. explicit memory management. ACM SIGPLAN Notices, 40(10):313–326, October 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shankar:2005:RSO

[1473] Ajeet Shankar, S. Subramanya Sastry, Rastislav Bodík, and James E. Smith. Runtime specialization with optimistic heap analysis. ACM SIGPLAN Notices, 40(10):327–343, October 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Allan:2005:ATM

[1474] Chris Allan, Pavel Avgustinov, Aske Simon Christensen, Laurie Hendren, Sascha Kuzins, Ondřej Lhoták, Oege de Moor, Damien Sereni, Ganesh Sittampalam, and Julian Tibble. Adding trace matching with free variables to AspectJ. ACM SIGPLAN Notices, 40 (10):345–364, October 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Martin:2005:FAE

[1475] Michael Martin, Benjamin Livshits, and Monica S. Lam. Finding application errors and security flaws using PQL: a program query language. ACM SIGPLAN Notices, 40(10):365– 383, October 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Goldsmith:2005:RQP

[1476] Simon Goldsmith, Robert O’Callahan, and Alex Aiken. Relational queries over program traces. ACM SIGPLAN Notices, 40(10):385–402, October 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ahern:2005:FJR

[1477] Alexander Ahern and Nobuko Yoshida.

Formalising Java RMI with explicit code mobility. ACM SIGPLAN Notices, 40(10):403–422, October 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gregor:2005:LSG

[1478] Douglas Gregor and Andrew Lumsdaine. Lifting sequential graph algorithms for distributed-memory parallel computation. ACM SIGPLAN Notices, 40(10):423–437, October 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Welc:2005:SFJ

[1479] Adam Welc, Suresh Jagannathan, and Antony Hosking. Safe futures for Java. ACM SIGPLAN Notices, 40(10):439– 453, October 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vanDooren:2005:CRC

[1480] Marko van Dooren and Eric Steegmans. Combining the robustness of checked exceptions with the flexibility of unchecked exceptions using anchored exception declarations. ACM SIG-PLAN Notices, 40(10):455–471, October 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Liu:2005:IAO

[1481] Yanhong A. Liu, Scott D. Stoller, Michael Gorbovitski, Tom Rothamel, and Yanni Ellen Liu. Incrementalization across object abstraction. ACM SIGPLAN Notices, 40(10):473– 486, October 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cunei:2005:PFD

[1482] Antonio Cunei and Jan Vitek. PolyD: a flexible dispatching framework. ACM SIGPLAN Notices, 40(10):487–503, October 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Edwards:2005:SUS

[1483] Jonathan Edwards. Subtext: uncovering the simplicity of programming. ACM SIGPLAN Notices, 40(10):505– 518, October 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Charles:2005:XOO

[1484] Philippe Charles, Christian Grothoff, Vijay Saraswat, Christopher Donawa, Allan Kielstra, Kemal Ebcioglu, Christoph von Praun, and Vivek Sarkar. X10: an object-oriented approach to non-uniform cluster computing. ACM SIGPLAN Notices, 40(10): 519–538, October 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vanderburg:2005:SMA

[1485] Glenn Vanderburg. A simple model of agile software processes — or — extreme programming annealed. ACM SIGPLAN Notices, 40(10):539– 545, October 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:2005:FRT

[1486] Paul Frenger. Forth report: Ten years of Forth in ACM Sigplan Notices: part 1. ACM SIGPLAN Notices, 40(11): 4–16, November 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Behera:2005:IAL

[1487] Chandan Kumar Behera and Pawan Kumar. An improved algorithm for loop dead optimization. ACM SIG-PLAN Notices, 40(11):18–28, November 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Li:2005:MBC

[1488] Bixin Li, Ying Zhou, Yancheng Wang, and Junhui Mo. Matrix-based component dependence representation and its applications in software quality assurance. ACM SIGPLAN Notices, 40(11): 29–36, November 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Siegfried:2005:ESL

[1489] Robert M. Siegfried, Denis Diakoniarakis, Kenneth G. Franqueiro, and Amol Jain. Extending a scripting language for Visual Basic forms. ACM SIGPLAN Notices, 40(11):37– 40, November 2005. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Savidis:2005:MDI

[1490] A. Savidis. More dynamic imperative languages. ACM SIGPLAN Notices, 40(12):6–13, December 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kumar:2005:PCO

[1491] Rajeev Kumar, Amit Gupta, B. S. Pankaj, Mrinmoy Ghosh, and P. P. Chakrabarti. Post-compilation optimization for multiple gains with pattern matching. ACM SIGPLAN Notices, 40(12):14–23, December 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Loka:2005:HCC

[1492] Raghavendra Rao Loka. Hybrid compilation: a case study. ACM SIGPLAN Notices, 40(12):24–27, December 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sierra:2005:IDO

[1493] José L. Sierra, Antonio Navarro, Baltasar Fernández-Manjón, and Alfredo Fernández-Valmayor. Incremental definition and operationalization of domain-specific markup languages in ADDS. ACM SIGPLAN Notices, 40 (12):28–37, December 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McKinna:2006:WDT

[1494] James McKinna. Why dependent types matter. ACM SIGPLAN Notices, 41 (1):1, January 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fisher:2006:NDD

[1495] Kathleen Fisher, Yitzhak Mandelbaum, and David Walker. The next 700 data description languages. ACM SIGPLAN Notices, 41(1):2–15, January 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:2006:HMD

[1496] Chengliang Zhang, Chen Ding, Mitsunori Ogihara, Yutao Zhong, and Youfeng Wu. A hierarchical model of data locality. ACM SIGPLAN Notices, 41(1):16–29, January 2006.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gautam:2006:SR

[1497] Gautam and S. Rajopadhye. Simplifying reductions. ACM SIGPLAN Notices, 41(1):30–41, January 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Leroy:2006:FCC

[1498] Xavier Leroy. Formal certification of a compiler back-end or: programming a compiler with a proof assistant. ACM SIGPLAN Notices, 41(1):42–54, January 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bishop:2006:ELH

[1499] Steve Bishop, Matthew Fairbairn, Michael Norrish, Peter Sewell, Michael Smith, and Keith Wansbrough. Engineering with logic: HOL specification and symbolic-evaluation testing for TCP implementations. ACM SIG-PLAN Notices, 41(1):55–66, January 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dam:2006:DPS

[1500] Mads Dam. Decidability and proof systems for language-based noninterference relations. ACM SIGPLAN Notices, 41(1):67–78, January 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hunt:2006:FSS

[1501] Sebastian Hunt and David Sands. On flow-sensitive security types. ACM SIGPLAN Notices, 41(1):79–90, January 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Amtoft:2006:LIF

[1502] Torben Amtoft, Sruthi Bandhakavi, and Anindya Banerjee. A logic for information flow in object-oriented programs. ACM SIGPLAN Notices, 41(1): 91–102, January 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vouillon:2006:PRT

[1503] Jérôme Vouillon. Polymorphic regular tree types and patterns. ACM SIG-PLAN Notices, 41(1):103–114, January 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lahiri:2006:VPW

[1504] Shuvendu K. Lahiri and Shaz Qadeer.

Verifying properties of well-founded linked lists. ACM SIGPLAN Notices, 41(1):115–126, January 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Might:2006:EAD

[1505] Matthew Might and Olin Shivers. Environment analysis via ΔCFA. ACM SIGPLAN Notices, 41(1):127–140, January 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Koutavas:2006:SBR

[1506] Vasileios Koutavas and Mitchell Wand. Small bisimulations for reasoning about higher-order imperative programs. ACM SIGPLAN Notices, 41(1): 141–152, January 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Alur:2006:FCL

[1507] Rajeev Alur, Swarat Chaudhuri, and P. Madhusudan. A fixpoint calculus for local and global program flows. ACM SIGPLAN Notices, 41(1):153–165, January 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Odersky:2006:SEC

[1508] Martin Odersky. The Scala experiment: can we provide better language support for component systems? ACM SIG-PLAN Notices, 41(1):166–167, January 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Danner:2006:ATS

[1509] Norman Danner and James S. Royer. Adventures in time and space. ACM SIGPLAN Notices, 41(1):168–179, January 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cohen:2006:SKN

[1510] Albert Cohen, Marc Duranton, Christine Eisenbeis, Claire Pagetti, Florence Plateau, and Marc Pouzet. Nsynchronous Kahn networks: a relaxed model of synchrony for real-time systems. ACM SIGPLAN Notices, 41(1):

180–193, January 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:2006:CDC

[1511] Guangyu Chen, Feihui Li, and Mahmut Kandemir. Compiler-directed channel allocation for saving power in onchip networks. ACM SIGPLAN Notices, 41(1):194–205, January 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Danielsson:2006:FLR

[1512] Nils Anders Danielsson, John Hughes, Patrik Jansson, and Jeremy Gibbons. Fast and loose reasoning is morally. ACM SIGPLAN Notices, 41(1):206– 217, January 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meunier:2006:MSB

[1513] Philippe Meunier, Robert Bruce Findler, and Matthias Felleisen. Modular set-based analysis from contracts. ACM SIGPLAN Notices, 41(1):218– 231, January 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pottier:2006:STI

[1514] François Pottier and Yann Régis-Gianas. Stratified type inference for generalized algebraic data types. ACM SIGPLAN Notices, 41(1):232–244, January 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Flanagan:2006:HTC

[1515] Cormac Flanagan. Hybrid type checking. ACM SIGPLAN Notices, 41(1): 245–256, January 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kim:2006:PMT

[1516] Ik-Soon Kim, Kwangkeun Yi, and Cristiano Calcagno. A polymorphic modal type system for Lisp-like multistaged languages. ACM SIGPLAN Notices, 41(1):257–268, January 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sweeney:2006:NMP

[1517] Tim Sweeney. The next mainstream programming language: a game developer’s perspective. ACM SIGPLAN Notices, 41(1):269, January 2006. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ernst:2006:VCC

[1518] Erik Ernst, Klaus Ostermann, and William R. Cook. A virtual class calculus. ACM SIGPLAN Notices, 41(1): 270–282, January 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Liu:2006:II

[1519] Jed Liu, Aaron Kimball, and Andrew C. Myers. Interruptible iterators. ACM SIGPLAN Notices, 41(1): 283–294, January 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reis:2006:SCC

[1520] Gabriel Dos Reis and Bjarne Stroustrup. Specifying C++ concepts. ACM SIGPLAN Notices, 41(1):295–308, January 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Thielecke:2006:FRA

[1521] Hayo Thielecke. Frame rules from answer types for code pointers. ACM SIGPLAN Notices, 41(1):309–319, January 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ni:2006:CAP

[1522] Zhaozhong Ni and Zhong Shao. Certified assembly programming with embedded code pointers. ACM SIGPLAN Notices, 41(1):320–333, January 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vaziri:2006:ASC

[1523] Mandana Vaziri, Frank Tip, and Julian Dolby. Associating synchronization constraints with data in an objectoriented language. ACM SIGPLAN Notices, 41(1):334–345, January 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McCloskey:2006:ASI

[1524] Bill McCloskey, Feng Zhou, David Gay, and Eric Brewer. Autolocker: synchronization inference for atomic sections. ACM SIGPLAN Notices, 41(1): 346–358, January 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lu:2006:PRE

[1525] Yi Lu and John Potter. Protecting representation with effect encapsulation. ACM SIGPLAN Notices, 41(1): 359–371, January 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Su:2006:ECI

[1526] Zhendong Su and Gary Wassermann. The essence of command injection attacks in Web applications. ACM SIG-PLAN Notices, 41(1):372–382, January 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dantas:2006:HA

[1527] Daniel S. Dantas. Harmless advice. ACM SIGPLAN Notices, 41(1):383– 396, January 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Menon:2006:VSP

[1528] Vijay S. Menon, Neal Glew, Brian R. Murphy, Andrew McCreight, Tatiana Shpeisman, Ali-Reza Adl-Tabatabai, and Leaf Petersen. A verifiable SSA program representation for aggressive compiler optimization. ACM SIG-PLAN Notices, 41(1):397–408, January 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Olinsky:2006:SAC

[1529] Reuben Olinsky, Christian Lindig, and Norman Ramsey. Staged allocation: a compositional technique for specifying and implementing procedure calling conventions. ACM SIGPLAN Notices, 41(1):409–421, January 2006.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xue:2006:DSL

[1530] Deyi Xue. Developing a superset language of Smalltalk for modeling distributed mechanical design systems. ACM SIGPLAN Notices, 41(2): 8–17, February 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Borah:2006:IOO

[1531] Kallol Borah. Indus: an object oriented language for Ubiquitous computing. ACM SIGPLAN Notices, 41(2): 18–24, February 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2006:MPG

[1532] Anonymous. MxYacc: parser generator for mixed languages. ACM SIG-PLAN Notices, 41(2):25–28, February 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Murphy:2006:SSU

[1533] Mike Murphy. Some suggestions for the use of tables in compilers. ACM SIG-PLAN Notices, 41(2):29–33, February 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bailey:2006:PPA

[1534] Mark W. Bailey, John Hatcliff, and Frank Tip. PEPM 2006 paper abstracts. ACM SIGPLAN Notices, 41 (3):4–8, March 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Janakiram:2006:SCL

[1535] D. Janakiram, R. Venkateswarlu, A. Vijay Srinivas, and A. Uday Kumar. Soma: a compositional language for distributed systems. ACM SIGPLAN Notices, 41(3):10–19, March 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ning:2006:MGW

[1536] Zhang Ning and Guangze Xiong. Minimizing GC work by analysis of live objects. ACM SIGPLAN Notices, 41 (3):20–29, March 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dekker:2006:LFP

[1537] Anthony H. Dekker. Lazy functional programming in Java. ACM SIG-PLAN Notices, 41(3):30–39, March 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chang:2006:SCA

[1538] Byeong-Mo Chang. Static check analysis for Java stack inspection. ACM SIGPLAN Notices, 41(3):40–48, March 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:2006:TYF

[1539] Paul Frenger. Ten years of Forth in ACM Sigplan Notices: part 2. ACM SIGPLAN Notices, 41(4):3–14, April 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mann:2006:TBG

[1540] Paul B. Mann. A translational BNF grammar notation (TBNF). ACM SIGPLAN Notices, 41(4):16–23, April 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lenka:2006:SML

[1541] Dillip Kumar Lenka and Pawan Kumar. States merging in LR parser. ACM SIGPLAN Notices, 41(4):24–29, April 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jiang:2006:AAR

[1542] Shujuan Jiang, Baowen Xu, and Liang Shi. An approach to analyzing recursive programs with exception handling constructs. ACM SIGPLAN Notices, 41(4):30–35, April 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tuya:2006:PGS

[1543] Javier Tuya, M. José Suárez-Cabal, and Claudio de la Riva. A practical guide to SQL white-box testing. ACM SIGPLAN Notices, 41(4):36–41, April 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Loka:2006:CRH

[1544] Raghavendra Rao Loka. Compilation reuse and hybrid compilation: an experiment. ACM SIGPLAN Notices, 41 (4):42–49, April 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Behera:2006:IAL

[1545] Chandan Kumar Behera and Pawan Kumar. An improved algorithm for loop dead optimization. ACM SIG-PLAN Notices, 41(5):11–20, May 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Moon:2006:TMS

[1546] Sewon Moon and Byeong-Mo Chang. A thread monitoring system for multithreaded Java programs. ACM SIG-PLAN Notices, 41(5):21–29, May 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:2006:MPS

[1547] Yingzhou Zhang, Jose Emilio Labra Gayo, and Agustín Cernuda del Río. A monadic program slicer. ACM SIG-PLAN Notices, 41(5):30–38, May 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Adams:2006:FAM

[1548] Michael D. Adams and David S. Wise. Fast additions on masked integers. ACM SIGPLAN Notices, 41(5):39–45, May 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frost:2006:NTP

[1549] Richard A. Frost and Rahmatullah Hafiz. A new top-down parsing algorithm to accommodate ambiguity and left recursion in polynomial time. ACM SIGPLAN Notices, 41(5):46–54, May 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Carlstrom:2006:ATP

[1550] Brian D. Carlstrom, Austen McDonald, Hassan Chafi, JaeWoong Chung, Chi Cao Minh, Christos Kozyrakis, and Kunle Olukotun. The ATOMOS transactional programming language. ACM SIGPLAN Notices, 41(6):1–13, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Harris:2006:OMT

[1551] Tim Harris, Mark Plesko, Avraham Shinnar, and David Tarditi. Optimizing memory transactions. ACM SIG-PLAN Notices, 41(6):14–25, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Adl-Tabatabai:2006:CRS

[1552] Ali-Reza Adl-Tabatabai, Brian T.

Lewis, Vijay Menon, Brian R. Murphy, Bratin Saha, and Tatiana Shpeisman. Compiler and runtime support for efficient software transactional memory. ACM SIGPLAN Notices, 41(6):26–37, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Grimm:2006:BET

[1553] Robert Grimm. Better extensibility through modular syntax. ACM SIG-PLAN Notices, 41(6):38–51, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ertl:2006:FFI

[1554] M. Anton Ertl, Kevin Casey, and David Gregg. Fast and flexible instruction selection with on-demand tree-parsing automata. ACM SIGPLAN Notices, 41 (6):52–60, June 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Triantafyllis:2006:FUW

[1555] Spyridon Triantafyllis, Matthew J. Bridges, Easwaran Raman, Guilherme Ottoni, and David I. August. A framework for unrestricted whole-program optimization. ACM SIGPLAN Notices, 41(6):61–71, June 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Neamtiu:2006:PDS

[1556] Iulian Neamtiu, Michael Hicks, Gareth Stoyle, and Manuel Oriol. Practical dynamic software updating for C. ACM SIGPLAN Notices, 41(6):72–83, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fei:2006:APR

[1557] Long Fei and Samuel P. Midkiff. Artemis: practical runtime monitoring of applications for execution anomalies. ACM SIGPLAN Notices, 41(6):84–95, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Acar:2006:EAS

[1558] Umut A. Acar, Guy E. Blelloch, Matthias Blume, and Kanat Tangwongsan. An experimental analysis of self-adjusting computation. ACM SIGPLAN Notices, 41(6):96–107, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Barton:2006:SMP

[1559] Christopher Barton, Călin Casçaval, George Almási, Yili Zheng, Montse Farreras, Siddhartha Chatterje, and José Nelson Amaral. Shared memory programming for large scale machines. ACM SIGPLAN Notices, 41 (6):108–117, June 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ren:2006:ODP

[1560] Gang Ren, Peng Wu, and David Padua. Optimizing data permutations for SIMD devices. ACM SIG-PLAN Notices, 41(6):118–131, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nuzman:2006:AVI

[1561] Dorit Nuzman, Ira Rosen, and Ayal Zaks. Auto-vectorization of interleaved data for SIMD. ACM SIG-PLAN Notices, 41(6):132–143, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dhurjati:2006:SEA

[1562] Dinakar Dhurjati, Sumant Kowshik, and Vikram Adve. SAFECode: enforcing alias analysis for weakly typed languages. ACM SIGPLAN Notices, 41 (6):144–157, June 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berger:2006:DPM

[1563] Emery D. Berger and Benjamin G. Zorn. DieHard: probabilistic memory safety for unsafe languages. ACM SIG-PLAN Notices, 41(6):158–168, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:2006:PDS

[1564] Xiangyu Zhang, Neelam Gupta, and Rajiv Gupta. Pruning dynamic slices with confidence. ACM SIGPLAN Notices, 41(6):169–180, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Johnson:2006:CSD

[1565] Troy A. Johnson and Rudolf Eigenmann. Context-sensitive domainindependent algorithm composition and selection. ACM SIGPLAN Notices, 41(6):181–192, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:2006:RNE

[1566] Guangyu Chen, Feihui Li, Mahmut Kandemir, and Mary Jane Irwin. Reducing NoC energy consumption through compiler-directed channel voltage scaling. ACM SIGPLAN Notices, 41(6):193–203, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Koes:2006:GPR

[1567] David Ryan Koes and Seth Copen Goldstein. A global progressive register allocator. ACM SIGPLAN Notices, 41 (6):204–215, June 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nakaike:2006:PBG

[1568] Takuya Nakaike, Tatsushi Inagaki, Hideaki Komatsu, and Toshio Nakatani. Profile-based global live-range splitting. ACM SIGPLAN Notices, 41(6): 216–227, June 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bridges:2006:AIS

[1569] Matthew J. Bridges, Neil Vachharajani, Guilherme Ottoni, and David I. August. Automatic instruction scheduler retargeting by reverse-engineering. ACM SIGPLAN Notices, 41(6):228– 238, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lau:2006:OPA

[1570] Jeremy Lau, Matthew Arnold, Michael Hind, and Brad Calder. Online performance auditing: using hot optimizations without getting burned. ACM SIGPLAN Notices, 41(6):239– 251, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chilimbi:2006:CCC

[1571] Trishul M. Chilimbi and Ran Shaham. Cache-conscious coallocation of hot data streams. ACM SIG-PLAN Notices, 41(6):252–262, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhuang:2006:AEA

[1572] Xiaotong Zhuang, Mauricio J. Serrano, Harold W. Cain, and Jong-Deok Choi. Accurate, efficient, and adaptive calling context profiling. ACM SIG-PLAN Notices, 41(6):263–271, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jarvi:2006:ASG

[1573] Jaakko Järvi, Douglas Gregor, Jeremiah Willcock, Andrew Lumsdaine, and Jeremy Siek. Algorithm specialization in generic programming: challenges of constrained generics in C++. ACM SIGPLAN Notices, 41(6):272– 282, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Spoonhower:2006:ESP

[1574] Daniel Spoonhower, Joshua Auerbach, David F. Bacon, Perry Cheng, and David Grove. Eventrons: a safe programming construct for high-frequency hard real-time applications. ACM SIG-PLAN Notices, 41(6):283–294, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shivers:2006:CTC

[1575] Olin Shivers and Matthew Might. Continuations and transducer composition. ACM SIGPLAN Notices, 41(6): 295–307, June 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Naik:2006:ESR

[1576] Mayur Naik, Alex Aiken, and John Whaley. Effective static race detection for Java. ACM SIGPLAN Notices, 41 (6):308–319, June 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pratikakis:2006:LCS

[1577] Polyvios Pratikakis, Jeffrey S. Foster, and Michael Hicks. LOCK-SMITH: context-sensitive correlation analysis for race detection. ACM SIG-PLAN Notices, 41(6):320–331, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:2006:PGP

[1578] Wen ke Chen, Sanjay Bhansali, Trishul Chilimbi, Xiaofeng Gao, and Weihaw Chuang. Profile-guided proactive garbage collection for locality optimization. ACM SIGPLAN Notices, 41(6): 332–340, June 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vechev:2006:CPD

[1579] Martin T. Vechev, Eran Yahav, and David F. Bacon. Correctnesspreserving derivation of concurrent garbage collection algorithms. ACM SIGPLAN Notices, 41(6):341–353, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kermany:2006:CCI

[1580] Haim Kermany and Erez Petrank. The Compressor: concurrent, incremental, and parallel compaction. ACM SIG-PLAN Notices, 41(6):354–363, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Guyer:2006:FMS

[1581] Samuel Z. Guyer, Kathryn S. McKinley, and Daniel Frampton. Free-Me: a static analysis for automatic individual object reclamation. ACM SIG-PLAN Notices, 41(6):364–375, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gulwani:2006:CAI

[1582] Sumit Gulwani and Ashish Tiwari. Combining abstract interpreters. ACM SIGPLAN Notices, 41(6):376–386, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sridharan:2006:RBC

[1583] Manu Sridharan and Rastislav Bodík. Refinement-based context-sensitive points-to analysis for Java. ACM SIG-PLAN Notices, 41(6):387–400, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Feng:2006:MVA

[1584] Xinyu Feng, Zhong Shao, Alexander Vaynberg, Sen Xiang, and Zhaozhong Ni. Modular verification of assembly code with stack-based control abstractions. ACM SIGPLAN Notices, 41(6): 401–414, June 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cook:2006:TPS

[1585] Byron Cook, Andreas Podelski, and Andrey Rybalchenko. Termination proofs for systems code. ACM SIG-PLAN Notices, 41(6):415–426, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Martonosi:2006:ESW

[1586] Margaret Martonosi. Embedded systems in the wild: ZebraNet software, hardware, and deployment experiences. ACM SIGPLAN Notices, 41(7):1, July 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vonPlaten:2006:FLO

[1587] Carl von Platen and Johan Eker.

Feedback linking: optimizing object code layout for updates. ACM SIG-PLAN Notices, 41(7):2–11, July 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:2006:MDS

[1588] Kun Zhang and Santosh Pande. Minimizing downtime in seamless migrations of mobile applications. ACM SIG-PLAN Notices, 41(7):12–21, July 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Spivak:2006:SPT

[1589] Michal Spivak and Sivan Toledo. Storing a persistent transactional object heap on flash memory. ACM SIG-PLAN Notices, 41(7):22–33, July 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Regehr:2006:DAT

[1590] John Regehr and Usit Duongsaa. Deriving abstract transfer functions for analyzing embedded software. ACM SIGPLAN Notices, 41(7):34–43, July 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cooprider:2006:PAD

[1591] Nathan Cooprider and John Regehr. Pluggable abstract domains for analyzing embedded software. ACM SIG-PLAN Notices, 41(7):44–53, July 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mine:2006:FSV

[1592] Antoine Miné. Field-sensitive value analysis of embedded C programs with union types and pointer arithmetics. ACM SIGPLAN Notices, 41(7):54–63, July 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kreahling:2006:RCC

[1593] William Kreahling, Stephen Hines, David Whalley, and Gary Tyson. Reducing the cost of conditional transfers of control by using comparison specifications. ACM SIGPLAN Notices, 41 (7):64–71, July 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhuang:2006:ETM

[1594] Xiaotong Zhuang and Santosh Pande. Effective thread management on network processors with compiler analysis. ACM SIGPLAN Notices, 41(7):72–82, July 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kulkarni:2006:SNO

[1595] Prasad A. Kulkarni, David B. Whalley, Gary S. Tyson, and Jack W. Davidson. In search of near-optimal optimization phase orderings. ACM SIG-PLAN Notices, 41(7):83–92, July 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Danne:2006:EST

[1596] Klaus Danne and Marco Platzner. An EDF schedulability test for periodic tasks on reconfigurable hardware devices. ACM SIGPLAN Notices, 41 (7):93–102, July 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sandberg:2006:FWF

[1597] Christer Sandberg, Andreas Ermedahl, Jan Gustafsson, and Björn Lisper. Faster WCET flow analysis by program slicing. ACM SIGPLAN Notices, 41 (7):103–112, July 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Prochnow:2006:SSS

[1598] Steffen Prochnow, Claus Traulsen, and Reinhard von Hanxleden. Synthesizing safe state machines from Esterel. ACM SIGPLAN Notices, 41(7):113–124, July 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Edwards:2006:ECG

[1599] Stephen A. Edwards and Olivier Tardieu. Efficient code generation from SHIM models. ACM SIGPLAN Notices, 41(7):125–134, July 2006.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rothamel:2006:GOC

[1600] Tom Rothamel, Yanhong A. Liu, Constance L. Heitmeyer, and Elizabeth I. Leonard. Generating optimized code from SCR specifications. ACM SIG-PLAN Notices, 41(7):135–144, July 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Farfeleder:2006:ECG

[1601] Stefan Farfeleder, Andreas Krall, Edwin Steiner, and Florian Brandner. Effective compiler generation by architecture description. ACM SIG-PLAN Notices, 41(7):145–152, July 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:2006:PLA

[1602] Jian-Jia Chen and Tei-Wei Kuo. Procrastination for leakage-aware ratemonotonic scheduling on a dynamic voltage scaling processor. ACM SIG-PLAN Notices, 41(7):153–162, July 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mutyam:2006:CDT

[1603] Madhu Mutyam, Feihui Li, Vijaykrishnan Narayanan, Mahmut Kandemir, and Mary Jane Irwin. Compilerdirected thermal management for VLIW functional units. ACM SIG-PLAN Notices, 41(7):163–172, July 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Park:2006:BAI

[1604] Sanghyun Park, Aviral Shrivastava, Nikil Dutt, Alex Nicolau, Yunheung Paek, and Eugene Earlie. Bypass aware instruction scheduling for register file power reduction. ACM SIG-PLAN Notices, 41(7):173–181, July 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yan:2006:ADE

[1605] Leipo Yan, Thambipillai Srikanthan, and Niu Gang. Area and delay estimation for FPGA implementation of coarse-grained reconfigurable architectures. ACM SIGPLAN Notices, 41 (7):182–188, July 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pandey:2006:BCB

[1606] Raju Pandey and Jeffrey Wu. BOTS:

a constraint-based component system for synthesizing scalable software systems. ACM SIGPLAN Notices, 41 (7):189–198, July 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:2006:OCS

[1607] Weihua Zhang, Xinglong Qian, Ye Wang, Binyu Zang, and Chuanqi Zhu. Optimizing compiler for shared-memory multiple SIMD architecture. ACM SIGPLAN Notices, 41(7):199–208, July 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Findler:2006:SFP

[1608] Robert Bruce Findler. Scheme and Functional Programming 2006: paper abstracts. ACM SIGPLAN Notices, 41 (8):6–9, August 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:2006:BRD

[1609] Paul Frenger. Book review: Designing Embedded Hardware. ACM SIG-PLAN Notices, 41(8):10–14, August 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Denckla:2006:MCB

[1610] Ben Denckla. Many cyclic block diagrams do not need parallel semantics. ACM SIGPLAN Notices, 41(8):16–20, August 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Das:2006:OSO

[1611] Dibyendu Das. Optimizing subroutines with optional parameters in F90 via function cloning. ACM SIG-PLAN Notices, 41(8):21–28, August 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:2006:ICT

[1612] Fuh-Gwo Chen and Ting-Wei Hou. Instruction-coated translation: an approach to restructure directly threaded interpreters with low cohesion. ACM SIGPLAN Notices, 41(8):29–33, August 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Velazquez-Iturbide:2006:PDL

[1613] J. Ángel Velázquez-Iturbide. Principled design of logical fisheye views of functional expressions. ACM SIG-PLAN Notices, 41(8):34–43, August 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dybvig:2006:DCS

[1614] R. Kent Dybvig. The development of Chez Scheme. ACM SIGPLAN Notices, 41(9):1–12, September 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Might:2006:IFA

[1615] Matthew Might and Olin Shivers. Improving flow analyses via ΓCFA: abstract garbage collection and counting. ACM SIGPLAN Notices, 41(9): 13–25, September 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kiselyov:2006:DDB

[1616] Oleg Kiselyov, Chung chieh Shan, and Amr Sabry. Delimited dynamic binding. ACM SIGPLAN Notices, 41(9): 26–37, September 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Walker:2006:STF

[1617] David Walker, Lester Mackey, Jay Ligatti, George A. Reis, and David I. August. Static typing for a faulty lambda calculus. ACM SIGPLAN Notices, 41(9):38–49, September 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jones:2006:SUB

[1618] Simon Peyton Jones, Dimitrios Vytiniotis, Stephanie Weirich, and Geoffrey Washburn. Simple unification-based type inference for GADTs. ACM SIG-PLAN Notices, 41(9):50–61, September 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nanevski:2006:PSH

[1619] Aleksandar Nanevski, Greg Morrisett, and Lars Birkedal. Polymorphism and separation in Hoare type theory. ACM SIGPLAN Notices, 41(9): 62–73, September 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nakata:2006:RMP

[1620] Keiko Nakata and Jacques Garrigue. Recursive modules for programming. ACM SIGPLAN Notices, 41(9):74– 86, September 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Owens:2006:SFM

[1621] Scott Owens and Matthew Flatt. From structures and functors to modules and units. ACM SIGPLAN Notices, 41(9): 87–98, September 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rossberg:2006:MLD

[1622] Andreas Rossberg. The missing link: dynamic components for ML. ACM SIGPLAN Notices, 41(9):99– 110, September 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fisher:2006:SAS

[1623] David Fisher and Olin Shivers. Static analysis for syntax objects. ACM SIG-PLAN Notices, 41(9):111–121, September 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shapiro:2006:PPC

[1624] Marc Shapiro. Practical proofs of concurrent programs. ACM SIGPLAN Notices, 41(9):123, September 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Donnelly:2006:TE

[1625] Kevin Donnelly and Matthew Fluet.

Transactional events. ACM SIG-PLAN Notices, 41(9):124–135, September 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ziarek:2006:SMC

[1626] Lukasz Ziarek, Philip Schatz, and Suresh Jagannathan. Stabilizers: a modular checkpointing abstraction for concurrent functional programs. ACM SIGPLAN Notices, 41(9):136– 147, September 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dowse:2006:MDC

[1627] Malcolm Dowse and Andrew Butterfield. Modelling deterministic concurrent I/O. ACM SIGPLAN Notices, 41(9):148–159, September 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chlipala:2006:MDC

[1628] Adam Chlipala. Modular development of certified program verifiers with a proof assistant. ACM SIGPLAN Notices, 41(9):160–171, September 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hickey:2006:MMR

[1629] Jason Hickey, Aleksey Nogin, Xin Yu, and Alexei Kopylov. Mechanized metareasoning using a hybrid HOAS/de Bruijn representation and reflection. ACM SIGPLAN Notices, 41(9):172– 183, September 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rosu:2006:ESC

[1630] Grigore Roşu. Equality of streams is a Π0 over 2-complete problem. ACM SIGPLAN Notices, 41(9):184– 191, September 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frisch:2006:OX

[1631] Alain Frisch. OCaml + XDuce. ACM SIGPLAN Notices, 41(9):192– 200, September 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kawanaka:2006:BBT

[1632] Shinya Kawanaka and Haruo Hosoya. biXid: a bidirectional transformation language for XML. ACM SIG-PLAN Notices, 41(9):201–214, September 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bird:2006:FYF

[1633] Richard S. Bird. Fifteen years of functional pearls. ACM SIGPLAN Notices, 41(9):215, September 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bringert:2006:PAC

[1634] Björn Bringert and Aarne Ranta. A pattern for almost compositional functions. ACM SIGPLAN Notices, 41(9): 216–226, September 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Katsumata:2006:AFF

[1635] Shin ya Katsumata and Susumu Nishimura. Algebraic fusion of functions with an accumulating parameter and its improvement. ACM SIG-PLAN Notices, 41(9):227–238, September 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Blume:2006:EPF

[1636] Matthias Blume, Umut A. Acar, and Wonseok Chae. Extensible programming with first-class cases. ACM SIG-PLAN Notices, 41(9):239–250, September 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vytiniotis:2006:BTI

[1637] Dimitrios Vytiniotis, Stephanie Weirich, and Simon Peyton Jones. Boxy types: inference for higher-rank types and impredicativity. ACM SIGPLAN Notices, 41(9):251–262, September 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Abadi:2006:ACC

[1638] Martín Abadi. Access control in a core calculus of dependency. ACM SIG-PLAN Notices, 41(9):263–273, September 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lin:2006:PMO

[1639] Chuan kai Lin. Programming monads operationally with Unimo. ACM SIG-PLAN Notices, 41(9):274–285, September 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Denielou:2006:APS

[1640] Pierre-Malo Deniélou and James J. Leifer. Abstraction preservation and subtyping in distributed languages. ACM SIGPLAN Notices, 41(9):286– 297, September 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hutchins:2006:EDCa

[1641] DeLesley Hutchins. Eliminating distinctions of class: using prototypes to model virtual classes. ACM SIG-PLAN Notices, 41(10):1–20, October 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nystrom:2006:JNIa

[1642] Nathaniel Nystrom, Xin Qi, and Andrew C. Myers. J&: nested intersection for scalable software composition. ACM SIGPLAN Notices, 41(10): 21–36, October 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Warth:2006:SSOa

[1643] Alessandro Warth, Milan Stanojević, and Todd Millstein. Statically scoped object adaptation with expanders. ACM SIGPLAN Notices, 41(10):37– 56, October 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Andreae:2006:FIP

[1644] Chris Andreae, James Noble, Shane Markstrum, and Todd Millstein. A framework for implementing pluggable type systems. ACM SIGPLAN Notices, 41(10):57–74, October 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fairbanks:2006:DFMa

[1645] George Fairbanks, David Garlan, and William Scherlis. Design fragments make using frameworks easier. ACM SIGPLAN Notices, 41(10):75–88, October 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cohen:2006:JJTa

[1646] Tal Cohen, Joseph (Yossi) Gil, and Itay Maman. JTL: the Java tools language. ACM SIGPLAN Notices, 41(10):89– 108, October 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bockisch:2006:AVMa

[1647] Christoph Bockisch, Matthew Arnold, Tom Dinkelaker, and Mira Mezini. Adapting virtual machine techniques for seamless aspect support. ACM SIG-PLAN Notices, 41(10):109–124, October 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bockisch:2006:ECFa

[1648] Christoph Bockisch, Sebastian Kanthak, Michael Haupt, Matthew Arnold, and Mira Mezini. Efficient control flow quantification. ACM SIGPLAN Notices, 41(10):125–138, October 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Eugster:2006:UPJa

[1649] Patrick Eugster. Uniform proxies for Java. ACM SIGPLAN Notices, 41(10): 139–152, October 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Maebe:2006:JSBa

[1650] Jonas Maebe, Dries Buytaert, Lieven Eeckhout, and Koen De Bosschere. Javana: a system for building customized Java program analysis tools. ACM SIGPLAN Notices, 41(10):153– 168, October 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Blackburn:2006:DBJa

[1651] Stephen M. Blackburn, Robin Garner, Chris Hoffmann, Asjad M. Khang, Kathryn S. McKinley, Rotem Bentzur, Amer Diwan, Daniel Feinberg, Daniel Frampton, Samuel Z. Guyer, Martin Hirzel, Antony Hosking, Maria Jump, Han Lee, J. Eliot B. Moss, B. Moss, Aashish Phansalkar, Darko Stefanović, Thomas VanDrunen, Daniel von Dincklage, and Ben Wiedermann. The Da-Capo benchmarks: Java benchmarking development and analysis. ACM SIG-PLAN Notices, 41(10):169–190, October 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Titzer:2006:VOHa

[1652] Ben L. Titzer. Virgil: objects on the head of a pin. ACM SIGPLAN Notices, 41(10):191–208, October 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bravenboer:2006:DFEa

[1653] Martin Bravenboer, Éric Tanter, and Eelco Visser. Declarative, formal, and extensible syntax definition for aspectJ. ACM SIGPLAN Notices, 41(10):209– 228, October 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cavazos:2006:MSDa

[1654] John Cavazos and Michael F. P.

O’Boyle. Method-specific dynamic compilation using logistic regression. ACM SIGPLAN Notices, 41(10):229– 240, October 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ogata:2006:RCIa

[1655] Kazunori Ogata, Tamiya Onodera, Kiyokuni Kawachiya, Hideaki Komatsu, and Toshio Nakatani. Replay compilation: improving debuggability of a just-in-time compiler. ACM SIG-PLAN Notices, 41(10):241–252, October 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Herlihy:2006:FFIa

[1656] Maurice Herlihy, Victor Luchangco, and Mark Moir. A flexible framework for implementing software transactional memory. ACM SIGPLAN Notices, 41(10):253–262, October 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Russell:2006:ESRa

[1657] Kenneth Russell and David Detlefs. Eliminating synchronization-related atomic operations with biased locking and bulk rebiasing. ACM SIGPLAN Notices, 41(10):263–272, October 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chin:2006:FBAa

[1658] Wei-Ngan Chin, Florin Craciun, Siau-Cheng Khoo, and Corneliu Popeea. A flow-based approach for variant parametric types. ACM SIGPLAN Notices, 41(10):273–290, October 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gregor:2006:CLSa

[1659] Douglas Gregor, Jaakko Järvi, Jeremy Siek, Bjarne Stroustrup, Gabriel Dos Reis, and Andrew Lumsdaine. Concepts: linguistic support for generic programming in C++. ACM SIG-PLAN Notices, 41(10):291–310, October 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Potanin:2006:GOGa

[1660] Alex Potanin, James Noble, Dave Clarke, and Robert Biddle. Generic ownership for generic Java. ACM SIG-PLAN Notices, 41(10):311–324, October 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Liu:2006:FFCa

[1661] Yu David Liu and Scott F. Smith.

A formal framework for component deployment. ACM SIGPLAN Notices, 41(10):325–344, October 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wasserrab:2006:OSTa

[1662] Daniel Wasserrab, Tobias Nipkow, Gregor Snelting, and Frank Tip. An operational semantics and type safety proof for multiple inheritance in C++. ACM SIGPLAN Notices, 41(10):345– 362, October 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Darga:2006:ESMa

[1663] Paul T. Darga and Chandrasekhar Boyapati. Efficient software model checking of data structure properties. ACM SIGPLAN Notices, 41(10):363– 382, October 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kit:2006:IRCa

[1664] Lo Kwun Kit, Chan Kwun Man, and Elisa Baniassad. Isolating and relating concerns in requirements using latent semantic analysis. ACM SIGPLAN Notices, 41(10):383–396, October 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baxter:2006:USJa

[1665] Gareth Baxter, Marcus Frean, James Noble, Mark Rickerby, Hayden Smith, Matt Visser, Hayden Melton, and Ewan Tempero. Understanding the shape of Java software. ACM SIG-PLAN Notices, 41(10):397–412, October 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sahavechaphan:2006:XMSa

[1666] Naiyana Sahavechaphan and Kajal Claypool. XSnippet: mining For sample code. ACM SIGPLAN Notices, 41(10):413–430, October 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gabriel:2006:CSa

[1667] Richard P. Gabriel and Ron Goldman. Conscientious software. ACM SIG-PLAN Notices, 41(10):433–450, October 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Simonyi:2006:ISa

[1668] Charles Simonyi, Magnus Christerson, and Shane Clifford. Intentional software. ACM SIGPLAN Notices, 41(10): 451–464, October 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Waldo:2006:SDa

[1669] Jim Waldo. On system design.

ACM SIGPLAN Notices, 41(10):467– 480, October 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Steimann:2006:PSAa

[1670] Friedrich Steimann. The paradoxical success of aspect-oriented programming. ACM SIGPLAN Notices, 41(10): 481–497, October 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rosenblum:2006:IVC

[1671] Mendel Rosenblum. Impact of virtualization on computer architecture and operating systems. ACM SIG-PLAN Notices, 41(11):1, November 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Adams:2006:CSH

[1672] Keith Adams and Ole Agesen. A comparison of software and hardware techniques for x86 virtualization. ACM SIGPLAN Notices, 41(11): 2–13, November 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jones:2006:GMB

[1673] Stephen T. Jones, Andrea C. Arpaci-Dusseau, and Remzi H. Arpaci-Dusseau. Geiger: monitoring the buffer cache in a virtual machine environment. ACM SIGPLAN Notices, 41(11): 14–24, November 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Crandall:2006:TSD

[1674] Jedidiah R. Crandall, Gary Wassermann, Daniela A. S. de Oliveira, Zhendong Su, S. Felix Wu, and Frederic T. Chong. Temporal search: detecting hidden malware timebombs with virtual machines. ACM SIGPLAN Notices, 41(11):25–36, November 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lu:2006:ADA

[1675] Shan Lu, Joseph Tucek, Feng Qin, and Yuanyuan Zhou. AVIO: detecting atomicity violations via access interleaving invariants. ACM SIGPLAN Notices, 41(11):37–48, November 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xu:2006:RTR

[1676] Min Xu, Mark D. Hill, and Rastislav Bodik. A regulated transitive reduction (RTR) for longer memory race recording. ACM SIGPLAN Notices, 41(11): 49–60, November 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bond:2006:BBE

[1677] Michael D. Bond and Kathryn S.

McKinley. Bell: bit-encoding online memory leak detection. ACM SIG-PLAN Notices, 41(11):61–72, November 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shyam:2006:ULC

[1678] Smitha Shyam, Kypros Constantinides, Sujay Phadke, Valeria Bertacco, and Todd Austin. Ultra low-cost defect protection for microprocessor pipelines. ACM SIGPLAN Notices, 41(11):73– 82, November 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reddy:2006:UPB

[1679] Vimal K. Reddy, Eric Rotenberg, and Sailashri Parthasarathy. Understanding prediction-based partial redundant threading for low-overhead, high- coverage fault tolerance. ACM SIGPLAN Notices, 41(11):83–94, November 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Parashar:2006:SSB

[1680] Angshuman Parashar, Anand Sivasubramaniam, and Sudhanva Gurumurthi. SlicK: slice-based locality exploitation for efficient redundant multithreading. ACM SIGPLAN Notices, 41(11):95– 105, November 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Heath:2006:MFT

[1681] Taliver Heath, Ana Paula Centeno, Pradeep George, Luiz Ramos, and Yogesh Jaluria. Mercury and Freon: temperature emulation and management for server systems. ACM SIGPLAN Notices, 41(11):106–116, November 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kgil:2006:PUS

[1682] Taeho Kgil, Shaun D’Souza, Ali Saidi, Nathan Binkert, Ronald Dreslinski, Trevor Mudge, Steven Reinhardt, and Krisztian Flautner. PicoServer: using 3D stacking technology to enable a compact energy efficient chip multiprocessor. ACM SIGPLAN Notices, 41(11):117–128, November 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Coons:2006:SPS

[1683] Katherine E. Coons, Xia Chen, Doug Burger, Kathryn S. McKinley, and Sundeep K. Kushwaha. A spatial path scheduling algorithm for EDGE architectures. ACM SIGPLAN Notices, 41(11):129–140, November 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mercaldi:2006:IST

[1684] Martha Mercaldi, Steven Swanson, Andrew Petersen, Andrew Putnam, Andrew Schwerin, Mark Oskin, and Susan J. Eggers. Instruction scheduling for a tiled dataflow architecture. ACM SIGPLAN Notices, 41(11):141– 150, November 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gordon:2006:ECG

[1685] Michael I. Gordon, William Thies, and Saman Amarasinghe. Exploiting coarse-grained task, data, and pipeline parallelism in stream programs. ACM SIGPLAN Notices, 41 (11):151–162, November 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mishra:2006:TES

[1686] Mahim Mishra, Timothy J. Callahan, Tiberiu Chelcea, Girish Venkataramani, Seth C. Goldstein, and Mihai Budiu. Tartan: evaluating spatial computation for whole program execution. ACM SIGPLAN Notices, 41(11):163– 174, November 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Eyerman:2006:PCA

[1687] Stijn Eyerman, Lieven Eeckhout, Tejas Karkhanis, and James E. Smith. A performance counter architecture for computing accurate CPI components. ACM SIGPLAN Notices, 41(11):175– 184, November 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lee:2006:AER

[1688] Benjamin C. Lee and David M.

Brooks. Accurate and efficient regression modeling for microarchitectural performance and power prediction. ACM SIGPLAN Notices, 41(11): 185–194, November 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ipek:2006:EEA

[1689] Engin Ïpek, Sally A. McKee, Rich Caruana, Bronis R. de Supinski, and Martin Schulz. Efficiently exploring architectural design spaces via predictive modeling. ACM SIGPLAN Notices, 41(11):195–206, November 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kharbutli:2006:CEP

[1690] Mazen Kharbutli, Xiaowei Jiang, Yan Solihin, Guru Venkataramani, and Milos Prvulovic. Comprehensively and efficiently protecting the heap. ACM SIGPLAN Notices, 41(11):207– 218, November 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chilimbi:2006:HIH

[1691] Trishul M. Chilimbi and Vinod Ganapathy. HeapMD: identifying heapbased bugs using anomaly detection. ACM SIGPLAN Notices, 41(11):219– 228, November 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Narayanasamy:2006:RSM

[1692] Satish Narayanasamy, Cristiano Pereira, and Brad Calder. Recording shared memory dependencies using strata. ACM SIGPLAN Notices, 41(11):229– 240, November 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Patwardhan:2006:DTS

[1693] Jaidev P. Patwardhan, Vijeta Johri, Chris Dwyer, and Alvin R. Lebeck.

A defect tolerant self-organizing nanoscale SIMD architecture. ACM SIGPLAN Notices, 41(11):241–251, November 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schuchman:2006:PTA

[1694] Ethan Schuchman and T. N. Vijaykumar. A program transformation and architecture support for quantum uncomputation. ACM SIGPLAN Notices, 41(11):252–263, November 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mysore:2006:IC

[1695] Shashidhar Mysore, Banit Agrawal, Navin Srivastava, Sheng-Chih Lin, Kaustav Banerjee, and Tim Sherwood. Introspective 3D chips. ACM SIGPLAN Notices, 41(11):264–273, November 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cantin:2006:SP

[1696] Jason F. Cantin, Mikko H. Lipasti, and James E. Smith. Stealth prefetching. ACM SIGPLAN Notices, 41(11):274– 282, November 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chakraborty:2006:CSE

[1697] Koushik Chakraborty, Philip M. Wells, and Gurindar S. Sohi. Computation spreading: employing hardware migration to specialize CMP cores on-the-fly. ACM SIGPLAN Notices, 41(11):283– 292, November 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Miller:2006:SBI

[1698] Jason E. Miller and Anant Agarwal. Software-based instruction caching for embedded processors. ACM SIGPLAN Notices, 41(11):293–302, November 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Li:2006:MEM

[1699] Xin Li, Marian Boldt, and Reinhard von Hanxleden. Mapping Esterel onto a multi-threaded embedded processor. ACM SIGPLAN Notices, 41(11):303– 314, November 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Binkert:2006:INI

[1700] Nathan L. Binkert, Ali G. Saidi, and Steven K. Reinhardt. Integrated network interfaces for high-bandwidth TCP/IP. ACM SIGPLAN Notices, 41 (11):315–324, November 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tarditi:2006:AUD

[1701] David Tarditi, Sidd Puri, and Jose Oglesby. Accelerator: using data parallelism to program GPUs for generalpurpose uses. ACM SIGPLAN Notices, 41(11):325–335, November 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Damron:2006:HTM

[1702] Peter Damron, Alexandra Fedorova, Yossi Lev, Victor Luchangco, Mark Moir, and Daniel Nussbaum. Hybrid transactional memory. ACM SIGPLAN Notices, 41(11):336–346, November 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chuang:2006:UPB

[1703] Weihaw Chuang, Satish Narayanasamy, Ganesh Venkatesh, Jack Sampson, Michael Van Biesbrouck, Gilles Pokam, Brad Calder, and Osvaldo Colavin. Unbounded page-based transactional memory. ACM SIGPLAN Notices, 41 (11):347–358, November 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Moravan:2006:SNT

[1704] Michelle J. Moravan, Jayaram Bobba, Kevin E. Moore, Luke Yen, Mark D. Hill, Ben Liblit, Michael M. Swift, and David A. Wood. Supporting nested transactional memory in logTM. ACM SIGPLAN Notices, 41(11):359– 370, November 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chung:2006:TTM

[1705] JaeWoong Chung, Chi Cao Minh, Austen McDonald, Travis Skare, Hassan Chafi, Brian D. Carlstrom, Christos Kozyrakis, and Kunle Olukotun. Tradeoffs in transactional memory virtualization. ACM SIGPLAN Notices, 41(11):371–381, November 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kawahito:2006:NIR

[1706] Motohiro Kawahito, Hideaki Komatsu, Takao Moriyama, Hiroshi Inoue, and Toshio Nakatani. A new idiom recognition framework for exploiting hardware-assist instructions. ACM SIGPLAN Notices, 41(11):382–393, November 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bansal:2006:AGP

[1707] Sorav Bansal and Alex Aiken. Automatic generation of peephole superoptimizers. ACM SIGPLAN Notices, 41(11):394–403, November 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Solar-Lezama:2006:CSF

[1708] Armando Solar-Lezama, Liviu Tancau, Rastislav Bodik, Sanjit Seshia, and Vijay Saraswat. Combinatorial sketching for finite programs. ACM SIGPLAN Notices, 41(11):404– 415, November 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

DaSilva:2006:PPA

[1709] Jeff Da Silva and J. Gregory Steffan. A probabilistic pointer analysis for speculative optimizations. ACM SIGPLAN Notices, 41(11):416–425, November 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:2006:FSS

[1710] Paul Frenger. Forth & small systems report: potential unrealized: the MicroMED advisor. ACM SIGPLAN Notices, 41(12):10–16, December 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wu:2006:FSP

[1711] Fangjun Wu. Formal specification of program slicing. ACM SIGPLAN Notices, 41(12):18–27, December 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mannarswamy:2006:TTR

[1712] Sandya Mannarswamy, Shruti Doval, Hariharan Sandanagobalane, and Mahesha Nanjundaiah. TRICK: tracking and reusing compiler’s knowledge. ACM SIGPLAN Notices, 41(12):28– 37, December 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wang:2006:DIG

[1713] B. Wang and C. Pronk. Design and implementation of a GUI for the TLC model checker. ACM SIGPLAN Notices, 41(12):38–43, December 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dubey:2006:GCP

[1714] Alpana Dubey. Goodness criteria for programming language grammar rules. ACM SIGPLAN Notices, 41(12):44– 53, December 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tang:2007:PRI

[1715] Audrey Tang. Perl 6: reconciling the irreconcilable. ACM SIGPLAN Notices, 42(1):1, January 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Matthews:2007:OSM

[1716] Jacob Matthews and Robert Bruce Findler. Operational semantics for multi-language programs. ACM SIG-PLAN Notices, 42(1):3–10, January 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Avgustinov:2007:SSP

[1717] Pavel Avgustinov, Elnar Hajiyev, Neil Ongkingco, Oege de Moor, Damien Sereni, Julian Tibble, and Mathieu Verbaere. Semantics of static pointcuts in aspectJ. ACM SIGPLAN Notices, 42 (1):11–23, January 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:2007:TIL

[1718] Juan Chen. A typed intermediate language for compiling multiple inheritance. ACM SIGPLAN Notices, 42(1): 25–30, January 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jump:2007:CDM

[1719] Maria Jump and Kathryn S. McKinley. Cork: dynamic memory leak detection for garbage-collected languages. ACM SIGPLAN Notices, 42(1):31–38, January 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Polishchuk:2007:DHT

[1720] Marina Polishchuk, Ben Liblit, and Chloë W. Schulze. Dynamic heap type inference for program understanding and debugging. ACM SIGPLAN Notices, 42(1):39–46, January 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Godefroid:2007:CDT

[1721] Patrice Godefroid. Compositional dynamic test generation. ACM SIG-PLAN Notices, 42(1):47–54, January 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shen:2007:LAU

[1722] Xipeng Shen, Jonathan Shaw, Brian Meeker, and Chen Ding. Locality approximation using time. ACM SIG-PLAN Notices, 42(1):55–61, January 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dreyer:2007:MTC

[1723] Derek Dreyer, Robert Harper, Manuel M. T. Chakravarty, and Gabriele Keller. Modular type classes. ACM SIGPLAN Notices, 42(1):63–70, January 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Siskind:2007:FCN

[1724] Jeffrey Mark Siskind and Barak A.

Pearlmutter. First-class nonstandard interpretations by opening closures. ACM SIGPLAN Notices, 42(1):71–76, January 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mandelbaum:2007:PMF

[1725] Yitzhak Mandelbaum, Kathleen Fisher, David Walker, Mary Fernandez, and Artem Gleyzer. PADS/ML: a functional data description language. ACM SIGPLAN Notices, 42(1):77–83, January 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pitts:2007:GUN

[1726] Andrew M. Pitts and Mark R. Shinwell. Generative unbinding of names. ACM SIGPLAN Notices, 42(1):85–95, January 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tuch:2007:TBS

[1727] Harvey Tuch, Gerwin Klein, and Michael Norrish. Types, bytes, and separation logic. ACM SIGPLAN Notices, 42(1):97–108, January 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Appel:2007:VMM

[1728] Andrew W. Appel, Paul-André Melliès, Christopher D. Richards, and Jérôme Vouillon. A very modal model of a modern, major, general type system. ACM SIGPLAN Notices, 42(1): 109–122, January 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Calcagno:2007:CLM

[1729] Cristiano Calcagno, Philippa Gardner, and Uri Zarfaty. Context logic as modal logic: completeness and parametric inexpressivity. ACM SIGPLAN Notices, 42(1):123–134, January 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Batory:2007:ITP

[1730] Don Batory. From implementation to theory in product synthesis. ACM SIG-PLAN Notices, 42(1):135–136, January 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lammel:2007:SYB

[1731] Ralf Lämmel. Scrap your boilerplate with XPath-like combinators. ACM SIGPLAN Notices, 42(1):137–142, January 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ohori:2007:LFF

[1732] Atsushi Ohori and Isao Sasano.

Lightweight fusion by fixed point promotion. ACM SIGPLAN Notices, 42 (1):143–154, January 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pearlmutter:2007:LMH

[1733] Barak A. Pearlmutter and Jeffrey Mark Siskind. Lazy multivariate higher-order forward-mode AD. ACM SIGPLAN Notices, 42(1):155–160, January 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stovring:2007:CCI

[1734] Kristian Støvring and Soren B. Lassen. A complete, co-inductive syntactic theory of sequential control and state. ACM SIGPLAN Notices, 42(1):161– 172, January 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lee:2007:TMM

[1735] Daniel K. Lee, Karl Crary, and Robert Harper. Towards a mechanized metatheory of standard ML. ACM SIGPLAN Notices, 42(1):173–184, January 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Might:2007:LFA

[1736] Matthew Might. Logic-flow analysis of higher-order programs. ACM SIG-PLAN Notices, 42(1):185–198, January 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wiedermann:2007:EQS

[1737] Ben Wiedermann and William R. Cook. Extracting queries by static analysis of transparent persistence. ACM SIGPLAN Notices, 42(1):199– 210, January 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berdine:2007:VAI

[1738] Josh Berdine, Aziem Chawdhary, Byron Cook, Dino Distefano, and Peter O’Hearn. Variance analyses from invariance analyses. ACM SIGPLAN Notices, 42(1):211–224, January 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Malacaria:2007:AST

[1739] Pasquale Malacaria. Assessing security threats of looping constructs. ACM SIGPLAN Notices, 42(1):225–235, January 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yu:2007:JIB

[1740] Dachuan Yu, Ajay Chander, Nayeem Islam, and Igor Serikov. JavaScript instrumentation for browser security. ACM SIGPLAN Notices, 42(1):237– 249, January 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bugliesi:2007:SIT

[1741] Michele Bugliesi and Marco Giunti. Secure implementations of typed channel abstractions. ACM SIGPLAN Notices, 42(1):251–262, January 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Murthy:2007:APL

[1742] Chet Murthy. Advanced programming language design in enterprise software: a lambda-calculus theorist wanders into a datacenter. ACM SIG-PLAN Notices, 42(1):263–264, January 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cook:2007:PPE

[1743] Byron Cook, Alexey Gotsman, Andreas Podelski, Andrey Rybalchenko, and Moshe Y. Vardi. Proving that programs eventually do something good. ACM SIGPLAN Notices, 42(1):265– 276, January 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gulwani:2007:PVP

[1744] Sumit Gulwani and Nebojsa Jojic. Program verification as probabilistic inference. ACM SIGPLAN Notices, 42(1): 277–289, January 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Emmi:2007:LA

[1745] Michael Emmi, Jeffrey S. Fischer, Ranjit Jhala, and Rupak Majumdar. Lock allocation. ACM SIGPLAN Notices, 42 (1):291–296, January 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Parkinson:2007:MVN

[1746] Matthew Parkinson, Richard Bornat, and Peter O’Hearn. Modular verification of a non-blocking stack. ACM SIG-PLAN Notices, 42(1):297–302, January 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kahlon:2007:AIP

[1747] Vineet Kahlon and Aarti Gupta. On the analysis of interacting pushdown systems. ACM SIGPLAN Notices, 42 (1):303–314, January 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reppy:2007:SCM

[1748] John Reppy and Yingqi Xiao. Specialization of CML message-passing primitives. ACM SIGPLAN Notices, 42(1): 315–326, January 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Naik:2007:CMA

[1749] Mayur Naik and Alex Aiken. Conditional must not aliasing for static race detection. ACM SIGPLAN Notices, 42(1):327–338, January 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jhala:2007:IAA

[1750] Ranjit Jhala and Rupak Majumdar. Interprocedural analysis of asynchronous programs. ACM SIGPLAN Notices, 42 (1):339–350, January 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vaswani:2007:PPP

[1751] Kapil Vaswani, Aditya V. Nori, and Trishul M. Chilimbi. Preferential path profiling: compactly numbering interesting paths. ACM SIGPLAN Notices, 42(1):351–362, January 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ghica:2007:GSS

[1752] Dan R. Ghica. Geometry of synthesis: a structured approach to VLSI design. ACM SIGPLAN Notices, 42(1): 363–375, January 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Preda:2007:SBA

[1753] Mila Dalla Preda, Mihai Christodorescu, Somesh Jha, and Saumya Debray. A semantics-based approach to malware detection. ACM SIGPLAN Notices, 42(1):377–388, January 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:2007:TPB

[1754] Hsiang-Yang Chen, Ting-Wei Hou, and Chun-Liang Lin. Tamper-proofing basis path by using oblivious hashing on Java. ACM SIGPLAN Notices, 42(2): 9–16, February 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kumar:2007:PST

[1755] Rajeev Kumar and Soham S. Chakraborty. Precise static type analysis for object oriented programs. ACM SIGPLAN Notices, 42(2):17–26, February 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tanter:2007:DSC

[1756] Éric Tanter. On dynamically-scoped crosscutting mechanisms. ACM SIG-PLAN Notices, 42(2):27–33, February 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mahoney:2007:UCS

[1757] William R. Mahoney and William L. Sousan. Using common off-the-shelf tools to implement dynamic aspects. ACM SIGPLAN Notices, 42(2):34– 41, February 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:2007:FSS

[1758] Paul Frenger. Forth & small systems report. ACM SIGPLAN Notices, 42 (3):2–5, March 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lin:2007:SEA

[1759] Chin-Yang Lin and Ting-Wei Hou. A simple and efficient algorithm for cycle collection. ACM SIGPLAN Notices, 42 (3):7–13, March 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pareja-Flores:2007:WIF

[1760] Cristóbal Pareja-Flores, Jamie Urquiza-Fuentes, and J. Ángel Velázquez-Iturbide. WinHIPE: an IDE for functional programming based on rewriting and visualization. ACM SIG-PLAN Notices, 42(3):14–23, March 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jiang:2007:AAT

[1761] Shujuan Jiang and Yuanpeng Jiang.

An analysis approach for testing exception handling programs. ACM SIG-PLAN Notices, 42(4):3–8, April 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hou:2007:AIU

[1762] Ting-Wei Hou and Fuh-Gwo Chen.

An anomaly in an interpreter using GCC source-code-level register allocation. ACM SIGPLAN Notices, 42 (4):9–13, April 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Oh:2007:PEU

[1763] Minkyoung Oh, Jiyeon Lee, Byeong-Mo Chang, Joonseon Ahn, and Kyung-Goo Doh. A programming environment for ubiquitous computing environment. ACM SIGPLAN Notices, 42(4):14–22, April 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smirnov:2007:RPA

[1764] Igor B. Smirnov. Raw pointers in application classes of C++ considered harmful. ACM SIGPLAN Notices, 42 (4):23–31, April 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shi:2007:ILL

[1765] Huihui Shi, Yi Wang, Haibing Guan, and Alei Liang. An intermediate language level optimization framework for dynamic binary translation. ACM SIG-PLAN Notices, 42(5):3–9, May 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Banerjee:2007:SID

[1766] Jayasri Banerjee, Anup Kumar Bandyopadhyay, and Ajit Kumar Mandal. Some investigations on deadlock freedom issues of a cyclically connected system using Dijkstra’s weakest precondition calculus. ACM SIGPLAN Notices, 42(5):10–15, May 2007. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Novark:2007:EAC

[1767] Gene Novark, Emery D. Berger, and Benjamin G. Zorn. Exterminator: automatically correcting memory errors with high probability. ACM SIG-PLAN Notices, 42(6):1–11, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Burckhardt:2007:CCC

[1768] Sebastian Burckhardt, Rajeev Alur, and Milo M. K. Martin. CheckFence: checking consistency of concurrent data types on relaxed memory models. ACM SIGPLAN Notices, 42(6):12–21, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Narayanasamy:2007:ACB

[1769] Satish Narayanasamy, Zhenghao Wang, Jordan Tigani, Andrew Edwards, and Brad Calder. Automatically classifying benign and harmful data races all using replay analysis. ACM SIG-PLAN Notices, 42(6):22–31, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wassermann:2007:SPA

[1770] Gary Wassermann and Zhendong Su.

Sound and precise analysis of Web applications for injection vulnerabilities. ACM SIGPLAN Notices, 42(6):32–41, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Perry:2007:FTT

[1771] Frances Perry, Lester Mackey, George A.

Reis, Jay Ligatti, David I. August, and David Walker. Fault-tolerant typed assembly language. ACM SIG-PLAN Notices, 42(6):42–53, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chlipala:2007:CTP

[1772] Adam Chlipala. A certified typepreserving compiler from lambda calculus to assembly language. ACM SIG-PLAN Notices, 42(6):54–65, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cai:2007:CSM

[1773] Hongxu Cai, Zhong Shao, and Alexander Vaynberg. Certified self-modifying code. ACM SIGPLAN Notices, 42 (6):66–77, June 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shpeisman:2007:EIO

[1774] Tatiana Shpeisman, Vijay Menon, Ali-Reza Adl-Tabatabai, Steven Balensiefer, Dan Grossman, Richard L. Hudson, Katherine F. Moore, and Bratin Saha. Enforcing isolation and ordering in STM. ACM SIGPLAN Notices, 42(6):78–88, June 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nethercote:2007:VFH

[1775] Nicholas Nethercote and Julian Seward. Valgrind: a framework for heavyweight dynamic binary instrumentation. ACM SIGPLAN Notices, 42(6): 89–100, June 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ha:2007:IER

[1776] Jungwoo Ha, Christopher J. Rossbach, Jason V. Davis, Indrajit Roy, Hany E. Ramadan, Donald E. Porter, David L. Chen, and Emmett Witchel. Improved error reporting for software that uses black-box components. ACM SIG-PLAN Notices, 42(6):101–111, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sridharan:2007:TS

[1777] Manu Sridharan, Stephen J. Fink, and Rastislav Bodik. Thin slicing. ACM SIGPLAN Notices, 42(6): 112–122, June 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ramanathan:2007:SSI

[1778] Murali Krishna Ramanathan, Ananth Grama, and Suresh Jagannathan. Static specification inference using predicate mining. ACM SIGPLAN Notices, 42(6):123–134, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Scherpelz:2007:AIO

[1779] Erika Rice Scherpelz, Sorin Lerner, and Craig Chambers. Automatic inference of optimizer flow functions from semantic meanings. ACM SIG-PLAN Notices, 42(6):135–145, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Morita:2007:AIG

[1780] Kazutaka Morita, Akimasa Morihata, Kiminori Matsuzaki, Zhenjiang Hu, and Masato Takeichi. Automatic inversion generates divide-and-conquer parallel programs. ACM SIGPLAN Notices, 42(6):146–155, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wang:2007:EAP

[1781] Perry H. Wang, Jamison D. Collins, Gautham N. Chinya, Hong Jiang, Xinmin Tian, Milind Girkar, Nick Y. Yang, Guei-Yuan Lueh, and Hong Wang. EX-OCHI: architecture and programming environment for a heterogeneous multicore multithreaded system. ACM SIG-PLAN Notices, 42(6):156–166, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Solar-Lezama:2007:SS

[1782] Armando Solar-Lezama, Gilad Arnold, Liviu Tancau, Rastislav Bodik, Vijay Saraswat, and Sanjit Seshia. Sketching stencils. ACM SIGPLAN Notices, 42 (6):167–178, June 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Killian:2007:MLS

[1783] Charles Edwin Killian, James W. Anderson, Ryan Braud, Ranjit Jhala, and Amin M. Vahdat. Mace: language support for building distributed systems. ACM SIGPLAN Notices, 42(6): 179–188, June 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Li:2007:CET

[1784] Peng Li and Steve Zdancewic. Combining events and threads for scalable network services implementation and evaluation of monadic, application-level concurrency primitives. ACM SIG-PLAN Notices, 42(6):189–199, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kothari:2007:REP

[1785] Nupur Kothari, Ramakrishna Gummadi, Todd Millstein, and Ramesh Govindan. Reliable and efficient programming abstractions for wireless sensor networks. ACM SIGPLAN Notices, 42(6):200–210, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kulkarni:2007:OPR

[1786] Milind Kulkarni, Keshav Pingali, Bruce Walter, Ganesh Ramanarayanan, Kavita Bala, and L. Paul Chew. Optimistic parallelism requires abstractions. ACM SIGPLAN Notices, 42(6): 211–222, June 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ding:2007:SBO

[1787] Chen Ding, Xipeng Shen, Kirk Kelsey, Chris Tice, Ruke Huang, and Chengliang Zhang. Software behavior oriented parallelization. ACM SIG-PLAN Notices, 42(6):223–234, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Krishnamoorthy:2007:EAP

[1788] Sriram Krishnamoorthy, Muthu Baskaran, Uday Bondhugula, J. Ramanujam, Atanas Rountev, and P. Sadayappan.

Effective automatic parallelization of stencil computations. ACM SIGPLAN Notices, 42(6):235–244, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Elmas:2007:GRT

[1789] Tayfun Elmas, Shaz Qadeer, and Serdar Tasiran. Goldilocks: a race and transaction-aware Java runtime. ACM SIGPLAN Notices, 42(6):245– 255, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Guo:2007:SAI

[1790] Bolei Guo, Neil Vachharajani, and David I. August. Shape analysis with inductive recursion synthesis. ACM SIGPLAN Notices, 42(6): 256–265, June 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gotsman:2007:TMS

[1791] Alexey Gotsman, Josh Berdine, Byron Cook, and Mooly Sagiv. Threadmodular shape analysis. ACM SIG-PLAN Notices, 42(6):266–277, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lattner:2007:MCS

[1792] Chris Lattner, Andrew Lenharth, and Vikram Adve. Making contextsensitive points-to analysis with heap cloning practical for the real world. ACM SIGPLAN Notices, 42(6):278– 289, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hardekopf:2007:AGF

[1793] Ben Hardekopf and Calvin Lin. The ant and the grasshopper: fast and accurate pointer analysis for millions of lines of code. ACM SIGPLAN Notices, 42(6): 290–299, June 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Beyer:2007:PI

[1794] Dirk Beyer, Thomas A. Henzinger, Rupak Majumdar, and Andrey Rybalchenko. Path invariants. ACM SIG-PLAN Notices, 42(6):300–309, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shankar:2007:DAI

[1795] Ajeet Shankar and Rastislav Bodík.

DITTO: automatic incrementalization of data structure invariant checks (in Java). ACM SIGPLAN Notices, 42 (6):310–319, June 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cook:2007:PTT

[1796] Byron Cook, Andreas Podelski, and Andrey Rybalchenko. Proving thread termination. ACM SIGPLAN Notices, 42(6):320–330, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kodumal:2007:RAS

[1797] John Kodumal and Alex Aiken. Regularly annotated set constraints. ACM SIGPLAN Notices, 42(6):331–341, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Geneves:2007:ESA

[1798] Pierre Genevès, Nabil Layaïda, and Alan Schmitt. Efficient static analysis of XML paths and types. ACM SIG-PLAN Notices, 42(6):342–351, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Titzer:2007:ESA

[1799] Ben L. Titzer, Joshua Auerbach, David F. Bacon, and Jens Palsberg. The ExoVM system for automatic VM and application reduction. ACM SIG-PLAN Notices, 42(6):352–362, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cooprider:2007:OCC

[1800] Nathan Dean Cooprider and John David Regehr. Offline compression for on-chip RAM. ACM SIGPLAN Notices, 42 (6):363–372, June 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schneider:2007:OOD

[1801] Florian T. Schneider, Mathias Payer, and Thomas R. Gross. Online optimizations driven by hardware performance monitoring. ACM SIG-PLAN Notices, 42(6):373–382, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Li:2007:UUC

[1802] Weijia Li, Youtao Zhang, Jun Yang, and Jiang Zheng. UCC: updateconscious compilation for energy efficiency in wireless sensor networks.

ACM SIGPLAN Notices, 42(6):383– 393, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Li:2007:PDE

[1803] Feihui Li, Guangyu Chen, Mahmut Kandemir, and Ibrahim Kolcu. Profiledriven energy reduction in network-on-chips. ACM SIGPLAN Notices, 42(6): 394–404, June 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Renganarayanan:2007:PTL

[1804] Lakshminarayanan Renganarayanan, DaeGon Kim, Sanjay Rajopadhye, and Michelle Mills Strout. Parameterized tiled loops for free. ACM SIG-PLAN Notices, 42(6):405–414, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:2007:TLE

[1805] Xiangyu Zhang, Sriraman Tallam, Neelam Gupta, and Rajiv Gupta. Towards locating execution omission errors. ACM SIGPLAN Notices, 42(6): 415–424, June 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lerner:2007:STE

[1806] Benjamin S. Lerner, Matthew Flower, Dan Grossman, and Craig Chambers. Searching for type-error messages. ACM SIGPLAN Notices, 42 (6):425–434, June 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dillig:2007:SED

[1807] Isil Dillig, Thomas Dillig, and Alex Aiken. Static error detection using semantic inconsistency inference. ACM SIGPLAN Notices, 42(6):435– 445, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Musuvathi:2007:ICB

[1808] Madanlal Musuvathi and Shaz Qadeer. Iterative context bounding for systematic testing of multithreaded programs. ACM SIGPLAN Notices, 42 (6):446–455, June 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vechev:2007:CSA

[1809] Martin T. Vechev, Eran Yahav, David F. Bacon, and Noam Rinetzky. CGCExplorer: a semi-automated search procedure for provably correct concurrent collectors. ACM SIGPLAN Notices, 42(6):456–467, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McCreight:2007:GFC

[1810] Andrew McCreight, Zhong Shao, Chunxiao Lin, and Long Li. A general framework for certifying garbage collectors and their mutators. ACM SIG-PLAN Notices, 42(6):468–479, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cherem:2007:PML

[1811] Sigmund Cherem, Lonnie Princehouse, and Radu Rugina. Practical memory leak detection using guarded value-flow analysis. ACM SIGPLAN Notices, 42 (6):480–491, June 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Permandla:2007:TSP

[1812] Pratibha Permandla, Michael Roberson, and Chandrasekhar Boyapati. A type system for preventing data races and deadlocks in the Java Virtual Machine language: 1. ACM SIG-PLAN Notices, 42(7):10, July 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Huang:2007:AFL

[1813] Tai-Yi Huang, Pin-Chuan Chou, Cheng-Han Tsai, and Hsin-An Chen. Automated fault localization with statistically suspicious program states. ACM SIGPLAN Notices, 42(7):11–20, July 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhong:2007:FAE

[1814] Xiliang Zhong and Cheng-Zhong Xu.

Frequency-aware energy optimization for real-time periodic and aperiodic tasks. ACM SIGPLAN Notices, 42 (7):21–30, July 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhu:2007:DCL

[1815] Yifan Zhu and Frank Mueller. DVSleak:

combining leakage reduction and voltage scaling in feedback EDF scheduling. ACM SIGPLAN Notices, 42 (7):31–40, July 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

AbouGhazaleh:2007:ICL

[1816] Nevine AbouGhazaleh, Alexandre Ferreira, Cosmin Rusu, Ruibin Xu, Frank Liberato, Bruce Childers, Daniel Mosse, and Rami Melhem. Integrated CPU and l2 cache voltage scaling using machine learning. ACM SIG-PLAN Notices, 42(7):41–50, July 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Auerbach:2007:JTF

[1817] Joshua Auerbach, David F. Bacon, Daniel T. Iercan, Christoph M. Kirsch, V. T. Rajan, Harald Roeck, and Rainer Trummer. Java takes flight: timeportable real-time programming with exotasks. ACM SIGPLAN Notices, 42 (7):51–62, July 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hong:2007:JCA

[1818] SungHyun Hong, Jin-Chul Kim, Jin Woo Shin, Soo-Mook Moon, Hyeong-Seok Oh, Jaemok Lee, and Hyung-Kyu Choi. Java client ahead-of-time compiler for embedded systems. ACM SIGPLAN Notices, 42(7):63–72, July 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Alias:2007:BIL

[1819] Christophe Alias, Fabrice Baray, and Alain Darte. Bee+Cl@k: an implementation of lattice-based array contraction in the source-to-source translator rose. ACM SIGPLAN Notices, 42(7):73–82, July 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bennett:2007:CSS

[1820] Richard Vincent Bennett, Alastair Colin Murray, Björn Franke, and Nigel Topham. Combining source-to-source transformations and processor instruction set extensions for the automated design-space exploration of embedded systems. ACM SIGPLAN Notices, 42(7):83–92, July 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ahn:2007:OCH

[1821] Minwook Ahn, Jooyeon Lee, and Yunheung Paek. Optimistic coalescing for heterogeneous register architectures. ACM SIGPLAN Notices, 42(7):93–102, July 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bouchez:2007:CSE

[1822] Florent Bouchez, Alain Darte, and Fabrice Rastello. On the complexity of spill everywhere under SSA form. ACM SIGPLAN Notices, 42(7):103–112, July 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xu:2007:TNR

[1823] Weifeng Xu and Russell Tessier. Tetris:

a new register pressure control technique for VLIW processors. ACM SIGPLAN Notices, 42(7):113–122, July 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pizlo:2007:HRT

[1824] Filip Pizlo, Antony L. Hosking, and Jan Vitek. Hierarchical real-time garbage collection. ACM SIGPLAN Notices, 42(7):123–133, July 2007. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

OBrien:2007:ICC

[1825] Kathryn O’Brien. Issues and challenges in compiling for the CBEA. ACM SIG-PLAN Notices, 42(7):134, July 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kaestner:2007:SWC

[1826] Daniel Kaestner. Safe worst-case execution time analysis by abstract interpretation of executable code. ACM SIGPLAN Notices, 42(7):135, July 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Engelsma:2007:ESM

[1827] Jonathan Engelsma. Enabling seamless mobility: an enablers, experiences and tools perspective. ACM SIG-PLAN Notices, 42(7):136, July 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Huang:2007:JTE

[1828] Po-Kuan Huang, Matin Hashemi, and Soheil Ghiasi. Joint throughput and energy optimization for pipelined execution of embedded streaming applications. ACM SIGPLAN Notices, 42 (7):137–139, July 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cho:2007:ISH

[1829] Hansu Cho, Samar Abdi, and Daniel Gajski. Interface synthesis for heterogeneous multi-core systems from transaction level models. ACM SIG-PLAN Notices, 42(7):140–142, July 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Viskic:2007:AGE

[1830] Ines Viskic, Samar Abdi, and Daniel D.

Gajski. Automatic generation of embedded communication SW for heterogeneous MPSoC platforms. ACM SIGPLAN Notices, 42(7):143–145, July 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:2007:ECF

[1831] Chung-Kai Chen, Ling-Hua Tseng, Shih-Chang Chen, Young-Jia Lin, Yi-Ping You, Chia-Han Lu, and Jenq-Kuen Lee. Enabling compiler flow for embedded VLIW DSP processors with distributed register files. ACM SIGPLAN Notices, 42(7):146–148, July 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Walther:2007:ART

[1832] Karsten Walther, René Herzog, and Jörg Nolte. Analyzing the real-time behaviour of deeply embedded event driven systems. ACM SIGPLAN Notices, 42(7):149–151, July 2007. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Coffman:2007:GPT

[1833] Joel Coffman, Christopher Healy, Frank Mueller, and David Whalley. Generalizing parametric timing analysis. ACM SIGPLAN Notices, 42(7): 152–154, July 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:2007:CDA

[1834] Guangyu Chen, Feihui Li, and Mahmut Kandemir. Compiler-directed application mapping for NoC based chip multiprocessors. ACM SIGPLAN Notices, 42(7):155–157, July 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yan:2007:SEW

[1835] Shan Yan and Bill Lin. Stream execution on wide-issue clustered VLIW architectures. ACM SIGPLAN Notices, 42(7):158–160, July 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chu:2007:CDP

[1836] Michael L. Chu and Scott A. Mahlke. Code and data partitioning for finegrain parallelism. ACM SIGPLAN Notices, 42(7):161–164, July 2007. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hines:2007:AIF

[1837] Stephen Roderick Hines, Gary Tyson, and David Whalley. Addressing instruction fetch bottlenecks by using an instruction register file. ACM SIG-PLAN Notices, 42(7):165–174, July 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yan:2007:WAI

[1838] Jun Yan and Wei Zhang. WCET analysis of instruction caches with prefetching. ACM SIGPLAN Notices, 42(7): 175–184, July 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ning:2007:EMP

[1839] Ke Ning and David Kaeli. External memory page remapping for embedded multimedia systems. ACM SIGPLAN Notices, 42(7):185–194, July 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cho:2007:DDS

[1840] Hyungmin Cho, Bernhard Egger, Jaejin Lee, and Heonshik Shin. Dynamic data scratchpad memory management for a memory subsystem with an MMU. ACM SIGPLAN Notices, 42 (7):195–206, July 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Li:2007:SAD

[1841] Lian Li, Quan Hoang Nguyen, and Jingling Xue. Scratchpad allocation for data aggregates in superperfect graphs. ACM SIGPLAN Notices, 42 (7):207–216, July 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

In:2007:SSW

[1842] Jihyun In, Ilhoon Shin, and Hyojun Kim. SWL: a search-while-load demand paging scheme with NAND flash memory. ACM SIGPLAN Notices, 42 (7):217–226, July 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wang:2007:OSC

[1843] Qin Wang, Junpu Chen, Weihua Zhang, Min Yang, and Binyu Zang. Optimizing software cache performance of packet processing applications. ACM SIGPLAN Notices, 42(7):227–236, July 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ravindran:2007:CMP

[1844] Rajiv Ravindran, Michael Chu, and Scott Mahlke. Compiler-managed partitioned data caches for low power. ACM SIGPLAN Notices, 42(7):237– 247, July 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dunlavey:2007:PTI

[1845] Michael Dunlavey. Performance tuning with instruction-level cost derived from call-stack sampling. ACM SIGPLAN Notices, 42(8):4–8, August 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gabay:2007:CJR

[1846] Yarom Gabay and Assaf J. Kfoury. A calculus for Java’s reference objects. ACM SIGPLAN Notices, 42(8):9–17, August 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sewell:2007:OET

[1847] Peter Sewell, Francesco Zappa Nardelli, Scott Owens, Gilles Peskine, Thomas Ridge, Susmit Sarkar, and Rok Strniša. Ott: effective tool support for the working semanticist. ACM SIGPLAN Notices, 42(9):1–12, September 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sozeau:2007:PIF

[1848] Matthieu Sozeau. Programming finger trees in Coq. ACM SIGPLAN Notices, 42(9):13–24, September 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wazny:2007:ERF

[1849] Jeremy Wazny. Experience report:

functional programming in c-rules. ACM SIGPLAN Notices, 42(9):25– 28, September 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Syme:2007:EPM

[1850] Don Syme, Gregory Neverov, and James Margetson. Extensible pattern matching via a lightweight language extension. ACM SIGPLAN Notices, 42(9):29–40, September 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Danvy:2007:BSC

[1851] Olivier Danvy and Michael Spivey. On Barron and Strachey’s Cartesian product function. ACM SIGPLAN Notices, 42(9):41–46, September 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Matsuda:2007:BTB

[1852] Kazutaka Matsuda, Zhenjiang Hu, Keisuke Nakano, Makoto Hamana, and Masato Takeichi. Bidirectionalization transformation based on automatic derivation of view complement functions. ACM SIGPLAN Notices, 42(9): 47–58, September 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Elliott:2007:TFP

[1853] Conal M. Elliott. Tangible functional programming. ACM SIGPLAN Notices, 42(9):59–70, September 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sereni:2007:TAC

[1854] Damien Sereni. Termination analysis and call graph construction for higherorder functional programs. ACM SIG-PLAN Notices, 42(9):71–84, September 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

VanHorn:2007:RCP

[1855] David Van Horn and Harry G. Mairson. Relating complexity and precision in control flow analysis. ACM SIG-PLAN Notices, 42(9):85–96, September 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Filinski:2007:IRA

[1856] Andrzej Filinski and Kristian Støvring.

Inductive reasoning about effectful data types. ACM SIGPLAN Notices, 42(9):97–110, September 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Leijen:2007:TDT

[1857] Daan Leijen. A type directed translation of MLF to system F. ACM SIG-PLAN Notices, 42(9):111–122, September 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lloyd:2007:DPA

[1858] John W. Lloyd. Declarative programming for artificial intelligence applications. ACM SIGPLAN Notices, 42(9): 123–124, September 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fredlund:2007:MMC

[1859] Lars-Åke Fredlund and Hans Svensson. McErlang: a model checker for a distributed functional programming language. ACM SIGPLAN Notices, 42(9):125–136, September 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sims:2007:ERR

[1860] Steve Sims and Daniel C. DuVarney.

Experience report: the reactis validation tool. ACM SIGPLAN Notices, 42(9):137–140, September 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Plasmeijer:2007:IES

[1861] Rinus Plasmeijer, Peter Achten, and Pieter Koopman. iTasks: executable specifications of interactive work flow systems for the web. ACM SIG-PLAN Notices, 42(9):141–152, September 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Welsh:2007:ERS

[1862] Noel Welsh and David Gurnell. Experience report: scheme in commercial Web application development. ACM SIGPLAN Notices, 42(9):153– 156, September 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Herman:2007:FPG

[1863] David Herman. Functional pearl: the great escape or, how to jump the border without getting caught. ACM SIG-PLAN Notices, 42(9):157–164, September 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Flatt:2007:ADC

[1864] Matthew Flatt, Gang Yu, Robert Bruce Findler, and Matthias Felleisen. Adding delimited and composable control to a production programming environment. ACM SIGPLAN Notices, 42 (9):165–176, September 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kennedy:2007:CCC

[1865] Andrew Kennedy. Compiling with continuations, continued. ACM SIG-PLAN Notices, 42(9):177–190, September 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Park:2007:TSH

[1866] Sungwoo Park. Type-safe higherorder channels in ML-like languages. ACM SIGPLAN Notices, 42(9):191– 202, September 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nystrom:2007:EHL

[1867] Jan Nyström, Phil Trinder, and David King. Evaluating high-level distributed language constructs. ACM SIG-PLAN Notices, 42(9):203–212, September 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Beshers:2007:ERU

[1868] Clifford Beshers, David Fox, and Jeremy Shaw. Experience report: using functional programming to manage a Linux distribution. ACM SIG-PLAN Notices, 42(9):213–218, September 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pfenning:2007:SIT

[1869] Frank Pfenning. Subtyping and intersection types revisited. ACM SIG-PLAN Notices, 42(9):219, September 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenzel:2007:ERB

[1870] Leif Frenzel. Experience report: building an eclipse-based IDE for Haskell. ACM SIGPLAN Notices, 42(9):220– 222, September 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ennals:2007:UFF

[1871] Rob Ennals and David Gay. Userfriendly functional programming for Web mashups. ACM SIGPLAN Notices, 42(9):223–234, September 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hage:2007:GUA

[1872] Jurriaan Hage, Stefan Holdermans, and Arie Middelkoop. A generic usage analysis with subeffect qualifiers. ACM SIGPLAN Notices, 42(9):235– 246, September 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cooper:2007:ERH

[1873] Jonathan Cooper and Steve McKeever. Experience report: a Haskell interpreter for cellML. ACM SIGPLAN Notices, 42(9):247–250, September 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Harris:2007:FDI

[1874] Tim Harris and Satnam Singh. Feedback directed implicit parallelism. ACM SIGPLAN Notices, 42(9):251– 264, September 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brassel:2007:LCV

[1875] Bernd Braßel, Michael Hanus, Sebastian Fischer, Frank Huch, and Germán Vidal. Lazy call-by-value evaluation. ACM SIGPLAN Notices, 42(9):265– 276, September 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marlow:2007:FLU

[1876] Simon Marlow, Alexey Rodriguez Yakushev, and Simon Peyton Jones. Faster laziness using dynamic pointer tagging. ACM SIGPLAN Notices, 42 (9):277–288, September 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dreyer:2007:TSR

[1877] Derek Dreyer. A type system for recursive modules. ACM SIGPLAN Notices, 42(9):289–302, September 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ghuloum:2007:IPR

[1878] Abdulaziz Ghuloum and R. Kent Dybvig. Implicit phasing for R6RS libraries. ACM SIGPLAN Notices, 42 (9):303–314, September 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Coutts:2007:SFL

[1879] Duncan Coutts, Roman Leshchinskiy, and Don Stewart. Stream fusion: from lists to streams to nothing at all. ACM SIGPLAN Notices, 42(9):315– 326, September 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jones:2007:CPS

[1880] Simon Peyton Jones. Call-pattern specialisation for Haskell programs. ACM SIGPLAN Notices, 42(9):327– 337, September 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ekman:2007:JEJ

[1881] Torbjörn Ekman and Görel Hedin. The JastAdd extensible Java compiler. ACM SIGPLAN Notices, 42(10): 1–18, October 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hirzel:2007:JGJ

[1882] Martin Hirzel and Robert Grimm. Jeannie: granting Java Native Interface developers their wishes. ACM SIG-PLAN Notices, 42(10):19–38, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tan:2007:IIL

[1883] Gang Tan and Greg Morrisett. Ilea: inter-language analysis across Java and C. ACM SIGPLAN Notices, 42(10): 39–56, October 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Georges:2007:SRJ

[1884] Andy Georges, Dries Buytaert, and Lieven Eeckhout. Statistically rigorous Java performance evaluation. ACM SIGPLAN Notices, 42(10):57–76, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xian:2007:MAP

[1885] Feng Xian, Witawas Srisa-an, and Hong Jiang. Microphase: an approach to proactively invoking garbage collection for improved performance. ACM SIGPLAN Notices, 42(10):77–96, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bond:2007:PCC

[1886] Michael D. Bond and Kathryn S.

McKinley. Probabilistic calling context. ACM SIGPLAN Notices, 42(10): 97–112, October 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Igarashi:2007:VPT

[1887] Atsushi Igarashi and Mirko Viroli.

Variant path types for scalable extensibility. ACM SIGPLAN Notices, 42(10): 113–132, October 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gasiunas:2007:DC

[1888] Vaidas Gasiunas, Mira Mezini, and Klaus Ostermann. Dependent classes. ACM SIGPLAN Notices, 42(10):133– 152, October 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sasitorn:2007:CNS

[1889] James Sasitorn and Robert Cartwright.

Component NEXTGEN: a sound and expressive component framework for Java. ACM SIGPLAN Notices, 42(10):

153–170, October 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ducasse:2007:UCV

[1890] Stéphane Ducasse, Roel Wuyts, Alexandre Bergel, and Oscar Nierstrasz. User-changeable visibility: resolving unanticipated name clashes in traits. ACM SIGPLAN Notices, 42 (10):171–190, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smaragdakis:2007:TIC

[1891] Yannis Smaragdakis, Anthony Kay, Reimer Behrends, and Michal Young. Transactions with isolation and cooperation. ACM SIGPLAN Notices, 42(10): 191–210, October 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Spring:2007:SHT

[1892] Jesper H. Spring, Jean Privat, Rachid Guerraoui, and Jan Vitek. Stream-Flex: high-throughput stream programming in Java. ACM SIGPLAN Notices, 42(10):211–228, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kuehne:2007:CPL

[1893] Thomas Kuehne and Daniel Schreiber. Can programming be liberated from the two-level style: multi-level programming with DeepJava. ACM SIG-PLAN Notices, 42(10):229–244, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mitchell:2007:CBL

[1894] Nick Mitchell and Gary Sevitsky. The causes of bloat, the limits of health. ACM SIGPLAN Notices, 42(10):245– 260, October 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dekel:2007:NRC

[1895] Uri Dekel and James D. Herbsleb. Notation and representation in collaborative object-oriented design: an observational study. ACM SIGPLAN Notices, 42(10):261–280, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Leff:2007:WEV

[1896] Avraham Leff and James T. Rayfield.

WebRB: evaluating a visual domainspecific language for building relational web-applications. ACM SIGPLAN Notices, 42(10):281–300, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bierhoff:2007:MTC

[1897] Kevin Bierhoff and Jonathan Aldrich.

Modular typestate checking of aliased objects. ACM SIGPLAN Notices, 42 (10):301–320, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Greenfieldboyce:2007:TQI

[1898] David Greenfieldboyce and Jeffrey S.

Foster. Type qualifier inference for Java. ACM SIGPLAN Notices, 42(10):

321–336, October 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fahndrich:2007:EOI

[1899] Manuel Fahndrich and Songtao Xia. Establishing object invariants with delayed types. ACM SIGPLAN Notices, 42(10):337–350, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shaner:2007:MVH

[1900] Steve M. Shaner, Gary T. Leavens, and David A. Naumann. Modular verification of higher-order methods with mandatory calls specified by model programs. ACM SIGPLAN Notices, 42(10):351–368, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rinard:2007:UEP

[1901] Martin C. Rinard. Using early phase termination to eliminate load imbalances at barrier synchronization points. ACM SIGPLAN Notices, 42(10):369– 386, October 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Elkarablieh:2007:SSA

[1902] Bassem Elkarablieh, Sarfraz Khurshid, Duy Vu, and Kathryn S. McKinley. STARC: static analysis for efficient repair of complex data. ACM SIGPLAN Notices, 42(10):387–404, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bond:2007:TBA

[1903] Michael D. Bond, Nicholas Nethercote, Stephen W. Kent, Samuel Z. Guyer, and Kathryn S. McKinley. Tracking bad apples: reporting the origin of null and undefined value errors. ACM SIG-PLAN Notices, 42(10):405–422, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ma:2007:IAE

[1904] Kin-Keung Ma and Jeffrey S. Foster. Inferring aliasing and encapsulation properties for Java. ACM SIG-PLAN Notices, 42(10):423–440, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cameron:2007:MO

[1905] Nicholas R. Cameron, Sophia Drossopoulou, James Noble, and Matthew J. Smith. Multiple ownership. ACM SIGPLAN Notices, 42(10):441–460, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Muller:2007:OTU

[1906] Peter Müller and Arsenii Rudich.

Ownership transfer in universe types. ACM SIGPLAN Notices, 42(10):461– 478, October 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bierman:2007:LTF

[1907] Gavin M. Bierman, Erik Meijer, and Mads Torgersen. Lost in translation: formalizing proposed extensions to C#.

ACM SIGPLAN Notices, 42(10):479– 498, October 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Strnisa:2007:JMS

[1908] Rok Strniša, Peter Sewell, and Matthew Parkinson. The Java module system: core design and semantic definition. ACM SIGPLAN Notices, 42(10):499–514, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kojarski:2007:AAC

[1909] Sergei Kojarski and David H. Lorenz. Awesome: an aspect co-weaving system for composing multiple aspectoriented extensions. ACM SIGPLAN Notices, 42(10):515–534, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pothier:2007:SOD

[1910] Guillaume Pothier, Éric Tanter, and José Piquer. Scalable omniscient debugging. ACM SIGPLAN Notices, 42 (10):535–552, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Buytaert:2007:UHS

[1911] Dries Buytaert, Andy Georges, Michael Hind, Matthew Arnold, Lieven Eeckhout, and Koen De Bosschere. Using HPM-sampling to drive dynamic compilation. ACM SIGPLAN Notices, 42(10):553–568, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:2007:MEG

[1912] Feng Chen and Grigore Roşu. MOP:

an efficient and generic runtime verification framework. ACM SIGPLAN Notices, 42(10):569–588, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Avgustinov:2007:MTM

[1913] Pavel Avgustinov, Julian Tibble, and Oege de Moor. Making trace monitors feasible. ACM SIGPLAN Notices, 42(10):589–608, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rinard:2007:LCZ

[1914] Martin C. Rinard. Living in the comfort zone. ACM SIGPLAN Notices, 42(10):611–622, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McDirmid:2007:LIL

[1915] Sean McDirmid. Living it up with a live programming language. ACM SIG-PLAN Notices, 42(10):623–638, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Edwards:2007:NIA

[1916] Jonathan Edwards. No ifs, ands, or buts: uncovering the simplicity of conditionals. ACM SIGPLAN Notices, 42(10):639–658, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fleissner:2007:EAA

[1917] Sebastian Fleissner and Elisa L. A. Baniassad. Epi-aspects: aspect-oriented conscientious software. ACM SIG-PLAN Notices, 42(10):659–674, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meijer:2007:CUP

[1918] Erik Meijer. Confessions of a used programming language salesman. ACM SIGPLAN Notices, 42(10):677–694, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Grossman:2007:TMG

[1919] Dan Grossman. The transactional memory /garbage collection analogy. ACM SIGPLAN Notices, 42(10):695– 706, October 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:2007:HYM

[1920] Paul Frenger. Hacking your mouse: errata for Mar. 2007. ACM SIG-PLAN Notices, 42(11):3, November 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sarma:2007:SSS

[1921] M. Sarma and R. Mall. Synthesis of system state models. ACM SIGPLAN Notices, 42(11):5–14, November 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Svensson:2007:FXD

[1922] Kurt Svensson. Faster XML data validation in a programming language with XML datatypes. ACM SIGPLAN Notices, 42(11):15–21, November 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xia:2007:FSE

[1923] Feng Xia, Guosong Tian, and Youxian Sun. Feedback scheduling: an event-driven paradigm. ACM SIG-PLAN Notices, 42(12):7–14, December 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Syropoulos:2007:PMV

[1924] Apostolos Syropoulos. Π machines:

virtual machines realizing graph structured transition P systems. ACM SIG-PLAN Notices, 42(12):15–22, December 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fontana:2008:SBM

[1925] Walter Fontana. Systems biology, models, and concurrency. ACM SIG-PLAN Notices, 43(1):1–2, January 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Aydemir:2008:EFM

[1926] Brian Aydemir, Arthur Charguéraud, Benjamin C. Pierce, Randy Pollack, and Stephanie Weirich. Engineering formal metatheory. ACM SIG-PLAN Notices, 43(1):3–15, January 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tristan:2008:FVT

[1927] Jean-Baptiste Tristan and Xavier Leroy. Formal verification of translation validators: a case study on instruction scheduling optimizations. ACM SIGPLAN Notices, 43(1):17–27, January 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Voigtlander:2008:MAA

[1928] Janis Voigtländer. Much ado about two (pearl): a pearl on parallel prefix computation. ACM SIGPLAN Notices, 43 (1):29–35, January 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Neamtiu:2008:CEV

[1929] Iulian Neamtiu, Michael Hicks, Jeffrey S. Foster, and Polyvios Pratikakis. Contextual effects for version-consistent dynamic software updating all and safe concurrent programming. ACM SIG-PLAN Notices, 43(1):37–49, January 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Moore:2008:HLS

[1930] Katherine F. Moore and Dan Grossman. High-level small-step operational semantics for transactions. ACM SIG-PLAN Notices, 43(1):51–62, January 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Abadi:2008:STM

[1931] Martín Abadi, Andrew Birrell, Tim Harris, and Michael Isard. Semantics of transactional memory and automatic mutual exclusion. ACM SIG-PLAN Notices, 43(1):63–74, January 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Parkinson:2008:SLA

[1932] Matthew J. Parkinson and Gavin M.

Bierman. Separation logic, abstraction and inheritance. ACM SIGPLAN Notices, 43(1):75–86, January 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chin:2008:EMO

[1933] Wei-Ngan Chin, Cristina David, Huu Hai Nguyen, and Shengchao Qin. Enhancing modular OO verification with separation logic. ACM SIGPLAN Notices, 43(1):87–99, January 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brotherston:2008:CPP

[1934] James Brotherston, Richard Bornat, and Cristiano Calcagno. Cyclic proofs of program termination in separation logic. ACM SIGPLAN Notices, 43(1): 101–112, January 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Asperti:2008:ICR

[1935] Andrea Asperti. The intensional content of Rice’s theorem. ACM SIG-PLAN Notices, 43(1):113–119, January 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gaboardi:2008:LAP

[1936] Marco Gaboardi, Jean-Yves Marion, and Simona Ronchi Della Rocca. A logical account of pspace. ACM SIG-PLAN Notices, 43(1):121–131, January 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Danielsson:2008:LST

[1937] Nils Anders Danielsson. Lightweight semiformal time complexity analysis for purely functional data structures. ACM SIGPLAN Notices, 43(1):133– 144, January 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McMillan:2008:RHP

[1938] K. L. McMillan. Relevance heuristics for program analysis. ACM SIGPLAN Notices, 43(1):145–146, January 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gupta:2008:PNT

[1939] Ashutosh Gupta, Thomas A. Henzinger, Rupak Majumdar, Andrey Rybalchenko, and Ru-Gang Xu. Proving non-termination. ACM SIGPLAN Notices, 43(1):147–158, January 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chaudhuri:2008:SAR

[1940] Swarat Chaudhuri. Subcubic algorithms for recursive state machines.

ACM SIGPLAN Notices, 43(1):159– 169, January 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lahiri:2008:BFR

[1941] Shuvendu Lahiri and Shaz Qadeer.

Back to the future: revisiting precise program verification using SMT solvers. ACM SIGPLAN Notices, 43 (1):171–182, January 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Unkel:2008:AIS

[1942] Christopher Unkel and Monica S.

Lam. Automatic inference of stationary fields: a generalization of Java’s final fields. ACM SIGPLAN Notices, 43(1): 183–195, January 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zheng:2008:DDA

[1943] Xin Zheng and Radu Rugina. Demanddriven alias analysis for C. ACM SIG-PLAN Notices, 43(1):197–208, January 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nita:2008:TPD

[1944] Marius Nita, Dan Grossman, and Craig Chambers. A theory of platformdependent low-level software. ACM SIGPLAN Notices, 43(1):209–220, January 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yorsh:2008:GPC

[1945] Greta Yorsh, Eran Yahav, and Satish Chandra. Generating precise and concise procedure summaries. ACM SIG-PLAN Notices, 43(1):221–234, January 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gulwani:2008:LAI

[1946] Sumit Gulwani, Bill McCloskey, and Ashish Tiwari. Lifting abstract interpreters to quantified logical domains. ACM SIGPLAN Notices, 43(1):235– 246, January 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chang:2008:RIS

[1947] Bor-Yuh Evan Chang and Xavier Rival. Relational inductive shape analysis. ACM SIGPLAN Notices, 43(1): 247–260, January 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Castagna:2008:TCW

[1948] Giuseppe Castagna, Nils Gesbert, and Luca Padovani. A theory of contracts for Web services. ACM SIGPLAN Notices, 43(1):261–272, January 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Honda:2008:MAS

[1949] Kohei Honda, Nobuko Yoshida, and Marco Carbone. Multiparty asynchronous session types. ACM SIG-PLAN Notices, 43(1):273–284, January 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Minsky:2008:CT

[1950] Yaron M. Minsky. Caml trading. ACM SIGPLAN Notices, 43(1):285, January 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McBride:2008:CLM

[1951] Conor McBride. Clowns to the left of me, jokers to the right (pearl): dissecting data structures. ACM SIG-PLAN Notices, 43(1):287–295, January 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Johann:2008:FSP

[1952] Patricia Johann and Neil Ghani. Foundations for structured programming with GADTs. ACM SIGPLAN Notices, 43(1):297–308, January 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Acar:2008:ISA

[1953] Umut A. Acar, Amal Ahmed, and Matthias Blume. Imperative selfadjusting computation. ACM SIG-PLAN Notices, 43(1):309–322, January 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fournet:2008:CSI

[1954] Cédric Fournet and Tamara Rezk.

Cryptographically sound implementations for typed information-flow security. ACM SIGPLAN Notices, 43(1): 323–335, January 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Laud:2008:CSC

[1955] Peeter Laud. On the computational soundness of cryptographically masked flows. ACM SIGPLAN Notices, 43(1): 337–348, January 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Alavi:2008:EET

[1956] Hamed Seiied Alavi, Seth Gilbert, and Rachid Guerraoui. Extensible encoding of type hierarchies. ACM SIG-PLAN Notices, 43(1):349–358, January 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zeilberger:2008:FHO

[1957] Noam Zeilberger. Focusing and higherorder abstract syntax. ACM SIGPLAN Notices, 43(1):359–369, January 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pientka:2008:TTF

[1958] Brigitte Pientka. A type-theoretic foundation for programming with higher-order abstract syntax and firstclass substitutions. ACM SIGPLAN Notices, 43(1):371–382, January 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Herbelin:2008:ACN

[1959] Hugo Herbelin and Silvia Ghilezan. An approach to call-by-name delimited continuations. ACM SIGPLAN Notices, 43(1):383–394, January 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tobin-Hochstadt:2008:DIT

[1960] Sam Tobin-Hochstadt and Matthias Felleisen. The design and implementation of typed scheme. ACM SIG-PLAN Notices, 43(1):395–406, January 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bohannon:2008:BRL

[1961] Aaron Bohannon, J. Nathan Foster, Benjamin C. Pierce, Alexandre Pilkiewicz, and Alan Schmitt. Boomerang: resourceful lenses for string data. ACM SIGPLAN Notices, 43(1):407–419, January 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fisher:2008:DSF

[1962] Kathleen Fisher, David Walker, Kenny Q. Zhu, and Peter White. From dirt to shovels: fully automatic tool generation from ad hoc data. ACM SIGPLAN Notices, 43(1):421–434, January 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fitzgerald:2008:VAS

[1963] John Fitzgerald, Peter Gorm Larsen, and Shin Sahara. VDMTools: advances in support for formal modeling in VDM. ACM SIGPLAN Notices, 43 (2):3–11, February 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lu:2008:CIE

[1964] Jia Lu, Jing Ying, Ming hu Wu, and Cang Hong Jin. Crosscutting invariant and an efficient checking algorithm using program slicing. ACM SIGPLAN Notices, 43(2):12–20, February 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Qian:2008:PSU

[1965] Ju Qian and Baowen Xu. Program slicing under UML scenario models. ACM SIGPLAN Notices, 43(2):21–24, February 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:2008:ASA

[1966] Danfeng Zhang, Yao Guo, and Xiangqun Chen. AspectC2C: a symmetric aspect extension to the C language. ACM SIGPLAN Notices, 43(2): 25–32, February 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Winfree:2008:TMP

[1967] Erik Winfree. Toward molecular programming with DNA. ACM SIG-PLAN Notices, 43(3):1, March 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:2008:OVB

[1968] Xiaoxin Chen, Tal Garfinkel, E. Christopher Lewis, Pratap Subrahmanyam, Carl A. Waldspurger, Dan Boneh, Jeffrey Dwoskin, and Dan R. K. Ports. Overshadow: a virtualization-based approach to retrofitting protection in commodity operating systems. ACM SIGPLAN Notices, 43(3):2–13, March 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McCune:2008:HLC

[1969] Jonathan M. McCune, Bryan Parno, Adrian Perrig, Michael K. Reiter, and Arvind Seshadri. How low can you go?: recommendations for hardwaresupported minimal TCB code execution. ACM SIGPLAN Notices, 43(3): 14–25, March 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bhargava:2008:ATD

[1970] Ravi Bhargava, Benjamin Serebrin, Francesco Spadini, and Srilatha Manne. Accelerating two-dimensional page walks for virtualized systems. ACM SIGPLAN Notices, 43(3):26–35, March 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lee:2008:ETL

[1971] Benjamin C. Lee and David Brooks. Efficiency trends and limits from comprehensive microarchitectural adaptivity. ACM SIGPLAN Notices, 43(3):36–47, March 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Raghavendra:2008:NPS

[1972] Ramya Raghavendra, Parthasarathy Ranganathan, Vanish Talwar, Zhikui Wang, and Xiaoyun Zhu. No ’power’ struggles: coordinated multi-level power management for the data center. ACM SIGPLAN Notices, 43(3):48–59, March 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ballapuram:2008:EAS

[1973] Chinnakrishnan S. Ballapuram, Ahmad Sharif, and Hsien-Hsin S. Lee. Exploiting access semantics and program behavior to reduce snoop power in chip multiprocessors. ACM SIG-PLAN Notices, 43(3):60–69, March 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mallik:2008:PMU

[1974] Arindam Mallik, Jack Cosgrove, Robert P. Dick, Gokhan Memik, and Peter Dinda. PICSEL: measuring userperceived performance to control dynamic frequency scaling. ACM SIG-PLAN Notices, 43(3):70–79, March 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Joao:2008:IPO

[1975] Jose A. Joao, Onur Mutlu, Hyesoon Kim, Rishi Agarwal, and Yale N. Patt. Improving the performance of objectoriented languages with dynamic predication of indirect jumps. ACM SIG-PLAN Notices, 43(3):80–90, March 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wegiel:2008:MCV

[1976] Michal Wegiel and Chandra Krintz. The mapping collector: virtual memory support for generational, parallel, and concurrent compaction. ACM SIG-PLAN Notices, 43(3):91–102, March 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Devietti:2008:HAS

[1977] Joe Devietti, Colin Blundell, Milo M. K. Martin, and Steve Zdancewic. Hardbound: architectural support for spatial safety of the C programming language. ACM SIGPLAN Notices, 43(3):103–114, March 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lvin:2008:ATA

[1978] Vitaliy B. Lvin, Gene Novark, Emery D. Berger, and Benjamin G. Zorn. Archipelago: trading address space for reliability and security. ACM SIGPLAN Notices, 43(3): 115–124, March 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Choi:2008:ABP

[1979] Bumyong Choi, Leo Porter, and Dean M. Tullsen. Accurate branch prediction for short threads. ACM SIG-PLAN Notices, 43(3):125–134, March 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Srikantaiah:2008:ASP

[1980] Shekhar Srikantaiah, Mahmut Kandemir, and Mary Jane Irwin. Adaptive set pinning: managing shared caches in chip multiprocessors. ACM SIG-PLAN Notices, 43(3):135–144, March 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tuck:2008:SSE

[1981] James Tuck, Wonsun Ahn, Luis Ceze, and Josep Torrellas. SoftSig: softwareexposed hardware signatures for code analysis and optimization. ACM SIG-PLAN Notices, 43(3):145–156, March 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Burcea:2008:PV

[1982] Ioana Burcea, Stephen Somogyi, Andreas Moshovos, and Babak Falsafi. Predictor virtualization. ACM SIG-PLAN Notices, 43(3):157–167, March 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ganapathy:2008:DIM

[1983] Vinod Ganapathy, Matthew J. Renzelmann, Arini Balakrishnan, Michael M. Swift, and Somesh Jha. The design and implementation of microdrivers. ACM SIGPLAN Notices, 43(3):168– 178, March 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Weinsberg:2008:TFC

[1984] Yaron Weinsberg, Danny Dolev, Tal Anker, Muli Ben-Yehuda, and Pete Wyckoff. Tapping into the fountain of CPUs: on operating system support for programmable devices. ACM SIG-PLAN Notices, 43(3):179–188, March 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shen:2008:HCD

[1985] Kai Shen, Ming Zhong, Sandhya Dwarkadas, Chuanpeng Li, Christopher Stewart, and Xiao Zhang. Hardware counter driven on-the-fly request signatures. ACM SIGPLAN Notices, 43(3):189–200, March 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

VanErtvelde:2008:DPA

[1986] Luk Van Ertvelde and Lieven Eeckhout. Dispersing proprietary applications as benchmarks through code mutation. ACM SIGPLAN Notices, 43(3): 201–210, March 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mysore:2008:UVF

[1987] Shashidhar Mysore, Bita Mazloom, Banit Agrawal, and Timothy Sherwood. Understanding and visualizing full systems with data flow tomography. ACM SIGPLAN Notices, 43(3): 211–221, March 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ottoni:2008:COG

[1988] Guilherme Ottoni and David I. August. Communication optimizations for global multi-threaded instruction scheduling. ACM SIGPLAN Notices, 43(3):222–232, March 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kulkarni:2008:OPB

[1989] Milind Kulkarni, Keshav Pingali, Ganesh Ramanarayanan, Bruce Walter, Kavita Bala, and L. Paul Chew. Optimistic parallelism benefits from data partitioning. ACM SIGPLAN Notices, 43(3):233–243, March 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cox:2008:XEO

[1990] Russ Cox, Tom Bergan, Austin T. Clements, Frans Kaashoek, and Eddie Kohler. Xoc, an extension-oriented compiler for systems programming. ACM SIGPLAN Notices, 43(3):244– 254, March 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wells:2008:AIF

[1991] Philip M. Wells, Koushik Chakraborty, and Gurindar S. Sohi. Adapting to intermittent faults in multicore systems. ACM SIGPLAN Notices, 43(3): 255–264, March 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Li:2008:UPH

[1992] Man-Lap Li, Pradeep Ramachandran, Swarup Kumar Sahoo, Sarita V. Adve, Vikram S. Adve, and Yuanyuan Zhou. Understanding the propagation of hard errors to software and implications for resilient system design. ACM SIG-PLAN Notices, 43(3):265–276, March 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Suleman:2008:FDT

[1993] M. Aater Suleman, Moinuddin K. Qureshi, and Yale N. Patt. Feedbackdriven threading: power-efficient and high-performance execution of multithreaded workloads on CMPs. ACM SIGPLAN Notices, 43(3):277–286, March 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Linderman:2008:MPM

[1994] Michael D. Linderman, Jamison D.

Collins, Hong Wang, and Teresa H. Meng. Merge: a programming model for heterogeneous multi-core systems. ACM SIGPLAN Notices, 43(3):287– 296, March 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gummaraju:2008:SPG

[1995] Jayanth Gummaraju, Joel Coburn, Yoshio Turner, and Mendel Rosenblum. Streamware: programming general-purpose multicore processors using streams. ACM SIGPLAN Notices, 43(3):297–307, March 2008. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nightingale:2008:PSC

[1996] Edmund B. Nightingale, Daniel Peek, Peter M. Chen, and Jason Flinn. Parallelizing security checks on commodity hardware. ACM SIGPLAN Notices, 43(3):308–318, March 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Castro:2008:BBR

[1997] Miguel Castro, Manuel Costa, and Jean-Philippe Martin. Better bug reporting with better privacy. ACM SIG-PLAN Notices, 43(3):319–328, March 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lu:2008:LMC

[1998] Shan Lu, Soyeon Park, Eunsoo Seo, and Yuanyuan Zhou. Learning from mistakes: a comprehensive study on real world concurrency bug characteristics. ACM SIGPLAN Notices, 43(3): 329–339, March 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ce:2008:PVM

[1999] Yu Ce, Xu Zhen, Sun Ji-zhou, Meng Xiao-jing, Huang Yan-yan, and Wu Hua-bei. ParaModel: a visual modeling and code skeleton generation system for programming parallel applications. ACM SIGPLAN Notices, 43 (4):4–10, April 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Campanoni:2008:PDC

[2000] Simone Campanoni, Giovanni Agosta, and Stefano Crespi Reghizzi. A parallel dynamic compiler for CIL bytecode. ACM SIGPLAN Notices, 43(4):11–20, April 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sui:2008:ART

[2001] Yang Yi Sui, Jun Lin, and Xiao Tuo Zhang. An automated refactoring tool for dataflow visual programming language. ACM SIGPLAN Notices, 43 (4):21–28, April 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hook:2008:IDC

[2002] James Hook and Michael Sperber. The ICFP developer conference 2008. ACM SIGPLAN Notices, 43(5):3–4, May 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:2008:HJ

[2003] Paul Frenger. Hard Java. ACM SIG-PLAN Notices, 43(5):5–9, May 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Falsafi:2008:I

[2004] Babak Falsafi and Pascal Felber. Introduction. ACM SIGPLAN Notices, 43(5):10, May 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bailey:2008:TPA

[2005] Mark W. Bailey. TRANSACT 2008 paper abstracts. ACM SIGPLAN Notices, 43(5):11–14, May 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Menon:2008:SGL

[2006] Vijay Menon, Steven Balensiefer, Tatiana Shpeisman, Ali-Reza Adl-Tabatabai, Richard L. Hudson, Bratin Saha, and Adam Welc. Single global lock semantics in a weakly atomic STM. ACM SIGPLAN Notices, 43 (5):15–26, May 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Terauchi:2008:CRF

[2007] Tachio Terauchi. Checking race freedom via linear programming. ACM SIGPLAN Notices, 43(6):1–10, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sen:2008:RDR

[2008] Koushik Sen. Race directed random testing of concurrent programs. ACM SIGPLAN Notices, 43(6):11–21, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Blackburn:2008:IMR

[2009] Stephen M. Blackburn and Kathryn S. McKinley. Immix: a mark-region garbage collector with space efficiency, fast collection, and mutator performance. ACM SIGPLAN Notices, 43 (6):22–32, June 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pizlo:2008:SCR

[2010] Filip Pizlo, Erez Petrank, and Bjarne Steensgaard. A study of concurrent real-time garbage collectors. ACM SIGPLAN Notices, 43(6):33–44, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wang:2008:CCA

[2011] Xi Wang, Zhilei Xu, Xuezheng Liu, Zhenyu Guo, Xiaoge Wang, and Zheng Zhang. Conditional correlation analysis for safe region-based memory management. ACM SIGPLAN Notices, 43 (6):45–55, June 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Amin:2008:AVM

[2012] Ahmed M. Amin, Mithuna Thottethodi, T. N. Vijaykumar, Steven Wereley, and Stephen C. Jacobson. Automatic volume management for programmable microfluidics. ACM SIG-PLAN Notices, 43(6):56–67, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boehm:2008:FCC

[2013] Hans-J. Boehm and Sarita V. Adve.

Foundations of the C++ concurrency memory model. ACM SIGPLAN Notices, 43(6):68–78, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Huang:2008:ESS

[2014] Shan Shan Huang and Yannis Smaragdakis. Expressive and safe static reflection with MorphJ. ACM SIG-PLAN Notices, 43(6):79–89, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pouchet:2008:IOP

[2015] Louis-Noël Pouchet, Cédric Bastoul, Albert Cohen, and John Cavazos. Iterative optimization in the polyhedral model: part ii, multidimensional time. ACM SIGPLAN Notices, 43(6):90–100, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bondhugula:2008:PAP

[2016] Uday Bondhugula, Albert Hartono, J. Ramanujam, and P. Sadayappan. A practical automatic polyhedral parallelizer and locality optimizer. ACM SIGPLAN Notices, 43 (6):101–113, June 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kudlur:2008:OES

[2017] Manjunath Kudlur and Scott Mahlke. Orchestrating the execution of stream programs on multicore platforms. ACM SIGPLAN Notices, 43(6):114–124, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vechev:2008:DLF

[2018] Martin Vechev and Eran Yahav. Deriving linearizable fine-grained concurrent objects. ACM SIGPLAN Notices, 43 (6):125–135, June 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Solar-Lezama:2008:SCD

[2019] Armando Solar-Lezama, Christopher Grant Jones, and Rastislav Bodik. Sketching concurrent data structures. ACM SIGPLAN Notices, 43(6):136– 148, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anderson:2008:SCD

[2020] Zachary Anderson, David Gay, Rob Ennals, and Eric Brewer. SharC: checking data sharing strategies for multithreaded C. ACM SIGPLAN Notices, 43(6):149–158, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rondon:2008:LT

[2021] Patrick M. Rondon, Ming Kawaguci, and Ranjit Jhala. Liquid types.

ACM SIGPLAN Notices, 43(6):159– 169, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Feng:2008:CLL

[2022] Xinyu Feng, Zhong Shao, Yuan Dong, and Yu Guo. Certifying low-level programs with hardware interrupts and preemptive threads. ACM SIGPLAN Notices, 43(6):170–182, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:2008:TPC

[2023] Juan Chen, Chris Hawblitzel, Frances Perry, Mike Emmi, Jeremy Condit, Derrick Coetzee, and Polyvios Pratikaki. Type-preserving compilation for large-scale optimizing objectoriented compilers. ACM SIGPLAN Notices, 43(6):183–192, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McCamant:2008:QIF

[2024] Stephen McCamant and Michael D.

Ernst. Quantitative information flow as network flow capacity. ACM SIG-PLAN Notices, 43(6):193–205, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Godefroid:2008:GBW

[2025] Patrice Godefroid, Adam Kiezun, and Michael Y. Levin. Grammar-based whitebox fuzzing. ACM SIGPLAN Notices, 43(6):206–215, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pereira:2008:RAP

[2026] Fernando Magno Quintão Pereira and Jens Palsberg. Register allocation by puzzle solving. ACM SIGPLAN Notices, 43(6):216–226, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hack:2008:CCG

[2027] Sebastian Hack and Gerhard Goos. Copy coalescing by graph recoloring. ACM SIGPLAN Notices, 43(6): 227–237, June 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xin:2008:EPE

[2028] Bin Xin, William N. Sumner, and Xiangyu Zhang. Efficient program execution indexing. ACM SIGPLAN Notices, 43(6):238–248, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kahlon:2008:BTS

[2029] Vineet Kahlon. Bootstrapping: a technique for scalable flow and contextsensitive pointer alias analysis. ACM SIGPLAN Notices, 43(6):249–259, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vonDincklage:2008:EFP

[2030] Daniel von Dincklage and Amer Diwan. Explaining failures of program analyses. ACM SIGPLAN Notices, 43 (6):260–269, June 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dillig:2008:SCS

[2031] Isil Dillig, Thomas Dillig, and Alex Aiken. Sound, complete and scalable path-sensitive analysis. ACM SIG-PLAN Notices, 43(6):270–280, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gulwani:2008:PAC

[2032] Sumit Gulwani, Saurabh Srivastava, and Ramarathnam Venkatesan. Program analysis as constraint solving. ACM SIGPLAN Notices, 43(6):281– 292, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Flanagan:2008:VSC

[2033] Cormac Flanagan, Stephen N. Freund, and Jaeheon Yi. Velodrome: a sound and complete dynamic atomicity checker for multithreaded programs. ACM SIGPLAN Notices, 43 (6):293–303, June 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cherem:2008:ILA

[2034] Sigmund Cherem, Trishul Chilimbi, and Sumit Gulwani. Inferring locks for atomic sections. ACM SIGPLAN Notices, 43(6):304–315, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chugh:2008:DAC

[2035] Ravi Chugh, Jan W. Voung, Ranjit Jhala, and Sorin Lerner. Dataflow analysis for concurrent programs using datarace detection. ACM SIG-PLAN Notices, 43(6):316–326, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wegiel:2008:XTS

[2036] Michal Wegiel and Chandra Krintz. XMem: type-safe, transparent, shared memory for cross-runtime communication and coordination. ACM SIG-PLAN Notices, 43(6):327–338, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Halbwachs:2008:DPA

[2037] Nicolas Halbwachs and Mathias Péron. Discovering properties about arrays in simple programs. ACM SIGPLAN Notices, 43(6):339–348, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zee:2008:FFV

[2038] Karen Zee, Viktor Kuncak, and Martin Rinard. Full functional verification of linked data structures. ACM SIG-PLAN Notices, 43(6):349–361, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Musuvathi:2008:FSM

[2039] Madanlal Musuvathi and Shaz Qadeer. Fair stateless model checking. ACM SIGPLAN Notices, 43(6):362–371, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Guerraoui:2008:MCT

[2040] Rachid Guerraoui, Thomas A. Henzinger, Barbara Jobstmann, and Vasu Singh. Model checking transactional memories. ACM SIGPLAN Notices, 43 (6):372–382, June 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Auerbach:2008:FTG

[2041] Joshua Auerbach, David F. Bacon, Rachid Guerraoui, Jesper Honig Spring, and Jan Vitek. Flexible task graphs: a unified restricted thread programming model for Java. ACM SIG-PLAN Notices, 43(7):1–11, July 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lee:2008:EHS

[2042] Seong-Won Lee, Soo-Mook Moon, and Seong-Moo Kim. Enhanced hot spot detection heuristics for embedded Java just-in-time compilers. ACM SIG-PLAN Notices, 43(7):13–22, July 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Badea:2008:IJS

[2043] Carmen Badea, Alexandru Nicolau, and Alexander V. Veidenbaum. Impact of JVM superoperators on energy consumption in resource-constrained embedded systems. ACM SIGPLAN Notices, 43(7):23–30, July 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ebner:2008:GIS

[2044] Dietmar Ebner, Florian Brandner, Bernhard Scholz, Andreas Krall, Peter Wiedermann, and Albrecht Kadlec. Generalized instruction selection using ssa-graphs. ACM SIGPLAN Notices, 43(7):31–40, July 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cho:2008:CDD

[2045] Doosan Cho, Sudeep Pasricha, Ilya Issenin, Nikil Dutt, Yunheung Paek, and SunJun Ko. Compiler driven data layout optimization for regular/irregular array access patterns. ACM SIG-PLAN Notices, 43(7):41–50, July 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reineke:2008:RCA

[2046] Jan Reineke and Daniel Grund. Relative competitive analysis of cache replacement policies. ACM SIGPLAN Notices, 43(7):51–60, July 2008. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anand:2008:RSS

[2047] Madhukar Anand and Insup Lee. Robust and sustainable schedulability analysis of embedded software. ACM SIGPLAN Notices, 43(7):61–70, July 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Homayoun:2008:IPR

[2048] Houman Homayoun, Sudeep Pasricha, Mohammad Makhzan, and Alex Veidenbaum. Improving performance and reducing energy-delay with adaptive resource resizing for out-of-order embedded processors. ACM SIGPLAN Notices, 43(7):71–78, July 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rajopadhye:2008:DSI

[2049] Sanjay Rajopadhye, Gautam Gupta, and Lakshminarayanan Renganarayanan. A domain specific interconnect for reconfigurable computing. ACM SIG-PLAN Notices, 43(7):79–88, July 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lee:2008:FFC

[2050] Jaejin Lee, Junghyun Kim, Choonki Jang, Seungkyun Kim, Bernhard Egger, Kwangsub Kim, and SangYong Han. FaCSim: a fast and cycleaccurate architecture simulator for embedded systems. ACM SIGPLAN Notices, 43(7):89–100, July 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Delaval:2008:TSA

[2051] Gwenaël Delaval, Alain Girault, and Marc Pouzet. A type system for the automatic distribution of higher-order synchronous dataflow programs. ACM SIGPLAN Notices, 43(7):101–110, July 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cohen:2008:EEP

[2052] Norman H. Cohen and Karl Trygve Kalleberg. EventScript: an eventprocessing language based on regular expressions with actions. ACM SIGPLAN Notices, 43(7):111–120, July 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Biernacki:2008:CDM

[2053] Dariusz Biernacki, Jean-Louis Colaço, Gregoire Hamon, and Marc Pouzet. Clock-directed modular code generation for synchronous data-flow languages. ACM SIGPLAN Notices, 43 (7):121–130, July 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Newton:2008:DEC

[2054] Ryan R. Newton, Lewis D. Girod, Michael B. Craig, Samuel R. Madden, and John Gregory Morrisett. Design and evaluation of a compiler for embedded stream programs. ACM SIGPLAN Notices, 43(7):131–140, July 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bachir:2008:PPP

[2055] Mounira Bachir, Sid-Ahmed-Ali Touati, and Albert Cohen. Post-pass periodic register allocation to minimise loop unrolling degree. ACM SIG-PLAN Notices, 43(7):141–150, July 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

DeSutter:2008:PRB

[2056] Bjorn De Sutter, Paul Coene, Tom Vander Aa, and Bingfeng Mei. Placement-and-routing-based register allocation for coarse-grained reconfigurable arrays. ACM SIGPLAN Notices, 43(7):151–160, July 2008. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wang:2008:OSA

[2057] Li Wang, Xuejun Yang, Jingling Xue, Yu Deng, Xiaobo Yan, Tao Tang, and Quan Hoang Nguyen. Optimizing scientific application loops on stream processors. ACM SIGPLAN Notices, 43 (7):161–170, July 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McKinley:2008:IPQ

[2058] Kathryn S. McKinley. Improving publication quality by reducing bias with double-blind reviewing and author response. ACM SIGPLAN Notices, 43 (8):5–9, August 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jones:2008:ISM

[2059] Richard Jones and Steve Blackburn.

International Symposium of Memory Management (ISMM 2008). ACM SIG-PLAN Notices, 43(8):12–14, August 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bailey:2008:IPA

[2060] Mark W. Bailey. ISMM 2008 paper abstracts. ACM SIGPLAN Notices, 43 (8):15–19, August 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lampson:2008:LSE

[2061] Butler W. Lampson. Lazy and speculative execution in computer systems. ACM SIGPLAN Notices, 43(9): 1–2, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cheney:2008:FFU

[2062] James Cheney. FLUX: functional updates for XML. ACM SIGPLAN Notices, 43(9):3–14, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Castagna:2008:TIX

[2063] Giuseppe Castagna and Kim Nguyen. Typed iterators for XML. ACM SIG-PLAN Notices, 43(9):15–26, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jia:2008:APL

[2064] Limin Jia, Jeffrey A. Vaughan, Karl Mazurak, Jianzhou Zhao, Luke Zarko, Joseph Schorr, and Steve Zdancewic. AURA: a programming language for authorization and audit. ACM SIG-PLAN Notices, 43(9):27–38, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Oury:2008:PP

[2065] Nicolas Oury and Wouter Swierstra. The power of Pi. ACM SIGPLAN Notices, 43(9):39–50, September 2008.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schrijvers:2008:TCO

[2066] Tom Schrijvers, Simon Peyton Jones, Manuel Chakravarty, and Martin Sulzmann. Type checking with open type functions. ACM SIGPLAN Notices, 43(9):51–62, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Remy:2008:MMG

[2067] Didier Rémy and Boris Yakobowski.

From ML to ML^{F}: graphic type constraints with efficient type inference. ACM SIGPLAN Notices, 43(9): 63–74, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Guillemette:2008:TPC

[2068] Louis-Julien Guillemette and Stefan Monnier. A type-preserving compiler in Haskell. ACM SIGPLAN Notices, 43(9):75–86, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jones:2008:ERP

[2069] Mark P. Jones. Experience report:

playing the DSL card. ACM SIGPLAN Notices, 43(9):87–90, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Henglein:2008:GDS

[2070] Fritz Henglein. Generic discrimination:

sorting and partitioning unshared data in linear time. ACM SIGPLAN Notices, 43(9):91–102, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Effinger-Dean:2008:TEM

[2071] Laura Effinger-Dean, Matthew Kehrt, and Dan Grossman. Transactional events for ML. ACM SIGPLAN Notices, 43(9):103–114, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Convey:2008:ERE

[2072] Christian Convey, Andrew Fredricks, Christopher Gagner, Douglas Maxwell, and Lutz Hamel. Experience report: Erlang in acoustic ray tracing. ACM SIGPLAN Notices, 43(9):115– 118, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fluet:2008:ITP

[2073] Matthew Fluet, Mike Rainey, John Reppy, and Adam Shaw. Implicitlythreaded parallelism in Manticore. ACM SIGPLAN Notices, 43(9):119– 130, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Danvy:2008:DIP

[2074] Olivier Danvy. Defunctionalized interpreters for programming languages. ACM SIGPLAN Notices, 43(9):131– 142, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chlipala:2008:PHO

[2075] Adam Chlipala. Parametric higherorder abstract syntax for mechanized semantics. ACM SIGPLAN Notices, 43 (9):143–156, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ahmed:2008:TCC

[2076] Amal Ahmed and Matthias Blume.

Typed closure conversion preserves observational equivalence. ACM SIG-PLAN Notices, 43(9):157–168, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Morihata:2008:WIR

[2077] Akimasa Morihata, Kiminori Matsuzaki, and Masato Takeichi. Write it recursively: a generic framework for optimal path queries. ACM SIG-PLAN Notices, 43(9):169–178, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Adams:2008:ENE

[2078] Michael D. Adams and R. Kent Dybvig. Efficient nondestructive equality checking for trees and graphs. ACM SIGPLAN Notices, 43(9):179– 188, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hinze:2008:FPS

[2079] Ralf Hinze. Functional pearl: streams and unique fixed points. ACM SIG-PLAN Notices, 43(9):189–200, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fischer:2008:DFT

[2080] Sebastian Fischer and Herbert Kuchen. Data-flow testing of declarative programs. ACM SIGPLAN Notices, 43 (9):201–212, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chargueraud:2008:FTC

[2081] Arthur Charguéraud and François Pottier. Functional translation of a calculus of capabilities. ACM SIGPLAN Notices, 43(9):213–224, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Augustsson:2008:PTS

[2082] Lennart Augustsson, Howard Mansell, and Ganesh Sittampalam. Paradise: a two-stage DSL embedded in Haskell. ACM SIGPLAN Notices, 43(9):225– 228, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nanevski:2008:YDT

[2083] Aleksandar Nanevski, Greg Morrisett, Avraham Shinnar, Paul Govereau, and Lars Birkedal. Ynot: dependent types for imperative programs. ACM SIG-PLAN Notices, 43(9):229–240, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fluet:2008:SFG

[2084] Matthew Fluet, Mike Rainey, and John Reppy. A scheduling framework for general-purpose parallel languages. ACM SIGPLAN Notices, 43(9):241– 252, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Spoonhower:2008:SPP

[2085] Daniel Spoonhower, Guy E. Blelloch, Robert Harper, and Phillip B. Gibbons. Space profiling for parallel functional programs. ACM SIGPLAN Notices, 43(9):253–264, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jones:2008:PPT

[2086] Mark P. Jones. Polymorphism and page tables: systems programming from a functional programmer’s perspective. ACM SIGPLAN Notices, 43 (9):265–266, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Krauss:2008:PMP

[2087] Alexander Krauss. Pattern minimization problems over recursive data types. ACM SIGPLAN Notices, 43 (9):267–274, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

VanHorn:2008:DCC

[2088] David Van Horn and Harry G. Mairson. Deciding k CFA is complete for EXPTIME. ACM SIGPLAN Notices, 43(9):275–282, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Leijen:2008:HST

[2089] Daan Leijen. HMF: simple type inference for first-class polymorphism. ACM SIGPLAN Notices, 43(9):283– 294, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vytiniotis:2008:FFC

[2090] Dimitrios Vytiniotis, Stephanie Weirich, and Simon Peyton Jones. FPH: first-class polymorphism for Haskell. ACM SIGPLAN Notices, 43(9):295– 306, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dreyer:2008:MMM

[2091] Derek Dreyer and Andreas Rossberg. Mixin’ up the ML module system. ACM SIGPLAN Notices, 43(9):307– 320, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ley-Wild:2008:CSA

[2092] Ruy Ley-Wild, Matthew Fluet, and Umut A. Acar. Compiling selfadjusting programs with continuations. ACM SIGPLAN Notices, 43(9):321– 334, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mainland:2008:FSF

[2093] Geoffrey Mainland, Greg Morrisett, and Matt Welsh. Flask: staged functional programming for sensor networks. ACM SIGPLAN Notices, 43 (9):335–346, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nanavati:2008:ERP

[2094] Ravi Nanavati. Experience report: a pure shirt fits. ACM SIGPLAN Notices, 43(9):347–352, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Park:2008:FN

[2095] Sungwoo Park, Jinha Kim, and Hyeonseung Im. Functional netlists. ACM SIGPLAN Notices, 43(9):353–366, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dolstra:2008:NPF

[2096] Eelco Dolstra and Andres Löh. NixOS:

a purely functional Linux distribution. ACM SIGPLAN Notices, 43(9):367– 378, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Duke:2008:ERV

[2097] David J. Duke, Rita Borgo, Colin Runciman, and Malcolm Wallace. Experience report: visualizing data through functional pipelines. ACM SIGPLAN Notices, 43(9):379–382, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Foster:2008:QL

[2098] J. Nathan Foster, Alexandre Pilkiewicz, and Benjamin C. Pierce. Quotient lenses. ACM SIGPLAN Notices, 43 (9):383–396, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dolstra:2008:RTI

[2099] Eelco Dolstra, Jurriaan Hage, Bastiaan Heeren, Stefan Holdermans, Johan Jeuring, Andres Löh, Clara Löh, Arie Middelkoop, Alexey Rodriguez, and John van Schie. Report on the Tenth ICFP Programming Contest. ACM SIGPLAN Notices, 43(9):397– 408, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Willis:2008:CIJ

[2100] Darren Willis, David J. Pearce, and James Noble. Caching and incrementalisation in the Java Query Language. ACM SIGPLAN Notices, 43(10):1– 18, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wiedermann:2008:IQE

[2101] Ben Wiedermann, Ali Ibrahim, and William R. Cook. Interprocedural query extraction for transparent persistence. ACM SIGPLAN Notices, 43(10): 19–36, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tatlock:2008:DTR

[2102] Zachary Tatlock, Chris Tucker, David Shuffelton, Ranjit Jhala, and Sorin Lerner. Deep typechecking and refactoring. ACM SIGPLAN Notices, 43 (10):37–52, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Russo:2008:JPV

[2103] Claudio V. Russo. Join patterns for Visual Basic. ACM SIGPLAN Notices, 43(10):53–72, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gil:2008:WIS

[2104] Joseph Gil and Itay Maman. Whiteoak:

introducing structural typing into Java. ACM SIGPLAN Notices, 43(10):73– 90, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kats:2008:MSB

[2105] Lennart C. L. Kats, Martin Bravenboer, and Eelco Visser. Mixing source and bytecode: a case for compilation by normalization. ACM SIGPLAN Notices, 43(10):91–108, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bond:2008:TML

[2106] Michael D. Bond and Kathryn S.

McKinley. Tolerating memory leaks. ACM SIGPLAN Notices, 43(10):109– 126, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shankar:2008:JLD

[2107] Ajeet Shankar, Matthew Arnold, and Rastislav Bodik. Jolt: lightweight dynamic analysis and removal of object churn. ACM SIGPLAN Notices, 43(10):127–142, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Arnold:2008:QER

[2108] Matthew Arnold, Martin Vechev, and Eran Yahav. QVM: an efficient runtime for detecting defects in deployed systems. ACM SIGPLAN Notices, 43(10):143–162, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xian:2008:CAS

[2109] Feng Xian, Witawas Srisa-an, and Hong Jiang. Contention-aware scheduler: unlocking execution parallelism in multithreaded Java programs. ACM SIGPLAN Notices, 43(10):163–180, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schneider:2008:DOE

[2110] Florian T. Schneider, Vijay Menon, Tatiana Shpeisman, and Ali-Reza Adl-Tabatabai. Dynamic optimization for efficient strong atomicity. ACM SIGPLAN Notices, 43(10):181–194, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ni:2008:DIT

[2111] Yang Ni, Adam Welc, Ali-Reza Adl-Tabatabai, Moshe Bach, Sion Berkowits, James Cownie, Robert Geva, Sergey Kozhukow, Ravi Narayanaswamy, Jeffrey Olivier, Serguei Preis, Bratin Saha, Ady Tal, and Xinmin Tian. Design and implementation of transactional constructs for C/C++. ACM SIGPLAN Notices, 43(10):195–212, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Distefano:2008:JTP

[2112] Dino Distefano and Matthew J. Parkinson J. jStar: towards practical verification for Java. ACM SIGPLAN Notices, 43(10):213–226, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Beckman:2008:VCU

[2113] Nels E. Beckman, Kevin Bierhoff, and Jonathan Aldrich. Verifying correct usage of atomic blocks and typestate. ACM SIGPLAN Notices, 43(10):227– 244, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gopinathan:2008:EOP

[2114] Madhu Gopinathan and Sriram K. Rajamani. Enforcing object protocols by combining static and runtime analysis. ACM SIGPLAN Notices, 43(10):245– 260, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chern:2008:ISD

[2115] Rick Chern and Kris De Volder. The impact of static-dynamic coupling on remodularization. ACM SIGPLAN Notices, 43(10):261–276, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schafer:2008:SER

[2116] Max Schäfer, Torbjörn Ekman, and Oege de Moor. Sound and extensible renaming for Java. ACM SIGPLAN Notices, 43(10):277–294, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tansey:2008:ARI

[2117] Wesley Tansey and Eli Tilevich. Annotation refactoring: inferring upgrade transformations for legacy applications. ACM SIGPLAN Notices, 43(10):295– 312, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dagenais:2008:ESA

[2118] Barthélémy Dagenais and Laurie Hendren. Enabling static analysis for partial Java programs. ACM SIGPLAN Notices, 43(10):313–328, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ferrara:2008:SUC

[2119] Pietro Ferrara, Francesco Logozzo, and Manuel Fahndrich. Safer unsafe code for .NET. ACM SIGPLAN Notices, 43(10):329–346, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Naeem:2008:TLA

[2120] Nomair A. Naeem and Ondrej Lhotak. Typestate-like analysis of multiple interacting objects. ACM SIGPLAN Notices, 43(10):347–366, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Georges:2008:JPE

[2121] Andy Georges, Lieven Eeckhout, and Dries Buytaert. Java performance evaluation through rigorous replay compilation. ACM SIGPLAN Notices, 43 (10):367–384, September 2008. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kawachiya:2008:ARM

[2122] Kiyokuni Kawachiya, Kazunori Ogata, and Tamiya Onodera. Analysis and reduction of memory inefficiencies in Java strings. ACM SIGPLAN Notices, 43(10):385–402, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Prokopski:2008:APC

[2123] Gregory B. Prokopski and Clark Verbrugge. Analyzing the performance of code-copying virtual machines. ACM SIGPLAN Notices, 43(10):403–422, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Moors:2008:GHK

[2124] Adriaan Moors, Frank Piessens, and Martin Odersky. Generics of a higher kind. ACM SIGPLAN Notices, 43(10): 423–438, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Oliveira:2008:VPR

[2125] Bruno C.d.S. Oliveira, Meng Wang, and Jeremy Gibbons. The visitor pattern as a reusable, generic, typesafe component. ACM SIGPLAN Notices, 43(10):439–456, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nystrom:2008:CTO

[2126] Nathaniel Nystrom, Vijay Saraswat, Jens Palsberg, and Christian Grothoff. Constrained types for object-oriented languages. ACM SIGPLAN Notices, 43(10):457–474, September 2008. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bellamy:2008:ELT

[2127] Ben Bellamy, Pavel Avgustinov, Oege de Moor, and Damien Sereni. Efficient local type inference. ACM SIGPLAN Notices, 43(10):475–492, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Roberson:2008:ESM

[2128] Michael Roberson, Melanie Harries, Paul T. Darga, and Chandrasekhar Boyapati. Efficient software model checking of soundness of type systems. ACM SIGPLAN Notices, 43(10):493– 504, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smith:2008:JTI

[2129] Daniel Smith and Robert Cartwright. Java type inference is broken: can we fix it? ACM SIGPLAN Notices, 43(10):505–524, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schippers:2008:DBS

[2130] Hans Schippers, Dirk Janssens, Michael Haupt, and Robert Hirschfeld. Delegationbased semantics for modularizing crosscutting concerns. ACM SIGPLAN Notices, 43(10):525–542, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baldi:2008:TAL

[2131] Pierre F. Baldi, Cristina V. Lopes, Erik J. Linstead, and Sushil K. Bajracharya. A theory of aspects as latent topics. ACM SIGPLAN Notices, 43(10):543–562, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Muschevici:2008:MDP

[2132] Radu Muschevici, Alex Potanin, Ewan Tempero, and James Noble. Multiple dispatch in practice. ACM SIGPLAN Notices, 43(10):563–582, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bouillet:2008:TBA

[2133] Eric Bouillet, Mark Feblowitz, Zhen Liu, Anand Ranganathan, and Anton Riabov. A tag-based approach for the design and composition of information processing applications. ACM SIGPLAN Notices, 43(10):585– 602, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Simpkins:2008:TAP

[2134] Christopher Simpkins, Sooraj Bhat, Charles Isbell, Jr., and Michael Mateas.

Towards adaptive programming: integrating reinforcement learning into a programming language. ACM SIGPLAN Notices, 43(10):603–614, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gabriel:2008:DD

[2135] Richard P. Gabriel. Designed as designer. ACM SIGPLAN Notices, 43 (10):617–632, September 2008. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bailey:2008:SPLa

[2136] Mark W. Bailey. SIGPLAN programming language curriculum workshop: Workshop organization. ACM SIG-PLAN Notices, 43(11):1–6, November 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bailey:2008:SPLb

[2137] Mark W. Bailey. SIGPLAN programming language curriculum workshop: Discussion summaries and recommendations. ACM SIGPLAN Notices, 43(11):6–29, November 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bailey:2008:SPLc

[2138] Mark W. Bailey. SIGPLAN programming language curriculum workshop: Workshop report summary. ACM SIG-PLAN Notices, 43(11):29–30, November 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Allen:2008:STC

[2139] Eric Allen. Some things that computer science majors should know. ACM SIGPLAN Notices, 43(11):32– 35, November 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bailey:2008:IPL

[2140] Mark W. Bailey. Injecting programming language concepts throughout the curriculum: an inclusive strategy. ACM SIGPLAN Notices, 43(11):36– 38, November 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bodik:2008:SLU

[2141] Rastislav Bodik. Small languages in an undergraduate PL/Compiler course. ACM SIGPLAN Notices, 43(11):39– 44, November 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bruce:2008:PLL

[2142] Kim Bruce and Stephen N. Freund.

Programming languages in a liberal arts education. ACM SIGPLAN Notices, 43(11):45–49, November 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bruce:2008:PLP

[2143] Kim Bruce and Stephen N. Freund.

Programming languages as part of core computer science. ACM SIGPLAN Notices, 43(11):50–54, November 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cook:2008:HLP

[2144] William R. Cook. High-level problems in teaching undergraduate programming languages. ACM SIGPLAN Notices, 43(11):55–58, November 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Felleisen:2008:WTP

[2145] Matthias Felleisen. Why teach programming languages in this day and age and how to go about it. ACM SIG-PLAN Notices, 43(11):59–61, November 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fisher:2008:WNM

[2146] Kathleen Fisher. We need more than one: why students need a sophisticated understanding of programming languages. ACM SIGPLAN Notices, 43(11):62–65, November 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fisler:2008:IDS

[2147] Kathi Fisler. Implementing domainspecific languages as the foundation of an honors intro CS course. ACM SIG-PLAN Notices, 43(11):66–70, November 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Harper:2008:PPP

[2148] Robert Harper. Position paper: practical foundations for programming languages. ACM SIGPLAN Notices, 43 (11):71–73, November 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hind:2008:ADB

[2149] Michael Hind. Addressing the disconnect between the good and the popular. ACM SIGPLAN Notices, 43(11): 74–76, November 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hughes:2008:ETF

[2150] John Hughes. Experiences from teaching functional programming at Chalmers. ACM SIGPLAN Notices, 43 (11):77–80, November 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Krishnamurthi:2008:TPL

[2151] Shriram Krishnamurthi. Teaching programming languages in a post-Linnaean age. ACM SIGPLAN Notices, 43(11):81–83, November 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Larus:2008:PRC

[2152] James Larus. PL research and its consequences on PL curriculum. ACM SIGPLAN Notices, 43(11):84– 86, November 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lea:2008:LPE

[2153] Doug Lea, David F. Bacon, and David Grove. Languages and performance engineering: method, instrumentation, and pedagogy. ACM SIGPLAN Notices, 43(11):87–92, November 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Leavens:2008:UCP

[2154] Gary T. Leavens. Use concurrent programming models to motivate teaching of programming languages. ACM SIG-PLAN Notices, 43(11):93–98, November 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pollock:2008:RPT

[2155] Lori Pollock. Rethinking pedagogy for teaching PL with more than PL concepts in mind. ACM SIGPLAN Notices, 43(11):99–103, November 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reges:2008:MPL

[2156] Stuart Reges. Marketing the programming languages course. ACM SIGPLAN Notices, 43(11):104–107, November 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reynolds:2008:STT

[2157] John C. Reynolds. Some thoughts on teaching programming and programming languages. ACM SIGPLAN Notices, 43(11):108–110, November 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rinard:2008:UPL

[2158] Martin Rinard. Using programming language concepts to teach general thinking skills. ACM SIGPLAN Notices, 43(11):111–118, November 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sestoft:2008:PLC

[2159] Peter Sestoft. Programming language concepts for software developers. ACM SIGPLAN Notices, 43(11):119– 123, November 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sheldon:2008:AOA

[2160] Mark A. Sheldon and Franklyn Turbak.

An aspect-oriented approach to the undergraduate programming language curriculum. ACM SIGPLAN Notices, 43(11):124–129, November 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shivers:2008:WTP

[2161] Olin Shivers. Why teach programming languages. ACM SIGPLAN Notices, 43(11):130–132, November 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wand:2008:PLF

[2162] Mitchell Wand and Daniel P. Friedman.

Programming languages: fundamental concepts for expanding and disciplining the mind. ACM SIGPLAN Notices, 43(11):133–135, November 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pistoia:2008:WPL

[2163] Marco Pistoia and Úlfar Erlingsson. Workshop on Programming Languages and Analysis for Security (PLAS 2008). ACM SIGPLAN Notices, 43(12):3– 4, December 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pistoia:2008:PPA

[2164] Marco Pistoia and Úlfar Erlingsson. PLAS 2008 paper abstracts. ACM SIG-PLAN Notices, 43(12):5–8, December 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chaudhuri:2008:TSD

[2165] Avik Chaudhuri, Prasad Naldurg, and Sriram Rajamani. A type system for data-flow integrity on Windows Vista. ACM SIGPLAN Notices, 43(12):9– 20, December 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Swamy:2008:VES

[2166] Nikhil Swamy and Michael Hicks. Verified enforcement of stateful information release policies. ACM SIGPLAN Notices, 43(12):21–31, December 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pistoia:2008:PLP

[2167] Marco Pistoia and Úlfar Erlingsson. Programming languages and program analysis for security: a three-year retrospective. ACM SIGPLAN Notices, 43(12):32–39, December 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Harris:2009:LCT

[2168] Tim Harris. Language constructs for transactional memory. ACM SIG-PLAN Notices, 44(1):1, January 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Elmas:2009:CAA

[2169] Tayfun Elmas, Shaz Qadeer, and Serdar Tasiran. A calculus of atomic actions. ACM SIGPLAN Notices, 44(1): 2–15, January 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gotsman:2009:PNB

[2170] Alexey Gotsman, Byron Cook, Matthew Parkinson, and Viktor Vafeiadis. Proving that non-blocking algorithms don’t block. ACM SIGPLAN Notices, 44(1): 16–28, January 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Abadi:2009:MCT

[2171] Martin Abadi and Gordon Plotkin. A model of cooperative threads. ACM SIGPLAN Notices, 44(1):29–40, January 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xu:2009:SCC

[2172] Dana N. Xu, Simon Peyton Jones, and Koen Claessen. Static contract checking for Haskell. ACM SIGPLAN Notices, 44(1):41–52, January 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Qi:2009:MTS

[2173] Xin Qi and Andrew C. Myers. Masked types for sound object initialization. ACM SIGPLAN Notices, 44(1):53–65, January 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Leijen:2009:FTR

[2174] Daan Leijen. Flexible types: robust type inference for first-class polymorphism. ACM SIGPLAN Notices, 44(1): 66–77, January 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lublinerman:2009:MCG

[2175] Roberto Lublinerman, Christian Szegedy, and Stavros Tripakis. Modular code generation from synchronous block diagrams: modularity vs. code size. ACM SIGPLAN Notices, 44(1):78–89, January 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Barthe:2009:FCC

[2176] Gilles Barthe, Benjamin Grégoire, and Santiago Zanella Béguelin. Formal certification of code-based cryptographic proofs. ACM SIGPLAN Notices, 44(1): 90–101, January 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ganty:2009:VLA

[2177] Pierre Ganty, Rupak Majumdar, and Andrey Rybalchenko. Verifying liveness for asynchronous programs. ACM SIGPLAN Notices, 44(1):102–113, January 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brunel:2009:FFB

[2178] Julien Brunel, Damien Doligez, René Rydhof Hansen, Julia L. Lawall, and Gilles Muller. A foundation for flow-based program matching: using temporal logic and model checking. ACM SIGPLAN Notices, 44(1):114– 126, January 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gulwani:2009:SPE

[2179] Sumit Gulwani, Krishna K. Mehra, and Trishul Chilimbi. SPEED: precise and efficient static estimation of program computational complexity. ACM SIG-PLAN Notices, 44(1):127–139, January 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Monniaux:2009:AMA

[2180] David P. Monniaux. Automatic modular abstractions for linear constraints. ACM SIGPLAN Notices, 44(1):140– 151, January 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Barker:2009:WCO

[2181] Chris Barker. Wild control operators. ACM SIGPLAN Notices, 44(1): 152, January 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Garcia:2009:LED

[2182] Ronald Garcia, Andrew Lumsdaine, and Amr Sabry. Lazy evaluation and delimited control. ACM SIGPLAN Notices, 44(1):153–164, January 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Voigtlander:2009:BFP

[2183] Janis Voigtländer. Bidirectionalization for free! (pearl). ACM SIGPLAN Notices, 44(1):165–176, January 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Morihata:2009:THT

[2184] Akimasa Morihata, Kiminori Matsuzaki, Zhenjiang Hu, and Masato Takeichi. The third homomorphism theorem on trees: downward & upward lead to divide-and-conquer. ACM SIG-PLAN Notices, 44(1):177–185, January 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ley-Wild:2009:CSS

[2185] Ruy Ley-Wild, Umut A. Acar, and Matthew Fluet. A cost semantics for self-adjusting computation. ACM SIG-PLAN Notices, 44(1):186–199, January 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tozawa:2009:CWP

[2186] Akihiko Tozawa, Michiaki Tatsubori, Tamiya Onodera, and Yasuhiko Minamide. Copy-on-write in the PHP language. ACM SIGPLAN Notices, 44(1):

200–212, January 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bronson:2009:FDB

[2187] Nathan G. Bronson, Christos Kozyrakis, and Kunle Olukotun. Feedbackdirected barrier optimization in a strongly isolated STM. ACM SIG-PLAN Notices, 44(1):213–225, January 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hardekopf:2009:SSF

[2188] Ben Hardekopf and Calvin Lin. Semisparse flow-sensitive pointer analysis. ACM SIGPLAN Notices, 44(1):226– 238, January 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gulwani:2009:CFT

[2189] Sumit Gulwani, Tal Lev-Ami, and Mooly Sagiv. A combination framework for tracking partition sizes. ACM SIGPLAN Notices, 44(1):239–251, January 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wang:2009:TDA

[2190] Yin Wang, Stéphane Lafortune, Terence Kelly, Manjunath Kudlur, and Scott Mahlke. The theory of deadlock avoidance via discrete control. ACM SIGPLAN Notices, 44(1):252–263, January 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tate:2009:ESN

[2191] Ross Tate, Michael Stepp, Zachary Tatlock, and Sorin Lerner. Equality saturation: a new approach to optimization. ACM SIGPLAN Notices, 44(1):264–276, January 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jonsson:2009:PSH

[2192] Peter A. Jonsson and Johan Nordlander. Positive supercompilation for a higher order call-by-value language. ACM SIGPLAN Notices, 44(1):277– 288, January 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Calcagno:2009:CSA

[2193] Cristiano Calcagno, Dino Distefano, Peter O’Hearn, and Hongseok Yang. Compositional shape analysis by means of bi-abduction. ACM SIGPLAN Notices, 44(1):289–300, January 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Simpson:2009:LTC

[2194] Alex Simpson. Linear types for computational effects. ACM SIGPLAN Notices, 44(1):301, January 2009. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Condit:2009:UTC

[2195] Jeremy Condit, Brian Hackett, Shuvendu K. Lahiri, and Shaz Qadeer. Unifying type checking and property checking for low-level code. ACM SIG-PLAN Notices, 44(1):302–314, January 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Feng:2009:LRG

[2196] Xinyu Feng. Local rely-guarantee reasoning. ACM SIGPLAN Notices, 44(1): 315–327, January 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brotherston:2009:CBL

[2197] James Brotherston and Cristiano Calcagno. Classical BI: a logic for reasoning about dualising resources. ACM SIGPLAN Notices, 44(1):328–339, January 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ahmed:2009:SDR

[2198] Amal Ahmed, Derek Dreyer, and Andreas Rossberg. State-dependent representation independence. ACM SIG-PLAN Notices, 44(1):340–353, January 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Montagu:2009:MAT

[2199] Benoît Montagu and Didier Rémy.

Modeling abstract types in modules with open existential types. ACM SIG-PLAN Notices, 44(1):354–365, January 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Krishnaswami:2009:FPM

[2200] Neelakantan R. Krishnaswami. Focusing on pattern matching. ACM SIG-PLAN Notices, 44(1):366–378, January 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sarkar:2009:SXC

[2201] Susmit Sarkar, Peter Sewell, Francesco Zappa Nardelli, Scott Owens, Tom Ridge, Thomas Braibant, Magnus O. Myreen, and Jade Alglave. The semantics of x86-CC multiprocessor machine code. ACM SIGPLAN Notices, 44(1):379– 391, January 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boudol:2009:RMM

[2202] Gérard Boudol and Gustavo Petri. Relaxed memory models: an operational approach. ACM SIGPLAN Notices, 44(1):392–403, January 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Guerraoui:2009:SPL

[2203] Rachid Guerraoui and Michal Kapalka. The semantics of progress in lock-based transactional memory. ACM SIG-PLAN Notices, 44(1):404–415, January 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kobayashi:2009:THO

[2204] Naoki Kobayashi. Types and higherorder recursion schemes for verification of higher-order programs. ACM SIG-PLAN Notices, 44(1):416–428, January 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ridge:2009:VDS

[2205] Thomas Ridge. Verifying distributed systems: the operational approach. ACM SIGPLAN Notices, 44(1):429– 440, January 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hawblitzel:2009:AVP

[2206] Chris Hawblitzel and Erez Petrank.

Automated verification of practical garbage collectors. ACM SIGPLAN Notices, 44(1):441–453, January 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kiselyov:2009:LMR

[2207] Oleg Kiselyov and Chung chieh Shan.

Lightweight monadic regions. ACM SIGPLAN Notices, 44(2):1–12, February 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Russo:2009:LLW

[2208] Alejandro Russo, Koen Claessen, and John Hughes. A library for light-weight information-flow security in haskell. ACM SIGPLAN Notices, 44(2):13– 24, February 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pucella:2009:HST

[2209] Riccardo Pucella and Jesse A. Tov.

Haskell session types with (almost) no class. ACM SIGPLAN Notices, 44(2): 25–36, February 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Runciman:2009:SLS

[2210] Colin Runciman, Matthew Naylor, and Fredrik Lindblad. SmallCheck and Lazy SmallCheck: automatic exhaustive testing for small values. ACM SIG-PLAN Notices, 44(2):37–48, February 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mitchell:2009:APE

[2211] Neil Mitchell and Colin Runciman. Not all patterns, but enough: an automatic verifier for partial but sufficient pattern matching. ACM SIGPLAN Notices, 44(2):49–60, February 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bernardy:2009:YEH

[2212] Jean-Philippe Bernardy. Yi: an editor in Haskell for Haskell. ACM SIG-PLAN Notices, 44(2):61–62, February 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Viera:2009:HDY

[2213] Marcos Viera, S. Doaitse Swierstra, and Eelco Lempsink. Haskell, do you read me?: constructing and composing efficient top-down parsers at runtime. ACM SIGPLAN Notices, 44(2): 63–74, February 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ahn:2009:SSS

[2214] Ki Yung Ahn and Tim Sheard. Shared subtypes: subtyping recursive parametrized algebraic data types.

ACM SIGPLAN Notices, 44(2):75– 86, February 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jones:2009:LPD

[2215] Mark P. Jones and Iavor S. Diatchki.

Language and program design for functional dependencies. ACM SIGPLAN Notices, 44(2):87–98, February 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kariotis:2009:MMF

[2216] Pericles S. Kariotis, Adam M. Procter, and William L. Harrison. Making monads first-class with Template Haskell. ACM SIGPLAN Notices, 44(2):99– 110, February 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rodriguez:2009:CLG

[2217] Alexey Rodriguez, Johan Jeuring, Patrik Jansson, Alex Gerdes, Oleg Kiselyov, and Bruno C. d. S. Oliveira. Comparing libraries for generic programming in haskell. ACM SIGPLAN Notices, 44(2):111–122, February 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Allwood:2009:CCL

[2218] Tristan O. R. Allwood and Susan Eisenbach. CLASE: cursor library for a structured editor. ACM SIGPLAN Notices, 44(2):123–124, February 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Coutts:2009:HBI

[2219] Duncan Coutts, Isaac Potoczny-Jones, and Don Stewart. Haskell: batteries included. ACM SIGPLAN Notices, 44(2):125–126, February 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gebhart:2009:ETC

[2220] Mark Gebhart, Bertrand A. Maher, Katherine E. Coons, Jeff Diamond, Paul Gratz, Mario Marino, Nitya Ranganathan, Behnam Robatmili, Aaron Smith, James Burrill, Stephen W. Keckler, Doug Burger, and Kathryn S. McKinley. An evaluation of the TRIPS computer system. ACM SIGPLAN Notices, 44(3):1–12, March 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pistol:2009:AIN

[2221] Constantin Pistol, Wutichai Chongchitmate, Christopher Dwyer, and Alvin R. Lebeck. Architectural implications of nanoscale integrated sensing and computing. ACM SIGPLAN Notices, 44(3):13–24, March 2009. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Park:2009:CEA

[2222] Soyeon Park, Shan Lu, and Yuanyuan Zhou. CTrigger: exposing atomicity violation bugs from their hiding places. ACM SIGPLAN Notices, 44(3):25–36, March 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sidiroglou:2009:AAS

[2223] Stelios Sidiroglou, Oren Laadan, Carlos Perez, Nicolas Viennot, Jason Nieh, and Angelos D. Keromytis. ASSURE: automatic software self-healing using rescue points. ACM SIGPLAN Notices, 44(3):37–48, March 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lenharth:2009:RDO

[2224] Andrew Lenharth, Vikram S. Adve, and Samuel T. King. Recovery domains: an organizing principle for recoverable operating systems. ACM SIGPLAN Notices, 44(3):49–60, March 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dimitrov:2009:ABB

[2225] Martin Dimitrov and Huiyang Zhou.

Anomaly-based bug prediction, isolation, and validation: an automated approach for software debugging. ACM SIGPLAN Notices, 44(3):61–72, March 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Montesinos:2009:CSH

[2226] Pablo Montesinos, Matthew Hicks, Samuel T. King, and Josep Torrellas. Capo: a software-hardware interface for practical deterministic multiprocessor replay. ACM SIGPLAN Notices, 44 (3):73–84, March 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Devietti:2009:DDS

[2227] Joseph Devietti, Brandon Lucia, Luis Ceze, and Mark Oskin. DMP: deterministic shared memory multiprocessing. ACM SIGPLAN Notices, 44 (3):85–96, March 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Olszewski:2009:KED

[2228] Marek Olszewski, Jason Ansel, and Saman Amarasinghe. Kendo: efficient deterministic multithreading in software. ACM SIGPLAN Notices, 44(3): 97–108, March 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tiwari:2009:CIF

[2229] Mohit Tiwari, Hassan M. G. Wassel, Bita Mazloom, Shashidhar Mysore, Frederic T. Chong, and Timothy Sherwood. Complete information flow tracking from the gates up. ACM SIG-PLAN Notices, 44(3):109–120, March 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tam:2009:RAL

[2230] David K. Tam, Reza Azimi, Livio B. Soares, and Michael Stumm. RapidMRC: approximating L2 miss rate curves on commodity systems for online optimizations. ACM SIGPLAN Notices, 44(3):121–132, March 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Eyerman:2009:PTC

[2231] Stijn Eyerman and Lieven Eeckhout. Per-thread cycle accounting in SMT processors. ACM SIGPLAN Notices, 44(3):133–144, March 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hofmann:2009:MBM

[2232] Owen S. Hofmann, Christopher J.

Rossbach, and Emmett Witchel. Maximum benefit from a minimal HTM. ACM SIGPLAN Notices, 44(3):145– 156, March 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dice:2009:EEC

[2233] Dave Dice, Yossi Lev, Mark Moir, and Daniel Nussbaum. Early experience with a commercial hardware transactional memory implementation. ACM SIGPLAN Notices, 44(3):157– 168, March 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wells:2009:MMM

[2234] Philip M. Wells, Koushik Chakraborty, and Gurindar S. Sohi. Mixed-mode multicore reliability. ACM SIGPLAN Notices, 44(3):169–180, March 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rajamani:2009:IDE

[2235] Sriram Rajamani, G. Ramalingam, Venkatesh Prasad Ranganath, and Kapil Vaswani. ISOLATOR: dynamically ensuring isolation in concurrent programs. ACM SIGPLAN Notices, 44(3):181–192, March 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tucek:2009:EOV

[2236] Joseph Tucek, Weiwei Xiong, and Yuanyuan Zhou. Efficient online validation with delta execution. ACM SIG-PLAN Notices, 44(3):193–204, March 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meisner:2009:PES

[2237] David Meisner, Brian T. Gold, and Thomas F. Wenisch. PowerNap: eliminating server idle power. ACM SIG-PLAN Notices, 44(3):205–216, March 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Caulfield:2009:GUF

[2238] Adrian M. Caulfield, Laura M. Grupp, and Steven Swanson. Gordon: using flash memory to build fast, powerefficient clusters for data-intensive applications. ACM SIGPLAN Notices, 44(3):217–228, March 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gupta:2009:DFT

[2239] Aayush Gupta, Youngjae Kim, and Bhuvan Urgaonkar. DFTL: a flash translation layer employing demandbased selective caching of page-level address mappings. ACM SIGPLAN Notices, 44(3):229–240, March 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Aleen:2009:CAS

[2240] Farhana Aleen and Nathan Clark.

Commutativity analysis for software parallelization: letting program transformations see the big picture. ACM SIGPLAN Notices, 44(3):241–252, March 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Suleman:2009:ACS

[2241] M. Aater Suleman, Onur Mutlu, Moinuddin K. Qureshi, and Yale N. Patt. Accelerating critical section execution with asymmetric multi-core architectures. ACM SIGPLAN Notices, 44(3): 253–264, March 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mytkowicz:2009:PWD

[2242] Todd Mytkowicz, Amer Diwan, Matthias Hauswirth, and Peter F. Sweeney. Producing wrong data without doing anything obviously wrong! ACM SIGPLAN Notices, 44(3):265– 276, March 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bond:2009:LP

[2243] Michael D. Bond and Kathryn S.

McKinley. Leak pruning. ACM SIG-PLAN Notices, 44(3):277–288, March 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wegiel:2009:DPC

[2244] Michal Wegiel and Chandra Krintz.

Dynamic prediction of collection yield for managed runtimes. ACM SIG-PLAN Notices, 44(3):289–300, March 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Menon:2009:TSA

[2245] Aravind Menon, Simon Schubert, and Willy Zwaenepoel. TwinDrivers: semiautomatic derivation of fast and safe hypervisor network drivers from guest OS drivers. ACM SIGPLAN Notices, 44(3):301–312, March 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Burcea:2009:PBV

[2246] Ioana Burcea and Andreas Moshovos. Phantom-BTB: a virtualized branch target buffer design. ACM SIGPLAN Notices, 44(3):313–324, March 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ramani:2009:SSF

[2247] Karthik Ramani, Christiaan P. Gribble, and Al Davis. StreamRay: a stream filtering architecture for coherent ray tracing. ACM SIGPLAN Notices, 44(3):325–336, March 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cameron:2009:ASS

[2248] Robert D. Cameron and Dan Lin. Architectural support for SWAR text processing with parallel bit streams: the inductive doubling principle. ACM SIGPLAN Notices, 44(3):337–348, March 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Blelloch:2009:PT

[2249] Guy E. Blelloch. Parallel thinking.

ACM SIGPLAN Notices, 44(4):1–2, April 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kulkarni:2009:HMP

[2250] Milind Kulkarni, Martin Burtscher, Rajeshkar Inkulu, Keshav Pingali, and Calin Casçaval. How much parallelism is there in irregular applications? ACM SIGPLAN Notices, 44(4):3–14, April 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kang:2009:ETM

[2251] Seunghwa Kang and David A. Bader.

An efficient transactional memory algorithm for computing minimum spanning forest of sparse graphs. ACM SIGPLAN Notices, 44(4):15–24, April 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zyulkyarov:2009:AQU

[2252] Ferad Zyulkyarov, Vladimir Gajinov, Osman S. Unsal, Adrián Cristal, Eduard Ayguadé, Tim Harris, and Mateo Valero. Atomic quake: using transactional memory in an interactive multiplayer game server. ACM SIGPLAN Notices, 44(4):25–34, April 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yau:2009:AAM

[2253] Siu Yau, Vijay Karamcheti, Denis Zorin, Kostadin Damevski, and Steven G. Parker. Application-aware management of parallel simulation collections. ACM SIGPLAN Notices, 44 (4):35–44, April 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Michael:2009:IWS

[2254] Maged M. Michael, Martin T. Vechev, and Vijay A. Saraswat. Idempotent work stealing. ACM SIGPLAN Notices, 44(4):45–54, April 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hiraishi:2009:BBL

[2255] Tasuku Hiraishi, Masahiro Yasugi, Seiji Umatani, and Taiichi Yuasa. Backtracking-based load balancing. ACM SIGPLAN Notices, 44(4):55–64, April 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Li:2009:ESM

[2256] Tong Li, Dan Baumberger, and Scott Hahn. Efficient and scalable multiprocessor fair scheduling using distributed weighted round-robin. ACM SIGPLAN Notices, 44(4):65–74, April 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wang:2009:MPM

[2257] Zheng Wang and Michael F. P. O’Boyle. Mapping parallelism to multicores: a machine learning based approach. ACM SIGPLAN Notices, 44 (4):75–84, April 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Allen:2009:SSD

[2258] Matthew D. Allen, Srinath Sridharan, and Gurindar S. Sohi. Serialization sets: a dynamic dependencebased parallel execution model. ACM SIGPLAN Notices, 44(4):85–96, April 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hill:2009:OBS

[2259] Mark D. Hill. Opportunities beyond single-core microprocessors. ACM SIG-PLAN Notices, 44(4):97, April 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Patt:2009:MCD

[2260] Yale N. Patt. Multi-core demands multi-interfaces. ACM SIGPLAN Notices, 44(4):99–100, April 2009. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lee:2009:OGC

[2261] Seyong Lee, Seung-Jai Min, and Rudolf Eigenmann. OpenMP to GPGPU: a compiler framework for automatic translation and optimization. ACM SIGPLAN Notices, 44(4): 101–110, April 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yang:2009:CGC

[2262] Xuejun Yang, Li Wang, Jingling Xue, Yu Deng, and Ying Zhang. Comparability graph coloring for optimizing utilization of stream register files in stream processors. ACM SIG-PLAN Notices, 44(4):111–120, April 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Quintana-Orti:2009:SDL

[2263] Gregorio Quintana-Ortí, Francisco D. Igual, Enrique S. Quintana-Ortí, and Robert A. van de Geijn. Solving dense linear systems on platforms with multiple hardware accelerators. ACM SIG-PLAN Notices, 44(4):121–130, April 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schneider:2009:CPM

[2264] Scott Schneider, Jae-Seung Yeom, Benjamin Rose, John C. Linford, Adrian Sandu, and Dimitrios S. Nikolopoulos. A comparison of programming models for multiprocessors with explicitly managed memory hierarchies. ACM SIGPLAN Notices, 44(4):131– 140, April 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Spear:2009:CSC

[2265] Michael F. Spear, Luke Dalessandro, Virendra J. Marathe, and Michael L. Scott. A comprehensive strategy for contention management in software transactional memory. ACM SIG-PLAN Notices, 44(4):141–150, April 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Agrawal:2009:SON

[2266] Kunal Agrawal, I-Ting Angelina Lee, and Jim Sukha. Safe open-nested transactions through ownership. ACM SIG-PLAN Notices, 44(4):151–162, April 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ramadan:2009:CCT

[2267] Hany E. Ramadan, Indrajit Roy, Maurice Herlihy, and Emmett Witchel. Committing conflicting transactions in an STM. ACM SIGPLAN Notices, 44 (4):163–172, April 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ratanaworabhan:2009:DTA

[2268] Paruj Ratanaworabhan, Martin Burtscher, Darko Kirovski, Benjamin Zorn, Rahul Nagpal, and Karthik Pattabiraman. Detecting and tolerating asymmetric races. ACM SIGPLAN Notices, 44(4): 173–184, April 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Abadi:2009:TMS

[2269] Martín Abadi, Tim Harris, and Mojtaba Mehrara. Transactional memory with strong atomicity using offthe-shelf memory protection hardware. ACM SIGPLAN Notices, 44(4):185– 196, April 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ranganathan:2009:IPP

[2270] Parthasarathy Ranganathan. Industrial perspectives panel. ACM SIG-PLAN Notices, 44(4):197, April 2009.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nicolau:2009:TEP

[2271] Alexandru Nicolau, Guangqiang Li, and Arun Kejariwal. Techniques for efficient placement of synchronization primitives. ACM SIGPLAN Notices, 44 (4):199–208, April 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Son:2009:CDD

[2272] Seung Woo Son, Mahmut Kandemir, Mustafa Karakoy, and Dhruva Chakrabarti. A compiler-directed data prefetching scheme for chip multiprocessors. ACM SIGPLAN Notices, 44 (4):209–218, April 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baskaran:2009:CAD

[2273] Muthu Manikandan Baskaran, Nagavijayalakshmi Vydyanathan, Uday Kumar Reddy Bondhugula, J. Ramanujam, Atanas Rountev, and P. Sadayappan. Compiler-assisted dynamic scheduling for effective parallelization of loop nests on multicore processors. ACM SIGPLAN Notices, 44(4): 219–228, April 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tallent:2009:EPM

[2274] Nathan R. Tallent and John M. Mellor-Crummey. Effective performance measurement and analysis of multithreaded applications. ACM SIGPLAN Notices, 44(4):229–240, April 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kistler:2009:PCA

[2275] Michael Kistler, John Gunnels, Daniel Brokenshire, and Brad Benton. Petascale computing with accelerators. ACM SIGPLAN Notices, 44(4):241–250, April 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xue:2009:MSR

[2276] Ruini Xue, Xuezheng Liu, Ming Wu, Zhenyu Guo, Wenguang Chen, Weimin Zheng, Zheng Zhang, and Geoffrey Voelker. MPIWiz: subgroup reproducible replay of MPI applications. ACM SIGPLAN Notices, 44(4):251– 260, April 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vo:2009:FVP

[2277] Anh Vo, Sarvani Vakkalanka, Michael DeLisi, Ganesh Gopalakrishnan, Robert M. Kirby, and Rajeev Thakur. Formal verification of practical MPI programs. ACM SIGPLAN Notices, 44 (4):261–270, April 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bikshandi:2009:EPI

[2278] Ganesh Bikshandi, Jose G. Castanos, Sreedhar B. Kodali, V. Krishna Nandivada, Igor Peshansky, Vijay A. Saraswat, Sayantan Sur, Pradeep Varma, and Tong Wen. Efficient, portable implementation of asynchronous multi-place programs. ACM SIGPLAN Notices, 44(4):271– 282, April 2009. CODEN SINODQ.

ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dennis:2009:HBP

[2279] Jack B. Dennis. How to build programmable multi-core chips. ACM SIGPLAN Notices, 44(4):283–284, April 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Long:2009:ASC

[2280] Guoping Long, Dongrui Fan, and Junchao Zhang. Architectural support for Cilk computations on many-core architectures. ACM SIGPLAN Notices, 44 (4):285–286, April 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ma:2009:CRS

[2281] Wenjing Ma and Gagan Agrawal. A compiler and runtime system for enabling data mining applications on GPUs. ACM SIGPLAN Notices, 44 (4):287–288, April 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Huang:2009:EGO

[2282] Lei Huang, Deepak Eachempati, Marcus W. Hervey, and Barbara Chapman. Exploiting global optimizations for OpenMP programs in the OpenUH compiler. ACM SIGPLAN Notices, 44 (4):289–290, April 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Volos:2009:NDI

[2283] Haris Volos, Adam Welc, Ali-Reza Adl-Tabatabai, Tatiana Shpeisman, Xinmin Tian, and Ravi Narayanaswamy.

NePalTM: design and implementation of nested parallelism for transactional memory systems. ACM SIGPLAN Notices, 44(4):291–292, April 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kejariwal:2009:PSA

[2284] Arun Kejariwal and Calin Casçaval.

Parallelization spectroscopy: analysis of thread-level parallelism in HPC programs. ACM SIGPLAN Notices, 44 (4):293–294, April 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ha:2009:PRN

[2285] Phuong Hoai Ha, Philippas Tsigas, and Otto J. Anshus. Preliminary results on NB-FEB, a synchronization primitive for parallel programming. ACM SIG-PLAN Notices, 44(4):295–296, April 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dash:2009:STD

[2286] Alokika Dash and Brian Demsky. Software transactional distributed shared memory. ACM SIGPLAN Notices, 44 (4):297–298, April 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yang:2009:SBP

[2287] Ke Yang, Bingsheng He, Qiong Luo, Pedro V. Sander, and Jiaoying Shi. Stack-based parallel recursion on graphics processors. ACM SIG-PLAN Notices, 44(4):299–300, April 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bhatele:2009:TAT

[2288] Abhinav Bhatelé, Eric Bohm, and Laxmikant V. Kalé. Topology aware task mapping techniques: an API and case study. ACM SIGPLAN Notices, 44 (4):301–302, April 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Markstrum:2009:TCR

[2289] Shane A. Markstrum, Robert M. Fuhrer, and Todd D. Millstein. Towards concurrency refactoring for X10. ACM SIGPLAN Notices, 44(4):303– 304, April 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Scott:2009:THR

[2290] Stephen L. Scott, Christian Engelmann, Geoffroy R. Vallée, Thomas Naughton, Anand Tikotekar, George Ostrouchov, Chokchai Leangsuksun, Nichamon Naksinehaboon, Raja Nassar, Mihaela Paun, Frank Mueller, Chao Wang, Arun B. Nagarajan, and Jyothish Varma. A tunable holistic resiliency approach for high-performance computing systems. ACM SIGPLAN Notices, 44(4):305–306, April 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kulkarni:2009:TBT

[2291] Chinmay Eishan Kulkarni, Osman Unsal, Adrián Cristal, Eduard Ayguadé, and Mateo Valero. Turbocharging boosted transactions or: how I learnt to stop worrying and love longer transactions. ACM SIGPLAN Notices, 44 (4):307–308, April 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Alglave:2009:SPA

[2292] Jade Alglave, Anthony Fox, Samin Ishtiaq, Magnus O. Myreen, Susmit Sarkar, Peter Sewell, and Francesco Zappa Nardelli. The semantics of power and ARM multiprocessor machine code (abstract only). ACM SIGPLAN Notices, 44(5):8, May 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ziarek:2009:SWB

[2293] Lukasz Ziarek, Suresh Jagannathan, Matthew Fluet, and Umut A. Acar. Speculative N-way barriers (abstract only). ACM SIGPLAN Notices, 44(5): 8, May 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zain:2009:LPH

[2294] Abdallah Deeb I. Al Zain, Kevin Hammond, Jost Berthold, Phil Trinder, Greg Michaelson, and Mustafa Aswad. Low-pain, high-gain multicore programming in Haskell: coordinating irregular symbolic computations on multicore architectures (abstract only). ACM SIGPLAN Notices, 44(5):8–9, May 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Budimlic:2009:DAM

[2295] Zoran Budimlic, Aparna M. Chandramowlishwaran, Kathleen Knobe, Geoff N. Lowney, Vivek Sarkar, and Leo Treggiari. Declarative aspects of memory management in the concurrent collections parallel programming model (abstract only). ACM SIGPLAN Notices, 44(5):9, May 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sulzmann:2009:CPCa

[2296] Martin Sulzmann, Edmund S. L. Lam, and Simon Marlow. Comparing the performance of concurrent linked-list implementations in Haskell (abstract only). ACM SIGPLAN Notices, 44(5): 9, May 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Herhut:2009:CCS

[2297] Stephan Herhut, Sven-Bodo Scholz, and Clemens Grelck. Controlling chaos: on safe side-effects in data-parallel operations (abstract only). ACM SIG-PLAN Notices, 44(5):9–10, May 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sulzmann:2009:CPCb

[2298] Martin Sulzmann, Edmund S. L. Lam, and Simon Marlow. Comparing the performance of concurrent linked-list implementations in Haskell. ACM SIG-PLAN Notices, 44(5):11–20, May 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Subramanian:2009:DSU

[2299] Suriya Subramanian, Michael Hicks, and Kathryn S. McKinley. Dynamic software updates: a VM-centric approach. ACM SIGPLAN Notices, 44 (6):1–12, June 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Neamtiu:2009:STU

[2300] Iulian Neamtiu and Michael Hicks. Safe and timely updates to multi-threaded programs. ACM SIGPLAN Notices, 44 (6):13–24, June 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hammer:2009:CCB

[2301] Matthew A. Hammer, Umut A. Acar, and Yan Chen. CEAL: a C-based language for self-adjusting computation. ACM SIGPLAN Notices, 44(6):25–37, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ansel:2009:PLC

[2302] Jason Ansel, Cy Chan, Yee Lok Wong, Marek Olszewski, Qin Zhao, Alan Edelman, and Saman Amarasinghe. PetaBricks: a language and compiler for algorithmic choice. ACM SIG-PLAN Notices, 44(6):38–49, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chugh:2009:SIF

[2303] Ravi Chugh, Jeffrey A. Meister, Ranjit Jhala, and Sorin Lerner. Staged information flow for JavaScript. ACM SIGPLAN Notices, 44(6):50–62, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Roy:2009:LPF

[2304] Indrajit Roy, Donald E. Porter, Michael D. Bond, Kathryn S. McKinley, and Emmett Witchel. Laminar: practical fine-grained decentralized information flow control. ACM SIG-PLAN Notices, 44(6):63–74, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Livshits:2009:MSI

[2305] Benjamin Livshits, Aditya V. Nori, Sriram K. Rajamani, and Anindya Banerjee. Merlin: specification inference for explicit information flow problems. ACM SIGPLAN Notices, 44(6):75–86, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tripp:2009:TET

[2306] Omer Tripp, Marco Pistoia, Stephen J. Fink, Manu Sridharan, and Omri Weisman. TAJ: effective taint analysis of web applications. ACM SIG-PLAN Notices, 44(6):87–97, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anderson:2009:LAC

[2307] Zachary R. Anderson, David Gay, and Mayur Naik. Lightweight annotations for controlling sharing in concurrent data structures. ACM SIG-PLAN Notices, 44(6):98–109, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Joshi:2009:RDP

[2308] Pallavi Joshi, Chang-Seo Park, Koushik Sen, and Mayur Naik. A randomized dynamic program analysis technique for detecting real deadlocks. ACM SIG-PLAN Notices, 44(6):110–120, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Flanagan:2009:FEP

[2309] Cormac Flanagan and Stephen N. Freund. FastTrack: efficient and precise dynamic race detection. ACM SIG-PLAN Notices, 44(6):121–133, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marino:2009:LES

[2310] Daniel Marino, Madanlal Musuvathi, and Satish Narayanasamy. LiteRace: effective sampling for lightweight data-race detection. ACM SIGPLAN Notices, 44(6):134–143, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Petrank:2009:PGP

[2311] Erez Petrank, Madanlal Musuvathi, and Bjarne Steensgaard. Progress guarantee for parallel programs via bounded lock-freedom. ACM SIG-PLAN Notices, 44(6):144–154, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dragojevic:2009:STM

[2312] Aleksandar Dragojević, Rachid Guerraoui, and Michal Kapalka. Stretching transactional memory. ACM SIG-PLAN Notices, 44(6):155–165, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mehrara:2009:PSA

[2313] Mojtaba Mehrara, Jeff Hao, Po-Chun Hsu, and Scott Mahlke. Parallelizing sequential applications on commodity hardware using a low-cost software transactional memory. ACM SIG-PLAN Notices, 44(6):166–176, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tournavitis:2009:THA

[2314] Georgios Tournavitis, Zheng Wang, Björn Franke, and Michael F. P. O’Boyle. Towards a holistic approach to auto-parallelization: integrating profile-driven parallelism detection and machine-learning based mapping. ACM SIGPLAN Notices, 44(6): 177–187, June 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hooimeijer:2009:DPS

[2315] Pieter Hooimeijer and Westley Weimer. A decision procedure for subset constraints over regular languages. ACM SIGPLAN Notices, 44(6):188–198, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schwerdfeger:2009:VCD

[2316] August C. Schwerdfeger and Eric R. Van Wyk. Verifiable composition of deterministic grammars. ACM SIG-PLAN Notices, 44(6):199–210, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Torre:2009:ARP

[2317] Salvatore La Torre, Madhusudan Parthasarathy, and Gennaro Parlato. Analyzing recursive programs using a fixed-point calculus. ACM SIG-PLAN Notices, 44(6):211–222, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Srivastava:2009:PVU

[2318] Saurabh Srivastava and Sumit Gulwani. Program verification using templates over predicate abstraction. ACM SIGPLAN Notices, 44(6):223– 234, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Aftandilian:2009:GAU

[2319] Edward E. Aftandilian and Samuel Z.

Guyer. GC assertions: using the garbage collector to check heap properties. ACM SIGPLAN Notices, 44 (6):235–244, June 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nagarakatte:2009:SHC

[2320] Santosh Nagarakatte, Jianzhou Zhao, Milo M. K. Martin, and Steve Zdancewic. SoftBound: highly compatible and complete spatial memory safety for C. ACM SIGPLAN Notices, 44(6):245–258, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Oiwa:2009:IMS

[2321] Yutaka Oiwa. Implementation of the memory-safe full ANSI-C compiler. ACM SIGPLAN Notices, 44(6): 259–269, June 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rubio-Gonzalez:2009:EPA

[2322] Cindy Rubio-González, Haryadi S. Gunawi, Ben Liblit, Remzi H. Arpaci-Dusseau, and Andrea C. Arpaci-Dusseau. Error propagation analysis for file systems. ACM SIG-PLAN Notices, 44(6):270–280, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Qi:2009:SCB

[2323] Xin Qi and Andrew C. Myers. Sharing classes between families. ACM SIG-PLAN Notices, 44(6):281–292, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rendel:2009:TSR

[2324] Tillman Rendel, Klaus Ostermann, and Christian Hofer. Typed selfrepresentation. ACM SIGPLAN Notices, 44(6):293–303, June 2009. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kawaguchi:2009:TBD

[2325] Ming Kawaguchi, Patrick Rondon, and Ranjit Jhala. Type-based data structure verification. ACM SIGPLAN Notices, 44(6):304–315, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tristan:2009:VVL

[2326] Jean-Baptiste Tristan and Xavier Leroy. Verified validation of lazy code motion. ACM SIGPLAN Notices, 44 (6):316–326, June 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kundu:2009:POC

[2327] Sudipta Kundu, Zachary Tatlock, and Sorin Lerner. Proving optimizations correct using parameterized program equivalence. ACM SIGPLAN Notices, 44(6):327–337, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zee:2009:IPL

[2328] Karen Zee, Viktor Kuncak, and Martin C. Rinard. An integrated proof language for imperative programs. ACM SIGPLAN Notices, 44(6):338– 351, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ravitch:2009:AGL

[2329] Tristan Ravitch, Steve Jackson, Eric Aderhold, and Ben Liblit. Automatic generation of library bindings using static analysis. ACM SIG-PLAN Notices, 44(6):352–362, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chandra:2009:SPA

[2330] Satish Chandra, Stephen J. Fink, and Manu Sridharan. Snugglebug: a powerful approach to weakest preconditions. ACM SIGPLAN Notices, 44(6): 363–374, June 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gulwani:2009:CFR

[2331] Sumit Gulwani, Sagar Jain, and Eric Koskinen. Control-flow refinement and progress invariants for bound analysis. ACM SIGPLAN Notices, 44(6): 375–385, June 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Inoue:2009:SMM

[2332] Hiroshi Inoue, Hideaki Komatsu, and Toshio Nakatani. A study of memory management for Web-based applications on multicore processors. ACM SIGPLAN Notices, 44(6):386– 396, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Novark:2009:EPL

[2333] Gene Novark, Emery D. Berger, and Benjamin G. Zorn. Efficiently and precisely locating memory leaks and bloat. ACM SIGPLAN Notices, 44(6): 397–407, June 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shacham:2009:CAS

[2334] Ohad Shacham, Martin Vechev, and Eran Yahav. Chameleon: adaptive selection of collections. ACM SIG-PLAN Notices, 44(6):408–418, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xu:2009:GFP

[2335] Guoqing Xu, Matthew Arnold, Nick Mitchell, Atanas Rountev, and Gary Sevitsky. Go with the flow: profiling copies to find runtime bloat. ACM SIG-PLAN Notices, 44(6):419–430, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Saha:2009:PMH

[2336] Bratin Saha, Xiaocheng Zhou, Hu Chen, Ying Gao, Shoumeng Yan, Mohan Rajagopalan, Jesse Fang, Peinan Zhang, Ronny Ronen, and Avi Mendelson. Programming model for a heterogeneous x86 platform. ACM SIGPLAN Notices, 44(6):431–440, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tallent:2009:BAM

[2337] Nathan R. Tallent, John M. Mellor-Crummey, and Michael W. Fagan. Binary analysis for measurement and attribution of program performance. ACM SIGPLAN Notices, 44(6):441– 452, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hoffman:2009:SAT

[2338] Kevin J. Hoffman, Patrick Eugster, and Suresh Jagannathan. Semanticsaware trace analysis. ACM SIG-PLAN Notices, 44(6):453–464, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gal:2009:TBJ

[2339] Andreas Gal, Brendan Eich, Mike Shaver, David Anderson, David Mandelin, Mohammad R. Haghighat, Blake Kaplan, Graydon Hoare, Boris Zbarsky, Jason Orendorff, Jesse Ruderman, Edwin W. Smith, Rick Reitmaier, Michael Bebenita, Mason Chang, and Michael Franz. Trace-based just-in-time type specialization for dynamic languages. ACM SIGPLAN Notices, 44 (6):465–478, June 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stotzer:2009:MSO

[2340] Eric J. Stotzer and Ernst L. Leiss. Modulo scheduling without overlapped lifetimes. ACM SIGPLAN Notices, 44 (7):1–10, July 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Caspi:2009:SOS

[2341] Paul Caspi, Jean-Louis Colaço, Léonard Gérard, Marc Pouzet, and Pascal Raymond. Synchronous objects with scheduling policies: introducing safe shared memory in Lustre. ACM SIG-PLAN Notices, 44(7):11–20, July 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Oh:2009:RCA

[2342] Taewook Oh, Bernhard Egger, Hyunchul Park, and Scott Mahlke. Recurrence cycle aware modulo scheduling for coarse-grained reconfigurable architectures. ACM SIGPLAN Notices, 44 (7):21–30, July 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zou:2009:PFT

[2343] Jia Zou, Joshua Auerbach, David F.

Bacon, and Edward A. Lee. PTIDES on flexible task graph: real-time embedded system building from theory to practice. ACM SIGPLAN Notices, 44 (7):31–40, July 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lee:2009:COR

[2344] Jongeun Lee and Aviral Shrivastava.

A compiler optimization to reduce soft errors in register files. ACM SIG-PLAN Notices, 44(7):41–49, July 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Leather:2009:RPE

[2345] Hugh Leather, Michael O’Boyle, and Bruce Worton. Raced profiles: efficient selection of competing compiler optimizations. ACM SIGPLAN Notices, 44(7):50–59, July 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yang:2009:ECS

[2346] Xuejun Yang, Nathan Cooprider, and John Regehr. Eliminating the call stack to save RAM. ACM SIGPLAN Notices, 44(7):60–69, July 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Blazy:2009:LRU

[2347] Sandrine Blazy and Benoit Robillard.

Live-range unsplitting for faster optimal coalescing. ACM SIGPLAN Notices, 44(7):70–79, July 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sarkar:2009:PAM

[2348] Abhik Sarkar, Frank Mueller, Harini Ramaprasad, and Sibin Mohan. Pushassisted migration of real-time tasks in multi-core processors. ACM SIG-PLAN Notices, 44(7):80–89, July 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mankin:2009:STM

[2349] Jennifer Mankin, David Kaeli, and John Ardini. Software transactional memory for multicore embedded systems. ACM SIGPLAN Notices, 44 (7):90–98, July 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Udupa:2009:SES

[2350] Abhishek Udupa, R. Govindarajan, and Matthew J. Thazhuthaveetil. Synergistic execution of stream programs on multicores with accelerators. ACM SIGPLAN Notices, 44(7):99–108, July 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Heinz:2009:TDE

[2351] Thomas Heinz and Reinhard Wilhelm. Towards device emulation code generation. ACM SIGPLAN Notices, 44 (7):109–118, July 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hines:2009:GIF

[2352] Stephen Roderick Hines, Yuval Peress, Peter Gavin, David Whalley, and Gary Tyson. Guaranteeing instruction fetch behavior with a Lookahead Instruction Fetch Engine (LIFE). ACM SIGPLAN Notices, 44(7):119–128, July 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McKechnie:2009:DFB

[2353] Paul E. McKechnie, Michaela Blott, and Wim A. Vanderbauwhede. Debugging FPGA-based packet processing systems through transactionlevel communication-centric monitoring. ACM SIGPLAN Notices, 44(7): 129–136, July 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gracioli:2009:TIE

[2354] Giovani Gracioli and Sebastian Fischmeister. Tracing interrupts in embedded software. ACM SIGPLAN Notices, 44(7):137–146, July 2009. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Moore:2009:ACD

[2355] Ryan W. Moore, José A. Baiocchi, Bruce R. Childers, Jack W. Davidson, and Jason D. Hiser. Addressing the challenges of DBT for the ARM architecture. ACM SIGPLAN Notices, 44 (7):147–156, July 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fidge:2009:IHS

[2356] Colin J. Fidge and Diane Corney. Integrating hardware and software information flow analyses. ACM SIG-PLAN Notices, 44(7):157–166, July 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Andre:2009:SVT

[2357] Charles André and Frédéric Mallet. Specification and verification of time requirements with CCSL and Esterel. ACM SIGPLAN Notices, 44(7):167– 176, July 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chaudhuri:2009:LBS

[2358] Avik Chaudhuri. Language-based security on Android (abstract only). ACM SIGPLAN Notices, 44(8):4, August 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

DeVries:2009:ABV

[2359] Brian W. DeVries, Gopal Gupta, Kevin W. Hamlen, Scott Moore, and Meera Sridhar. ActionScript bytecode verification with co-logic programming (abstract only). ACM SIG-PLAN Notices, 44(8):4, August 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jia:2009:EIFa

[2360] Limin Jia and Steve Zdancewic. Encoding information flow in AURA (abstract only). ACM SIGPLAN Notices, 44(8):4, August 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wasserrab:2009:PBN

[2361] Daniel Wasserrab, Denis Lohner, and Gregor Snelting. On PDG-based noninterference and its modular proof (abstract only). ACM SIGPLAN Notices, 44(8):4–5, August 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Askarov:2009:CMI

[2362] Aslan Askarov and Andrei Sabelfeld.

Catch me if you can: permissive yet secure error handling (abstract only). ACM SIGPLAN Notices, 44(8):5, August 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Balliu:2009:WPA

[2363] Musard Balliu and Isabella Mastroeni.

A weakest precondition approach to active attacks analysis (abstract only). ACM SIGPLAN Notices, 44(8):5, August 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Newsome:2009:MCC

[2364] James Newsome, Stephen McCamant, and Dawn Song. Measuring channel capacity to distinguish undue influence (abstract only). ACM SIG-PLAN Notices, 44(8):5, August 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Austin:2009:EPDa

[2365] Thomas H. Austin and Cormac Flanagan. Efficient purely-dynamic information flow analysis (abstract only). ACM SIGPLAN Notices, 44(8):6, August 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Birgisson:2009:IST

[2366] Arnar Birgisson and Úlfar Erlingsson. An implementation and semantics for transactional memory introspection in Haskell (abstract only). ACM SIG-PLAN Notices, 44(8):6, August 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Broberg:2009:FSS

[2367] Niklas Broberg and David Sands. Flowsensitive semantics for dynamic information flow policies (abstract only). ACM SIGPLAN Notices, 44(8):6, August 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shinnar:2009:LIF

[2368] Avraham Shinnar, Marco Pistoia, and Anindya Banerjee. A language for information flow: dynamic tracking in multiple interdependent dimensions (abstract only). ACM SIGPLAN Notices, 44(8):6, August 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jia:2009:EIFb

[2369] Limin Jia and Steve Zdancewic. Encoding information flow in AURA. ACM SIGPLAN Notices, 44(8):7–19, August 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Austin:2009:EPDb

[2370] Thomas H. Austin and Cormac Flanagan. Efficient purely-dynamic information flow analysis. ACM SIG-PLAN Notices, 44(8):20–31, August 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Steele:2009:OFC

[2371] Guy L. Steele, Jr. Organizing functional code for parallel execution or, foldl and foldr considered slightly harmful. ACM SIGPLAN Notices, 44 (9):1–2, September 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL https://dl.acm.org/
citation.cfm?id=1596551.

Hinze:2009:FPT

[2372] Ralf Hinze. Functional pearl: la tour d’Hanoï. ACM SIGPLAN Notices, 44 (9):3–10, September 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL https://dl.acm.org/
citation.cfm?id=1596555.

Fischer:2009:PFL

[2373] Sebastian Fischer, Oleg Kiselyov, and Chung chieh Shan. Purely functional lazy non-deterministic programming. ACM SIGPLAN Notices, 44 (9):11–22, September 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL https://dl.acm.org/
citation.cfm?id=1596556.

Sculthorpe:2009:SFR

[2374] Neil Sculthorpe and Henrik Nilsson.

Safe functional reactive programming through dependent types. ACM SIG-PLAN Notices, 44(9):23–34, September 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.

acm.org/citation.cfm?id=1596558.

Liu:2009:CCA

[2375] Hai Liu, Eric Cheng, and Paul Hudak. Causal commutative arrows and their optimization. ACM SIGPLAN Notices, 44(9):35–46, September 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.acm.

org/citation.cfm?id=1596559.

Felleisen:2009:FSF

[2376] Matthias Felleisen, Robert Bruce Findler, Matthew Flatt, and Shriram Krishnamurthi. A functional I/O system or, fun for freshman kids. ACM SIG-PLAN Notices, 44(9):47–58, September 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.

acm.org/citation.cfm?id=1596561.

Newton:2009:ERE

[2377] Ryan R. Newton and Teresa Ko. Experience report: embedded, parallel computer-vision with a functional DSL. ACM SIGPLAN Notices, 44 (9):59–64, September 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL https://dl.acm.org/
citation.cfm?id=1596562.

Marlow:2009:RSM

[2378] Simon Marlow, Simon Peyton Jones, and Satnam Singh. Runtime support for multicore Haskell. ACM SIGPLAN Notices, 44(9):65–78, September 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.acm.

org/citation.cfm?id=1596563.

Chlipala:2009:EIP

[2379] Adam Chlipala, Gregory Malecha, Greg Morrisett, Avraham Shinnar, and Ryan Wisnesky. Effective interactive proofs for higher-order imperative programs. ACM SIGPLAN Notices, 44(9):79–90, September 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.acm.

org/citation.cfm?id=1596565.

Klein:2009:ERS

[2380] Gerwin Klein, Philip Derrin, and Kevin Elphinstone. Experience report: seL4: formally verifying a high-performance microkernel. ACM SIGPLAN Notices, 44(9):91–96, September 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.acm.

org/citation.cfm?id=1596566.

Benton:2009:BSI

[2381] Nick Benton and Chung-Kil Hur.

Biorthogonality, step-indexing and compiler correctness. ACM SIGPLAN Notices, 44(9):97–108, September 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.acm.

org/citation.cfm?id=1596567.

Flatt:2009:SCB

[2382] Matthew Flatt, Eli Barzilay, and Robert Bruce Findler. Scribble: closing the book on ad hoc documentation tools. ACM SIGPLAN Notices, 44(9):109–120, September 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.acm.

org/citation.cfm?id=1596569.

Pierce:2009:LUU

[2383] Benjamin C. Pierce. Lambda, the ultimate TA: using a proof assistant to teach programming language foundations. ACM SIGPLAN Notices, 44(9):121–122, September 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.acm.

                                        org/citation.cfm?id=1596552.

Licata:2009:UBC

[2384] Daniel R. Licata and Robert Harper. A universe of binding and computation. ACM SIGPLAN Notices, 44(9): 123–134, September 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL https://dl.acm.org/
                                        citation.cfm?id=1596571.

Neis:2009:NPP

[2385] Georg Neis, Derek Dreyer, and Andreas Rossberg. Non-parametric parametricity. ACM SIGPLAN Notices, 44 (9):135–148, September 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL https://dl.acm.org/
                                        citation.cfm?id=1596572.

Claessen:2009:FRC

[2386] Koen Claessen, Michal Palka, Nicholas Smallbone, John Hughes, Hans Svensson, Thomas Arts, and Ulf Wiger. Finding race conditions in Erlang with QuickCheck and PULSE. ACM SIGPLAN Notices, 44(9):149–160, September 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL

                                        https://dl.acm.org/citation.cfm?
                                        id=1596574.

Ziarek:2009:PMC

[2387] Lukasz Ziarek, KC Sivaramakrishnan, and Suresh Jagannathan. Partial memoization of concurrency and communication. ACM SIGPLAN Notices, 44(9):161–172, September 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.acm.

org/citation.cfm?id=1596575.

VoigtlAnder:2009:FTI

[2388] Janis Voigtländer. Free theorems involving type constructor classes: functional pearl. ACM SIGPLAN Notices, 44(9):173–184, September 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.acm.

org/citation.cfm?id=1596577.

Sampson:2009:ERH

[2389] Curt J. Sampson. Experience report:

Haskell in the ‘real world’: writing a commercial application in a lazy functional language. ACM SIGPLAN Notices, 44(9):185–190, September 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.acm.

org/citation.cfm?id=1596578.

Elliott:2009:BD

[2390] Conal M. Elliott. Beautiful differentiation. ACM SIGPLAN Notices, 44 (9):191–202, September 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL https://dl.acm.org/
citation.cfm?id=1596579.

Gazagnaire:2009:OEH

[2391] Thomas Gazagnaire and Vincent Hanquez. OXenstored: an efficient hierarchical and transactional database using functional programming with reference cell comparisons. ACM SIGPLAN Notices, 44(9):203–214, September 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.acm.

org/citation.cfm?id=1596581.

Pagano:2009:ERU

[2392] Bruno Pagano, Olivier Andrieu, Thomas Moniot, Benjamin Canou, Emmanuel Chailloux, Philippe Wang, Pascal Manoury, and Jean-Louis Colaço. Experience report: using objective caml to develop safety-critical embedded tools in a certification framework. ACM SIGPLAN Notices, 44(9): 215–220, September 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL https://dl.acm.org/
citation.cfm?id=1596582.

GenevAs:2009:IQI

[2393] Pierre Genevès, Nabil Layaïda, and Vincent Quint. Identifying query incompatibilities with evolving XML schemas. ACM SIGPLAN Notices, 44 (9):221–230, September 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL https://dl.acm.org/
citation.cfm?id=1596583.

Piponi:2009:CMD

[2394] Dan P. Piponi. Commutative monads, diagrams and knots. ACM SIGPLAN Notices, 44(9):231–232, September 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.

acm.org/citation.cfm?id=1596553.

Yakushev:2009:GPF

[2395] Alexey Rodriguez Yakushev, Stefan Holdermans, Andres Löh, and Johan Jeuring. Generic programming with fixed points for mutually recursive datatypes. ACM SIGPLAN Notices, 44(9):233–244, September 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.acm.

org/citation.cfm?id=1596585.

Viera:2009:AGF

[2396] Marcos Viera, S. Doaitse Swierstra, and Wouter Swierstra. Attribute grammars fly first-class: how to do aspect oriented programming in Haskell. ACM SIGPLAN Notices, 44 (9):245–256, September 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL https://dl.acm.org/
citation.cfm?id=1596586.

Reppy:2009:PCM

[2397] John Reppy, Claudio V. Russo, and Yingqi Xiao. Parallel concurrent ML. ACM SIGPLAN Notices, 44(9): 257–268, September 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL https://dl.acm.org/
citation.cfm?id=1596588.

Chaudhuri:2009:CML

[2398] Avik Chaudhuri. A concurrent ML library in concurrent Haskell. ACM SIGPLAN Notices, 44(9):269–280, September 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL

https://dl.acm.org/citation.cfm?
id=1596589.

Cuoq:2009:ERO

[2399] Pascal Cuoq, Julien Signoles, Patrick Baudin, Richard Bonichon, Géraud Canet, Loïc Correnson, Benjamin Monate, Virgile Prevosto, and Armand Puccetti. Experience report: OCaml for an industrial-strength static analysis framework. ACM SIGPLAN Notices, 44(9):281–286, September 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.acm.

org/citation.cfm?id=1596591.

Midtgaard:2009:CFA

[2400] Jan Midtgaard and Thomas P. Jensen. Control-flow analysis of function calls and returns by abstract interpretation. ACM SIGPLAN Notices, 44(9): 287–298, September 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL https://dl.acm.org/
citation.cfm?id=1596592.

McCarthy:2009:ARW

[2401] Jay A. McCarthy. Automatically RESTful web applications: marking modular serializable continuations. ACM SIGPLAN Notices, 44(9): 299–310, September 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL https://dl.acm.org/
citation.cfm?id=1596594.

Balat:2009:ERO

[2402] Vincent Balat, Jérôme Vouillon, and Boris Yakobowski. Experience report: ocsigen, a web programming framework. ACM SIGPLAN Notices, 44(9):311–316, September 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.acm.

org/citation.cfm?id=1596595.

Rompf:2009:IFC

[2403] Tiark Rompf, Ingo Maier, and Martin Odersky. Implementing firstclass polymorphic delimited continuations by a type-directed selective CPS-transform. ACM SIGPLAN Notices, 44(9):317–328, September 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.acm.

org/citation.cfm?id=1596596.

Swamy:2009:TTC

[2404] Nikhil Swamy, Michael Hicks, and Gavin M. Bierman. A theory of typed coercions and its applications. ACM SIGPLAN Notices, 44(9):329–340, September 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL

https://dl.acm.org/citation.cfm?
id=1596598.

Schrijvers:2009:CDT

[2405] Tom Schrijvers, Simon Peyton Jones, Martin Sulzmann, and Dimitrios Vytiniotis. Complete and decidable type inference for GADTs. ACM SIGPLAN Notices, 44(9):341–352, September 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.

                                        acm.org/citation.cfm?id=1596599.

Meyerovich:2009:FPL

[2406] Leo A. Meyerovich, Arjun Guha, Jacob Baskin, Gregory H. Cooper, Michael Greenberg, Aleks Bromfield, and Shriram Krishnamurthi. Flapjax: a programming language for Ajax applications. ACM SIGPLAN Notices, 44(10): 1–20, October 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Malayeri:2009:CMI

[2407] Donna Malayeri and Jonathan Aldrich. CZ: multiple inheritance without diamonds. ACM SIGPLAN Notices, 44 (10):21–40, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ducournau:2009:EAO

[2408] Roland Ducournau, Floréal Morandat, and Jean Privat. Empirical assessment of object-oriented implementations with multiple inheritance and static typing. ACM SIGPLAN Notices, 44(10):41–60, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lublinerman:2009:PPO

[2409] Roberto Lublinerman, Swarat Chaudhuri, and Pavol Cerny. Parallel programming with object assemblies. ACM SIGPLAN Notices, 44(10):61– 80, October 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berger:2009:GSM

[2410] Emery D. Berger, Ting Yang, Tongping Liu, and Gene Novark. Grace: safe multithreaded programming for C/ C++. ACM SIGPLAN Notices, 44(10): 81–96, October 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bocchino:2009:TES

[2411] Robert L. Bocchino, Jr., Vikram S.

Adve, Danny Dig, Sarita V. Adve, Stephen Heumann, Rakesh Komuravelli, Jeffrey Overbey, Patrick Simmons, Hyojin Sung, and Mohsen Vakilian. A type and effect system for deterministic parallel Java. ACM SIGPLAN Notices, 44(10):97–116, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bloom:2009:TRC

[2412] Bard Bloom, John Field, Nathaniel Nystrom, Johan Östlund, Gregor Richards, Rok Strniša, Jan Vitek, and Tobias Wrigstad. Thorn: robust, concurrent, extensible scripting on the JVM. ACM SIGPLAN Notices, 44(10): 117–136, October 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Inoue:2009:HJV

[2413] Hiroshi Inoue and Toshio Nakatani.

How a Java VM can get more from a hardware performance monitor. ACM SIGPLAN Notices, 44(10): 137–154, October 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ha:2009:CDA

[2414] Jungwoo Ha, Matthew Arnold, Stephen M. Blackburn, and Kathryn S. McKinley. A concurrent dynamic analysis framework for multicore hardware. ACM SIGPLAN Notices, 44(10):155– 174, October 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mytkowicz:2009:ICP

[2415] Todd Mytkowicz, Devin Coughlin, and Amer Diwan. Inferred call path profiling. ACM SIGPLAN Notices, 44(10): 175–190, October 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Charles:2009:ACC

[2416] Philippe Charles, Robert M. Fuhrer, Stanley M. Sutton, Jr., Evelyn Duesterwald, and Jurgen Vinju. Accelerating the creation of customized, language-specific IDEs in Eclipse. ACM SIGPLAN Notices, 44(10):191– 206, October 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lee:2009:DAY

[2417] Byeongcheol Lee, Martin Hirzel, Robert Grimm, and Kathryn S. McKinley. Debug all your code: portable mixed-environment debugging. ACM SIGPLAN Notices, 44(10): 207–226, October 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Leijen:2009:DTP

[2418] Daan Leijen, Wolfram Schulte, and Sebastian Burckhardt. The design of a task parallel library. ACM SIGPLAN Notices, 44(10):227–242, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bravenboer:2009:SDS

[2419] Martin Bravenboer and Yannis Smaragdakis. Strictly declarative specification of sophisticated points-to analyses. ACM SIGPLAN Notices, 44(10): 243–262, October 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Saito:2009:STC

[2420] Chieri Saito and Atsushi Igarashi. Self type constructors. ACM SIGPLAN Notices, 44(10):263–282, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Furr:2009:PGS

[2421] Michael Furr, Jong hoon (David) An, and Jeffrey S. Foster. Profile-guided static typing for dynamic scripting languages. ACM SIGPLAN Notices, 44 (10):283–300, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Song:2009:ESL

[2422] Myoungkyu Song and Eli Tilevich.

Enhancing source-level programming tools with an awareness of transparent program transformations. ACM SIG-PLAN Notices, 44(10):301–320, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Abi-Antoun:2009:SEC

[2423] Marwan Abi-Antoun and Jonathan Aldrich. Static extraction and conformance analysis of hierarchical runtime architectural structure using annotations. ACM SIGPLAN Notices, 44(10): 321–340, October 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nanda:2009:DMT

[2424] Mangala Gowri Nanda, Senthil Mani, Vibha Singhal Sinha, and Saurabh Sinha. Demystifying model transformations: an approach based on automated rule inference. ACM SIGPLAN Notices, 44(10):341–360, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhao:2009:AWL

[2425] Yi Zhao, Jin Shi, Kai Zheng, Haichuan Wang, Haibo Lin, and Ling Shao. Allocation wall: a limiting factor of Java applications on emerging multicore platforms. ACM SIGPLAN Notices, 44(10):361–376, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ogasawara:2009:NAM

[2426] Takeshi Ogasawara. NUMA-aware memory manager with dominantthread-based copying GC. ACM SIG-PLAN Notices, 44(10):377–390, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pluquet:2009:ECP

[2427] Frédéric Pluquet, Stefan Langerman, and Roel Wuyts. Executing code in the past: efficient in-memory object graph versioning. ACM SIGPLAN Notices, 44(10):391–408, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vonDincklage:2009:OPI

[2428] Daniel von Dincklage and Amer Diwan. Optimizing programs with intended semantics. ACM SIGPLAN Notices, 44(10):409–424, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tsafrir:2009:MDW

[2429] Dan Tsafrir, Robert W. Wisniewski, David F. Bacon, and Bjarne Stroustrup. Minimizing dependencies within generic classes for faster and smaller programs. ACM SIGPLAN Notices, 44 (10):425–444, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kats:2009:PRF

[2430] Lennart C. L. Kats, Maartje de Jonge, Emma Nilsson-Nyman, and Eelco Visser. Providing rapid feedback in generated modular language environments: adding error recovery to scannerless generalized-LR parsing. ACM SIGPLAN Notices, 44(10):445–464, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Riehle:2009:DPD

[2431] Dirk Riehle. Design pattern density defined. ACM SIGPLAN Notices, 44(10): 469–480, October 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mullen:2009:WCO

[2432] Thomas Mullen. Writing code for other people: cognitive psychology and the fundamentals of good software design principles. ACM SIGPLAN Notices, 44(10):481–492, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Overbey:2009:RLR

[2433] Jeffrey L. Overbey and Ralph E. Johnson. Regrowing a language: refactoring tools allow programming languages to evolve. ACM SIGPLAN Notices, 44(10):493–502, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Knoll:2009:PL

[2434] Roman Knöll and Mira Mezini. π: a pattern language. ACM SIGPLAN Notices, 44(10):503–522, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mahoney:2009:SEM

[2435] Mark Mahoney. Software evolution and the moving picture metaphor. ACM SIGPLAN Notices, 44(10):525– 528, October 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Quillien:2009:AAA

[2436] Jenny Quillien, Pam Rostal, and Dave West. Agile anthropology and Alexander’s architecture: an essay in three voices. ACM SIGPLAN Notices, 44 (10):529–546, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baniassad:2009:EPL

[2437] Elisa Baniassad and Clayton Myers.

An exploration of program as language. ACM SIGPLAN Notices, 44(10):547– 556, October 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cook:2009:UDA

[2438] William R. Cook. On understanding data abstraction, revisited. ACM SIG-PLAN Notices, 44(10):557–572, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Allen:2009:OOP

[2439] Jim Allen, Zena Ariola, Pierre-Louis Curien, Matthew Fluet, Jeff Foster, Dan Grossman, Robert Harper, Hugo Herbelin, Yannis Smaragdakis, David Walker, and Steve Zdancewic. An overview of the Oregon programming languages summer school. ACM SIG-PLAN Notices, 44(11):1–3, November 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Danish:2009:OSD

[2440] Matthew Danish and Hongwei Xi. Operating system development with ATS: work in progress: (abstract only). ACM SIGPLAN Notices, 44(11):6, November 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Monnier:2009:STH

[2441] Stefan Monnier and David Haguenauer. Singleton types here, singleton types there, singleton types everywhere: (abstract only). ACM SIG-PLAN Notices, 44(11):6, November 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Balzer:2009:MRA

[2442] Stephanie Balzer and Thomas R. Gross. Modular reasoning about invariants over shared state with interposed data members: (abstract only). ACM SIGPLAN Notices, 44(11):6–7, November 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Christiansen:2009:FTF

[2443] Jan Christiansen, Daniel Seidel, and Janis Voigtländer. Free theorems for functional logic programs: (abstract only). ACM SIGPLAN Notices, 44 (11):7, November 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stump:2009:RTG

[2444] Aaron Stump and Evan Austin. Resource typing in Guru: (abstract only). ACM SIGPLAN Notices, 44 (11):7, November 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Weirich:2009:AGD

[2445] Stephanie Weirich and Chris Casinghino. Arity-generic datatype-generic programming: (abstract only). ACM SIGPLAN Notices, 44(11):7, November 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kalibera:2009:CBV

[2446] Tomas Kalibera, Pavel Parizek, Ghaith Haddad, Gary T. Leavens, and Jan Vitek. Challenge benchmarks for verification of real-time programs: (abstract only). ACM SIGPLAN Notices, 44(11): 7–8, November 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vitek:2009:SPT

[2447] Jan Vitek. Of scripts and programs: tall tales, urban legends, and future prospects. ACM SIGPLAN Notices, 44(12):1–2, December 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tanter:2009:BSD

[2448] Éric Tanter. Beyond static and dynamic scope. ACM SIGPLAN Notices, 44(12):3–14, December 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Axelsen:2009:GPT

[2449] Eyvind W. Axelsen and Stein Krogdahl. Groovy package templates: supporting reuse and runtime adaption of class hierarchies. ACM SIGPLAN Notices, 44(12):15–26, December 2009.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Strickland:2009:CFC

[2450] T. Stephen Strickland and Matthias Felleisen. Contracts for first-class modules. ACM SIGPLAN Notices, 44(12): 27–38, December 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Riehl:2009:LEO

[2451] Jonathan Riehl. Language embedding and optimization in Mython. ACM SIGPLAN Notices, 44(12):39–48, December 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Khoo:2009:DJA

[2452] Yit Phang Khoo, Michael Hicks, Jeffrey S. Foster, and Vibha Sazawal. Directing JavaScript with arrows. ACM SIGPLAN Notices, 44(12):49–58, December 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Costanza:2009:COS

[2453] Pascal Costanza, Charlotte Herzeel, and Theo D’Hondt. Context-oriented software transactional memory in Common Lisp. ACM SIGPLAN Notices, 44(12):59–68, December 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pluquet:2009:FTR

[2454] Frédéric Pluquet, Antoine Marot, and Roel Wuyts. Fast type reconstruction for dynamically typed programming languages. ACM SIGPLAN Notices, 44(12):69–78, December 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yermolovich:2009:ODL

[2455] Alexander Yermolovich, Christian Wimmer, and Michael Franz. Optimization of dynamic languages using hierarchical layering of virtual machines. ACM SIGPLAN Notices, 44 (12):79–88, December 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Furr:2009:RIL

[2456] Michael Furr, Jong hoon (David) An, Jeffrey S. Foster, and Michael Hicks. The Ruby intermediate language. ACM SIGPLAN Notices, 44(12):89–98, December 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ungar:2009:HOH

[2457] David Ungar and Sam S. Adams. Hosting an object heap on manycore hardware: an exploration. ACM SIG-PLAN Notices, 44(12):99–110, December 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Griswold:2001:MSP

[2458] William G. Griswold. Making slicing practical: the final mile. In ACM [2481], page 1. ISBN 1-58113-413-4.

LCCN ???? URL http://www.acm.
org/pubs/articles/proceedings/soft/
379605/p1-griswold/p1-griswold.
pdf;   http://www.acm.org/pubs/

                                        citations/proceedings/soft/379605/

p1-griswold/. Invited talk.

Reiss:2001:OB

[2459] Steven P. Reiss. An overview of BLOOM. In ACM [2481], pages 2–5. ISBN 1-58113-413-4. LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/soft/
                                        379605/p2-reiss/p2-reiss.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/soft/379605/p2-reiss/

. Supplement to ACM SIGPLAN Notices.

Agrawal:2001:EEC

[2460] Gagan Agrawal and Liang Guo. Evaluating explicitly context-sensitive program slicing. In ACM [2481], pages 6–12. ISBN 1-58113-413-4. LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/soft/
                                        379605/p6-agrawal/p6-agrawal.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/soft/379605/p6-agrawal/

. Supplement to ACM SIGPLAN Notices.

Souter:2001:CDU

[2461] Amie L. Souter and Lori L. Pollock. Contextual def-use associations for object aggregation. In ACM [2481], pages 13–19. ISBN 1-58113-413-4. LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/soft/
                                        379605/p13-souter/p13-souter.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/soft/379605/p13-souter/

. Supplement to ACM SIGPLAN Notices.

Palsberg:2001:TBA

[2462] Jens Palsberg. Type-based analysis and applications. In ACM [2481], pages 20–27. ISBN 1-58113-413-4. LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/soft/
379605/p20-palsberg/p20-palsberg.
pdf;   http://www.acm.org/pubs/
citations/proceedings/soft/379605/
p20-palsberg/. Invited talk.

Kadayif:2001:VVE

[2463] I. Kadayif, T. Chinoda, M. Kandemir, N. Vijaykirsnan, M. J. Irwin, and A. Sivasubramaniam. vEC: virtual energy counters. In ACM [2481], pages 28–31. ISBN 1-58113-413-4. LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/soft/
379605/p28-kadayif/p28-kadayif.
pdf;   http://www.acm.org/pubs/
citations/proceedings/soft/379605/

p28-kadayif/. Supplement to ACM SIGPLAN Notices.

Ancourt:2001:ARS

[2464] Corinne Ancourt and Thi Viet Nga Nguyen. Array resizing for scientific code debugging, maintenance and reuse. In ACM [2481], pages 32– 37. ISBN 1-58113-413-4. LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/soft/
379605/p32-ancourt/p32-ancourt.
pdf;   http://www.acm.org/pubs/
citations/proceedings/soft/379605/

p32-ancourt/. Supplement to ACM SIGPLAN Notices.

Braband:2001:SVD

[2465] Claus Braband, Anders Møller, and Michael Schwartzbach. Static validation of dynamically generated HTML. In ACM [2481], pages 38– 45. ISBN 1-58113-413-4. LCCN

????  URL http://www.acm.org/
pubs/citations/proceedings/soft/

                                        379605/p38-braband/. Supplement to

ACM SIGPLAN Notices.

Ryder:2001:CIA

[2466] Barbara G. Ryder and Frank Tip. Change impact analysis for objectoriented programs. In ACM [2481], pages 46–53. ISBN 1-58113-413-4.

                                        LCCN ???? URL http://www.acm.
                                        org/pubs/articles/proceedings/soft/
                                        379605/p46-ryder/p46-ryder.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/soft/379605/p46-ryder/

. Supplement to ACM SIGPLAN Notices.

Hind:2001:PAH

[2467] Michael Hind. Pointer analysis: haven’t we solved this problem yet? In ACM [2481], pages 54– 61. ISBN 1-58113-413-4. LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/soft/
                                        379605/p54-hind/p54-hind.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/soft/379605/p54-hind/

. Invited talk.

Maessen:2001:PAS

[2468] Jan-Willem Maessen, Vivek Sarkar, and David Grove. Program analysis for safety guarantees in a Java virtual machine written in Java. In ACM [2481], pages 62– 65. ISBN 1-58113-413-4. LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/soft/
                                        379605/p62-maessen/p62-maessen.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/soft/379605/

p62-maessen/. Supplement to ACM SIGPLAN Notices.

Mock:2001:DPS

[2469] Markus Mock, Manuvir Das, Craig Chambers, and Susan J. Eggers. Dynamic points-to sets: a comparison with static analyses and potential applications in program understanding and optimization. In ACM [2481], pages 66–72. ISBN 1-58113-413-4.

LCCN ???? URL http://www.acm.
org/pubs/articles/proceedings/soft/
379605/p66-mock/p66-mock.pdf;
http://www.acm.org/pubs/citations/
proceedings/soft/379605/p66-mock/

. Supplement to ACM SIGPLAN Notices.

Liang:2001:EEF

[2470] Donglin Liang, Maikel Pennings, and Mary Jean Harrold. Extending and evaluating flow-insensitive and context-insensitive points-to analyses for Java. In ACM [2481], pages 73– 79. ISBN 1-58113-413-4. LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/soft/
379605/p73-liang/p73-liang.pdf;
 http://www.acm.org/pubs/citations/
proceedings/soft/379605/p73-liang/

. Supplement to ACM SIGPLAN Notices.

Holzmann:2001:ESV

[2471] Gerard J. Holzmann. Economics of software verification. In ACM [2481], pages 80–89. ISBN 1-58113-413-4.

LCCN ???? URL http://www.acm.
org/pubs/articles/proceedings/soft/
379605/p80-holzmann/p80-holzmann.
pdf;   http://www.acm.org/pubs/
citations/proceedings/soft/379605/
p80-holzmann/. Invited talk.

Webber:2001:WCI

[2472] Adam Brooks Webber. What is a class invariant? In ACM [2481], pages 86–89. ISBN 1-58113-413-4. LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/soft/
                                        379605/p86-webber/p86-webber.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/soft/379605/p86-webber/

. Supplement to ACM SIGPLAN Notices.

Flanagan:2001:DRC

[2473] Cormac Flanagan and Stephen N. Freund. Detecting race conditions in large programs. In ACM [2481], pages 90–96. ISBN 1-58113-413-4. LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/soft/
                                        379605/p90-flanagan/p90-flanagan.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/soft/379605/

p90-flanagan/. Supplement to ACM SIGPLAN Notices.

Ball:2001:BPS

[2474] Thomas Ball and Sriram K. Rajamani. Bebop: a path-sensitive interprocedural dataflow engine. In ACM [2481], pages 97–103. ISBN 1-58113-413-4.

                                        LCCN ???? URL http://www.acm.
                                        org/pubs/articles/proceedings/soft/
                                        379605/p97-ball/p97-ball.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/soft/379605/p97-ball/

. Supplement to ACM SIGPLAN Notices.

Anonymous:2001:AIb

[2475] Anonymous. Author index. In ACM [2481], page 104. ISBN 1-58113-413-4. LCCN ???? Supplement to ACM SIGPLAN Notices.

ACM:2000:AIP

[2476] ACM, editor. ASPLOS-IX Proceedings: The Ninth International Conference on Architectural Support for Programming Languages and Operating Systems, Cambridge, MA, November 12–15, 2000, volume 35(11) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, November 2000. ISBN 1-58113-317-0.

ACM:2000:COO

[2477] ACM, editor. Conference on Objectoriented programming, systems, languages, and applications: October 15– 19, 2000, Minneapolis, MN, USA, volume 35(10) of ACM SIGPLAN Notices. Addison-Wesley, Reading, MA, USA, 2000. CODEN SINODQ. ISBN 0-201-48565-6. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN ????

ACM:2000:PAS

[2478] ACM, editor. Proceedings of the ACM SIGPLAN ’00 conference on Programming language design and implementation: June 18–21, 2000, Vancouver, BC, Canada. ACM Press, New York, NY, USA, 2000. CODEN SINODQ. ISBN 1-58113-199-2. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN ???? ACM order number 548000.

ACM:2000:PFA

[2479] ACM, editor. Proceedings of the fifth ACM SIGPLAN international conference on Functional programming: September 18–21, 2000, Montréal, Canada, volume 35(9) of ACM SIG-PLAN Notices. ACM Press, New York, NY, USA, 2000. CODEN SINODQ.

ISBN 1-58113-202-6. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN ???? ACM order number 565000.

Sreedhar:2000:PAS

[2480] Vugranam C. Sreedhar and Jong-Deok Choi, editors. Proceedings of the ACM SIGPLAN Workshop on Dynamic and Adaptive Compilation and Optimization (Dynamo ’00). January 18, 2000, Boston, Massachusetts, volume 35(7) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, 2000. ISBN 1-58113-241-7. ACM order number 549000.

ACM:2001:ASS

[2481] ACM, editor. ACM SIGPLAN– SIGSOFT workshop on Program analysis for software tools and engineering: June 18–19, 2001, Snowbird, Utah, USA: PASTE’01. ACM Press, New York, NY, USA, 2001. ISBN 1-58113-413-4. LCCN ???? Supplement to ACM SIGPLAN Notices.

ACM:2001:PAS

[2482] ACM, editor. POPL 2001: The 28th ACM SIGPLAN-SIGACT Symposium on Principles of Programming Languages: January 17–19, 2001, The Royal Society, London, United Kingdom. ACM Press, New York, NY, USA, 2001. ISBN 1-58113-336-7. LCCN ???? ACM order number 549011.

ACM:2001:PEA

[2483] ACM, editor. Proceedings of the Eighth ACM SIGPLAN Symposium on Principles and Practices of Parallel Programming: June 18–20, 2001, Snowbird, Utah, USA (PPoPP’01), volume 36(7) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, July 2001. ISBN 1-58113-346-4. LCCN ????

ACM:2002:PAS

[2484] ACM, editor. Proceedings of the 2002 ACM SIGPLAN Workshop Partial Evaluation and Semantics-Based Program Manipulation (PEPM’02), volume 37(3) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, March 2002. ISBN 1-58113-455-X. LCCN ???? ACM order number 551021.

McKinley:2003:YAS

[2485] Kathryn S. McKinley, editor. 20 Years of the ACM/SIGPLAN Conference on Programming Language Design and Implementation (1979–1999): A Selection, volume 39(4) of ACM SIG-PLAN Notices. ACM Press, New York, NY, USA, 2003. ISBN 1-58113-623-4. LCCN ????